A Complete Bibliography of the Journal of Time Series Analysis

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

28 May 2024 Version 1.11

Title word cross-reference

(0, 1) [AD23]. (INAR(p)) [JGY91]. (SINAR(p)) [KT11]. 2 [ZM01]. 2 × 2 [ZC19b]. $36.99 [Hal14]. $71.46 [Omb13]. α [CT06b, GGSW20, GSTW20, LDH19]. AR(1) [AB86, AF91, HW95b, HCH00, HB05, KLM16, Lim92, Log04, PZC14, Pet86, PPS14, SS96b, Wei85, Zie99]. AR(2) [MA93]. AR(p) [KS08a, ZL12b]. ARCH(1) [Aud05]. ARCH(∞) [HP17]. ARCH(p) [KS08a]. ARCH(q) [WSS04, CPR18]. ARIMA(p, 1, q) [MN95]. ARIMA(p, d, q) [Rei94]. ARMA(1, 1) [Oke98]. ARMA(p, q) [Bar87, Che06, RB13]. b [HV08, HI15, ILT14]. β [LDH19]. Bl(p, 0, p, 1) [SR91]. Cogarch(p, q) [IMR18]. Ear(p) [BM91]. F [WS20]. I(0) [KLN04b]. I(1) [CL97, KLN04b, WR08]. I(2) [NV96]. INAR(1) [AOA87, AK10, BS15a, DO04, PK13]. INAR(p) [DVW08, RN12, SO05, ZWZ11]. ∞ [NN21]. k [Pro03, WCG98]. L [BCT15]. L^{p} [LXT20]. l_{1} [BCK23]. M [AD99, AC18, Ber07, CN86, Giu17, HT86, LPS23, LLG09, SL04b]. NEAR(2)

[Per04]. NLAR(p) [ZW08]. p [CWDL97, Hog19, KT11, LPS23, ZBD06]. q [HK14, NB83]. R [BCT15]. R^{2} [Bha93]. S [Sib01]. SαS [CH15]. T [Wes19a, KT16, MMOV24, SBS23, TCCG19]. U [VP12]. U_{p} [Dah85]. VARMA(p, q) [WJM11]. X_{t} = A_{t}X_{t−1} + ϵ_{t} [Pou88]. Z [XZ22].

* [AT86].

’-Approach [Hog19]. -Conditioned [NB83]. -dependence [HK14]. -divergence [CT06b]. -Estimates [CN86]. -Estimation [Ber07, Giu17, SL04b, Sib01, LLG09]. -estimator [LPS23]. -Estimators [AD99, AC18]. -Factor [WCG98]. -Order [KT11]. -out [Pro03]. -Penalized [BCK23]. -Smoothing [HT86]. -Stable [GGSW20, GSTW20]. -Statistic [Dah85]. -Statistics [BCT15, VP12]. -valued [AD23, XZ22].

/1 [DS91].

0 [Aue22]. 0172 [Cao19]. 0172-7397 [Cao19].

1 [McL17, Pou17, Wil16]. 1-4822-1959-X [Pou17]. 10.1111 [KPT21]. 10.1111/jtsa.12460 [KPT21]. 11 [Pfe94].

2 [Lat17, Omb13, Rao16b]. 2007 [Ano07]. 2021 [Tay22a]. 2022 [Tay23c]. 2023 [Tay24b]. 24.99£ [Kar16]. 2nd [Bos09, Bos16, Che09, Lu18].

35th [NH19].

4 [Hal14]. 40 [KPT21].

5 [McC15, Wan21]. 531 [Rao14]. 5th [Wil16].

6 [Bos16, Cha16a, Kil18, Leo13, Pou16]. 60.00£ [Kar16].

7 [Nea13, Tur18]. 7397 [Cao19].

8 [Rao17].

978 [Aue22, Bos16, Cha16a, Hal14, Kil18, Lat17, Leo13, McC15, McL17, Nea13, Omb13, Pou16, Rao16b, Rao17, Tur18, Wan21, Wil16]. 978-0-12-803768-3 [Rao16a]. 978-0-19-968366-6 [Cha16a]. 978-0-19-969560-7 [Nea13]. 978-0-367-22109-6 [Nun20]. 978-0-4716-8717-7 [Tur18]. 978-0-521-17561-6 [Leo13]. 978-0-691-16627-8 [Neš16]. 978-1-107-09733-9 [Kar16]. 978-1-107-48250-0 [Kar16]. 978-1-107-63002-4 [Hal14]. 978-1-118-32655-8 [Rao17]. 978-1-118-67502-1 [Wil16].

978-1-118-74495-6 [Bos16]. 978-1-118-74511-3 [Bos16]. 978-1-118-74515-1 [Bos16]. 978-1-119-11518-2 [Rao16b]. 978-1-138-30386-7 [Yao20]. 978-1-4200-9460-2 [Omb13]. 978-1-42001-006-0 [Qui15]. 978-1-4398-4940-8 [Zha13]. 978-1-4398-7651-0 [Aue22]. 978-1-466-51084-5 [Wan21]. 978-1-4665-0225-3 [Rao14]. 978-1-4665-7773-2 [Lat17]. 978-1-4665-8322-1 [Mcn15]. 978-1-4822-2500-6 [Pou16]. 978-1-4822-5383-2 [Lu18]. 978-1-4987-3422-6 [Kil18]. 978-1-58488-176-6 [Qui15]. 978-1-58488-650-1 [McL17]. 978-3-642-31741-5 [McC15]. 978-3-642-35511-0 [Ter14]. 978-3-642-35512-7 [Ter14].

abrupt [LK98]. Absolute [ZZL20, Li12b, WD10]. absolutely [DS04]. Academic [Rao16a]. Academy [Ano94]. access [Ano20g]. account [KH04a]. Accounting [HN80, Kil98]. Accumulated [Kab87, LB11]. Accuracy [Gor81]. Accurate [ZP20b]. Acknowledgement [HB94a]. Acronyms [Gra82]. Adaptive [AC18, Büh96, DP10, FL00, Hid92, RT17c, XHN17, FK99, May22, Sch98]. Additive [BT06, Kab94, Whi85, mWK96, Zha24, AHT13, PR03, Vog99, YHN99]. Adequacy [AY96]. Adjacent [HQ89]. Adjusted [New80, Pfe94]. Adjusting [MO02]. Adjustment [Aka80, Ish84, JS24, MP87, PD02, SS01, Jan05b, KH04a, Kil11, PR09, Rod13, SL00, Sol04]. ADL [LL10]. Admissibility [Hal94b]. Advanced [Rao05]. Advances [WH19]. advantages [PS07]. Affected [MP84]. affine [PF21]. affine-transformed [PF21]. Against [BK03, HIP87, KS05, LKB15, Mar20, RB92, Tay03, AN08, AES06, BP12, DK13, Hid07, Kap05, Xia01]. aggregate [TW02]. Aggregated [CZ19, GM15]. Aggregates [PV98]. Aggregating [BC97]. Aggregation [dBCRT19, Elt94, Eng84, HN93, HV99, KOV94, MW16, OV04, SF11, Sou07, SW86b, TC05b, AVF98, BS02, Has13, PPS14, Sbr11, Zaf07]. Ahead [Ton82, KGY18]. AIC [Fin85, HB90, Qui88, Wan93a]. air [LZZ22]. Akaike [De 98b, HH93a, HT93, Ioa11, Maï12, WL98]. Akaike-Type [HH93a]. Alan [Kil18]. Alarm [BHL90]. Alexander [Neš16, Zha13]. Alexandre [Pou17]. Alfredo [Pou17]. algebra [ZM06]. Algorithm [BC95, FRR17, IJ99, KT94b, KS08b, KT01, LL95, MS07, MB97, PBT00, Rig92, SS82, CF24, Cra03, FK99, JN14, Kil98, KM04, Mau02, SBS23]. Algorithms [AC93, De 91, Gae00, Hua90a, Zho92, BP11, CGM08, EMNR09, UT12]. Almost [DM96, DD15, GL21, DH13, Len16, LL06, Kar16]. Alternative [Car85, PF95, ZG88, Kap05, PW05, Xia01]. Alternatives [BK03, Mar20, RB92, AN08, DK13]. Ambiguities [Fin84]. Amplitude [Has82, TM93]. Amplitude-Modulated [TM93]. Anal [Ano21a, Fei20]. analogue [Hid07, SP12]. Analysis [BC95, Bos16, Car85, Cav14a, Cha95, CK22, Dah83, FT85, Fei20, Fra05, FM04, Gab88, GCK99, Gra82, GSO^{+}17, HHP84, HN80, KPT21, KKJ18, Kil18, Kra16, LL92, Led90, LB11, LT95, LO16, MT94a, MT94b, MMNT20, NBQ16, Nun20, Oza82, PZ04, PW89, Pos20, Pou88, Pou17, Pri80, Pri96, Rao10c, Rig92, Rig96, Sto87, Sto90, TK93, Ter85, TM93, Tsa88, TW89, VWR87, Wal87b, Wil16, Yak87, ZT94, Bos09, Bro07, DE07, FL00, HK17, Jan10, KL09, KFS02, KP13, KOW22, KXS^{+}12, Kur11, Li98, Li12b, Mar12, Mau11, NS03, PR98, Par13, Rao08, Rao12b, RT17b, RSW08, SdJ22, Seo24, SY11, Spe10, SO12, SR07, SP18, VPWD11, VVD18, Ano07, Cao19, Cox94, Tay22a, Tay23c, Ano97b, Ano99c, Ano02, Ano03, Ano04, Ano05b]. Analysis [Ano06a, Ano20g, Che09, HKK23, LT18, NH19, Tay20, Tay21b, Tay24b]. Analytic [PS99, AG24]. Analytical [Pes07, Sbr11, WNS22]. analyzing [EP17]. Andrews [HH81]. angle [BEvdW12]. Anniversary [NH19]. Announcement [ACN23, Ano86, Ano05e, Tay13b, Tay18a, Tay19, Tay20, Tay21a, Tay22a, Tay22b, Tay23c, Tay23b, Ano23a, Tay21b, Tay23a, Tay24b]. Antedependence [Fok10]. Antipersistent [BP07]. Application [BC95, GPH83, GJ02, HHP84, HN21, McL94, McL95, Mil19, PT81b, Rob87, SS95, SS96a, Tia88, AK10, Cra03, GA16, KL09, KPRN03, LLOS08, Len16, LL06, LZZ22, Mil10, MRT07, NSK^{+}11, NK24, PC23, PW05, TvV02, VN17, WWG09, WCK12, ZPZ21]. Applications [BCT15, GR81, Hal14, McC15, Pou17, Rao14, Rig96, Tha90, Yu07, BB07a, BDL08, Bos09, CP17, DdM13, EM08, FP12, GSO^{+}17, HKVW22, HST23, HWBD11, JT11, Kri09, Qia14, RMSF10, Leo13, Kar16, Nea13]. Applied [CR90, Lu18, Mcn15, MVS87, PZ17, ZC19a, PS99, Kil18]. Approach [ATT03, CGN15, CV06, Duo84, Eng84, GT19a, Gra95, Hog19, JC17, KP90, KJ85, LH83, LWL93, LT83, MW05, MP84, Nun20, PW84, Pri80, SS82, TK93, Tua86, mWK96, WTSL17, ABT18a, CSD12, DC01, DRY21, FNV08, GKL11, GKD21, Jen12, Kim15, KMX17, LZ18, LV00, LLT14, Lie12, Lie05, MLS97, PT23, Rao11, RDB14, Ste05, WS02, WLC12, WNS22, Zho12, Zhu13]. approaches [ZP20a]. Approximate [Abr87, And93b, Azz82, HR02, HR93, JO06, Tua92, Che06, GS13]. Approximating [Ali83, Fin85]. Approximation [IMR18, PT05, Pöt90, Str96, AK24, FK99, HV08, YLC21]. Approximations [Wah89]. AR-GARCH [LLZZ22, MS08b]. Arch [Wei84, ZLSY20, LL97, MY02, BB12, BM03, BM09, CT06b, Cli07, GLP10, GSS17, IP08, KFS02, Kim15, MO02, NS13, NN21]. ARCH-M [MO02]. ARCH/GARCH [GLP10]. Arctic [ZC19b]. ARFIMA [DGPHS19, Nan14, STY97]. ARIMA [BM04, KT01, LW91, PRR04, PP88, Pic90, PD02, SW86b, TV83, TvV02, Wri95, Yaj85]. ARIMA-Model-Based [PD02]. Arising [Cha15a]. Arma [LM88, Wei84, AG95, AB09, AD99, BL01, BC12, Bro95, BB99, BFK12, CG07, CT87a, DZ18, DdM04, EMNR09, Eng84, FHK20, Fin84, Fin16, FG04, Gae00, Gir07, Had04, HS05, HW89, HZZGH83, JWW99, Kab83, Kak96, Kan81, Kar01, KHS03, KT94c, LM04, LM08, Lou08, LB00, Ma02, MA24, ML83, McL93, NLL12, NR07, NP90, Peñ84, PRC03, PW84, PS95, RMT90, Sak91b, SL04a, SL04b, SF05, Str96, ST91, Swi90, Tan87, Tig85, TC07, Tua87, Tua88, Ver87, Vol12, Wal95, WS02, Wan08, Whi85, YB06, Yu07, ZG85, ZPZ21, Zhu13]. ARMA-GARCH [AB09, Zhu13]. ARMAX [Pos05]. array [PPS14]. arrays [LSSC16]. article [Sza23]. Aruma [HA93]. Arup [Yao20]. ARX [Duc05]. Aspects [BL21, JA81, MS92, PT86]. Assessing [KH99, MMT05, Psa08, BFK19]. Assessment [Zha04]. Asset [CHS17, SCW19, TY10]. assigned [Efr19]. assisted [CR99]. associate [Ano22g]. Associated [Cha91, Fin84, Wri95]. Assumptions [Psa08, ADL18]. Asymmetric [AV05, SLN99, Sol04, TZ22, BM04, HA21, LZ20, MKS22]. asymmetry [SY20]. Asymptotes [CA99]. Asymptotic [AT87b, AV93, AY96, BB07b, CL06, CT96, Cha87, Cho91, Dah85, Deo97, DS91, EF06, FRS11, HP17, Kab83, KT94a, Kak96, KLN04b, KS18, KS19, Kur21, LTT18, LP19, MPR91, MW22, Och83, Por87, Ray88, RT09, Sai83, Sai86, SH87, SW21, ST85, ST05, ST91, Tan87, TAM11, Tom87, TvV02, Ver87, Wal95, WS20, Yaj85, Zha92, ZLP19, ZG85, BGT21, DZ17, DVW08, Erc11, HV08, Ioa10, Joh03, Kak99a, KM04, KS23b, LP10a, LP04, PZC14, Tak24, Wal00]. Asymptotics [AG08, HB93, HB94a, HB05, Muk99, Nie15, HI15, XLT23]. at-most-m [HPW17]. ATSA [Gra82]. Augmentation [FS94]. Augmented [LTZ20]. Australian [Ano94]. Authors [Tay20, Tay21b, Tay22a, Tay23c, Tay24b]. Auto [Mur85]. Auto-Regressive [Mur85]. Autocorrelated [USMS83]. Autocorrelation [And92, Cho91, Fas00, HT99, Hid92, Kan87, KPS04, Li84, Peñ84, SS89, HK14, LT17, Mar99, MZ06, RSW08, Sch98]. Autocorrelations [AC96, BC01, EF06, HR93, LM94a, ML83, PS92, Pap94, SH87, SL04a, SL96b, Yaj85, BFZ02, Deb11, MJ12, PRW04]. Autocorrelograms [GJ02]. Autocovariance [BEvdW12, CS15, Deg87, LL05, MG00, MP18, MIN^{+}16, Yao20, ZS01, BB14, BLL09, LLS08, LLBM^{+}11, MP10, PPS20, WP21]. Autocovariances [AC96, Bat83, Hal95, Kee97, RG89, Tia88, AK24, BC02, BKM21, DdM13, LBV09, MJ12, VY16]. autodependogram [BPN12]. Autogressive [QN81]. Automatic [Cam87, CGM08, Hen01b, HB94b]. autopersistence [WL11]. autoregessive [JP99]. autorégressif [Mok87]. Autoregression [AM18, CP16b, Dav91, GP06, HH93a, HK86, Hua90a, Kav89, KP93, McL94, McL95, MT90, PX06, SP01, Tha90, XA99, BB07b, BM10, FM04, Kak99b, KLN04b, Lie12, LZ24, MZ06, SR07, TZ22, YHN99, Zho13]. Autoregressions [BF96, BDD95, CV06, DFR21, HLHT94, Kni87, LH96, LS06, MWM97, PT82, Pol94, Pop90, Qui88, Hil13, Kil98, LR02, LP14, MMM22, PH02, TP03]. Autoregressive [AB99, AT87a, AT87b, AD23, AOA87, AQL89, AH92, And89, AM80, And92, AV93, AK90b, AHS06, AHP17, Azz81, Bai93, BCCR19, BKS97, BM89, BMY99, Bha83, Bha86, Bha89, Bha93, BM81, BLL05, Bol88, Bos96, BR06, BLT92, CT86, CG19, Cha95, CL95a, Che95, Chi91, Cho91, Chu96, CT87b, DGJ06, DSW80, DH22, DS91, Duo84, FT85, Fei20, GT93, GS20, Gor18, GL19, GJ06, HO84, Hal94a, HK90, HK22, Hög86, HZF93, Hz92, HR93, HN93, HHI18, Hua90b, HT93, Huz81, Hyn93, IY03, JL83, Jas03, JGY91, Kab93, KH99, KMS15, KOD09, Kan87, KM90, Kni87, KP89, KP90, KP95, KJ85, KF92, Kun97, LL92, LH83, Li93, Lju88, Lüt85, MP90, MS92, MM93, MT94a, MA20, MW05, MMT98, MK15, MM91, Mil95, Nas93, NQ80]. Autoregressive [Och83, OT98, PS92, Pap05, PF95, Pau84, PT81a, Pem87, Pet89, Pic82, PP97, Por87, Pöt90, QN82, Rai96, Ray88, RBY92, RA92, RB22, Sai86, SS16, SS90, SH87, SH88a, SK96, Sch16, SPA20, Sha08, SY11, She88, ST04, Shi93, SS95, SL96b, SS96a, SPM19, SR17, Spa93a, Spa93b, SH88b, SHLL96, TY10, TP85, Tom87, Ton82, Tua84a, Tua84b, Tua86, Tua88, Tua92, TT82, Ula93, Vel94, Vil01, Wah89, Wan93b, mWK96, XZL20, YL20, YR95, YL91, Zha92, ZT97, vG99, AN08, ADL18, AAD22, AE06, And97, AMZ13, And08, AS21, BBC16, BG00, BHL09, BBK23, Bha97, Bla14, BPT02, BKM21, BFZ02, BMH08, BH08, CS08, CSD12, Cam04, CZ12, CL01, CD09, CD12, CLL14, CWDL97, CS11, DA14, DdM13, FF13, FK13, FP18, FR07, FKMN02, FB21]. autoregressive [GMRO11, GH03, GG07, GL20, GJ23, GB98, HLZ23, HP14, HL11b, Huz07, HS11, Ioa10, Ioa11, JR22, Joh03, JS24, KY09, KT11, KL09, KK12, KR13, KGY18, KP21, KR98, Lat98, LL12, LJ23, LB11, LLY14, LK21, Lie05, LP19, Lug06, Mau02, MS23, Mok87, MTW04, NLR16, NS03, PR98, Psa01, PS03, PS06b, SL00, SO97a, Sch98, SM13a, SL97, Shi98, SS98, SF98, SLL97, ST05, Swe22, TK08, TS14, TC13, Tri12, UD09, UT12, VADG04, Vid09, WL11, WP21, WP14, Wie13, WL98, WD10, Xia01, YP06, ZHHH22, ZB02, ZM06, ZC12, ZKP22, Zha24, ZJN24, ZBD06, ZXC22, ZB05]. Autoregressive-Moving [CT87b, JL83, KM90, Pic82, Mau02]. Average [AH92, AM80, And92, AV93, AK90b, Bai93, BKS97, BM89, BMY99, BH94, Bha83, Bha89, Bos96, Bre94, BLT92, CT96, CG19, Cha95, Chi91, Cho91, Chu96, CT87b, DJM86, DSW80, DS91, GZ15, Had95, HR93, HN93, JL83, KM90, KP89, KP90, KF92, Lju88, MS92, MA20, Nas93, PS92, Pap05, Pic82, Por87, Pöt90, RBY92, Sai86, She88, Shi93, SS95, SS96a, SM06, SHLL96, Tua84a, Tua84b, Tua86, Ula93, Vel94, Wah89, Wan93b, YR95, AMZ13, And08, BFK13, De 98a, DA14, FF13, GG07, Huz07, KM21, Li12a, LZ20, Mau02, Mon98, PR98, SS98, SF98, ST05, kTR98, TC13, WD10, XPZL10, ZL24a, ZXC22]. Averaged [Hen01a, Lob97, SH12]. Averages [AT86, CH15, BEvdW12, Chu12, MR18, MS00a, MS01].

B [RW17]. Backdating [Mar07a]. Backtesting [DP20, Hog19]. backward [SO97a]. Balanced [BW18]. Band [Gor81]. Band-Limited [Gor81]. Banded [MP10]. Bands [BW18, SS90, Tom87]. Bandwidth [BB87, FRP99, Hen01b, HI19, KC96, Sou07, Vel00, AO09, BB07b, RS17, Sko01]. barely [KP08, ZL12a]. Bartlett [BF97, CCY16, Dah85, FZ09, Lar98, PI22]. Based [BT94, BDH^{+}18, DN95, DL15, Dit00, EF06, Hal92, Hal95, Hal94b, HD96, JLMB20, KLM16, LTK07, Mar07a, MW16, PD02, Sch96, ST97, TCCG19, Vel94, VN00, WTSL17, YL20, AS24, And08, AO09, AV08, BKM21, Bro07, BH10, CK13, DHT14, DR11, HKV24, Kur11, LLMR08, LT03, Lie12, LLL22, Mau11, MMOV24, Nag03, Olm23, Pes07, PV15, RCLM^{+}11, SF98, Sko01, TM98, Tsa07, THN24, Wan24, WCK12, YLC21, ZT18, ZJ06, FA03]. basic [Bra21]. Basis [CN17]. Bayes [LH83]. Bayesian [Rao17, Aka80, Ano21a, ATT03, BKS97, CSD12, Cam04, CJ82, CL95a, CV06, CS84, CS87, DF11, DGPHS19, GT93, GH03, HCH00, HW99, Ish84, Jen04, KFS02, KC10, LSSC16, LS06, MS92, MFM05, MT94a, MW05, MS23, MWM97, NBQ16, PR98, Pet19, QR98, QS00, RT02, SM13a, SPM19, Spe10, SR07, Unn04, VADG04, Vil01, mWK96, XPZL10, YP06, ZP20a, Nea13]. Behavior [Kur21, ZLP19, BGT21, KS23b]. Behaviour [KLN04a, ST85, BB07b, HV08, PP16a]. Belzunce [Rao16a]. Beran [Ter14]. Berlin [McC15]. Bernoulli [RNI13]. Bernstein [Kak06]. Bertram [RTW14]. beta [BWH19, PI22]. beta-Bartlett [PI22]. betas [CS11]. Between [AM07, DF80, LX96, Mar95, New80, PS95, PR88, PM92, RS19, VY90, CL97, Ger21, GS20, GOP^{+}12, HS05, HL18, LNVK02, Li06, MGRM10, PI22, SP12, YD12]. Beveridge [GB99, NV96]. beyond [NN21]. Bhattacharjee [Yao20]. Bias [ANW93, AC18, GMP15, KS19, LH96, PR95, SCW19, Zha92, Li24, Sko01, STY97, Sto19, ZM06]. Bias-Corrected [PR95]. Biases [Pop90]. Bidimensional [GGSW20]. Bifrequency [Bri80]. bifurcating [ZB05]. big [BWH19]. Bilinear [ARS86, Cha91, DB98, GR81, Gab88, Gra95, Gue87, KBB90, Kum86, Li84, Liu89a, Liu89b, Qui82, RRW83, SR88, SR91, SR92, SRHZT83, ST87, Ton81, AR10, DB03, Kri09, LPZ15, TT97b, Wan05]. Binary [Li93, JR22, WL11, WC14]. Binomial [ABT18b, DH17, GL19, WP14, CF14, ZJ06, Zhu11]. Biological [FP12, SO12]. biomechanical [BB07a]. Bispectral [Gab88, LT95]. Bispectrum [Sak91a, TM98]. Bivariate [GSTW20, LJ23, PT86, Rig96, KP20, PBSO23]. Blangiardo [Rao17]. Blind [Li93]. Block [DD15, JPP15, McE18, Tew18, DLPP14, GLN15, KH04b, LLRR^{+}21, PS21]. Block-Nested [McE18]. Blockwise [Bra05]. Blowfly [Tsa88, Bri12]. BLUE [Tan87]. Boca [Lat17, Nun20, Yao20]. Book [Ano82, Ano95, Ano97a, Ano98a, Ano99a, Ano99b, Ano00a, Ano01a, Ano01b, Ano01c, Ano05a, Ano05d, Aue22, Bas98, Bos09, Bos10, Bos16, Cao19, Cha16a, Cha05b, Che09, Cox94, Dod05, Erc08, Fea05, Hal12, Hal14, Jan05a, Jan05b, Kar16, Kil18, Kok13, Lat17, Lay98, Leo13, Lu18, McC15, McL17, Mcn15, Mil04, Mil05, Nea13, Neš16, Nun20, Omb13, Pou05, Pou16, Pou17, Pri04, Qui05, Qui15, Rao04a, Rao04b, Rao05, Rao10a, Rao14, Rao16a, Rao16b, Rao17, Shu05, Ter05, Ter14, Tur98, Tur11, Tur12, Tur18, Wan21, Whi05, Wil16, Yao20, Zha13, Pap22]. Bootstrap [AO09, AC18, Aue22, BDD95, BW18, CRT15, CPR18, CP03, CR99, CGN15, DD15, DGP15, DLLN15, FK13, Kab93, KPT19, KPT21, KYP20, MP18, MK15, PSU08, PS92, PP99, PRR04, PT05, Psa01, RR09, Sme15, WW15, ZZL20, AHT13, BKM21, CG07, CPR15, CHLT15, CNR17, DLPP14, FP18, FKMN02, GLN15, HK17, Hid07, JPP15, Kak99b, Kil98, KH04b, LLT14, MP10, MR12, PS21, Par13, Pos08, Tew18]. Bootstrap-assisted [CR99]. Bootstrap-based [AO09]. Bootstrapping [AD99, BCT15, BM09, HK08, KF92, LX01, SP08, Swe03, HS11, WS02]. Bose [Yao20]. Both [Mar20]. Bound [Kab87]. Boundary [BP18, CNR17, CD09]. Bounded [Wal87b]. Bounds [CISG16, KJ85, AK24, LP04, Tak24]. Box [Wil16, HC04]. brain [GOP^{+}12]. Break [DLRY08, TSL08, CP16a, Fos13, HOS15, HLZ23, ILT14, NR11, San18, Sen07, SK09, Wri98]. Breaks [AES06, CISG16, Sch22, WT19, Wes19a, AH13, BEL06, BH01, BH03, CT06a, CHLT15, GA04, GA08, HM03, Kap05, KP10, Rod13, WE07]. Breaks* [HR04]. Brillinger [MVS87]. broad [PBSO23]. Broadband [HB01, NM11, Hur01]. Broken [NVS06]. Broken-Trend [NVS06]. Brownian [BB07a, Hua12, KM99, PT02, YFL^{+}14, ZT06]. Bubble [Kur21, KS23b]. Bubbles [AHL^{+}18, FFGM15, Tay24a]. Busetti [HM03]. Business [Jan05a].

C [Bos16, Hal14, Kil18, Wil16]. calculating [BC02]. Calculation [Cha15a]. California [NSK^{+}11]. Call [ACN23, Ano09a, Tay24a]. Cambridge [Hal14, Kar16, Leo13]. Cameletti [Rao17]. Can [KT01]. Canadian [Lim87, Lüt82]. Canonical [PM92, VY90, VWR87, DHJ12]. CAPM [Pra82]. Carlo [BM89, Che93, Dit00, Mil06, PT81b, VADG04]. CARMA [BL13, BL19, BL21]. Carolina [Rao16a]. Casals [Pou17]. Case [LT83, PR88, Bau05, Kei03]. cases [PZC14]. Catastrophes [SHLL96]. Categorical [FK87, DE07, KXS^{+}12, ME98]. Cauchy [AB86]. causal [BL21, GB98, Had04, JP99, Ken12, RB13, SPH18, Swe22]. Causality [ADD19, GH19a, HS05, Kan81, Lay84, YK06, Bra13, BS02, Hos01, Tau23, THN24]. CCA [Bau05]. Censored [CP16b, HVS15]. Central [BT94, Cha91, Kee97, KL10, Mor83, Sto85, Yaj89, RB13, WCK12]. Centred [Kni87]. certain [Deo97]. CH [Cav14a]. Chain [FT85, Fei20, VADG04, ZKG22]. Chains [FR97, BS15b, Bra21, FRZ01]. challenges [DMHF12]. Change [ADL18, BK19, CT20, DK17, Ger18, HH12, KH98b, LTK07, NLL12, PZ17, Rai96, SPH18, AMS^{+}17, BP12, BHLS11, Bet16, BFK19, CF24, Ger21, GKD21, HK08, JFML13, KL09, KH98a, KA07, Kur18, LJ23, NAJ12, PZC14, PY22, RT02, Roz01, WC10, Yam11, YD12, ZL12b]. Change-Point [ADL18, Ger18, KH98b, HH12, BFK19, HK08, RT02, WC10, YD12]. Change-Points [Rai96, Bet16, JFML13, Roz01]. Changepoints [FKK12]. Changes [IL19, TZ19, ZC19b, AMS^{+}17, DH22, DQ23, HPW17, LK98, Sha11, Tay05]. Changing [And93a, Joy87, Wie13, WX18]. Channel [SS90]. chaos [OJHO00]. Chaotic [HW95a, KTL00, LC03, LLS97]. Chapman [Rao14]. Characteristic [JLMB20, DdRSK21]. Characteristics [BHL90]. Characterization [SL04a, GA16]. CHARME [SFK10]. Chebyshev [KTL00]. check [Jin18]. Checking [Hok83, ML83, McL94, McL95, MV03, JW16, LL97, UD09]. Checks [PT86].

Cheryl [Bos16]. Chi [Kat12]. Chi-squared [Kat12]. Chichester [Rao17]. Choi [Kar16]. Choice [Duo84, Gao97, HB90, Vel00, AO09, YP06]. Choices [Fok11]. Chun [Wan21]. Chun-Hui [Wan21]. Circle [AT87a, AT87b]. Circular [BG20, DPT12, HP23, TKOP20]. Class [BP07, BB99, CPR18, GZ88, Has82, JS90, Bra13, Bro07, Cha05a, Cle01, Cra03, DLRY08, FZ09, Gir07, GV10, HWBD11, Len16, LLS97, MKS22, MTW04, RZ10, TC05a, Wan09, ZL12a]. Classes [Car85]. classical [Par13]. Classics [Rao16b]. Classification [Rao11, YR92]. Classifying [Pic90]. Climate [Par13, PZ17, ASM21]. closed [Had04, JT11, ZJN24]. Closely [HQ89]. Clustering [DFF23, WT19, CT10, FP12, PT23, RB22, VP12]. Co [AT99, CRT15, GH19a, CT06a, Kil11]. Co- [AT99]. Co-Integration [CRT15, CT06a, Kil11]. Coefficient [AHS06, DC23, FT85, Fei20, GKY18, GL94, HB05, LPPS20, LH83, NQ80, QN81, Shi88, Wal87a, Wei85, Zie99, BHL09, CGP22, CL01, FK13, HP14, KL09, LJ23, Lie12, LP14, PPS14, Sch98, SY11, ZWZ11, ZBD06]. Coefficients [Bra16, Huz81, LL05, LTZ20, MB97, QN82, Ray88, UH95, BLL09, FL04, Hoy24, JS24, Kri22, MRT07, Nan14, TS14, Wan08]. Coherence [GOV19]. coherency [SH12]. Cohn [BC95]. Cointegrated [BSS17, Cha19, Che09, Dit04, Hoy20, JN19, Mar00, YK06, JPP15, Joh03, JS24, KPY22, PRC03, Seo24, Swe22]. Cointegrating [BK07, Mil10, Mil19, Sch22, TSL08, WW17, Kur11, KA07, SL00, Sun14]. Cointegration [AM07, BiS17, Cub95, Dav91, EP94, Has01, HI19, JN19, Joy92, Lev02, MW16, BDM98, BH13, dBCCO22, DHT14, Dit00, FL04, GM15, HMV08, HL11a, HI15, IC05, KA08, LNVK02, LL10, MLS97, PLNL22, Pes07, Vel03, WE07]. combination [Sol04]. Combined [OT98]. Combining [BH13, BFK19]. Comment [CS87]. comments [BH03, DS04]. commodity [KPRN03]. Common [BiS17, EP94, Wes19a]. Comparative [Car85, CG82, Kur11, Lim87, SO97b]. Compare [PT81b]. Comparing [Bau05, CD86, STY97, DL15, ZT18]. Comparison [ADSS18, Bri80, Elt94, Hög86, LLS02, Lüt85, PSU08, Pes07]. comparisons [MP16]. complete [AV08, DRY21]. Complex [AT87a, WS11]. Complexity [SR17]. Component [Elt94, KP93, MMNT20, RT17c, SH90, AE06, BM04, HW99, KP10, Seo24, Wal03]. Components [BHL90, CRT15, Cha16a, Cub95, FR83, Hot89, Joy92, KKJ18, MMNT20, Sin93, AV08, HLX10, Iac10, Iri02, ME98, PD12, Pou16]. Composite [BC97, LO16, SS89, ST03, PK13, SBS23]. Compound [DGJ06]. Computation [BM04, Bos96, CS84, McE18, PM92, SL04a, NN21, ZL24a]. Computational [CS87, HH93c]. Computationally [Ish84, SH09, Wes13]. Computer [ZM06]. computing [KM04, Vid09]. concentrations [DA14]. Concepts [Neš16, KP21]. Concerning [Wal95, Wal00]. Condition [HL06, Liu89b, DCCL03, HLT10, Sel10]. Conditional [AM18, AHP17, BH92, Bol88, Cha16b, Fin16, FRZ01, Gra95, HR02, Huz88, IY03, KOW22, KW12, LM94a, MS08a, Nie15, TT99, AAD22, AF22, AD23, AS21, Bla14, BWH19, CH14, GLP10, Hen01a, HL11a, Hon97, HL11b, LLOS08, LPS23, Lug06, RWZ20, VW15, WS02, Wes13, WR08]. Conditional-Sum-of-Squares [Nie15]. Conditionally [KH01, MMNT20, CS11, EM08]. Conditioned [NB83]. Conditioning [FM85]. Conditions [AK90a, Kan81, Mar92a, SS96c, Spa93a, Spa93b, CS11, DRS21, FKD10]. Confidence [BM81, BW18, CGN15, HW95b, KJ85, Kur18, SS90, Tom87, HK08, Kak99b, LP19, NSL07]. Connections [HB90]. Consinusoids [Has82]. Consistency [DS91, FRR17, JvS95, KK20, Lob97, Mas96, Pet86, ST97, CNR17]. Consistent [BH08, DGH06, EFT16, JP99, Kim91, KR98, MPR91, WW17, WP21, ZG88]. Constancy [Kur17, HM13, Ken12]. Constant [DC23, IP20]. Constraints [Arb08, BLL05, HT99, Gir07]. Constructing [Mar07b]. Contemporaneous [Mil19, PPS14, Zaf07, Bra13, PBSO23, Sun14]. context [BT13]. Continuous [Cha19, CT20, CT87a, Com96, Fin16, GS20, HW89, Hoy20, Hyn93, IP20, LM95, Sin93, Str96, TC13, Tho19, YL20, BP03, BFK12, BFK13, Cha99, DE07, Erc11, FHK20, FF13, GA16, Hoy24, Huz07, KM03, Maz12, SO97b, SCW19, SR07, TC05a, Vij06, Wal00]. continuous-discrete [Maz12]. Continuous-Parameter [HW89]. Continuous-Time [Fin16, Hyn93, Sin93, TC13, BFK12, BFK13, FHK20, FF13, GA16, Huz07, KM03, SCW19, SR07, TC05a]. contrast [CF24]. Contributions [MdS89]. control [DdRSK21, FK99, Jan10, Wil16]. Controlling [PD02]. convenient [Wes13]. Convergence [Kni87, KST95b, LDH19, FB13, Hua12, Jir16, OV04, PS99, RB13, San17]. Convolution [JT11]. Convolution-closed [JT11]. Conway [MA20]. Copula [BDH^{+}18, WTSL17, BS15b, KL13, NLL12]. Copula-Based [BDH^{+}18, WTSL17]. Copulas [Mcn15]. Corrected [BCK23, HT93, KS19, PR95, WL98]. Correcting [Li24]. Correction [Ano87, Ano92, Ano21a, AC18, Dit04, Fei20, GZW94, Spa93a, BDM98, CCY16, EM02, FL04, GMP15, Ioa11, Kil11, SAZ13, Sko01, SCW19]. corrections [Lar98]. Correlated [ACL01, BiS17, DM96, Fra84, KKJ18, LL05, MK93, Tia88, Wes19a, BLL09, GA16, KT16, LLS08, Len16, RGLA11, SA07]. Correlation [And93b, BB95, Bat88, Bha83, BHL94, Bol88, HS05, HG91, JLMB20, LS03b, PM92, BCFFT17, DGP15, Duc05, HST23, JW16, KF08, Li98, PZ04, PRC03, PBSO23, Zho12]. Correlational [KTL00]. Correlations [Cha87, MPR91, NB83, VY90, DHJ12, MY04, MS01]. Correlogram [Hal94b, PP12]. Correlogram-Based [Hal94b]. Corrigendum [Ano97c, Ano99d, Ano06b, Ano09b, Ano10, BMH09, JMP16, KPT21, Sza23, Tak24]. Cosinusoids [Wan93a]. Cosmological [Leo13]. Count [AF16, BS15a, MA20, SPA20, FM04, JT11, KL23, SM13b, ZJ06]. Counting [FHW94, RNI13]. Counts [MS08a, FKK12, JT03]. Course [Aue22]. Covariance [Had95, Hir06, Kro82, LH96, McE18, MPR91, Ott88, RS19, XL02, Yao20, ZC19a, DC01, Had04, HI15, Jen12, KL11, Li24, NHCLP08, RT17a, RS17, Tri11, WJM11]. Covariance-Stationary [Ott88]. Covariances [KT94a, Por87, Wal95, AK24, Kak99a, LR21]. covariate [KXS^{+}12]. Covariates [DC23, Fos13, KCW22, Wes13, ZKG22]. cover [Qui15]. Covered [LL95]. Cox [SB19]. Crashes [Tay24a]. CRC [Aue22, Kil18, Lat17, Lu18, McL17, Mcn15, Omb13, Pou16, Pou17, Qui15, Rao14, Zha13]. Cressie [Rao16b]. Crisis [WW17]. Criteria [AY96, De 01, Lüt85, TV83, HOS15, Maï12, PSSS09]. Criterion [BC97, Bha86, De 98b, HH93a, HP92, HT93, ST04, Ioa11, TH12, WL98]. Critical [CL95b]. Cross [De 01, Elt94, GGSW20, HB90, Kav89, Lob97, RS19, Vel00, Wes19a, AK24, Ioa22, LR21, Li24, MS01, Sko01]. Cross-Correlated [Wes19a]. cross-correlations [MS01]. Cross-Covariance [RS19, Li24]. cross-covariances [AK24, LR21]. Cross-Dependence [GGSW20]. Cross-Periodogram [Lob97]. Cross-Sectional [Elt94]. cross-spectrum [Ioa22]. Cross-Validation [Kav89, De 01, Vel00, Sko01]. Cross-Validatory [HB90]. Crossing [GT19b, HK90, CWDL97, Mar12]. Crossings [FHW94, Ked87, Mar99]. Cumulant [FP16, Kim91]. Cumulants [DO04, SR88, SR91, SO05]. cumulated [LLMR08]. cumulative [BFK19]. curve [SPH18]. curves [HKVW22]. CUSUM [LTK07, Wie13]. CUSUM-type [Wie13]. CVAR [JN18]. Cycle [Oza82]. Cycles [GA01, HS98, Jan05a, Bou08, CCGA13, Hid07, LP10b, MTW04, PR10, Tri06]. Cyclical [Kun97, Ton82, AR00, Art02].

D [Fei20, JMP16, Mar95, ZM01]. DAF [Maz12]. DAF-filters [Maz12]. Daily [TCCG19]. Damien [Nea13]. Data [ASJLZ19, BCT15, Bha82, BHL94, Bra16, BN92, Cao19, CZ19, Cha19, CWD00, CP16b, CK22, Dah83, Dun81, Fas00, Fok10, FS94, GL21, GT19a, GM85, GH19b, Hoy20, IJ99, JvS95, JO06, Kim16, LPPS20, LT95, Lim87, Lüt82, Ma02, Mar07a, MA20, MW16, Mil19, MS00b, Nun20, Rao12a, SPA20, SS96b, Sin93, Tho19, Ton82, Wes19a, ASM21, AK10, BJR17, BWH19, CCGA13, CPR15, Cha15b, Che06, Cle01, Efr19, FL00, FM04, HH12, HVS15, KOD09, KL09, KC11, KH98a, KLN04b, KH04b, KS23a, Kle08, LLL22, Mar99, MMOV24, Rao13, RT17b, RWZ20, RGLA11, Roz01, SAZ13, SS98, Sko01, SM13b, TKOP20, Tew18, VN17, Wan21, Wan24, ZJ06, Rao16b, Omb13]. Data-Dependent [BN92]. Data-Driven [ASJLZ19, KOD09]. date [Kur18, KS23b, Wri98]. David [Nea13, Nun20, Rao14]. Davies [Cra03]. Davies-Harte [Cra03]. Davis [Lat17]. DEA [CGN15]. Decay [McL98, Cha16b]. decaying [Deb11]. declining [TT97a]. Decomposition [CS84, CS87, OJ03, ZC19a, GB99]. decompositions [DP10]. Deconvolution [Li93, PS89, PT02, Com04, GJ23, LSSC16]. decreasing [May22]. Default [GH03]. defining [Hos01]. Definitions [AT99, KP15]. Deflation [BS15a]. deformation [Vij06, WWG09]. Degree [CAP94, Jan82]. Delay [DF80, HT88, Kur21, LT83]. Dellaportas [Nea13]. demand [RSVM15]. Densities [Abr87, BDH^{+}18, CD86, HH93c, PR88, Pöt90, ASK15, Jin18, Kom99, LL18]. Density [CMK05, CR90, FRP99, GM85, Kim91, KLN18, Kle08, KST95b, LQ20, Neu96, NP96, Rob87, Sak91b, Tom87, ZP20b, BP03, Dat97, Efr14, Efr19, Efr20, GPRV00, HV08, HK22, HT10, Kak06, KM99, Len16, LLS97, MS00b, MRT07, PP12, PP16a, SA07, WP14, ZKP22]. density-dependent [WP14]. departures [NK24]. Dependence [AD04, APH86, Ber07, BSG18, CT92, Cox91, Duf81, Ger18, GRS97, GJ01, GMR04, GGSW20, GSTW20, HW95a, HIP87, IL19, KLN18, LKB15, SBLS07, VDO95, WS20, AVF98, AG16, BP12, BT13, BCFFT17, CF24, GAHT01, Ger21, GOP^{+}12, HK14, Hid07, JT03, KP20, LRS21, LOS12, LG11, MORS21, MS00b, NKC15, Nie05, RT22, Sha11, SM13b, TT97a, VVD18, ZZL14, Zho12, Mcn15]. dependences [BPN12]. dependencies [SK09]. Dependent [BG20, BCK23, BN92, CT01, CN86, DLLN15, GOV19, GH19b, HHP84, Hid97, LDN19, McL17, PC05, Pou88, Pri80, Pri96, SSW15, TP85, TT82, XX18, Yaj89, BDL08, Bet16, Bra11, Bra22, CPR15, DHT14, HVS15, KM03, KP15, KH98a, KW04, KXS^{+}12, LLBM^{+}11, Li14, MA24, Mar07b, MMM22, NSL07, Psa01, PV15, Rao08, RT02, RNI13, Sko01, Sta20, ST05, Tew18, Wan24, WP14]. deposition [DF11]. Depth [KLM16]. derivation [ZM06]. Description [Hua90b, NSK^{+}11]. Design [NHCLP08, WC10]. Detect [TZ19, Tay05]. Detecting [DQ23, Hog18, HL11b, HPW17, LT92, Tay24a, ZG88, KYP20, TD11, WWW12]. Detection [AY88, BO05, BT06, IL19, Kab94, Ked87, Lay84, Mar99, SPM19, ADL18, AMS^{+}17, AE06, BBC16, BFK19, CF24, DLRY08, HH12, Lou08, NLL12, Rao12b, RT02, SPH18]. Determination [CRT15, LX96, Pau84, Qui88, Sak93, Tia88, TP85, DH13, LLY14, PS03, PS06b]. determine [KP10, Sme15]. Determining [BB87, Cav14b, HG91, Jan82, KH94, KR13, KP93, LK21]. Determinism [Bat83]. Deterministic [BHL90, CRT15, CT20, DN95, HS98, SH90, BGT21, HLX10, Iac10, McC13, YLC12]. Detrending [Aok91, Cha15a, Wes13]. Developments [Kar16, Tay24a, CPR15, Cox94]. Deviation [ZZL20, WD10, Yab12]. Deviations [Dun81, Li12b, Sto19]. device [BPN12]. Diagnosing [TT99]. Diagnostic [APH86, HP92, Hok83, LL97, ML83, McL94, McL95, MV03, PT86, UD09]. Diagnostics [GCK99, Led90, DZQ10, Pro03]. Diagonal [KBB90, MW22]. Dickey [FA03, AN92, CHLT15, Ioa10, KLN04a, LKN05, OT98]. Difference [ANW93, AN94, AP95, DO04, HP92, LT83, Rei94, SR88, SR91, SO05, CK13, CF98, JN14, Tsa07, WH11]. difference-based [Tsa07]. Difference-Delay [LT83]. Differenced [PP88]. Differences [Hog18]. Differencing [CAP94, GJ80, Jan82, KPT19, KPT21, Lüt82, SL96a]. Different [Gue87, dBCCO22, GB99, RMSF10]. Differential [RMT90, Zha13, Cle01, HLM03, Sim08]. Diffuse [AK90a, Mar92a, SS96c, FKD10, KD03]. Diffusion [JO06, CW99, SHL22]. Dimension [FHW94, LK21, PS06b]. Dimensional [BCK23, GT19a, GJ01, LHR82, XX18, BB14, CLY17, Cle01, GP02, HNT23, KP21, LTT18, MMM22, PA23, RCLM^{+}11, TNH23, XCG24]. dimensions [KM09]. Dimitris [Aue22]. direct [LNVK02]. Directed [BBK23]. Direction [BG20, ZT97]. disaggregate [GAP09]. Disaggregation [HV99, SW86a]. discontinuous [Arv14]. discount [Chu12]. Discrete [ACL01, Bro95, CT20, CT87a, Dic82, GS20, HW89, IMR18, JL83, JO06, Lat17, MMT05, ST85, Sto87, Tho19, Ton81, TC05b, VDO95, ASK15, Cle01, DR11, Huz07, JM04, Kei03, Maz12, NHCLP08, Sim08, SA07, TC13, Wal00, McC15]. Discrete-Parameter [HW89]. Discrete-Time [ACL01, Bro95, Dic82, IMR18, TC05b, Huz07, NHCLP08, SA07, Wal00]. Discrete-Valued [Lat17, Sto87]. Discriminant [DN99, Kra16, ZT94]. Discriminants [DNL81]. Discriminating [PR88, YD12]. Discrimination [Ala89, GQ17, Ger21]. Dispersed [SPA20]. Dispersion [SF05]. displacement [DZ17]. Distance [JLMB20, Pic90, AK24, DFF23, Zho12]. distances [DKV11]. Distinguished [Ano94, Tay20, Tay21b, Tay22a, Tay23c, Tay24b]. distortion [KA08, KT10]. distortions [LS03a]. distributed [AR10, Kle08]. Distribution [Ali83, Ber01, Cho91, Dah85, Del96, Fas00, GLML16, GGSW20, KST95a, KS19, LC03, Log04, Mar92b, Mic20, Och83, SL96b, SRHZT83, Spa93a, Spa93b, TCCG19, Tua92, Vel94, ZGH80, BF10, BWH19, CH14, DZ17, EFT16, HL17, HK22, HKV24, Jir16, KGY18, MMOV24, Nag03, Shi98, Wal00, WJM11, Yab12]. Distributional [LM94b, PS01]. Distributions [AT87a, AY96, AES06, CL06, Cha15a, Cha87, DG20, DdM13, Fin16, GMLS15, MT90, PT81a, Pet89, Ver87, KS18, LD04]. disturbance [Che06]. Disturbances [CT01, USMS83, Wri98]. Disturbed [Shi93]. divergence [CT06b]. Divisible [GMLS15]. Do [AT86]. DOI [KPT21]. Domain [Cha19, Cub95, Pra82, SR92, Tua86, Won97, BJR17, CR99, FP12, Has13, Jen12, Jin18, Kak13, KM03, Lev02, RDB14, RT17b, Tau23, Wan16, Yua00a]. Domenico [Leo13]. dominant [LZZ22]. dominating [BG00]. Double [CGN15, SHL22, CLL14, GL20, HL02, LZ24, TZ22]. Doubly [MG93, Pou86, Tjø86]. Douc [Rao14]. Douglas [Bos16]. Doukhan [Pap22]. Drift [CHS17, Ste05]. Driven [ASJLZ19, Azz82, DH17, FR97, Gor18, DRS21, FRZ01, KOD09, MS00b, Sim08, WC14, WHY22]. DSGE [Ano21a, Pet19]. Dual [GOV19]. Dual-Frequency [GOV19]. Duration [GG08, AAD22, HL11b]. Durations [GJ02]. Durbin [Ali83, KT01]. Durbin-Watson [Ali83]. during [GOP^{+}12]. Dynamic [AD84, Bra16, DF11, FM85, FS94, GJ23, Guo03, HN21, NBQ16, OJ03, Ott88, PW89, SBLS07, FZ23, GBY17, KM09, LOS12, PS21, TJL23, ZP20a, Cao19, Hal14]. Dynamical [BLL05, Sin93, GH11]. Dynamics [GS20, GJ06, IP08, KWPV12, San18, WLC12].

e-book [Bos16, Qui15]. Early [Ano05e]. Ecology [LT83]. Econometric [CZ19, Fra05]. Econometrics [Kok12, Mil06, Cha16a]. Economic [Hal12, Hal13, Hal14, Jan05a, MT94b, Wal87b, Pro23]. EDA [Bri12]. Edgeworth [Kak99b, Tan84]. Edited [Zha13, Cha16a, Lat17]. Edition [Bos16, Kil18, Lu18, Rao16b, Tur18, Wil16, Neš16, Bos09, Che09]. Editor [Cox94]. Editorial [Ano23a, CH11, LPR18, RW17, SO12, Tay13a, Tay13b, Tay18a, Tay18b, Tay18c, Tay19, Tay20, Tay21a, Tay21b, Tay22a, Tay22b, Tay23c, Tay23a, Tay23b, Tay24b]. editors [Ano22g, LT18, NH19]. eds [Bos16, Kil18, Nea13]. Edward [Pri94, Rob94]. Effect [Has01, Has13, HN93, Mar20, Fan05, Hos01, WCK12, YP06]. Effects [BiS17, CPR07, EV19, Kar01, Wes19b, CT10, GBY17, Sel10]. efficacy [HLM03]. Efficiency [CS87, EF06, Kab83, KT94a, Pra82, Sai83, ST88, Wal95, ZG85, Kak99a, LL12]. Efficient [BFZ02, Efr14, EMNR09, GLL06, GS21, HR88, Ish84, KOZ12, KA07, Lee16, LL06, MTJ14, NN21, PC05, Sch22, TS14, ZP20a, DVW08, HC00, KGY18, SH09, Wes13]. EGARCH [AA20, HL18, KS05]. EGARCH-M [HL18]. EGB2 [CH14]. Elasticity [IP20]. electrical [CT10]. Elements [Whi85]. Elimination [Hos01]. Elliott [Kil18]. elliptical [LP10b]. Elsevier [Rao16a]. EM-algorithm [SBS23]. embeddability [TC13]. Embeddable [BB99]. Embedding [Bro95, BL19, CT87a, HW89, Huz07]. embeddings [TJL23]. Embrechts [Neš16]. emissions [WCK12]. Empirical [AN92, BBC16, CR90, DH13, DLLN15, GLP10, JLMB20, LDN19, LH83, NSL07, TP82, Yau12, ZWZ11, ZPZ21, AK24, Bra05, CP17, CCY16, CH21, DdRSK21, HKV24, Jir16, Kak13, LLOS08, Tew18, YL22]. end [GKD21, HKV24, Kur18]. endogeneity [Sun14]. Endogenous [Kur17, Kil98]. Energy [LHR82, DFF23]. engineering [Rao11, Ter11]. Entropy [Pol94, SSX18, BLL09, BI12, Bra05, Gir07]. Envelope [Mar20]. Enveloping [KXS^{+}12]. environment [NLR16]. environmental [CH11, HWBD11, Rao12b]. epochs [RMSF10]. Equality [BG95, CF98, HR15, JW16, Jin18, LL18, LBV09, PPS20]. Equation [Gor18, IP08, LT83, BDM98, Cle01, WH11]. Equations [DO04, Qui82, SR88, SR91, SO05, TA88, BM03, CADF11, RY23, Sim08, TP03, Vol12, Zha13]. Equations* [HLM03]. Equidispersed [MA20]. equivalence [Ioa10]. Ergodic [sC88]. ergodicité [Mok87]. Ergodicity [ARS86, CG19, MG93, AF22, Kri09, LPS23, Lie05, Mok87, SFK10]. Eric [Rao14]. Erratum [Ano05c, Ano05d]. Error [AES06, BDM98, Bha93, Dit04, HH93a, IY03, Kab87, KH99, KS19, LR88, Li20, LP04, NP96, PS06a, Ray88, Sib01, Tak24, UH95, AD23, BGT21, BM04, EFT16, EM02, FL04, HDB98, Kil11, KLN04b, QS00, SAZ13, XCG24, YLC12]. Error-correction [BDM98]. Errors [AD84, ADD19, APH86, BCK23, Cha95, CN86, DH98, FR83, HH05, HZF93, JWW99, KK20, LPS99, RT17c, Sin93, TZ02, Wei84, XX18, CL97, CP16a, Deo97, FB13, FLL13, FR07, Ing01, KH04a, KT16, Kat12, KS18, KP08, LL12, LB11, LPS23, LLZZ22, LL97, Lub99, MA24, Psa01, Sch98, SL04b, Sta20, TD11, ZXC22, ZB05]. Errors-In-Variables [AD84, ADD19]. Escobar [Tur18]. Established [DN95]. Establishes [Ano94]. ESTAR [HMS13]. Estimability [Sel10]. Estimate [AF91, BB87, HH93a, HB90, Kul85, PV98, SS90, Tua87, ZW94, BDL08, Bha97, Gao97, KM03, Sko01]. Estimated [AT86, CD86, FR83, PT05, SH88a, Wat85, DGP15, DP20, HOS15, Joh03, LC03, QS00]. Estimates [AH92, AV05, BL01, Bha83, BM81, CN86, DS91, Gor81, Hyn93, Kni87, MP90, ST97, TP85, Cle01, DG98, DH22, Duc05, FF13, FG04, HD99, MP10, MY02, TvV02, WP21]. Estimating [BG20, BM03, CMK05, CT86, CL95b, HH93c, IY03, JC17, KM21, LPPS20, LHR82, Lüt85, MP18, NP96, Pfe94, Qui89, SH90, TA88, mWK96, Yam11, ZC19b, BFZ02, DKV11, FZ23, HLM03, Kom99, Sou07]. Estimation [AC93, AG95, AM05, Ano21a, AHS06, Bai94, BC01, BKS97, BM13, Bat88, BO05, BW00, BMY99, BT94, BG00, Ber07, Ber01, BHL09, Bha83, Bha89, Bha93, BB14, BM91, BLL05, Bri80, Büh96, BN92, CJ82, Cha19, CR90, CAP94, CWD00, Che91, Chi91, CT87b, Com96, DZ18, Deg87, DM96, Dun81, DH17, FRR17, FM96, Fer90, FHW94, FRS11, GR81, GL21, GBY17, Ger18, GPH83, GM85, GRS97, GM23, Giu17, GLML16, Gra95, Gri91, HT88, HH93b, HR88, HLZ23, HD96, Hid92, Hid97, HS04, HLHT94, HZZGH83, HZF93, HKK15, HHI18, HA93, Hua96, HB94b, HR95, Huz81, HB05, IMR18, IP20, JLMB20, JO06, KE88, Kav89, Kav93, KT94b, KBB90, Kim91, KLN18, KST95a, KST95b, KP89, KP90, KH98b, LH83, Li93, Li20, Lii85, LM95]. Estimation [LW19, LOS12, MG00, Ma02, MK93, May22, McC13, MM93, MPR91, MMT98, MS07, Mil95, MB97, MW97, Mur85, Neu96, NQ80, OM17, PC05, PS89, Pet19, PW84, PR95, PS95, Pos05, Pöt90, Pou89, QN81, RG89, RA92, Rei94, Rob87, RT17c, SSX18, SM13a, Sch22, SAZ13, SR92, Sha08, She87, Shi93, SS95, SS96a, SL04b, Sib01, SF93, SF05, SHL22, TNH23, TA88, Tua84a, Tua88, USMS83, UH95, VN17, Wah89, WSS04, Wes19b, Whi05, Wil17, YL22, YB06, YR95, ZM01, ZLSY20, ZL24a, ZG85, Zho92, ZW08, ZZL20, ZP20b, vS94, AG16, AG24, AB09, And08, AS21, BBC16, BDS12, BP03, BB12, BFK13, BMH08, BH08, CSD12, CPR07, CT06b, CL01, CLL14, DB03, DGH06, DC01, Dat97, DVW08, Efr14, Efr19, Efr20, EFT16, EK13, Fan05]. estimation [FHK20, FL04, FNV08, GAHT01, GLL06, GL20, GP02, HPY02, Hen01a, HL11a, HP14, HCT04, HI15, HB01, Hur01, Iac10, Iri02, JP99, KH04a, Kak06, KM09, KC11, KHS03, KL11, KL13, Kle08, KGY18, KP21, KR98, KW12, KOZ12, KA07, LM00, LJ23, LL06, LLMR08, LLBM^{+}11, LT03, LRS21, LLG09, Lug06, MMM22, Mon98, MRT07, NM11, Nie05, Nie11, OJHO00, PP12, PD12, PK13, Per04, PY22, PF21, Rao11, RDB14, RZ10, RS17, RB13, SY11, SF98, SO97b, SCW19, ST05, Sun14, TS14, Tri11, TC05a, Vel99, Vel03, VY16, Wal03, Wan05, WC10, Whi02, WD10, YFL^{+}14, ZC12, ZZL14, ZW12, ZB05, ZT06, Zhu13]. Estimator [ANW93, Cam87, GT19b, Hir06, KK20, KJ85, LH96, MM91, MVS87, Nie15, Och83, Qui00, SK96, ST91, Wan93a, XL02, Zie99, Arv14, Bar24, BM09, CF98, Erc11, HDB98, HS11, Ioa11, Ioa22, LR21, LPS23, Li24, MS00b, Sch98, SH12, TT97a]. Estimators [AT87a, AD99, AC18, BiS17, FRP99, Hal92, Hal95, Has93b, Kab83, Kak96, KS19, LDH19, PF95, Pet86, Pfe94, Pop90, RBY92, Rob83, Sai86, Tua92, Wri95, Zha92, BB07b, BM10, Bra11, Cha05a, HV08, Kak99b, KL10, KS18, KS23b, LL12, MTJ14, NHCLP08, Olm23, PP16a, Rao18, RT09, WHY22, XLT23, ZM06]. etext [Bos16]. Euler [Vij06]. Evaluating [Cli07, Smi08]. Evaluation [LB00, MT90, GB06]. Event [ATT03]. events [GKL11]. Evidence [AN92]. Evolutionary [Mél85, MdS89]. ex [MW22]. Exact [And93b, Fin85, Ma02, Mil95, MA93, NP96, Pem87, Shi88, Tho19, Tua87, vG99, GdSF13, JS24, Lug06, Mau02, Sim08, TP03, ZS17]. Exactly [VY16]. Exactly/ [VY16]. Examination [LKN05]. example [Arv14]. Examples [Mél85, Rao14]. Excel(R) [Mil06]. exchange [LG11]. exciting [CS08]. exhaustion [TY10]. Existence [BLT92, sC88, Cli07, Lat98, Liu89a, Liu89b, Qui82, RRW83]. Exogenous [Bau05, ZKG22]. Expansions [Cha87, Och83, Tan84, Kak99b, LP04, Olm23, Tak24]. Expectation [AC93]. Expectation-Maximization [AC93]. expectations [JS24]. Experiences [MVS87]. experiments [BEvdW12, Bri12, DZ17]. Explicit [AF91]. Exploring [GOP^{+}12]. Explosion [KLM16]. Explosive [AM18, AHL^{+}18, HB05, Mar20, Rai96, AN08, HLZ23]. Explosivity [FFGM15]. Exponent [Cli07, HS04]. Exponential [AQL89, Che91, Fer90, LH96, Mil84, SF05, AE06, HB01, NM11, NS03, ZB05, Zhu13]. exponentially [LZ20]. exponents [LC03]. Extended [Arb08, PI22]. Extending [FP18, Hog19]. Extension [ACL01, LL05]. Extensions [BGT21, PF21, PSU08]. Extracting [EP17, MMNT20, RCLM^{+}11]. extraction [BM04, MT15]. Extrapolation [LXT20]. Extreme [BT06, IC05, Mar12, MW22]. Extreme-Value [BT06]. Extremes [STA03, TT97b].

F [WCK12]. Factor [CLY17, GT19a, HNT23, HN21, HA93, LTZ20, TNH23, WCG98, ZC19a, BPT02, CS11, HKVW22, HST23, KM09, WNS22, ZS17]. Factor-Augmented [LTZ20]. Factorization [MP18]. factorizing [HT10]. Factors [EP94, DGP15]. facts [KS23a]. Family [AD04, Liu12, APS20, GdSF13]. Fan [Wan21]. Fast [KD00, KP89, Maz12, JN14, Rig92]. Favouring [PS06a]. feasible [Erc11]. features [Bra21, TD11]. feedback [Li12a]. Félix [Rao16a]. Feng [Ter14]. Ferrous [FFGM15]. Festivals [MP84]. Fields [Leo13, FPS21, Yua00a]. Filling [GS20]. Filter [BH91, Kit81, MB97, MP84, SS96c, Wat85, KOZ12, Nag03, SdJ22, WWG09]. Filtering [AK90a, AK90b, Azz82, GJ16, KT94b, KD03, Góm07, KD00, PT04]. filters [BM04, Maz12]. Finance [McC15, McC15]. Financial [AiL15, AHL^{+}18, CK22, Hal14, Kim16, Rao10a, SBLS07, CCGA13, KS23a, Ter11]. Finite [CL95b, DJM86, GJ01, Har81, IP08, Sto85, VY90, LV00]. Finiteness [FT85, Fei20]. First [AOA87, Aue22, BM81, CG19, Hoy20, HHI18, Jas03, JJL12, KY09, KBB90, MM91, Mur85, Och83, SH88b, Tua92, GH03, Hoy24, Kak99b, Lug06, NS03, RSW08, ZB05]. First- [Hoy20]. First-Order [AOA87, CG19, HHI18, Jas03, KBB90, MM91, Och83, SH88b, Tua92, JJL12, KY09, GH03, Kak99b, Lug06, NS03, RSW08, ZB05]. Fisher [KM90, KM04, Wan93b]. Fisz [FNV08]. Fit [AY96, CG82, EF06, Pap05, Sai83, Sch16, Vel94, And97, AS00, CR99, DdM04, Fan05, GF15, MMOV24, Nis09, Pre98]. fitted [Ioa10, KLN04b]. Fitting [APH86, Bha89, Bha93, CN17, Fin84, GS21, Huz88, KS08b, Kit81, PT82, PPT93, BP11, EF14, MR18, ZPZ21]. Fixed [CPR18, EV19, HV08, HI19, LL95, Wes19a, HI15, ILT14, KT16]. Fixed- [HV08, Wes19a, HI15, ILT14, KT16]. Flatness [Dro07]. Flexibility [LM94b]. Flexible [MV03, PBSO23, SPA20, SBS19, Qia14]. Flood [BHL90]. Florida [Lat17]. Flow [Hoy20, GSO^{+}17]. flows [EM08]. Following [FK04, FK99]. Forecast [PS06a, KLN04b, MP16, MO02]. Forecasting [AQL89, AMZ13, CJ82, Car85, GH11, GZF86, JA81, LWL93, MP84, NNC91, RA92, SS82, Ula93, Wil16, mWK96, ZC19a, CSD12, HW09, Jan10, TNH23, Vij06, Bos16]. Forecasts [ADSS18, KF08, PS07, WS02, Hal12]. foreign [LG11]. Forgetting [LH96, BPT02]. Form [Hid92, FL04, Had04, Lug06]. Forms [Abr87]. Formula [BF97, Kan87, FZ09]. Formulations [DN95]. forward [LT03]. Foundations [Kar16, Wan21]. Fourier [Rig92, DR11, FL00, HK17, HC20, KT94b, PLNL22, Sto85, Sto87, Sto90, TAM11, VVD18, Wal00, Yaj89, YLC21]. Fourier-PARMA [TAM11]. Fourth [FP16, Kim91]. Fourth-Order [FP16, Kim91]. FPE [De 98b]. Fractal [FHW94, PV98]. Fractional [ANW93, AN94, Che93, Fin16, GA08, GJ80, GZW89, GZW94, HMV08, Hos05, HI19, JN18, KPT19, KPT21, KT01, Nie15, Rei94, VDO95, Vil01, Wri95, BB07a, BG00, Dit00, FB13, GA04, Hua12, HI15, HB01, ILT14, JN14, KM99, NM11, PC23, PT02, Sim08, ST05, Vel03, Vel07, Vol12, ZT06]. Fractionally [CAP94, Chu96, DH98, Dit04, Has93b, HR95, JN19, KH98b, PP88, CP16a, GMP15, Nie11, PR98, Pos08, SH09, Tro13, Tsa07]. framework [BDM98]. Francis [Nun20, Wan21, Yao20, Zha13]. Free [HH16]. French [Mok87]. Frequencies [MW16, ZM01, ZG88, dBCCO22, DH13, HD99]. Frequency [CZ19, Cha19, CWD00, Cub95, GOV19, GH19a, HH93b, Hua96, HB93, HB94a, Kak13, KT94b, Kim16, LLMR08, PD12, Pra82, Qui00, SR92, ST97, Tho19, Tua86, Won97, ASM21, BG00, BFK12, BFK13, CR99, CK22, FF13, FP12, GM15, Has13, Jen12, Jin18, KS23a, LL06, Lev02, LK98, LLL22, Mil10, MTW04, RDB14, RT17b, SAZ13, Tau23, Wan16, Whi02, YFL^{+}14, Yua00a, Whi05]. Frequency-Domain [SR92, Jin18]. Frey [Neš16]. full [Kur11]. Fuller [FA03, KLN04a, AN92, CHLT15, Ioa10, LKN05, OT98]. Fuller-Type [AN92]. Function [AM80, BB95, Bat88, Bha83, Bha86, CS15, Cho91, Del96, HT99, Kan87, Kav93, Lii85, MG00, MA93, MW97, Peñ84, SS89, BM10, Cav14a, Che06, CGM08, DC01, DdRSK21, DQ23, Fos13, FZ23, HLZ23, HKV24, KH98a, KGY18, LL06, LLBM^{+}11, OJHO00, PLNL22, RT17a, TY10]. Functional [AHP17, CL01, CKR18, DC23, GHHK18, HKK15, KKJ18, LTZ20, PP97, RS19, RP20, Seo24, ZP20b, BCD18, BP18, HNT23, HR15, HKVW22, KR13, LRS21, PPS20, RS17, RWZ20, SK23, SY20, TNH23, Wan08, ZHHH22]. functionals [FMS02, GLN15, LP04, Tak24]. Functions [BW18, BN92, Eng84, Gri91, Huz88, JvS95, JLMB20, JC17, MR19, Pap94, EDD17, FKD10, GLL06, Jir16, WL11, ZSW22]. Functions-Based [JLMB20]. Further [BH03, IP20, Lay84]. Future [LX96, PM92, VY90, Li06].

G [Nea13]. gains [May22]. Gamma [GL21]. gap [GS20]. gappy [CF24]. gaps [Pro23]. GARCH [ZLY06, AB09, AS24, AD04, BC01, CPR07, CL00, CK15, DZQ10, FLO06, Fin16, GMLS15, GLML16, GLP10, HT99, HK20, HA21, Kar01, KPS04, Kim16, KOW22, Kri09, LLOS08, LKB15, LLL13, LLZZ22, Liu12, MKS22, MMOV24, MS08b, Mic20, NLL12, PS01, RZ10, SW21, SY20, XZ22, Zaf07, Zha04, ZL12a, ZW12, ZXC22, ZPZ21, Zhu11, Zhu13, ZZL20]. GARCH-in-Mean [CK15]. GARCH-SM [LLOS08]. GARCH-Type [Fin16]. Garcia [Pou17]. Garcia-Hiernaux [Pou17]. GARMA [WCG98]. Gaussian [AB99, APS20, And93b, AA20, ASK15, Bar00, BCFFT17, Bou08, Cha16b, CK13, CCY16, Cra03, DN99, DZ18, Dic82, FHW94, GdSF13, GRS97, GV10, GL20, GS13, HM13, HK22, HH93c, Huz07, JP99, KT94a, Kak96, Kak99a, KMS15, Kom99, Kul85, LT17, LDN19, LLBM^{+}11, LM88, LT92, Mar12, Nag03, Neu96, NS13, Pop90, Por87, PS95, Pou88, Rai96, SB19, SPM19, SA07, Spe10, SC97, ST91, TCCG19, Vel99, Vel03, Wal95, YB06]. Gaussianity [Hin82, TW02, Won97, Yua00a]. GDP [ZC19a]. General [ABT18b, And93b, CR90, DK17, FN97, Had95, HH93a, Hua90b, Liu89a, Pfe94, Pri80, Spa93a, Spa93b, BI12, CK13, FZ09, GV10, KP20, Ken12, LV00, WD10, ZL12a]. General-Lag [And93b]. Generalised [KPT19, KPT21]. Generalized [AH92, AHP17, Bol88, BR06, Cho91, Chu96, FK04, GZW89, GZW94, HL11a, HN21, JR22, KHS03, KP90, Len16, MN95, NKC15, Pap94, PS95, RBY92, TH12, TV83, ZLP19, ZXC22, Bra11, Bro07, DLPP14, FK99, HL17, Kak13, LL12, YL22]. generalized-risk-in-mean [YL22]. Generate [Gra88, KT01, KF08]. Generated [HW95a, JL83, KLN04b]. Generating [Bar87, SS89]. Generation [BI09, She88]. Genetic [Gae00, BP11, CGM08, UT12]. Geometric [BS15a, LPS23, MG93, Lie05, Mok87, RNI13, SFK10]. géométrique [Mok87]. Geometry [RMT90, vG99]. George [Wil16]. Gerald [Tur18]. Geweke [HDB98, NH19]. Ghosh [Ter14]. Gibbs [MT94a, PS99]. Giles [Cao19]. Gilles [McC15]. Gini [CS15]. Giovanni [Leo13]. Given [And87, BI09, Pöt90, SS89, ASK15]. GJR [XZ22]. Global [NNC91, EP17, VN17]. GLS [Cha15a]. GO [ZW12]. GO-GARCH [ZW12]. Goodness [AY96, And97, CG82, DdM04, EF06, MMOV24, Nis09, Pre98, Sch16, Vel94, AS00, CR99, Fan05, GF15]. Goodness-Of-Fit [AY96, EF06, Vel94, Sch16, And97, DdM04, MMOV24, Nis09, Pre98, AS00, CR99, Fan05, GF15]. GQARCH [YSX^{+}22]. GQL [MS08a]. Granger [ADD19, GH19a, Lay84, Tau23, THN24, YK06]. Granville [McL17]. Graphical [CV06, EDD17, HG91, BPN12]. graphs [BBK23, WL11, WR08]. Gray [Kil18]. GRCAR [LPS23]. Gregory [Wil16]. Greta [Wil16]. Group [DF80, Kra16, Nun20, Wan21, Yao20, Zha13]. grouping [WWW12]. Groups [Ala89]. Guest [LT18, NH19]. Guide [Tur18]. Gwilym [Wil16].

H [Lat17, Nun20]. Hahn [Tur18]. Hall [Lat17, Rao14]. Hall/CRC [Lat17]. Handbook [Hal12, Lat17, Rao10a]. Hannan [Ano94, Pri94, Rob94]. hard [Qui15]. Hardback [Lu18, Mcn15, Neš16, Nun20, Kar16, Yao20]. Hardcover [Bos16]. harmonic [Iri02, ME98, Wal03]. Harmonically [HH20]. Harmonizability [Joy87]. harmonizable [WS11]. Harry [Mcn15]. Harte [Cra03]. Harvey [Hal14, HM03]. Having [SS95, SS96a, Bar00, KS18]. Hawkes [EDD17]. Haywood [McL17]. hazard [GG08]. Heavy [GRT17, GZ15, Hal14, JMP16, YL20, BM13, GL20, HPY02, JMP12, KW04, Kri22, LPS23, MJ12]. heavy-tail [BM13]. Heavy-Tailed [GRT17, JMP16, YL20, GZ15, GL20, JMP12, KW04, Kri22, LPS23, MJ12]. HEGY [dBC07]. Heidelberg [McC15]. hemispheric [EP17]. Henry [Kil18]. Heterogeneity [CK22, EV19]. Heterogeneous [BCCR19, Jas03, Wes19b]. Heteroscedastic [GKY18, KH01, Wei86, CS11, EM08]. Heteroscedasticity [TT99, Hen01a, Lug06, RWZ20]. Heteroskedastic [Bol88, MMNT20, KYP20]. Heteroskedasticity [AM18, AHP17, BH92, HMW19, LM94a, PX06, FRZ01, HK14, HL11a, Hon97, PY22, Wes13]. Heteroskedasticity-Robust [HMW19]. Hidden [CWD00, FR97, ZM01, ZC19b, ZG88, BF10, FRZ01, Spe10, WWW12, Lu18]. hierarchical [DC01]. Hiernaux [Pou17]. High [BCK23, BFK12, BFK13, GT19a, Hua96, Kim16, PA23, Wah89, XX18, Yu07, ASM21, CLY17, CK22, FF13, HNT23, KS23a, KP21, LTT18, LLL22, MMM22, MTW04, TNH23, XCG24, YFL^{+}14]. High-Dimensional [BCK23, GT19a, XX18, PA23, CLY17, HNT23, KP21, LTT18, MMM22, TNH23, XCG24]. High-Frequency [Kim16, BFK12, BFK13, CK22, FF13, KS23a, YFL^{+}14]. High-Order [Wah89, MTW04]. Higher [AC96, DO04, Huz81, Ked87, SR88, SR91, SO05, ST91, XLT23, ZP20b, BFZ02, BMH08, Tri06]. higher-lag [BFZ02]. Higher-Order [DO04, Ked87, SR88, SR91, ST91, ZP20b, XLT23, BMH08]. Highly [MG00, PC23]. Hilbert [BSS17]. Hill [KK20]. history [Bri12]. HIV [WLC12]. Hoboken [Bos16, Wil16]. Hodrick [SdJ22]. Hodrick-Prescott [SdJ22]. Holan [Lat17]. homoscedasticity [CP17]. Honor [HKK23, RW17]. Honour [ACN23, NH19, LT18]. Hooker [Cao19]. Hopf [RY23]. hormonal [KWPV12]. Housing [WW17]. Hsin [Qui15]. Hudak [HDB98, NH19]. Hui [Wan21]. hybrid [LLT14]. Hyper [LP10b]. Hyper-spherical [LP10b]. Hyperbolic [McL98, Cha16b]. hyperbolically [Deb11]. Hypotheses [ST03, AR06, CL01, DKV11]. Hypothesis [BK03, Bre94, Eva80, Hal92, Kas82, Kim15, Kap05, KA08, LLZZ22, Sen07, Xia01].

i.i.d [MMOV24]. Iain [Lu18]. Ice [NNC91, ZC19b]. identically [AR10, Kle08]. Identifiability [AD84, BD92, FB21, DRS21]. Identification [BB07a, Bou08, DB98, DN95, Gae00, Hal94a, Hot89, KP95, Kum86, Li84, MW97, NP90, PS92, PW84, Tig85, TP82, Wal87a, CPR07, CGP22, CGM08, GJ23, HMS13, LLRR^{+}21, MS23, UT12]. Identify [GL94]. Identifying [CW82, Tsa89, Tua84b, GKL11]. idiosyncratic [ZS17]. ignored [DH22]. II [QN81, TNH23]. illustrations [MT15]. Impact [ZT06, HLT10, SPH18]. Implementation [Ish84]. Implementing [MW16]. Implicit [MFM05]. improper [WS11]. Improved [Bra11, BDD95, GV10, KS08a, MM12, MP16, RSW08, Vid04, HW09, MTW04, Vid09, WE07]. Improvement [CG07, De 98b, Kat08, LM04]. Improving [CS87]. Impulse [BW18, MR19, Wil17]. Inadequacy [HP92]. INARCH [EF14]. incidental [KT16]. Income [Lüt82]. Incomplete [Ma02, NP90]. increments [Bar00]. Independence [BR06, BD92, Del96, KL98, GS13, LT17, MGRM10, MW22, SP12, WR08]. Independent [AC96, Bar87, LLRR^{+}21, MMNT20, RT17c, She88, AR10, BC12, GJ23]. Indeterministic [HS98]. Index [Ano97b, Ano98b, Ano00b, Ano01d, Ano03, BC97, FHW94, PV98, XHN17, Ano99c, Ano02, Ano04, Ano05b, Ano06a, Ano07, BDS12]. indicators [BH10]. Indirect [DdRSK21, Arv14, FHK20]. Individual [Aok91, GBY17, Sme15]. infection [WLC12]. Inference [AT87b, ABT18b, AC18, BL19, CT96, CP16a, CL95a, DHJ12, DGPHS19, FK04, FP16, GKY18, GH19b, HR02, HKVW22, HI19, JMP16, JFML13, Jen04, JC17, Kab93, Kim16, LZ20, LPZ15, MFM05, PX06, PRW04, RS19, Vil01, Wes19b, XCG24, YL20, ZBD06, Zho13, ZZL20, AR00, BKM21, CLY17, CK13, CF14, DdRSK21, DGP15, FHK20, GB98, HOS15, HW99, JMP12, KA07, LZ18, Lug06, May22, MJ12, PRR04, QR98, RSW08, SB19, Ter11, Tsa07, VN17, Wan09, YD12, ZWZ11]. Inferences [MS08a, PZC14]. Infinite [BR06, Kni87, KT94c, KT01, Nas93, Pöt90, Bha97, BB14, CZ12, Hil13, KS18, KP08, LV00, LM08, MR12, ZL12a]. Infinite-Order [BR06]. Infinitely [GMLS15]. infinity [NN21]. inflated [JJL12]. Inflation [BS15a, MT15]. Influence [Bon05, DZQ10, JWW99, Zha04]. Info [Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f]. Information [Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano24a, Ano24b, Ano24c, Bra16, DNL81, GL94, HT93, KM90, LX96, LW19, Pon06, Wan93b, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Cha16b, GB06, GG07, HOS15, Ioa11, KM04, Li06, Maï12, NHCLP08, PS00, PSSS09, TH12, THN24, WL98]. INGARCH [FF10, PBSO23, SBS19]. Inhomogeneous [GL21]. Initial [AK90a, HL06, Mar92a, SS96c, Swi90, DCCL03, FKD10, HLT10, Sel10]. Initialization [SS96c]. Initializing [BH91]. INLA [Rao17]. Innovation [FP16, JO06, LWL93, KGY18, Sen07]. Innovations [AB86, BD92, GRT17, LM88, Mar00, MB97, BM13, BH08, DdM04, JJL12, Kri22, MMM22, NS13, NS03, QR98, ST05, TP03, TWVB00]. Input [OM17]. Inputs [Bau05, Li93, Tig85]. Insensitive [vS94]. instabilities [MP16]. Instability [Lee16]. Instantaneous [Lay84, BS02]. Instrumental [Hal92, Hal95]. Integer [ABT18b, AOA87, DK17, FLO06, GMLS15, Gor18, HA21, JGY91, Mic20, NR07, Sch16, AS24, BMH08, EMNR09, JJL12, KY09, KT11, KL09, Lat98, LJ23, MTJ14, NLR16, ZBD06, Zhu11]. Integer-Valued [ABT18b, AOA87, DK17, FLO06, GMLS15, Gor18, JGY91, NR07, Sch16, HA21, AS24, BMH08, EMNR09, KY09, KT11, KL09, Lat98, LJ23, NLR16, ZBD06, Zhu11]. integrals [San17]. Integrated [dBCRT19, Cha15a, CAP94, Che09, Chu96, Cor95, DC23, DH98, GH19a, GH91, Has93b, HN93, HR95, KH98b, LX01, BDL08, dBCO12, dBCCO22, Cha15b, CP16a, GMP15, JPP15, KP10, Nie11, PR98, Pos08, SH09, SL04b, ST05, Swe03, Tro13, Tsa07, Wan16]. Integration [BL13, CRT15, Che93, AT99, CT06a, GA04, GA08, HLX10, ILT14, Kil11, Sme15]. integro [WH11]. integro-difference [WH11]. Intensity [GL19]. Inter [GJ02]. Inter-Trade [GJ02]. Interaction [Ter85]. Interactions [ASJLZ19, HL18]. Interactive [EV19, Wes19b]. Intercepts [Lim92]. intercorrelated [HCT04]. intermittency [GAHT01]. Interpolation [Fra84, Kas82, LXT20, Mar07a, Pou89, Tan81]. Interpolators [BB95]. Interquartile [Yua00b]. Interval [ADSS18, HHI18, HP14, KGY18]. Intervals [BM81, BDD95, CGN15, HR93, Kab99, KH01, SLL^{+}20, Tur18, ZC19a, AMZ13, GLP10, HK08, KH04a, Kak99b, KCW22, MO02, NSL07, RR09, Vid04, Vid09]. Interventions [FF10]. Introduction [GJ80, KOC15, Lu18, NH19, Rao16a, LT18, Bos16, Tur11]. Introductory [Mil06, Bos10]. Invariance [Hos05, AS00, Gir07, PF21]. invariant [ZT18]. Inverse [AB99, BB95, Bat83, Bat88, Bha83, EF06, Had95, Kan87, RG89, Had04, MR18]. Invertibility [BH94, NR93, Pic82]. Invertible [Bar87, CT96, BL21, Pos08, RB13]. investigate [BPN12]. Investigation [BM89, Che93, HH81]. Irregular [KS08b, Mil19]. Irregularly [MVS87, SS96b]. Irreversibility [Cox91]. ISBN [Aue22, Bos16, Cha16a, Hal14, Kar16, Kil18, Lat17, Leo13, Lu18, McC15, McL17, Mcn15, Nea13, Neš16, Nun20, Omb13, Pou16, Pou17, Qui15, Rao14, Rao16a, Rao16b, Rao17, Ter14, Tur18, Wan21, Wil16, Yao20, Zha13]. ISBN-13 [Pou16]. Isotropic [LHR82]. ISSN [Cao19]. Issue [ACN23, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano24a, Ano24b, Ano24c, HKK23, NH19, RW17, Tay24a, BDP21, CH11, LT18, SO12, KOC15]. Issues [HN80, PS99]. Itô [Kim16, YSX^{+}22]. Iterative [Gri91, KT94b, MW97, CGM08]. IV [DHT14]. IV-based [DHT14].

J [ACN23, Ano21a, Fei20, Neš16, Rob94, Tur18]. Jackknife [ZL24b]. James [Cao19, Pri94]. Jan [Cha16a, Ter14]. January [Tay18b]. Jenkins [Wil16]. Jennings [Bos16]. Jerez [Pou17]. Jersey [Wil16]. Jianqing [Wan21]. Joe [Mcn15]. Johansen [PLNL22]. Johansen-type [PLNL22]. John [Bos16, KOC15, McL17, Rao16b, Rao17, Tur18, Wil16, RSVM15, Sav15]. Joint [BW18, Hal92, PY22, PS06b, Sen07, ST04, WW15, GA04, VW15]. Jointly [LK21]. Jose [Pou17]. Journal [KPT21, Ano97b, Ano99c, Ano02, Ano03, Ano04, Ano05b, Ano06a, Ano07, Ano20g, HKK23, LT18, NH19, Tay20, Tay21b, Tay22a, Tay23c, Tay24b]. JTSA [KOC15, BDP21]. jtsa.12460 [KPT21]. Julio [Rao16a]. Jump [LW19, Mar92b, SHL22, VADG04]. Jump-diffusion [SHL22]. jumps [LLL22, SCW19, YFL^{+}14]. just [KP08, ZL12a].

Kalman [BH91, MP84, SS96c, WWG09, Wat85]. Katerina [Ano21a]. Kay [Qui00]. Kernel [BB87, Com04, FRP99, HV92, KC96, KH04b, SSX18, BFK13, PS21, Sko01, Wan24]. Kessler [Zha13]. known [GA04]. Kolmogorov [Góm07]. Koopman [Cha16a]. KPSS [Has01, KT10, MW16]. kriging [RT17a]. Kronecker [Liu92]. Kulahci [Bos16]. Kulik [Ter14]. Kullback [MR18].

L [Bos16, Dat97, Fei20, Kil18, Lu18]. Lad [YL20]. Lad-Based [YL20]. Lag [ASJLZ19, And93b, Büh96, CAP94, DF80, GP02, TY00, Vil01, BI12, BFZ02, Kil98, LK21]. Lag-Window [Büh96]. Lagged [HKK15, Mar07b, RS19, RP20]. Lagrange [AN94, Lju88, OT98, Tua86, Tua87, ZS17]. Lags [GL94]. Langrock [Lu18]. Laplace [DGJ06]. LARCH [Bar24]. Large [BL01, Cha95, Cha05a, FG04, KP93, Zaf08, BBK23, GBY17, GP02, KM09, WJM11, WP21, Yao20]. Large-scale [Zaf08]. laser [LL06]. Lasso [AC18, SM13a, XX18]. law [SH12]. laws [TvV02]. leaf [VPWD11]. Learning [AiL15, GOP^{+}12, May22]. Least [AT87a, Bai94, BB12, Bri80, FRS11, Gra95, Hil13, Hua90a, HB05, KP90, KS19, LM00, LH96, LDH19, MW97, Pem87, Pet86, PS95, RBY92, Sto19, TP85, WD10, ZZL20, FG04, HOS15, HL11a, Ing01, Ioa11, KHS03, KF08, LL12, Li12b, TvV02, Wal03]. Least-Squares [Hua90a, HB05, KP90, LM00, Ioa11]. Leave [Pro03]. Leave- [Pro03]. Lecture [Leo13]. Leibler [MR18]. Length [Hua90b, Vil01, DG98, GP02, KP13, ZKG22]. lengths [DL15, Jin18, LL18]. Level [CHS17, FHW94, SH90, TSL08, ILT14, LLS02, LZZ22, Mar12, McC13, Tri11]. level-crossing [Mar12]. Levinson [Hua90a, KT01]. Levinson-Type [Hua90a]. Lévy [Fin16, EK13, WHY22]. Lévy-driven [WHY22]. Li [Qui15, Wan21]. Library [Rao16b]. light [EM08]. light-tailed [EM08]. Likelihood [ABT18b, AM80, Aud05, Bau05, Bro07, Deg87, FM96, FK04, FKD10, Huz88, IP20, JWW99, JC17, LM04, LDN19, Li20, LB00, Ma02, MS07, MM91, Mil95, MA93, NS03, RBY92, SK96, Shi93, ST99, Tua87, Ver87, YB06, YD12, ZL12b, AB09, AS21, BBC16, BMH08, CG07, Cav14a, Che06, CD09, CD12, CLL14, CCY16, CH21, CF14, Cle01, DB03, GdSF13, GLP10, GL20, HLM03, Kak13, Kur11, LD04, Mau02, NSL07, OJHO00, PS21, PK13, Per04, PF21, PW05, RSW08, SF98, SBS23, TP03, TC05a, Yau12, ZWZ11, ZC12, ZB05, ZPZ21, Zhu13, RA92]. Likelihood-based [Bro07, Kur11]. Limit [BT94, Cha91, Chu12, GP06, Kee97, Leo13, Mor83, Oza82, Sto85, Yaj89, ZL12a, Arv14, AA20, BT13, BDL08, BP18, FPS21, KL10, LP14, RB13, Sta20]. Limited [Gor81]. Limiting [Cha15a, LD04, Mic20, Tia88, Yab12, GPRV00, Shi98, WJM11]. Limits [Hog19, KS08a, GV10]. Lindner [Zha13]. Line [CJ82, HH93b, DA14]. linéaire [Mok87]. Linear [And87, ADD19, Bai94, BB95, Bat83, BSS17, CR90, Chi91, DN99, DH98, Fin85, FK04, FP16, FS94, Gor81, HHP84, HIP87, Has00, Has01, Hög86, JvS95, KOV94, KP89, KS19, LM94a, LLG09, LXT20, LM94b, MB97, MT90, NR93, OJ03, PS89, PT81a, PW89, Pri80, Rao05, SS89, SSX18, Sib01, ST85, TA88, Tho19, Tig85, TV83, Tsa88, TW89, USMS83, XX18, ZG85, BDS12, BF97, Bha97, BB14, BB07b, BM03, BM09, DGK14, Dat97, FK99, FPS21, FM98, GH11, HLT10, HP23, HMV08, HPW17, Jen12, KPRN03, KLN04b, KR98, Kri22, LL12, LLY14, LZ24, Mar12, MJ12, MP10, MZW09, Mok87, MWM97, NHCLP08, RT09, Sel10, Sim08, Sun14, kTR98, TZ22, VN17, XCG24, ZM06]. Linearity [Cox91, Hin82, Kil16, O’B87, PP97, RG80, Sak91a, TLG93, Won97, Yua00b, TM98, Yua00a]. Lines [HQ89]. Link [HK90, EDD17]. Ljung [Wil16]. LLC [Nun20, Yao20]. LM [Kil16, NR11]. LM-type [NR11, Kil16]. LME [FFGM15]. load [CT10]. Local [AG24, Aud05, BSG18, Bra16, CL06, DVW08, Iac10, KT16, Kur11, LT17, LRS21, LLG09, Mas96, Nie11, PP99, SP08, SH90, Vel00, Zha04, BCFFT17, BP18, DK13, FPS21, NK24, Tri11]. Locally [Büh96, CN17, HC20, JLMB20, PP16b, ST03, SP08, Tay03, WS11, DG98, GS13, HK17, Kle08, MKN22, Tam09, XCG24]. Location [Ger18, Li20, MM93, FZ23, HS04, Rao08]. location-dependent [Rao08]. location-scale [FZ23]. Log [Jan82, SP02, AO09, Ioa10, MRT07, PC23]. log-periodogram [AO09]. log-regression [MRT07]. log-stable [PC23]. logit [BWH19, WC14]. logit-beta [BWH19]. Lognormal [SF11]. Logspline [KST95a, KST95b]. Lomb [LLMR08]. Lomb-Scargle [LLMR08]. London [Leo13, Nun20, Ter14, Yao20, Zha13]. Long [AVF98, AV05, AM07, BCCR19, BT94, Ber07, BK19, Com96, Cox91, Fin16, GPH83, GRS97, GJ80, GSS17, GH19b, Has94, Hen01b, HH05, HB93, HB94a, HB94b, IL19, JMP16, Jen04, KPT19, KPT21, KPS04, KCW22, KE88, KLN18, KS19, LDN19, LKB15, LPPS20, LDH19, Lob97, Lub99, MORS21, Mar00, PC05, Ray93, ST88, Sib01, TZ19, VDO95, WCG98, YK06, AG16, AR00, Art02, BP12, BT13, BDS12, BC02, Bet16, Bou08, Bro07, CCGA13, CL97, Cra03, Deo97, DHJ12, Erc11, GAHT01, Ger21, Hen01a, HS04, HDB98, HD99, HC00, HB01, Hur01, JMP12, KP15, KP20, KP13, KS18, KW12, LV00, LLBM^{+}11, LRS21, LG11, McC13, MJ12, MS00b, MRT07, NM11, Nie05, NSL07, RT02, RMSF10, RS17, Sha11, SK09, Sou07, ST05, Tau23, TT97a, Tew18]. long [TC05a, WC10, Wri98, Yam11, Yau12, YD12, Ter14]. Long-Memory [BT94, BK19, GJ80, HB93, HB94a, HB94b, JMP16, Jen04, KPS04, KE88, LDH19, Lub99, Mar00, Ray93, ST88, HH05, Sib01, WCG98, BDS12, BC02, Bou08, Bro07, Hen01a, HDB98, HD99, HC00, HB01, JMP12, KP13, LV00, McC13, MJ12, MRT07, NM11, RMSF10, TC05a, Wri98, Yau12, YD12, Ter14]. Long-Range [Ber07, Cox91, GRS97, KLN18, LKB15, Ray93, VDO95, AVF98, AG16, BP12, BT13, Bet16, GAHT01, Ger21, KP15, KP20, LLBM^{+}11, LRS21, Nie05, NSL07, RT02, Sha11, SK09, ST05, TT97a, Tew18]. Long-Range-Dependent [PC05]. Long-Run [YK06, CCGA13, RS17, Tau23]. Long-term [KCW22]. Longitudinal [Fok10, IJ99]. Look [Lüt82]. Loop [Che95]. Louis [Tur18]. Low [And93b, Azz81, Hua96, HB93, HB94a, FM04, MTW04]. Low-Frequency [HB93, HB94a, MTW04]. Low-Lag [And93b]. Lower [Kab87, Wan05]. LSE [Tan87]. lstr [DE07]. Ltd [Rao16a]. Lund [Lat17]. Lyapounov [Cli07]. Lyapunov [LC03]. Lydia [Bha82]. Lying [FM96]. Lynx [Lim87].

M [AH92, Pou16, Wil16, HPW17, Kap05, Kee97, HL18, MO02]. MA [RTW14, Arv14, MP18, Wan08, Yab12]. MA-Sieve [MP18]. Macdonald [Lu18]. machina} [MW22]. Macroeconomic [GA01]. Magnitude [Wri95]. mail [RSVM15]. Management [Neš16]. Mandelbrot [Fin16]. Mann [Mau11]. Mann-Whitney [Mau11]. many [KCW22]. Map [HW95a]. maps [RCLM^{+}11]. Marco [McL17]. Margin [ZJN24]. Margin-closed [ZJN24]. Marginal [DG20, DH17, LQ20, Log04, Cha16b, GdSF13]. Marginals [SH88b]. Marinucci [Leo13]. Market [Pra82, WW17]. Markov [Fei20, Lu18, AF22, BF10, BBKL17, BS15b, Bra21, Cav14a, Cav14b, FT85, FRR17, FR97, FRZ01, HLHT94, KR98, LK21, Liu12, MT94b, PS03, PS06b, Smi08, Spe10, Ton81, VADG04, ZKG22, ZS01, ZJ06]. Markov-Switching [Liu12, Smi08, Cav14a, LK21, PS03]. Markovian [BCT15]. Marta [Rao17]. Martínez [Rao16a]. Martingale [HP92]. Masanobu [HKK23]. massive [KC11]. matching [KH04b]. Mathematical [Ano94, Leo13]. Mathieu [Zha13]. MATLAB [Rao11]. Matrices [Deg87, McE18, MIN^{+}16, BB14, HST23, KM21, MP10, SA07, WJM11, WP21, Yao20]. Matrix [CMK05, Had95, KM90, MP18, Sak91b, GB06, GG07, Had04, HT10, KL11, KM04, Len16, Zha24]. Matteo [Pou16]. Maurice [RTW14, RW17]. max [KOZ12]. max-stable [KOZ12]. Maxima [Kri22]. maximisation [CF24]. Maximization [AC93]. Maximum [Bau05, BLL09, BI12, BMH08, Cle01, DB03, Deg87, FM96, IP20, Li20, Ma02, MS07, MM91, Mil95, Per04, Pol94, RBY92, SK96, Shi93, TC05a, Tua87, YB06, ZC12, ZB05, AB09, AS21, CLL14, GL20, HLM03, LD04, PS21, PF21, SF98, ZB02, Zhu13]. Maxwell [MA20]. McAleer [Tay22b]. McElroy [Aue22]. MCMC [NR07]. McNeil [Neš16]. Mean [AD04, BG20, IP08, IY03, JMP16, Kab87, Kak96, Kar01, Ray88, RGLA11, ST88, SS01, AF22, BP12, BM04, Bla14, DH22, DQ23, DZ17, Ger21, HDB98, Ing01, JMP12, KH98a, KLN04b, KW04, NSL07, QS00, Sha11, VW15, VY16, YHN99, YL22, YLC12, CK15]. Mean-Square [Kab87]. Mean-Squared [IY03, Ing01, KLN04b]. Means [Dun81, Wes19a, Bet16, CGM08, HR15, TT97a]. Measure [Bat83, Bha93, Pic90, Wan16]. Measured [GJ06]. Measurement [BCK23, KK20, KS19, Sin93, WTSL17]. Measurements [AF91, GSO^{+}17]. Measures [GGSW20, GSTW20, BCFFT17, GMP15, Hos01, RCLM^{+}11, VVD18, Wyl08]. Measuring [PS07, Zho12, Jan05a]. mechanism [BDM98]. Medal [Ano94]. Median [Lug06, Zie99]. Median-Unbiased [Zie99, Lug06]. Meeker [Tur18]. Memorial [KOC15, BDP21]. Memory [AV05, AM07, BCCR19, BT94, BK19, Com96, CK15, Fin16, GPH83, GRS97, GJ80, GSS17, GH19b, Has94, Hen01b, HB93, HB94a, HB94b, HR95, JMP16, Jen04, KPT19, KPT21, KPS04, KE88, KS19, LPPS20, LDH19, Lob97, Lub99, Mar00, RW17, Ray93, ST88, TZ19, AR00, Art02, BDS12, BC02, Bou08, Bro07, CL97, CCY16, Cra03, DGH06, Deo97, Erc11, Hen01a, HS04, HH05, HDB98, HD99, HC00, HB01, Hur01, Iac10, JMP12, KP13, KS18, KW12, LV00, LLL13, McC13, MJ12, MRT07, NM11, RMSF10, RT09, Sib01, Sou07, TC05a, WC10, WCG98, Wri98, Yam11, Yau12, YD12, Ter14]. messy [Mil10]. Metals [FFGM15]. Method [Aok91, Bar87, Bau05, HH93c, Hok83, JO06, OM17, Pfe94, XY89, Bra11, HT10, Kak13, KM09, Kom99, MTJ14, MR18, PW05, RCLM^{+}11, ZW12, Jan05b]. Methodological [SW86a]. Methodology [MS92, WH19, YL91, AV08, GAP09]. Methods [Bau05, Bra16, Hal94b, HG91, KS08b, KP89, Neu96, PZ17, PT81b, Pou17, Rao14, Tay24a, Yak87, Zha13, CG07, CPR15, DRY21, GH11, GJ16, LL18, Lug06, Par13, PY22, Rao10b, Rao12b, RT17b, RT02, SH09, SO97b, Tsa07, Ter14]. Metric [GMR04]. Michael [Tay22b, Zha13]. Michela [Rao17]. microrheology [DMHF12]. Microsoft [Mil06]. Miguel [Pou17]. Mild [FFGM15]. Mildly [AM18]. Minimal [Psa08, DKV11]. minimax [XLT23]. Minimum [CT06b, DZ18, Hua90b, CHLT15]. Mink [Ter85]. Mis [Has94, LR88]. Mis-Specified [LR88]. Missing [Efr20, Nas94, Pou89, AK10, Bon05, BB12, Efr14, Efr19, SS98]. misspecification [STY97, KLN04a]. misspecifications [HL11b]. Misspecified [Bra22, Hal94a, EFT16, SL97]. Mixed [CZ19, Cha19, CH15, GH19a, Hoy20, KST95a, MP90, MW16, Qui15, SBS19, Tho19, WL05, Che06, GM15, Mil10, RN12, SAZ13, Sta20, WCK12, WWW12, Zhu13]. Mixed-Frequency [GH19a, GM15, Mil10, SAZ13]. Mixed-Norm [CH15]. Mixing [Arb08, VDO95, Lie05]. Mixture [Ber01, KMS15, AR10, BF10, CT10, HK22, LLL13, RB22, ZLY06]. Mixtures [DFR21, JL83]. Mode [CR90, Wan24]. mode-based [Wan24]. Model [ASJLZ19, AB86, And93a, AVW16, Ber01, Bha83, Bha89, Bha93, CJ82, CN17, CT87a, sC88, CT01, DO04, DB98, De 01, De 91, DS91, Elt94, Fer90, Gae00, GJ06, GSS17, GGSW20, HW95b, HCH00, HW89, HH81, HP92, HK20, HT93, Hur01, Huz88, IMR18, IP20, JGY91, JN18, JN19, KMS15, Kap01, KE88, KBB90, Kit81, LW91, LX96, Lii85, LS06, Ma02, MP87, McL93, MV03, MA20, MG93, MN95, MW97, Mur85, NBQ16, Pap05, PS89, Pet86, PW84, PD02, PW89, Pou86, Pra82, Ray88, RBY92, Rei94, SS89, SK96, SPA20, SR91, SH90, Shi88, Sib01, SO05, SHL22, TCCG19, Tan87, Tua84b, USMS83, WTSL17, WCG98, XHN17, YR95, ZW08, AR10, AD23, AS24, Arv14, AA20, BHLS11, Bra22, CSD12, Cha15b, CLY17, Che06]. model [CLL14, DE07, EFT16, EM08, HMS13, HC04, HS11, Joh03, KT11, Kil11, KC10, KR13, Kri09, LPS23, LPZ15, LP19, MKS22, Mau02, McC13, MW22, Mok87, NS03, Olm23, PZC14, PK13, Per04, PH02, PW05, Qia14, Rao08, RN12, RNI13, RB13, Sim08, SW21, SCW19, SY20, SBS23, Tri11, Unn04, UT12, VADG04, WLC12, Wie13, WC14, XZ22, YLC21, YL22, YP06, YLC12, YSX^{+}22, ZS01, ZS17, ZPZ21, Zhu11, ZL12b, ZLY06]. Model-Adaptive [XHN17]. modéle [Mok87]. Modeling [Aka80, BS98, DN95, Fok11, GT19a, GZF86, HL18, KP20, LZZ22, Mcn15, Ray93, SLL^{+}20, ASM21, EDD17, Hal13, HL17, HK22, HA21, Omb13, ZP20a, Tur11]. Modelling [Aok91, BS15a, CCGA13, CS15, CZ19, CV06, JA81, KWPV12, LV00, LM88, LT83, LM94b, Nan14, Ott88, Pou16, SBLS07, ZJ06, AV08, BPT02, BL13, Cha99, CLY17, CG11, LR02, SLL97, Tri12, UD09, Rao05]. Models [AC93, AB99, AG95, AF16, AT87b, ARS86, ABT18b, AQL89, AD99, AH92, AM80, AD84, And92, Ano21a, AK90a, AK90b, Arb08, AD04, AHS06, Bai93, BC01, BCCR19, BKS97, BL01, BM89, BBKL17, BH91, BMY99, BH94, BH92, BLL05, BI09, Bra16, CHS17, Car85, CRT15, CPR18, Cha15a, Cha19, CT20, CT86, Cha91, Cha95, CAP94, CL95a, Che95, Che91, CT87b, CG82, Cli07, Com96, CK15, CW82, CN86, DC23, DGJ06, Dav91, DZ18, Dit04, DH17, Duo84, DGPHS19, FK87, Fin84, Fin85, Fin16, FM85, FK04, Fok10, FRS11, FS94, GR81, GPH83, GT93, GKY18, Giu17, GMLS15, GS21, Gor18, GJ01, GJ80, Gra88, GL94, GSTW20, Gue87, HP17, HHP84, HO84, Hal94a, Hal95, Hal14, HK86, Hok83, HZF93, Hz92, HN21, HR93, Hot89]. Models [HN93, HV99, Hoy20, Hua90b, HA93, Hyn93, IP08, IY03, JS90, Jen04, KH99, Kan81, Kap01, Kar01, Kil16, KS08b, Kim16, KM90, KS05, KP89, KP90, KP95, KS19, KF92, Kum86, KLM16, LM04, LR88, LS03b, Li20, LTZ20, Lim87, LPS99, Lju88, Lu18, LB00, MS92, Mar92a, Mar92b, MMH88, MMT05, MT94b, ML83, McL94, McL95, MW05, MMT98, MS07, Mil84, MWM97, NLL12, NV96, NQ80, Nie15, OJ03, OM17, Oza82, PS92, Pap22, PP88, PT81a, Pem87, Peñ84, PPT93, Pet19, PT05, Pic90, PP97, PT86, PS95, Pöt90, Pre98, Pri80, QN81, RRW83, Rao17, RMT90, RA92, Sai83, Sai86, SS16, SH88a, SR88, SR92, SL04a, SBS19, Smi08, SF05, SH88b, ST87, Stu01, Swi90, TKOP20, TA88, Tho19, Tjø86, Ton82, Tsa89, TT99]. Models [Tua84a, Tua86, Tua87, Tua88, Tua92, Vel94, VWR87, Wah89, Wal87a, Wei84, Whi85, mWK96, WL05, XHN17, XZL20, YL20, YB06, YL91, Yu07, ZC19a, ZT97, ZLSY20, ZG85, ZZL20, dJ86, dJCCL94, vG99, AN08, ADL18, AF22, APS20, AE06, AMZ13, AK10, And08, AR06, AS21, Aud05, BF10, BP11, BM04, BG00, BHL09, BBK23, Bla14, BB12, BI12, BWH19, Bra13, BS02, Bro07, CS08, CH14, Cam04, CG07, CPR07, CT08, Cav14a, Cav14b, CGP22, CT06b, CL01, CK13, CGM08, CS11, CF14, Cle01, CT10, Com04, DB03, DLRY08, DF11, De 98a, Deo97, DRS21, DVW08, DdM04, Duc05, DdM13, EF14, EM02, Fan05, FHK20, FK13, FK99, FKD10, FG04, FR07, FZ23, FB21, GG08, GMRO11, GdSF13, GBY17, GH03, GV10, GB06, GG07, GLP10]. models [GL20, GMP15, Guo03, HOS15, HNT23, HLZ23, HP23, HP14, Hon97, HL11b, HPW17, HKVW22, HST23, HWBD11, HB01, Ing01, IC05, JS24, JT03, JT11, KM09, Kat12, KFS02, KHS03, KPY22, Ken12, KL13, Kim15, KOW22, KS23a, KK12, KM04, KD00, KD03, KS18, KP21, KR98, Kri09, KW12, LLOS08, LP10a, LD04, LJ23, Li12a, LLL13, LLY14, LZ20, LK21, LM08, Lou08, Lug06, Maï12, MA24, MTJ14, May22, MS08b, MS23, MO02, Mon98, MTW04, NM11, NS13, NN21, OJHO00, PS21, PT04, Pen07, PS99, PF21, PA23, PW05, Pro03, PS03, PS06b, PSSS09, QS00, RZ10, RB22, RR09, San18, Sbr11, Sch98, SM13a, SH09, Sel10, SAZ13, SS98, STY97, Spe10, SFK10, SR07, Sun14, TNH23, TWVB00, TvV02, TT97b, UD09, VW15, Vid04, Vid09, Wal03]. models [WS02, Wan05, Wan08, WC10, WCK12, WH11, WR08, WL98, WD10, WX18, XPZL10, XCG24, Yau12, YD12, Zaf08, ZHHH22, ZS01, ZM06, ZL12a, ZKP22, Zha24, ZJN24, ZL24a, ZW12, ZXC22, Zhu13, McL17]. moderate [NK24, Sto19, Yab12]. modifications [LKN05]. Modified [BS15a, Hir06, Hua12, Maï12, MTW04]. Modulated [Has82, TM93, GSO^{+}17]. Modulus [ZT97]. Moment [Gab88, Huz81, KP13, WHY22, Yu07, ADL18]. Moments [And87, And93b, AC96, AD04, Cli07, DO04, FT85, Fei20, SR88, SR91, SO05, Str96, Bra11, MTJ14, San17]. Money [Lüt82]. Monika [Yao20]. Monitoring [AHL^{+}18, Kur17, Kur21, WW17, DA14, GKD21, HKV24]. monkey [GOP^{+}12]. Monographs [Lu18, Mcn15]. Monro [FRR17]. Monte [BM89, Che93, Dit00, Mil06, PT81b, VADG04]. Montgomery [Bos16]. Monthly [MP84, Pon06, Tay98, ZC19a]. Morris [MP84]. most [HPW17]. motion [BB07a, Hua12, KM99, PT02, ZT06]. Moulines [Rao14]. movement [TY10]. Moving [AH92, AM80, AT86, And92, AV93, AK90b, Bai93, BKS97, BM89, BMY99, BH94, Bha83, Bha89, Bos96, Bre94, BLT92, CT96, CG19, Cha95, CH15, Chi91, Cho91, Chu96, CT87b, DJM86, DSW80, DS91, GZ15, Had95, HK17, HR93, HN93, JL83, KM90, KP89, KP90, KF92, Lju88, MS92, MS00a, MA20, Nas93, PS92, Pap05, Pic82, Por87, Pöt90, RBY92, Sai86, She88, Shi93, SS95, SS96a, SM06, SHLL96, Tua84a, Tua84b, Tua86, Ula93, Vel94, Wah89, Wan93b, YR95, AMZ13, And08, BFK13, Chu12, De 98a, DA14, FF13, GG07, Huz07, KM21, Li12a, LZ20, Mau02, MR18, MS01, Mon98, PR98, SS98, SF98, ST05, kTR98, TC13, WD10, XPZL10, ZL24a, ZXC22]. Moving-Average [AH92, And92, AV93, AK90b, BKS97, BM89, BMY99, BH94, Bha89, Bos96, Bre94, BLT92, CT96, CG19, Cha95, Chi91, Cho91, Chu96, DS91, Had95, HR93, HN93, KP89, KP90, KF92, Lju88, MS92, Nas93, PS92, Por87, Pöt90, RBY92, She88, SHLL96, Ula93, Vel94, Wah89, Wan93b, YR95, De 98a, GG07, Li12a, Mon98, PR98, SS98, kTR98]. MSE [PT05]. Mulero [Rao16a]. Multi [AMS^{+}17, Cai11, HKV24, Pem87, SS96a, TCCG19, TP03, Ton82, Tri12, BP11, EK13, HW09, KL11, KGY18, Nie11, ZSW22]. Multi-purpose [HKV24]. multi-regime [BP11]. Multi-Scale [AMS^{+}17]. Multi-Step [Pem87, HW09]. Multi-Step-Ahead [Ton82, KGY18]. Multi-Variate [SS96a, TCCG19, Cai11, TP03, Tri12, EK13, KL11, Nie11, ZSW22]. Multichannel [Sak93]. multidimensional [KWPV12, Sza22, Sza23]. Multimodal [Mar92b, RB92]. multinomial [BWH19, LOS12]. Multiple [CF24, CT87b, Liu89a, Lüt82, NBQ16, RT17c, SH87, ST87, Tig85, TP82, VWR87, WT19, BC12, BS02, CHLT15, Has13, JFML13, Jin18, Kat09, KYP20, Kim15, PR10, San18, WWW12]. Multiplicative [Li20, Mil84, Yaj85, AD23, AS24, GB06, GG07, YHN99]. multiplicative-error [AD23]. Multiplier [AN94, Lju88, OT98, Tua86, Tua87, ZS17]. multiplier-type [ZS17]. Multiscale [PV98, BB07a]. multistep [Bon01]. Multivariate [AG16, BT13, Bar87, BT94, CL00, CW82, DG20, DE07, Fer90, FR07, GKY18, GOV19, HS05, Hid97, Hog18, JA81, KOV94, Kan81, LR88, Mas96, MPR91, MB97, MR19, Nie15, NR93, Pau84, Pop90, PP16b, Qui88, Ray88, RBY92, SH88a, SBLS07, She87, SS95, SLL97, SM06, SH88b, Sto90, Stu01, Swi90, Tsa89, TT99, Ula93, Vil01, Wil17, Won97, WTSL17, APS20, ASK15, BS15b, BWH19, DZQ10, DFF23, EDD17, Góm07, GAP09, HKV24, Jen12, JPP15, KP15, KD00, LLRR^{+}21, LL97, MM12, MKS22, MY04, MT15, MMOV24, Mon98, PK13, PS07, PA23, San17, SH09, Spe10, ZP20a]. Murat [Bos16]. Murray [BDP21]. musical [Iri02]. Muskrat [Ter85]. Mutual [GL94, Cha16b, Li06].

N [JMP16, Kee97]. N. [Aue22]. Nankervis [KOC15]. NAR [FRR17]. Natural [Eva80]. Natural-Rate [Eva80]. Near [dBCRT19, Cha15a, sC88, Cha15b, IC05, WNS22]. near-cointegration [IC05]. Near-Integrated [dBCRT19, Cha15a, Cha15b]. Nearest [Yak87]. Nearest-Neighbour [Yak87]. Nearly [AF91, SP01, AG08, LR21, PZC14, VY16, ZC12]. Necessary [DRS21, Kan81]. Negative [ABT18b, And89, GL19, HZF93, Hz92, TC07, BM13, CF14, HL11b, Lat98, Zhu11]. Neglected [BK07]. Neighbour [Yak87]. Neil [Cha16a]. Nelson [GB99, NV96]. Nelson-Type [NV96]. Ness [Fin16]. Nested [McE18]. Network [TLG93, TD11]. Networks [PBT00, MS23, TJL23]. Neural [PBT00, TLG93, MS23, TD11]. Neurophysiological [Rig96]. neuroscience [Omb13]. Newbold [LT18]. next [Ano20g]. Nicholas [Nea13]. Nicholson [Bri12]. NJ [Bos16]. No [Ano21a, Bau05, Fei20, ADL18, Pes07]. no-cointegration [Pes07]. Noel [Rao16b]. Noise [And93b, FM96, Fra84, GSTW20, LT92, RB92, ST88, Spa93a, Spa93b, BCD18, KP10, MS23, PC23, Sim08, Vol12]. Noises [FR97]. Noisy [RT92, Tig85, LC03, RP20]. Non [And89, And92, BB95, BF97, BB99, Cam87, dBCO12, CT96, CZ12, CLL14, Cox91, DSW80, DZ18, DPT12, FK87, FRR17, FFGM15, GOV19, HHP84, Hal94b, HS05, Hid97, HH93c, Hög86, HZF93, Hz92, HA93, JS90, JvS95, KT94b, KLN18, Kok12, Kun97, LM88, LM94a, LT92, LM94b, MS08a, Mic20, Mil19, MT90, MWM97, Nag03, Nas93, Neu96, NBQ16, O’B87, PT81a, Pop90, Pou88, Pri80, SPM19, SHL22, Swi90, TCCG19, TK93, Tau23, TA88, TC07, Tsa88, TW89, VY90, VWR87, Wan93b, XY89, YK06, YFL^{+}14, ZZL14, dJCCL94, vS94, APS20, AK24, AHT13, Aud05, BM13, BH13, Bou08, Bra13, BL21, CT08, CHLT15, CCY16, DKV11, Efr14, FB13, FM98, GdSF13, GL20, GSO^{+}17, HM13, HLZ23, HL11b, HS11, Ioa22, JP99, Kei03, KR98]. non [Lat98, Lie12, LLG09, LOS12, MT15, Mok87, NS13, PZC14, PRC03, Pos08, Swe22, TWVB00, Vel99, WC10, WCK12, ZC12, ZSW22, Zho13]. Non-Admissibility [Hal94b]. non-causal [BL21, JP99, Swe22]. Non-Causality [HS05, YK06, Bra13]. non-central [WCK12]. non-cointegration [BH13]. Non-Contemporaneous [Mil19]. Non-Correlation [HS05, PRC03]. Non-Embeddable [BB99]. Non-Ferrous [FFGM15]. Non-Fourier [KT94b]. Non-Gaussian [DZ18, HH93c, LM88, LT92, Nag03, Neu96, Pop90, Pou88, SPM19, TCCG19, APS20, Bou08, CCY16, GdSF13, GL20, HM13, JP99]. Non-Invertible [CT96, BL21, Pos08]. Non-Linear [HHP84, Hög86, JvS95, LM94a, LM94b, MT90, PT81a, Pri80, TA88, Tsa88, TW89, BF97, MWM97, FM98, KR98, Mok87]. Non-Linearity [Cox91, O’B87]. Non-Minimum [DZ18]. Non-Negative [And89, HZF93, Hz92, TC07, BM13, HL11b, Lat98]. Non-Normal [DSW80, JS90, XY89, TWVB00]. Non-Parametric [Cam87, FRR17, Hid97, KLN18, Kok12, NBQ16, TK93, vS94, dBCO12, DPT12, Tau23, YFL^{+}14, ZZL14, Aud05, DKV11, Efr14, FB13, HS11, Ioa22, LLG09, NS13, WC10, ZSW22]. Non-Singularity [Wan93b]. non-smooth [Kei03]. Non-Stationarity [Kun97, VWR87, CLL14, AK24]. Non-Stationary [And92, BB95, FK87, GOV19, HA93, MS08a, Mic20, Nas93, Swi90, TW89, dJCCL94, CZ12, SHL22, AHT13, CT08, CHLT15, GSO^{+}17, HLZ23, Lie12, LOS12, MT15, PZC14, Vel99, ZC12, Zho13]. Non-Zero [VY90]. Noncausal [SS16, GJ16]. Noncausality [CL00]. nonconsecutive [BLL09]. nonexplosive [RZ10]. nonindependent [FR07]. Noninvertible [AT86, HR95]. Nonlinear [ASJLZ19, Arb08, APH86, BO05, BK03, Cor95, DFR21, EM02, FMS02, GT93, GJ02, GH91, GL94, GMR04, Has82, LR02, LTZ20, MV03, Neu96, Oza82, PP12, Pem87, Pol94, Rao10b, Wan08, WP21, Wan24, WTSL17, BCFFT17, Bla14, BM10, Cha05a, DGH06, DLRY08, DK13, DPT12, FZ09, HLX10, HS11, KWPV12, Kil11, KK12, KS18, LP10a, Lie05, LL97, MS08b, Muk99, Nis09, PS21, PSSS09, San18, SR07, Sun14, WH11, Zho12, Rao14]. Nonlinearities [AES06]. Nonlinearity [BK07, XPZL10]. Nonlinearly [GM85]. Nonnegative [MM93]. Nonnormal [HCH00]. Nonparametric [AVW16, BLL05, CT92, FLL13, Hir06, HCT04, MW05, PR95, Rao10b, Rob83, Roz01, TY00, mWK96, XL02, YHN99, DS04, Efr20, EDD17, FKMN02, HPY02, HV08, Hid07, KL10, MS23, MZW09, Olm23, PP16a, SY11].

Nonstationary [BH91, GZ88, HR88, HH93c, HR95, JN19, Kit81, Li93, MMT05, SP01, SL96b, Spa93a, Spa93b, TC05b, UH95, YR95, Zho92, BG00, BHL09, BH08, Li98, PR03, RZ10, SL97, Shi98, SP18, WWG09]. Norm [CH15]. Normal [Bar87, DSW80, JS90, MA93, Pet89, XY89, AK24, TWVB00]. normalcy [LZZ22]. normalcy-dominant [LZZ22]. Normality [DS91, GHHK18, DVW08, DdM04, RT09]. Normalization [ZLP19, HVS15, LZ18]. Normalized [TZ19, ABT18a, Bet16]. Norming [LP14]. Note [AT87b, Ali83, AD99, AB86, AK90b, Bro95, CT87a, CL95b, CS84, Cub95, GS21, HZZGH83, Kak99a, Lay84, McL93, Mor83, New80, PT81a, Pos05, Qui82, Qui88, She88, SW86a, Ton81, Ton82, TC07, TT99, Tua84b, VWR87, Wat85, BF10, BC02, Cha99, Dat97, GG07, HM03, Ing01, Li12a, NS13, PP16a, PT04, Wal03, WL98, YLC12, ZB02]. Notes [Leo13, Li06]. Nuisance [HH16, PW05]. null [CT06a, KA08, MLS97, Pes07, Xia01]. Number [FHW94, Hal14, KH94, KP93, KS94, Leo13, McC15, Nea13, Pöt90, Qui89, Sak93, Ter14, VY90, Wan93a, BEL06, Bos16, Cao19, Cav14b, Cha16a, HD99, Kar16, KP10, Kil18, Lat17, LM00, Lu18, McL17, Mcn15, Neš16, Omb13, Pou17, PS03, Qui15, Rao16a, Rao16b, Rao17, Spe10, Tur18, Wan21]. Numerical [JWW99, MS92, MT90, HT10].

Obituary [RTW14, Tay22b]. observation [DRS21]. observation-driven [DRS21]. Observational [JWW99]. Observations [Azz81, HT99, Mur85, Nas94, PR88, BB12, Bra11, Bra22, Efr14, HKV24, HR15, KW04, MMOV24, Pen07, PS00, RP20, YP06]. Observed [GM85, KK20, MVS87, SS96b, LSSC16]. Obtaining [CN86]. Occur [AT86]. oceanographic [GSO^{+}17]. On-Line [CJ82, HH93b, DA14]. One [DG20, Hon97, KT01, SH87, Cle01, ZB02]. one-dimensional [Cle01]. One-sided [Hon97]. Online [Ano05e, Ter14]. only [PW05]. Onset [Gil99, Ano99d]. Open [Che95, Ano20g, GKD21, HKV24]. open-end [GKD21, HKV24]. Open-Loop [Che95]. Operator [RS19]. operators [PPS20]. Optimal [AE06, BP03, FB13, FZ23, GRS97, Jir16, Mar07b, Ray93, SHLL96, Tay03, Ter11, ZLP19, Bra13, DG98, Mar12, TD11]. Optimality [BC97, Kul85, MM91]. Option [MFM05]. oracally [KGY18]. Oracle [AC18, Giu17, Sch22]. Order [AOA87, AQL89, AC96, Azz81, BS07, Bha83, BM81, BR06, Bri80, CG19, CL00, DO04, DG20, Duo84, FP16, Gab88, GH19b, Had95, HH93a, Hal94a, Hoy20, HHI18, Hua90b, Huz81, Jas03, KT11, Kak96, Kav89, Ked87, KBB90, Kim91, KP93, LX96, Lii85, Lüt85, MM91, Mur85, Och83, PS92, Pap94, Pau84, Pöt90, Qui88, SL96a, SS89, SR88, SR91, ST04, SO05, SH88b, ST91, Tan87, TP85, Tua84b, Tua88, Tua92, Wah89, YR92, ZP20b, BMH08, BH08, CWDL97, DR11, EMNR09, GH03, HLX10, HT99, Hoy24, JJL12, KY09, Kak99b, Kil98, KR13, LLY14, LK21, Lug06, MTW04, MS00b, NS03, PH02, PS06b, RSW08, Sme15, Tam09, TK08, Tri06, XLT23, ZB02, ZBD06, ZB05].

Ordered [Fok11, WSS04]. Orders [Swi90, Wri95, ZW94, IC05, Rao16a]. ordinal [LZZ22]. Ordinary [EF06, LDH19, LL12]. Ordinates [HB93, HB94a, ZGH80]. Origin [NP96]. Ornstein [OV04, WHY22]. Orthant [KS94]. Orthogonal [Rao18]. Other [Eng84, BCFFT17]. otoacoustic [WCK12]. Outlier [BO05, BT06, Kab94, Led90, Lou08, Sch96, SPM19, BBC16]. Outliers [AY88, AC93, AES06, AHT13, CPR07, PR03, Vog99]. Output [OM17]. Overdifferenced [CD94, SP01, HC00]. Overdifferencing [Har81]. Overdispersed [MA20]. Overdispersion [MS08a]. Overfitting [GJ06]. Overlapped [WWW12]. overlapping [TNZ04]. Overparametrized [Ver87]. Oxford [Cha16a, Hal12, Nea13]. Ozaki [Omb13]. Ozone [BHL94].

P [Wil16, sC88]. Packet [CN17]. Page [Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f]. Pages [Hal14, Leo13, McC15, Nea13, Ter14, Bos16, Cao19, Cha16a, Kar16, Kil18, Lat17, Lu18, McL17, Mcn15, Neš16, Omb13, Pou16, Pou17, Qui15, Rao14, Rao16a, Rao16b, Rao17, Tur18, Wan21, Zha13]. pair [LNVK02]. Paired [BG95, CF98]. Paleoclimate [Mil19]. Panel [BiS17, Bra16, LPPS20, Sta20, Wes19a, Wes19b, Che06, GBY17, HH12, KT16, Sme15]. Panels [EV19, HMW19, DHT14, HCT04, WNS22]. paper [Ano20g]. Paperback [Rao16a, Kar16]. Papers [ACN23, RSVM15, Sav15, Tay24a, Ano09a]. Parameter [AG95, ANW93, AM05, AD04, Azz82, AF91, BL01, BW00, BT94, BM81, CT20, CT87a, Che91, DK17, DH17, GAHT01, GRS97, GZF86, HW89, HT99, HB94b, HR95, KL09, Kul85, Kur17, Lee16, LJ23, Lii85, MM93, McL93, Mur85, NP90, Rei94, SS95, SS96a, SF05, TZ19, Tua92, Wan05, CNR17, DG98, DGH06, Erc11, FG04, HM13, HDB98, HD99, HB01, Iac10, KH04a, KPY22, Ken12, KL13, KK12, MRT07, Rao11, RT09, RB13, WC14, Yam11]. parameter-driven [WC14]. parameterization [MZ06, PS99]. Parameters [AT87a, And93a, BM91, Bos96, Chi91, DC23, FR83, HW95b, HH81, HH16, Kab83, Kni87, LW91, MP18, PT05, SH88a, Tua84a, Tua88, Wat85, ZW08, BM03, BM09, HL11a, HLM03, LOS12, McC13, Mon98, OV04, PW05, QS00, Rao08, RDB14, WSS04, Wie13, WHY22, ZT06]. Parametric [AV05, Cam87, DC23, DZ18, FRR17, GQ17, Hid97, KLN18, Kok12, Lee16, MLS97, NBQ16, Pre98, TK93, TZ19, XHN17, vS94, Aud05, BB07b, dBCO12, DKV11, DPT12, Efr14, FB13, HS11, Ioa22, KM09, KM03, Kom99, LS03a, LLG09, NM11, NS13, Sun14, Tau23, Vel03, WC10, YFL^{+}14, ZSW22, ZZL14]. PARCOR [ZP20a]. Paretian [TZ02]. PARMA [FRS11, TAM11, Wyl08]. Parsimonious [ASM21, DdM13]. Partial [And92, And93b, Bai93, Cho91, FK04, MZ06, NB83, SL04a, Yaj85, Yu07, Hos01, KPY22, Kur11]. Partially [AK90a, SS96c, LLY14]. Particle [KS08b, DZ17, KOZ12]. particular [LL06]. partitioning [Olm23]. partly [BB07b]. Parzen [Bha86]. Past [LX96, PM92, VY90, Li06]. patches [Pen07]. pathwise [DZ17]. Patterns [BS07, BI12]. Paul [LT18, Nea13, Neš16, Pap22]. Peak [vS94]. Peak-Insensitive [vS94]. Peaks [KS94, Pro23]. Pearson [Nag03, ZT18].

Peccati [Leo13]. Pelagatti [Pou16]. Peña [HC04]. Penalised [SR17]. Penalized [ASJLZ19, BCK23, PW05]. Percentage [PS06a]. Performance [BM89, PT86]. Period [LL95, MK93, Tia88, GA04]. Periodic [AG95, AAD22, AV93, BL01, BH94, BHL94, Bos96, BF96, DD15, GGSW20, HG91, LB00, McL94, McL95, Mur85, Sak91b, SL04a, Sha08, Ula93, UT12, AB09, AMZ13, DH13, DdM13, GLL06, GKL11, Jen12, LL06, LR02, MA24, Mar99, PD12, SO97a, TS14, UD09]. Periodically [ACL01, AM05, DM96, Dun81, KKJ18, LL05, MK93, Tia88, BLL09, dBCO12, GA16, LLS08, Len16, SA07]. Periodicities [Ked87, FL00, HW99, WWW12]. Periodicity [CT87b, GL21]. Periodogram [Bri80, Has93a, HD96, HB93, HB94a, JvS95, Kro82, LO16, Lob97, PP99, Rei94, SP08, SP02, ZGH80, AO09, BDL08, DRY21, FMS02, Hen01a, LLMR08, Li14, RMSF10, SH12, Vel07, WWW12]. periodograms [Wal00]. periods [GLL06, GA08]. Permutation [RT22]. Perron [LN99]. Persistence [EV19, Jas03, LTK07, MMT05, Rav89, GMP15, KYP20, SK09, Tay05]. persistent [Bou08]. perspective [GB99, RY23]. Perturbed [Oza82]. Petros [Nea13]. Petrova [Ano21a]. Phase [DZ18, LL95, KP20, PD12]. PhD [RTW14]. Phillips [LN99]. Pinkham [Bha82]. Pitman [Hal94b]. plane [GB98]. Plots [HH81]. Plug [HD99, RS17]. Plug-in [HD99, RS17]. Point [ADL18, GL21, Ger18, KH98b, Rig92, Rig96, BFK19, CF24, GKD21, HH12, HK08, LJ23, NLL12, NSK^{+}11, PZC14, RT02, SP12, WC10, Yam11, YD12]. Point-Process [GL21]. Points [Rai96, AMS^{+}17, Bet16, BH03, JFML13, NAJ12, PP16a, Roz01]. Poisson [AF16, DD15, DFR21, FM04, JJL12, MA20, SBS19]. Poisson-Sampled [DD15]. polio [KL09]. Politis [Aue22, JMP16]. pollution [DA14]. Polson [Nea13]. Polynomial [AM07, HH05, Mas96, WH11, Kak06]. polynomials [Bla14]. Polyvariograms [CA99]. Pooled [SP02]. Pooling [Mar07a]. Pooling-Based [Mar07a]. Population [Tsa88]. Porat [Wal95]. Porter [HDB98, NH19]. Porter-Hudak [HDB98, NH19]. Portfolio [LDH19, WTSL17, DP20]. Portmanteau [Arb08, BR06, HK14, Kat08, LM08, MKS22, MA24, WL05, XZL20, FR07, GF15, Kat09, Kat12, MM12, VW15, Zhu13]. Position [Kab94]. Positive [HHI18, AF22]. possible [CHLT15, PRW04, Sha11]. Possibly [GH19a, GMR04, KST95a, Zho92]. Postmodel [BM10]. Potentially [SHL22, HC00]. Power [CG82, GLML16, GL96, HL06, Mar20, TLG93, DK13, KT16, Kur11, MKS22, San17, SH12]. Powerful [HH16, Duc05, LKN05, WX18]. pp [Aue22, Wil16]. practice [Zaf08]. Practitioners [Tur18]. Pre [Oke98]. Pre-testing [Oke98]. Predictability [AiL15]. Prediction [ATT03, Bha93, BDD95, BN92, CH15, FR83, Fra84, GR81, HPY02, Har81, HN21, HN93, IY03, Kab87, KH99, Kab99, KH01, KS08a, Kar01, KGY18, LR88, LB00, O’B87, PP88, Pem87, PT05, PS00, Rav89, Ray88, Ray93, SLL^{+}20, SHLL96, Ton82, TV83, WW15, AMZ13, BGT21, Bon01, Bon05, DPT12, GV10, GKL11, GJ16, Ing01, KH04a, KCW22, Mar12, MZW09, QS00, RT17a, RY23, RR09, TP03, Vid04, Vid09, YLC12]. prediction-residual [GKL11]. Predictions [Stu01]. Predictive [BM89, Kab93, LTZ20, Olm23, PRR04]. Predictor [Sch96]. Predictors [SH88a, SP01, Ing01]. Preface [BDP21]. Preliminary [PW84, Sai86]. Prescott [SdJ22]. prescribed [SA07]. Presence [AC93, Fra84, Hid92, HG91, KS08b, Kra16, NP90, WS20, AMS^{+}17, Bar00, BEL06, BH01, CT06a, Efr14, Efr19, HL02, HMV08, Iac10, MP16, NKC15, NR11, PRW04, Psa01, Sha11, kTR98, TT97a]. Present [dBC07, PW05]. Press [Aue22, Cha16a, Hal14, Kar16, Kil18, Leo13, Lu18, McL17, Mcn15, Nea13, Neš16, Omb13, Pou16, Pou17, Qui15, Rao14, Rao16a, Zha13]. Prewhitened [Hir06, XL02]. Price [CHS17, Hal14, Leo13, Omb13, Kar16, TY10]. Prices [FFGM15, MFM05, KPRN03]. Priestley [RTW14, RW17]. primary [KPRN03]. Primer [Fok11]. Princeton [Neš16]. Principal [Cub95, Joy92, KKJ18, Seo24]. Principle [Hos05]. print [Ter14]. prior [SL00, TK08]. Priority [HB94a]. Priors [SR17, GH03]. Probabilistic [Ter14]. Probabilities [BBKL17, KS94]. Probability [Bos16, Gor18, HH93a, Lu18, Rob87, SH90, Tur18, Mcn15]. probit [GBY17]. Problem [ACL01]. Problems [Rig96]. Procedure [dBC07, Chi91, CN86, Ish84, ZG85, ZG88, KP10, KPY22, RS17, Vid09]. Procedures [AiL15, Bri80, TP85, HKV24, KYP20, KP21, Vog99]. Process [AT87a, AOA87, BC01, Bar87, BS15a, Bat83, Bha89, BM91, Bol88, CR90, Chu96, DLLN15, FLO06, FRR17, FHW94, FR83, GL21, GL19, GZF86, Had95, HHI18, Huz81, Kab83, KT94a, Kan87, LDN19, LH83, LQ20, Log04, Lüt85, Mic20, Och83, Oke98, PF95, Pic82, Rig96, SS89, Sak91b, SH87, She88, SRHZT83, SW86b, SC97, Ton81, TSL08, UH95, Wei86, Wei85, Yaj85, BB14, BFK12, EK13, Had04, HK22, HS04, Huz07, KY09, KLN04b, Lat98, MP10, NLR16, NSL07, PBSO23, RT17a, RMSF10, SL00, SLL97, SP12, TK08, Tew18, Vid04, Wal00, ZWZ11, ZB05]. processe [LL18]. Processes [ACL01, And87, And89, AV93, AM07, Azz82, Bai94, BSS17, BF97, BP07, Bos96, BLT92, Bro95, BB99, BL19, CN17, dBCRT19, CG19, CR90, CL95b, Chi91, Cho91, CL00, DJM86, DG20, DSW80, DF80, DM96, DD15, Dic82, Eng84, FT85, Fei20, Fin16, GS20, GLML16, GL96, GH19a, GZ15, GMR04, GZ88, GZW89, GZW94, HH20, HK90, HT99, Hög86, HR95, HB05, Jas03, JLMB20, JO06, Joy87, Kab93, KH01, KS08a, KOV94, Kak96, KPT19, KPT21, KPS04, KT94c, KH98b, Kun97, LL05, LX01, LM95, Liu12, MP90, Mar95, Mar92b, McC15, MM93, MM91, Mil95, MVS87, Mor83, Nas93, NP90, NR93, Ott88, Oza82, PC05, PS06a, PS01, PI22, Pet89, Por87, PP16b, Rai96, Ray93, Rig92, SS90, ST03, SSX18, Sch16, SP08, Shi93, SL96b]. Processes [SM06, SR17, Spa93a, Spa93b, ST85, Str96, SHLL96, ST91, Tay03, Ter14, Tia88, TP85, TV83, TC05b, TC07, TT82, Ula93, VY90, Ver87, VDO95, Vil01, Wal95, Wan93b, WT88, XY89, Yaj89, Yu07, ZM01, Zha04, AB09, AMS^{+}17, And97, AG08, AR00, ASK15, Bar24, BDS12, BC02, Bha97, BLL09, BC12, BFK13, BL13, BL21, BFZ02, BMH08, BH08, CL97, dBCO12, dBCCO22, Cha99, CZ12, Cha05a, CP17, CADF11, CD12, CWDL97, Cra03, DGK14, DZQ10, Dat97, Deb11, DS04, DKV11, EDD17, EMNR09, FF13, FF10, FZ09, GAHT01, GA16, GJ16, GS13, GB98, HK17, HL11b, JJL12, Jen12, JPP15, JR22, JP99, JM04, Kak99a, KOD09, KL09, KM03, KP13, KM21, Kri22, KOZ12, LLOS08, LLBM^{+}11, LB11, Lie05, MORS21, Mar12, MKN22, MZW09]. processes [MY02, Nan14, NR07, NSK^{+}11, Nie11, OV04, PR98, PRR04, PPS14, Pos08, PV15, RDB14, RT02, RT09, SO97a, SF11, Shi98, SB19, SO97b, SA07, SP18, Swe03, Tam09, TC13, Tri12, Tro13, TC05a, Tsa07, Vel07, VADG04, Vij06, WS11, Wan09, WP14, WHY22, XCG24, Zaf07, ZB02, ZC12, ZBD06, ZJ06]. product [LP04, Tak24]. Products [Eng84, Liu92]. professional [Zaf08]. Professor [HKK23, LT18, Pri94, Tay22b]. Profile [LW91]. Projection [XA99, ZT97, MS00b]. Properties [And92, AC18, BL01, BI09, Dav91, FRS11, FKMN02, HT99, Huz88, Kak96, KPS04, KTL00, MP90, Mar20, PS01, Por87, Pos08, Sai86, SH87, SH88a, SP01, SK96, ST91, Tan87, Ter14, Tri06, Yaj85, BCFFT17, Cha05a, Erc11, FG04, Gir07, LN99, Lie05, RMSF10, Tam09, WR08, Zaf08]. Property [CR90, Fin85, Had95, Kab99, Kul85, Pol94, Tia88, PF21]. proportion [AD23]. proportional [GG08]. proposal [Mon98]. prototype [NSK^{+}11]. Prototypical [MP87]. proving [Lie05]. Pseudo [Bau05, IP20]. Pseudo-Maximum [IP20]. Publication [NH19]. Publish [Ano20g]. Published [Aue22, Bos16, Cao19, Cha16a, Hal14, Kar16, Kil18, Lat17, Leo13, Lu18, McC15, McL17, Mcn15, Nea13, Neš16, Nun20, Pou16, Pou17, Qui15, Rao16b, Rao17, Ter14, Tur18, Wan21, Wil16, Yao20, Omb13]. Publishers [Zha13]. Pure [BK03, CL95b]. Purifying [RT92]. purpose [HKV24]. Pursuit [XA99].

Q [Tur18]. QMLE [AF16, GSS17, IP08, SW21]. Quadratic [Abr87, GSS17, HIP87, SF93, NHCLP08]. quality [LZZ22]. Quantifying [NAJ12]. Quantile [BCK23, CS08, CP16b, KMX17, Li14, LO16, MR19, BH10, CSD12, GMRO11, Kim15, KOW22, PV15]. quantile-based [PV15]. quantiles [Jir16, Muk99]. Quantitative [Neš16]. Quarterly [Pon06, ZC19a]. Quasi [AB09, ABT18b, AS21, CF14, Huz88, Li20, PS21, Pet19, TC05a, AG24, CG07, CLL14, GL20, PD12, Per04, VP12, Zhu13, Ano21a]. quasi- [AG24]. Quasi-Bayesian [Pet19, Ano21a]. Quasi-Likelihood [ABT18b, Huz88, CF14, CG07, Per04]. Quasi-Maximum [TC05a, AB09, AS21, PS21, CLL14, GL20, Zhu13].

R [Bos09, Lat17, Rao14, Che09, Kil18, Lu18, Nun20, Rao17, Bos10]. R-INLA [Rao17]. radial [BEvdW12]. radioactivity [DF11]. Radius [Liu92]. Rafal [Ter14]. rainfall [RCLM^{+}11]. Ramsay [Cao19]. Randall [Rao14]. Random [Abr87, AHS06, CL95b, CT10, DM96, Dic82, Elt94, FT85, GKY18, GZF86, Hal92, HB05, KOV94, LPPS20, LH83, LM95, NLR16, NQ80, Pou88, QN81, Ray88, SSW15, Ste05, XY89, BHL09, BH01, Efr20, FK13, FL04, FPS21, HP14, Hoy24, Ing01, KL09, KL10, Kri22, LJ23, MORS21, MS00a, OV04, PPS14, RDB14, RT17a, SLL97, TY10, WC10, Yua00a, ZWZ11, ZBD06, Fei20, Leo13]. Random-Coefficient [HB05, LPPS20, PPS14]. Randomly [LT95]. Randomness [HIP87, QN82]. Range [AES06, Ber07, Cox91, Ger18, GRS97, IL19, KLN18, LDN19, LKB15, PC05, Ray93, VDO95, Yua00b, AVF98, AG16, BP12, BT13, Bet16, FP18, GAHT01, Ger21, KP15, KP20, LLBM^{+}11, LRS21, LG11, MORS21, MS00b, Nie05, NSL07, PBSO23, RT02, Sha11, SK09, ST05, TT97a, Tew18]. ranges [KMX17]. Rank [And08, CMK05, CRT15, Duf81, FA03, HIP87, LLZZ22, RA92, TSL08, Kur11, Muk99, SL00]. Rank-based [And08]. Ranking [Duo84]. Rao [Cox94, Wil18]. rare [GKL11]. Rate [Eva80, GRS97, HK90, Kni87, KST95b, RB13, AMS^{+}17, CWDL97, Jir16, LL12, ZT06]. Rates [Mas96, FB13, LP14, PS99]. Ratio [DNL81, FM96, LM04, MN95, RA92, ST99, TZ02, CG07, CD09, CD12, NKC15, TNZ04, ZL12b]. rational [HT10, JS24]. Ratios [Ver87, VN17]. Raton [Lat17, Nun20, Yao20]. Ravishanker [Lat17]. ray [BEvdW12]. Real [AHL^{+}18, Tri11, ZC19a, ASK15, KL10, SW21, SA07]. Real-Time [AHL^{+}18, Tri11, KL10, SW21]. real-valued [SA07]. Reale [McL17]. Realizations [Bar87, She88]. Realized [BCCR19, LW19, KOW22]. Recognizing [CD94]. Reconsideration [Cor95]. Reconstruction [Pos20]. Records [DN95]. Recurrence [Gor18]. Recursive [AH92, Bon01, Bos96, Gri91, Had95, HLHT94, Hua90a, Kav89, Kit81, LH96, LB00, McE18, OM17, PT82, PPT93, Rod13, SS01]. Reduce [De 98b]. Reduced [RA92, KA08, THN24]. Reducing [KT10, LS03a]. Reduction [Zha92]. Redundancy [McL93]. Reference [PW89]. refined [LL12]. Regenerative [BCT15]. Regime [BBKL17, HP23, HLHT94, BP11, KOD09, KC10, KR98, PS06b, ZS01]. regime-switching [KC10]. regimes [Cav14b, PS03, Spe10]. Region [Pic82, LP19]. Regions [HW95b, WW15]. Regression [BSG18, BCK23, CT01, CN86, FK87, Hal95, HK86, HV92, HR88, Has82, Has93a, Has93b, HH05, HKK15, Kas82, KH94, KS19, LTZ20, LPS99, Mar00, Mas96, NVS06, Qui89, SL96a, Sak93, ST04, SP02, Shi88, Sib01, Stu01, Tan81, Tan87, TS94, USMS83, UH95, Wal87a, ABT18a, AO09, BP18, Deo97, FB13, FLL13, GH03, HPW17, HD99, KPY22, KMX17, KW12, LLG09, LOS12, MRT07, Nie05, Olm23, PW05, RP20, Sch98, Tsa07, Wan24, WC14]. Regressions [Has00, LX01, Mar95, Sch22, Wes19b, Bra05, CL97, DGP15, Kim15, KF08, KA07, Lub99, Mil10, SL04b]. Regressive [HZF93, Mur85]. Regressors [Kur17, Mil10, SL04b]. Regular [NP90, Sza22, Sza23, DS04]. Regularity [TV83]. Regularity/Singularity [TV83]. Regularized [MMM22]. regularly [MS00a, MS01]. Regulated [Tro13]. Reinsel [Wil16]. Related [HB05]. Relation [Sak91a, YL22]. Relations [New80, Bon01]. Relationship [DF80, PI22, PS95]. Relationships [Mil19, WW17]. Relationships* [BK07]. Relative [EF06, Gir07, GL96, Sai83, KPRN03]. Relevance [Kab99]. relevant [DQ23]. remainder [Che06]. Remodelled [Bha82]. remote [KC11]. remote-sensing [KC11]. renewal [AMS^{+}17]. Reparametrization [MS92]. Repeated [AF91]. Replicated [Azz81, Deg87, PR88]. Representation [Bha89, HNT23, Leo13, BC12, TC13]. Representations [GZ15, Gue87, SM06, Tho19, KP15]. resampling [ABT18a, Len16]. rescaled [FLL13]. Research [Ano94]. Researchers [Tur18]. Residual [CP17, Dit00, Fas00, Hal95, HZF93, HR93, LM94a, ML83, MMT98, MW16, SL96b, Tan87, DdM13, GKL11, LLOS08, Shi98]. Residual-Based [Dit00, MW16]. Residuals [Bai93, DSW80, Kav89, LL92, Pes07, Rob87, Vel94, Yu07, GH03, TS14]. Residuals-based [Pes07]. Resolution [HQ89, MTW04]. Response [BW18, MR19, Wil17]. restricted [CD09, CD12]. restriction [WSS04]. Restrictions [SS95, SS96a]. Results [AV93, IP20, Oke98, Wal95, YR95, BGT21, CG07, FL04, GB99, GB98, HNT23, LP10a, OV04, Sbr11, TAM11, Wal00]. Return [SLL^{+}20, SCW19, YL22]. Returns [AiL15, HL18, Wie13]. reverse [LT03]. Reversed [LL92]. Reversibility [BD92, Psa08, Pro23]. Reversible [VADG04, Bra21]. Review [Ano95, Ano98a, Ano99a, Ano99b, Ano00a, Ano01a, Ano01b, Ano05d, Aue22, Bas98, Bos09, Bos10, Bos16, Cao19, Cha16a, Che09, Cox94, Erc08, Fea05, Hal12, Hal14, Kar16, Kil18, Kok13, Lat17, Lay98, Leo13, Lu18, McC15, McL17, Mcn15, Nea13, Neš16, Nun20, Omb13, Pap22, Pou16, Pou17, Qui15, Rao05, Rao10a, Rao14, Rao16a, Rao16b, Rao17, Ter14, Tur98, Tur11, Tur12, Tur18, Wan21, Wil16, Yao20, Zha13, Jan05a, Jan05b, Mil04, Whi05]. Reviews [Ano82, Ano01c, Ano05a, Cha05b, Dod05, Mil05, Pou05, Pri04, Qui05, Rao04a, Rao04b, Shu05, Ter05, Ano97a]. Revised [Neš16, Rao16b]. Revisions [PD02]. revisiting [AVF98]. Richard [Fei20]. RINAR [KY09]. Risk [GLP10, Hog18, Neš16, WTSL17, DP20, Efr19, EFT16, TY10, YL22]. risk-return [YL22]. Rissanen [Kab87]. river [EM08]. Robbins [FRR17]. Robert [Nun20]. Robust [ABT18a, AES06, BKS97, BMY99, BCT15, BB07b, BM81, CH21, Duc05, EF14, FHK20, Fra84, Ger18, Ger21, Hen01b, HMW19, Kas82, KL11, KL13, LLBM^{+}11, LXT20, MG00, MY02, PBT00, Sha08, SBS19, Tan81, Art02, BDS12, HLT10, Li12b, McC13]. Robustness [Bha97, GT19b, HD96, HL02]. Roland [Lu18]. role [OJHO00]. Root [AK90b, AES06, Bea18, BK03, Bre94, CISG16, CP03, CL95b, GRT17, GL96, Hal95, HR02, HL06, HCH00, HR04, HMW19, HH16, KLN04a, OT98, PF95, SS16, SH87, SS96b, SS01, XX18, YR95, AG08, CT08, Cha15b, DK13, Fos13, GPRV00, HL02, Ioa10, Kap05, KT16, KP08, LLS02, Lar98, LD04, LT03, LLT14, LLZZ22, Lub99, MR12, NK24, Ott21, PSU08, PRW04, Psa01, Rod13, San18, Sen07, SL97, SS98, SF98, Sol04, Sto19, Swe03, kTR98, Tay05, Vog99, Wes13, WNS22, Xia01, Yab12, YLC12]. Roots [AT87a, AT87b, BF96, dBC07, GJ06, HK90, Kar16, Pau84, Pon06, RA92, ST99, SLN99, Tay98, Tay03, CHLT15, FM98, HL02, Joh03, LNV98, LG11, NR11, ZS17]. Rosenblatt [BGT21, BDP21]. Rotational [SS95, SS96a]. rounded [KY09]. rounding [LB11]. Rüdiger [Neš16]. Run [YK06, CCGA13, DHJ12, RS17, Tau23]. running [Mau11]. Runze [Wan21]. RUR [AES06].

S [Aue22, JMP16, Lat17, Nun20, Rao14]. Sample [AC96, BL01, Cha19, Cha95, CL95b, Del96, Har81, IP08, KT94a, KS94, MS01, Por87, SH87, ST88, San17, Tia88, Vel94, Wal95, Yaj85, Yua00b, Cha05a, DRY21, FG04, Kak99a, Kur18, Li24, Tay05, TNZ04, Wal00, WJM11]. Sampled [DD15, LT95, STA03, Sin93, Cha15b, FF13]. Sampler [MT94a, PS99]. samples [Rao18]. Sampling [BL19, DM96, KS08b, LM95, MW16, NP90, SS95, SS96a, BP03, BFK12, BFK13, HV08, WR08, ZT06]. Scalable [SB19]. scalar [AV08, PT23]. Scale [AMS^{+}17, BW00, FZ23, JM04, KS18, LLBM^{+}11, Zaf08]. Scargle [LLMR08]. scatter [ZW12]. scattering [BEvdW12]. scheme [KH04b, Len16]. Schemes [LM95]. Schur [BC95]. Schwarz [CF24]. Science [Rao14, Ano94, Wan21]. sciences [CH11, Rao12b, SO12, Ano94]. SCOMDY [KL13]. Score [AY88, SM13b, Yab12]. scores [Muk99]. Sea [ZC19b]. Search [LL95, Oke98]. Searching [PR03]. Seasonal [Aka80, BS98, Cub95, Fra05, Has94, HR04, Ish84, Joy92, KM90, KL23, KP95, LT03, MP87, NV96, PC05, Peñ84, PD02, Pon06, ST99, kTR98, Tay03, TV83, ZHHH22, AR00, Art02, Cha99, CGM08, DLPP14, GB06, GG07, KM04, LD04, NR11, PR09, TvV02, Jan05b]. Seasonality [FN97, Hal13]. Seasonally [dBCRT19, New80, Pfe94, dBCO12]. Second [CL00, HT99, Hoy20, Kil18, Pap94, Tam09, Tur18, YR92, DR11, TK08]. Second-Order [CL00, Hoy20, Pap94, YR92, HT99, Tam09, DR11, TK08]. Sectional [Elt94]. segment [DG98]. segments [HLZ23]. Seismic [DN95]. Selecting [BPT02, Hua90b, PBT00, PSSS09, YL91, ZT97]. Selection [CN17, Che95, De 01, Duo84, GH19b, HO84, HT93, Kap01, KC96, LX96, LM95, Maï12, ST04, TY00, Tua88, AR06, BBK23, BB07b, BM10, Bra22, BH08, Cam04, CLY17, CGP22, EMNR09, HD99, Hur01, KP13, LLY14, MS00b, PW05, RS17, Sko01, Sou07, Unn04, VADG04, Wan05, ZS01]. Self [TZ19, YL20, ZLP19, ABT18a, Bar00, Bet16, CS08, HVS15, LZ18, LPS23, RCLM^{+}11, ST05]. self-exciting [CS08]. Self-Normalization [ZLP19, HVS15, LZ18]. Self-Normalized [TZ19, ABT18a, Bet16]. self-similarity [Bar00, RCLM^{+}11, ST05]. Self-Weighted [YL20, LPS23]. Semi [AV05, DC23, DZ18, Lee16, Sun14, TZ19, BB07b, KM03, NM11, Vel03]. Semi-Parametric [AV05, DC23, DZ18, Lee16, TZ19, Sun14, BB07b, KM03, NM11, Vel03]. semimartingale [YFL^{+}14]. Semiparametric [ASJLZ19, AR00, Art02, BS98, GRS97, Hid92, HB94b, IL19, Jen04, LS03b, Nie05, Rao10b, TS94, Wan09, Bra22, GLL06, HB01, Hur01, Vel99]. sensing [KC11]. Sensitivity [ADD19]. Separability [CKR18]. separable [MY04, Wan05]. Separate [MMH88]. Separation [MIN^{+}16]. September [Tay18c]. Sequence [LL05]. Sequences [KTL00, MK93, Nas94, BGT21, PC23, Wyl08]. sequential [GKD21, KP10, Sme15, Ste05]. Sequentially [PBT00]. Serial [And93b, APH86, CT92, Cha87, Del96, Duf81, HIP87, LS03b, MPR91, BPN12, CF24, DGP15, Duc05, JT03, KF08, LT17, SM13b, VVD18]. Serially [CN86, KT16]. Serie [JMP16]. Series [AY88, AC93, AF16, ARS86, ABT18b, Ala89, And93a, AC96, AM05, Ano97b, Ano99c, Ano02, Ano03, Ano04, Ano05b, Ano06a, Ano07, Ano20g, Ano21a, AVW16, Aok91, AV05, APH86, AC18, Aue22, Azz81, BS07, BB95, BC97, BC95, BS15a, BO05, BH91, BH92, BG95, BG20, BK19, BLL05, BI09, Bos10, Bos16, BR06, BD92, BS98, CS15, Car85, Cha16a, CT96, Cha91, CD94, CAP94, Che09, CG82, CKR18, CW82, CS84, CS87, Cor95, Cox94, DN99, Deg87, DK17, Dit04, Dun81, DH17, Elt94, Eva80, FK87, Fei20, Fin84, Fin85, FK04, FP16, Fra84, Fra05, GR81, Gab88, GT19a, GCK99, GPH83, GT93, GA01, GRS97, Giu17, GHHK18, GOV19, Gra95, GJ80, Gra82, GH91, GZF86, HHP84, HO84, Hal94a, Hal95]. Series [HW95a, HS05, HKK23, HT86, HV92, HR88, Has82, Has93b, Has94, Hid97, Hin82, HH93c, Hog18, Hok83, HN80, HR93, HV99, HC20, HB93, HB94a, HB94b, Huz88, JL83, JWW99, Jan82, JS90, JA81, KMS15, KPT21, KE88, Kas82, KKJ18, Kil18, KBB90, KC96, KK20, Kit81, KT01, KP93, Kra16, Kul85, KL98, Kum86, LW91, LL92, Led90, LL95, Leo13, LR88, LT18, Li84, LM94a, LX96, LX01, LS03b, LWL93, Lim87, LPS99, Liu89a, Liu89b, LXT20, Lob97, Lu18, Lüt82, LM94b, MS08a, Mar00, Mar92b, Mas96, MMH88, MMT05, McC15, MT94a, MT94b, ML83, McL98, MV03, MG93, Mil84, Mil19, MA93, MP84, Neu96, Nie15, NH19, NBQ16, Nun20, O’B87, OJ03, OT98, Oza82, Pap94, Pau84, PS89, PBT00, PZ17, PT81b, PT86, Pos20, Pou89]. Series [PM92, Pou16, Pou17, Pre98, Pri80, Qui15, RG80, RRW83, RG89, Rao10c, Rao14, RS19, Rob83, Rob87, RT17c, Sai83, Sai86, SL96a, Sch96, STA03, SPA20, SBLS07, SR88, SR91, SR92, Sha08, She87, SS96b, Shi88, SS82, SPM19, Smi08, SH88b, ST87, Sto85, Sto87, SW86a, Stu01, Tan81, Tan84, TK93, Tay98, Tay03, Tay20, Tay21b, Tay22a, Tay23c, Tay24a, Tay24b, TA88, TP82, Tjø86, TM93, TS94, Tsa88, Tsa89, TY00, TW89, Tur11, Tur18, UH95, VWR87, Wal87b, WT19, WS20, Wil16, WL05, WTSL17, XA99, XHN17, Yak87, YB06, YL91, Yua00b, ZT94, ZW94, ZLY06, ZLP19, Zho92, ZP20b, ABT18a, AR10, AD23, APS20, AS00, ASM21, And08, Art02, AHT13, AH13, BCD18, BBC16, Bar00, BDL08, BM13, Bet16, Bon01, Bon05, Bos09]. series [Bou08, BC12, Bra05, BS02, Bro07, BFK19, BH01, BH03, CS08, Cai11, CSD12, CH14, CCGA13, CLY17, Cha16b, CD09, CCY16, CH21, CF14, CT10, CF98, CG11, CH11, DG98, DGH06, DC01, DRY21, DE07, DLRY08, DdRSK21, DFF23, DL15, DH13, DPT12, DP10, DA14, DLPP14, DR11, Fan05, FB13, FLL13, FK99, FP12, FL00, FKK12, FM04, GG08, GF15, GM15, GM23, GB06, Góm07, GAP09, GKL11, GLN15, GSO^{+}17, HPY02, Hal13, HNT23, HM13, HM03, HP23, Has13, HCT04, Hon97, HL11b, HR15, Hos01, HW99, HDB98, HD99, HC00, HB01, Hur01, HK08, JMP12, Jan10, JFML13, JW16, Jin18, Jir16, JT03, JT11, KP15, KYP20, Ken12, KH04b, KC10, KL11, Kom99, KGY18, KL23, KXS^{+}12, LC03, LLS02, LM00, LZ18, LD04, LV00, LLS08, Len16]. series [LLRR^{+}21, LNVK02, Li98, LLY14, LRS21, LL97, LPZ15, LTT18, LZZ22, LLS97, LOS12, Lub99, Mar99, MTJ14, MGRM10, MY04, Mau11, MJ12, MT15, ME98, MIN^{+}16, Mil10, MS00b, MRT07, Muk99, NM11, Nie05, Nis09, Omb13, PZ04, PD12, Par13, PT04, PT23, Pen07, PR03, PRC03, PPS20, PR10, PS00, Pro23, PSSS09, QR98, Rao10b, Rao18, RY23, RB22, RS17, RNI13, RT22, SK23, Sbr11, Seo24, SL97, Sme15, ST05, SO12, Sto19, Swe22, Sza22, Sza23, TKOP20, TNH23, TW02, TM98, TAM11, TWVB00, TJL23, TT97b, Unn04, UD09, VP12, Vel99, VW15, VY16, WWG09, WL11, WP21, WD10, WC14, WX18, XZ22, XLT23, Yau12, ZKP22, Zha24, ZJN24, ZP20a, Zho12, Zho13, ZJ06, Hal14, Lat17, McL17, Pap22, Bos09, Jan05a, Rao10a]. Series-Theory [Rao14]. Set [BC97, Fas00]. SETAR [DG20, De 01]. Sets [HT99, KC11, Kur18]. setting [Sta20]. shape [GLL06]. Shephard [Cha16a]. Shibor [SHL22]. Shift [Bai94, TSL08, RGLA11]. shifting [TT97a]. Shifts [LKB15, KYP20, LLS02, LM00, McC13]. Short [Deg87, Ger18, CCY16, KP13, LLBM^{+}11, Tau23]. short- [KP13, LLBM^{+}11, Tau23]. short-memory [CCY16]. Short-Range [Ger18]. Show [HH81]. shrinkage [Gao97]. Shrinked [PS06a]. Shumway [Nun20]. Sided [Sch96, Hon97]. Siem [Cha16a]. Sieve [CP03, MP18, MK15, BKM21, FP18, Pos08]. Sign [AS00]. Signal [FM96, MT15, BM04, Iri02, SP12]. Signal-To-Noise [FM96]. Signals [Gor81, RT92, GOP^{+}12, LSSC16]. signed [KT11]. Significance [BK03, HR93]. Significant [LLY14]. similarity [Bar00, Lie12, RCLM^{+}11, ST05]. similarity-based [Lie12]. Simple [Bre94, Has00, Liu89b, SRHZT83, Arv14, BCD18, LR21, Sha11, Vid09, Vog99]. Simplicial [KLM16]. Simulated [MVS87, HLM03]. Simulating [Cra03]. Simulation [ASK15, Com96, SA07, SC97, XY89, Cai11, GA16, GJ16, Mil06]. Simultaneous [BKM21, CGP22, HR93, SS90, SK96, Tom87]. Single [Kab94, OM17, XHN17, BP12, BDM98, BC12, DZ17, JFML13]. single-equation [BDM98]. Single-Index [XHN17]. Singular [HN21, Pos20, CADF11, FB21, KP13]. Singularity [TV83, Wan93b, HS04]. Sinusoidal [Qui89, Sak93]. Sinusoids [LT92]. SISO [KM04]. situations [TWVB00]. Size [Pic82, KA08, KT10, LS03a, LN99]. Skew [TCCG19]. Skew- [TCCG19]. skewed [PC23]. skip [Cha15b]. skip-sampled [Cha15b]. slope [Ioa10, VN17]. Slow [LDH19]. Slowly [Joy87, May22, Ott21, TT97a]. SM [LLOS08]. Small [HD96, HI15, BEvdW12, DRY21, TNZ04]. Small- [HI15]. small-angle [BEvdW12]. small-sample [TNZ04]. Smooth [LS06, SLN99, XZL20, ZL24a, BEL06, Bra16, GLN15, Kei03, Kil11, LNV98, San18, Sol04]. Smoothed [HD96, Rei94, SHL22]. smoother [Nag03]. Smoothing [AK90a, DCCL03, Fer90, HT86, HV92, KC96, SS82, DPT12, Góm07, KC11, KD00, KD03]. Society [Leo13]. Software [Pou17]. Solution [Pou86, Pou88]. Solutions [CADF11, Qui82, Vol12]. solving [BM03]. Some [Abr87, BCFFT17, Bre94, Cha15a, DS04, Fin84, FL04, Gab88, GB98, HT86, Hög86, Huz88, JA81, Kum86, Li84, Li06, Liu89b, MP90, NV96, Oke98, Por87, PT86, RRW83, Sai83, Sai86, SK96, STA03, Tjø86, TJL23, Tua84b, Wal00, Bra21, Bri12, CG07, Gir07, Kak99b, KS18, LKN05, LP14, OV04, PT04, Sbr11, Sol04, LP10a]. Sonia [Pou17]. Sons [Bos16, Rao17, Tur18, Wil16]. Sørensen [Zha13]. Sotoca [Pou17]. sound [Iri02]. sources [GJ23]. Space [AK90a, Aok91, BSS17, CG11, DB98, De 91, GT19a, IJ99, LWL93, Mar92a, MS07, PT05, dJ86, dJCCL94, CNR17, DB03, FKD10, GdSF13, Góm07, GAP09, GJ01, Guo03, KD00, KD03, Pen07, PS99, PF21, Pro03, Qia14, QS00, RR09, Sel10, SB19, WS02, WLC12, Pou17]. Space-Time [DB98, GT19a, CG11]. spaced [GM23]. Spaces [CH15, Ott88]. Sparse [XX18, PA23, RP20]. Sparsely [LL95]. Sparsity [KP21]. Spatial [Rao13, Rao16b, ZC19b, DF11, LLG09, Rao08, Wie13, Rao17]. Spatio [ASJLZ19, BWH19, GSTW20, HK20, KC11, Rao12a, Rao17, TCCG19, WH19, AR06, BJR17, Rao08, Rao13, RDB14, RT17b, RT17a, VPWD11, WH11]. Spatio-Temporal [ASJLZ19, GSTW20, HK20, Rao12a, Rao17, TCCG19, WH19, BWH19, KC11, AR06, BJR17, Rao08, Rao13, RDB14, RT17b, RT17a, VPWD11, WH11]. spatiotemporal [DQ23]. Special [ACN23, CH11, HKK23, LT18, NH19, RW17, SO12, Tay24a, BDP21, LPZ15]. Specification [CW99, Has94, MN95, Oke98, Pos05, Smi08, dJ86, DS04]. specifications [BS15b]. Specified [LR88]. Spectra [BDH^{+}18, MP90, YR92, IC05, LP04, ME98, Tak24, TKOP20, WWW12, ZT18, Qui15]. Spectral [Ala89, AY96, BDH^{+}18, BI09, Büh96, CMK05, CD86, Dah83, DNL81, DRY21, FF13, FM96, FRP99, GM85, Gor81, GQ17, HQ89, Has93b, Kim91, KLN18, KST95a, KST95b, KJ85, Kra16, LM95, LO16, Liu92, Mar00, MdS89, MVS87, Neu96, NP96, PS89, PR95, PR88, Pöt90, Pri96, RB92, Rig92, Rig96, SS90, Sak91b, TM93, Tom87, Vel94, WCK12, Wil17, Wyl08, Zha92, ZP20b, ASK15, BJR17, Bha97, Cha05a, DH22, Efr14, Efr19, Efr20, Gir07, HV08, Hen01a, HS04, HT10, HW99, Ioa11, Jin18, Kak06, KM99, Kom99, KXS^{+}12, Len16, Li12b, LL18, LLS97, MS00b, MRT07, PP12, PP16a, RCLM^{+}11, SdJ22, SA07, WJM11, WP21]. Spectral-based [WCK12]. Spectrum [BB87, Cam87, Dro07, HB90, Jan82, Kul85, Mél85, Pos20, RG89, Vel00, vS94, FNV08, Gao97, Ioa10, Ioa22, KP13, Wal00]. Speeds [TCCG19]. Sphere [Leo13]. spherical [LP10b]. sphericity [LTT18]. spot [BL13, YFL^{+}14]. spread [VPWD11]. Springer [Bos09, Cao19, McC15, Ter14]. Spurious [CL97, Mar95, NVS06, BS02]. SQMLE [AA20]. Square [AK90b, Kab87, Ray88, XX18, PZ04]. Square-Root [XX18]. Squared [HT99, IY03, LM94a, ML83, BM04, DZ17, HDB98, Ing01, Kat12, KLN04b, QS00, YLC12]. Squared-Residual [ML83]. Squares [AT87a, Bai94, BC01, Bri80, FRS11, Gra95, Hua90a, HB05, KP90, KS19, LTK07, LH96, LDH19, MW97, Nie15, Pem87, Pet86, PS95, RBY92, TP85, BB12, FG04, HOS15, HL11a, Hil13, Ing01, Ioa11, KHS03, KF08, LM00, LL12, Sto19, TvV02, Wal03]. Squares-Based [LTK07]. SRE [MW22]. Stability [CS11, Lim92, MS08b, Wei86, Wei85, KPY22, KK12]. Stable [De 91, GGSW20, GSTW20, KT94c, Nas94, TZ02, BH08, KOZ12, MORS21, PC23, QR98, ST05, ZC12]. Stage [CGN15, Wri95, Efr19, HOS15]. Standard [Dun81]. Standardized [Vel94]. Star [Kil16]. Starter [Aue22]. State [AK90a, Aok91, CJ82, CK22, De 91, GJ01, HHP84, IJ99, LWL93, Mar92a, MS07, Pen07, PT05, Pri80, dJ86, dJCCL94, FKD10, GdSF13, Góm07, GAP09, Guo03, KD00, KD03, LK21, PS99, PF21, Pro03, PS06b, Qia14, QS00, Rao11, RR09, Sel10, WS02, WLC12, Pou17]. State-Dependent [HHP84, Pri80]. State-Space [IJ99, MS07, PT05, dJ86, GJ01, GdSF13, Góm07, GAP09, Guo03, KD03, PF21, Pro03, RR09, Pou17]. static [GBY17]. Stationarity [AF22, AP95, CL06, Cha91, FT85, Fei20, Kun97, LP10a, Liu92, NVS06, Pic82, Pou86, Pou88, PP16b, VWR87, VN00, AK24, BJR17, BEL06, BFK19, BH03, BH10, CLL14, DR11, HM03, Jen12, Kri09, LS03a, Rao08, Xia01]. Stationary [Abr87, And92, AM05, AVW16, AM18, AM07, BB95, Bar87, Bat83, Bha89, BD92, BLT92, CN17, sC88, DG20, Dic82, FK87, FR83, GP06, GL96, GOV19, GZ15, Hin82, HK20, HA93, HC20, HS11, JL83, JLMB20, KT94a, Kei03, KF92, KS94, Liu89b, LXT20, Log04, MS08a, Mar20, MW05, Mic20, MIN^{+}16, Mil84, Mor83, Nas93, Neu96, Ott88, PS92, PS06a, Pet89, Pou89, PP16b, Qui82, RG80, RG89, Rig92, Rig96, SS89, ST03, SP08, SRHZT83, SM06, SR17, Sto85, SC97, SP18, Swi90, TC05b, TW89, VY90, WT88, Yaj89, Yua00b, ZT94, ZM01, ZL24b, dJCCL94, AN08, ABT18a, AG08, AHT13, ASK15, BGT21, Bar00, BS15b, BG00, Bon01, Bon05, Bra21, CT08, CHLT15, CZ12, Cha05a, Cra03, DG98, DKV11, DP10]. stationary [FHK20, Fos13, GSO^{+}17, GS13, HK17, HLZ23, Kak99a, KM03, KP10, Kle08, Lie12, LOS12, LBV09, Mau02, MT15, MKN22, PZC14, RDB14, RT17a, RMSF10, RS17, SM13a, SHL22, Sza22, Sza23, Tam09, Vel99, Vol12, WS11, Wal00, XCG24, ZC12, Zho13]. Statistic [Ali83, Dah85, Kat08, LM04, Tua87, LTT18, Yab12, ZT18]. Statistical [AiL15, BP12, DMHF12, Kim16, LT83, MT94b, Oza82, PT81b, Rao08, Rao12b, Rao14, Ter14, GLN15, NKC15, Par13, Ter11, TD11, Tur18, Wan21, Zha13]. Statistics [BCT15, Bos16, Dic82, Lu18, Mcn15, PS92, PP99, Rao12a, SF93, Tan84, Tua84b, Tur18, Bos09, CHLT15, GPRV00, Lar98, LD04, Rao13, SM13b, VP12, FA03, Mau11, Rao16b]. Steady [CJ82]. Step [Pem87, Ton82, ZLSY20, EK13, GL20, HW09, KPY22, KGY18]. Stephen [ACN23]. Stephens [Nea13]. steps [PY22]. Stepwise [Pos20, CGM08]. Stochastic [ACL01, BK19, CT01, DN95, DF80, Fin85, GA01, GS21, Gor18, GL19, Jen04, KS08b, KS05, LDH19, MG93, NK24, Ott88, PT81a, Pou86, Rao16a, SS89, Shi88, Tjø86, Ton81, Wri95, Zha13, Cle01, Com04, FK99, GAHT01, HWBD11, HLM03, Lat98, LP10b, McC13, NHCLP08, PA23, San17, SO97b, Sim08, SBS23, Tri12, Tri06, Vid04, WS11, Pap22]. Stock [Hoy20, Wie13]. Stoffer [Nun20, Rao14]. strategies [VADG04]. Stratospheric [BHL94]. Strength [HW95a]. strict [BH10]. Strictly [Qui82, Vol12, Bra21]. Strong [DS91, Mas96, Hid07, TD11, ZZL14]. Strongly [BG20, Hid97, Yaj89, KM03]. Structural [AES06, AH13, BW18, DGJ06, GT19a, HOS15, HR04, HV99, Ott88, PY22, Rav89, Sbr11, Sch22, Ter85, WT19, BHLS11, BH01, BH03, CGP22, CP16a, Fos13, FB21, GA04, GA08, HM03, HLZ23, Kap05, Kat12, KA07, Kur18, Rod13, WE07, WR08, Wri98, ZL12b]. Structural-Factor [GT19a]. Structure [AM80, Bol88, Gab88, Li84, MPR91, RZ10, RA92, SBLS07, Góm07, HW99, Lat98, PZ04, TD11, ZS01]. structures [BEvdW12, JW16]. Student [MMOV24, SBS23]. Student- [MMOV24, SBS23]. Studentizing [DGK14]. Studies [CG82, LT83]. Study [Eng84, HHP84, Lim87, Dit00, SL97, SO97b]. stylized [KS23a]. Sub [STA03, Tay05]. sub-sample [Tay05]. Sub-Sampled [STA03]. Subba [Cox94, Wil18]. Subclasses [DG20]. Subject [MW16]. Subordinated [LDN19]. Subordination [WT88]. Subprime [WW17]. Subsampling [ADD19, JMP12, KW04, LLS08, MJ12, JMP16]. Subset [GR81, Gae00, HO84, PT82, PPT93, PBT00, Tha90, YL91, ZT97, BPT02, MZ06, Unn04, Wan05]. Subspace [Bau05, GH11, SP18]. successive [TvV02]. Sucharita [Ter14]. Suddenly [And93a]. Sufficient [Dic82, Kan81, DRS21]. sugarcane [VPWD11]. Sum [DJM86, Nie15, Yu07, BFK19]. Summaries [LW91]. summary [NSK^{+}11]. Sums [Bai93, Eng84, CP17, Chu12, DGK14]. superharmonic [TK08]. superposition [GLL06]. support [PP16a]. Supwald [Kil16]. surface [GSO^{+}17]. surfaces [KXS^{+}12]. surveillance [AK10]. Survey [HN80]. Survival [Gor18]. Switching [BBKL17, FRR17, HLHT94, Liu12, MT94b, SK96, Smi08, AF22, Cav14a, Cav14b, HP23, KOD09, KFS02, KC10, LK21, PW05, PS03, PS06b, ZS01]. SWQMELE [ZPZ21]. symbolic [MGRM10]. Symmetric [Nas94, HST23, LZ18, TWVB00]. symmetry [LLOS08, PV15]. syndromic [AK10]. System [BHL90, MS23, Nag03, Sim08]. Systems [Li93, PBT00, Pos05, Sin93, Tig85, Wri95, XX18, YK06, GH11, GP02, KWPV12, Kur11, NHCLP08].

T [Cox94, JMP16, GL96, HL17, TP03]. Ta-Hsin [Qui15]. Tables [ZC19b]. Tail [Hog18, BM13, BDS12, Hil13]. tail-trimmed [Hil13]. Tailed [GRT17, JMP16, YL20, EM08, GL20, GZ15, JMP12, KW04, Kri22, LPS23, MJ12]. Tails [Hal14, HPY02, MS00a, MS01]. tangent [Kei03]. Taniguchi [HKK23]. taper [HC00]. Tapered [Dah83, JvS95, MP10, Roz01]. Tapering [AV05, Zha92]. Tapers [CWD00]. TAR [WTSL17]. Tata [Wil18]. Taylor [Nun20, Wan21, Yao20, Zha13, ACN23, Olm23]. technique [GPRV00]. Techniques [Neš16]. temperature [CG11, VN17]. temperatures [EP17]. Tempered [SK23]. Temporal [ASJLZ19, BS02, dBCRT19, GSTW20, HK20, MW16, Rao12a, Rao17, SF11, Sou07, SW86b, TCCG19, TC05b, WH19, AR06, BJR17, BWH19, Has13, KC11, Rao08, Rao13, RDB14, RT17b, RT17a, VPWD11, WH11]. Temporally [BCK23, CZ19, GM15]. term [KCW22]. Terms [KH94, Qui89, Sak93, Sib01, GJ06]. Test [AY88, Arb08, APH86, BH92, BR06, CP03, Dro07, Eva80, FN97, FA03, HL06, HCH00, Has01, HP92, KA08, LM04, Lju88, MN95, O’B87, PP97, Pre98, RG80, RA92, Sch96, Sch16, Shi88, TZ19, TLG93, TZ02, Tua86, Tua87, Vel94, VN00, WS20, XHN17, XZL20, AHT13, BCD18, BJR17, BEL06, Bet16, CG07, CD09, CD12, DR11, FR07, GA04, GPRV00, HW09, Hid07, ILT14, JW16, Jin18, KL09, KT10, Lar98, LD04, LJ23, LNVK02, LKN05, LLZZ22, LTT18, MM12, MKS22, MGRM10, Nis09, PV15, Rao08, Sha11, SP12, TM98, TNZ04, VW15, WWW12, Wes13, WX18, XPZL10, ZSW22, ZS17, ZPZ21, Zhu13]. Testing [AP95, AiL15, BiS17, Bar00, Bea18, BG95, BHLS11, Bet16, BK19, BK07, BH01, CT06a, CHLT15, Cha15b, CKR18, Cub95, DC23, Del96, DKV11, DK17, EV19, Fer90, FFGM15, FM98, GM15, GA01, Gil99, GHHK18, GH19a, HS05, HM13, HLX10, HMW19, Hin82, HR15, HST23, IP20, JN18, JT03, Kan81, Kas82, Ken12, Kil11, KPRN03, KK12, KS05, KP08, Kun97, LS03b, LW19, LLL22, LG11, LBV09, MMH88, Mil19, Pap05, PPS20, PR10, Pon06, PP16b, QN82, Rai96, RB92, SL96a, SS16, ST03, SS96b, SS98, SK09, Tay98, TT97a, TSL08, Wri98, Xia01, YR95, Yua00a, Yua00b, Ano99d, AR06, dBCO12, CNR17, CP17, CF98, DHT14, FL04, Fos13, HCT04, Hon97, Ioa10, Jen12, Kap05, KPY22, Kim15, LP10a, LLS08, MGRM10, MY04]. testing [MLS97, NS13, NK24, Oke98, Ott21, PT23, RGLA11, RT22, SL00, San18, SM13b, Tau23, TW02, Vog99, WSS04, Wie13, ZL24a]. Tests [AN92, AN94, AN08, AES06, BK03, Bre94, BH10, CL06, CISG16, CT92, Che93, CL95b, CG82, CD86, Duf81, EF06, GRT17, GL96, Hal92, Hal95, HIP87, HS98, HR04, HH16, Joy92, Kil16, KH98a, KLN04a, KL98, Kur21, KLM16, LKB15, Lee16, LTK07, LL18, Mar20, ME98, MW16, OT98, PF95, PRC03, PT81b, RWZ20, Sai83, SS01, ST99, Smi08, Tay03, Won97, WL05, YK06, ZT18, And97, AS00, Art02, BP12, BDM98, BH13, Bra13, Bra05, BFK19, BH03, CT08, CR99, CL01, CW99, DL15, DK13, DS04, Dit00, DdM04, Duc05, Fan05, GF15, HL02, HK14, HM03, HLT10, KT16, Kat12, KP13, Kur11, LT17, LLOS08, LLS02, LS03a, LN99, LT03, LL10, LLT14, LM08, MA24, Mar07b, MP16, MMOV24, MR12, NKC15, NR11, PSU08]. tests [PLNL22, PY22, Pes07, Psa01, Rod13, Roz01, Sen07, SF98, Sme15, Sol04, Swe03, kTR98, Tay05, THN24, WE07, ZL12b]. tests* [Kat09]. Texts [Rao14, Bos09]. Textures [YR92]. th [CWDL97, ZBD06]. th-order [CWDL97, ZBD06]. their [BDL08, CA99, PP16a, PC23, Yu07]. Theorem [BT94, Cha91, Kee97, Mor83, Yaj89, FPS21, RB13]. Theorems [Leo13, Sto85, BT13, BDL08, Chu12, KL10]. Theoretic [Bra16]. theoretical [Zaf08]. Theory [AY96, BT06, DNL81, GP06, HP17, HT86, IP08, Mél85, MdS89, Nea13, Pou17, Rao14, Wal87a, Arv14, AA20, BP18, Deo97, JPP15, LP14, LTT18, LP19, MW22, SW21, Sta20, ZL12a]. there [Sen07, Vog99]. thinning [AS24, WP14, ZJ06]. thinning-based [AS24]. Third [Bri80, Gab88, Kak96, SS89, Tan87, Hos01]. Third-Order [Bri80, Gab88, Kak96, SS89]. third-series [Hos01]. Three [BM89, LHR82, GL20]. Three-Dimensional [LHR82]. three-step [GL20]. Threshold [AB86, BLT92, CG19, CL95a, Che95, Cli07, GMRO11, GT93, GLML16, Kap01, Lim92, Log04, Mar92b, Pet86, Str96, Tha90, Ton82, WTSL17, YL20, YLC21, ZL12b, BHLS11, CS08, Cam04, De 98a, KMX17, LR02, LL10, SY20, WL98, XPZL10]. thresholding [PP12]. Thresholds [CT86, Gao97]. Time [AY88, AC93, AF16, ARS86, ABT18b, Ala89, ACL01, And93a, AC96, AM05, Ano97b, Ano99c, Ano02, Ano03, Ano04, Ano05b, Ano06a, Ano07, Ano20g, Ano21a, AVW16, Aok91, AV05, AD04, APH86, AHL^{+}18, AC18, Azz81, BS07, BB95, BC97, BC95, BS15a, BO05, BP11, BBKL17, BW00, BH91, BH92, BG95, BG20, BK19, BDH^{+}18, BLL05, BI09, Bos10, Bos16, BD92, Bro95, BS98, CH14, CS15, Car85, CT08, Cha16a, Cha19, CT20, CT96, Cha91, CD94, CAP94, Che09, CG82, Com96, CKR18, CW82, CS84, CS87, Cor95, Cox91, Cox94, DB98, DN99, DF80, Deg87, Dic82, DK17, Dit04, Dun81, DH17, Elt94, Eng84, Eva80, FK87, Fei20, Fin84, Fin85, Fin16, FK04, FP16, Fra84, Fra05, GR81, Gab88, GT19a, GCK99]. Time [GPH83, GT93, GA01, GRS97, GKY18, GS20, Giu17, GHHK18, GOV19, Gra95, GJ80, Gra82, GH91, GZF86, HHP84, HO84, Hal94a, Hal95, HW95a, Hal14, HS05, HKK23, HT88, HT86, HV92, HR88, Has82, Has00, Has93b, Has94, Has01, HH81, Hid97, Hin82, HH93c, Hog18, Hok83, HN80, HR93, HV99, Hoy20, HC20, HB93, HB94a, HB94b, Huz88, Hyn93, IMR18, IP20, JMP16, JL83, Jan82, JS90, Jan05a, Jan10, JA81, JM04, KMS15, KPT21, KE88, Kas82, KKJ18, Kil18, KBB90, KC96, KK20, Kit81, KT01, Kra16, Kul85, KL98, Kum86, LW91, Lat17, LL92, Led90, LL95, LR88, LT18, Li84, LM94a, LX96, Li98, LX01, LS03b, LWL93, LM95, Lim87, LPS99, Liu89a, Liu89b, LXT20, Lu18, Lüt82, LM94b, MS08a, Mar00, Mar92b, Mas96, Mau11, MMH88]. Time [MMT05, McC15, MT94a, MT94b, ML83, McL98, McL17, MV03, MG93, Mil84, MP84, Neu96, Nie15, NH19, NBQ16, O’B87, OJ03, OT98, Oza82, Pap94, Pap22, Pau84, PS89, PBT00, PZ17, Pet19, PT81b, PT86, Pos20, Pou89, PM92, Pou17, PH02, Pre98, Pri80, Pri96, Psa08, RG80, RRW83, RG89, Rao10a, Rao10c, Rao14, RS19, Rob83, Rob87, RT17c, Sai83, Sai86, SL96a, Sch96, STA03, SPA20, SBLS07, SR88, SR91, SR92, Sha08, She87, SS96b, Shi88, SS82, SPM19, Sin93, Smi08, SH88b, ST87, Sto85, Sto87, SW86a, Str96, Tan81, Tan84, TK93, Tay98, Tay03, Tay20, Tay21b, Tay22a, Tay23c, Tay24a, Tay24b, TA88, Tho19, TWVB00, TP82, TP85, Tjø86, TM93, Ton81, TS94, TC05b, Tsa88, Tsa89, TY00, TW89, Tur11]. Time [TT82, UH95, VP12, VWR87, VDO95, Vij06, Wal87b, WT19, WS20, WL05, WTSL17, XA99, XHN17, Yak87, YL91, Yua00b, ZT94, ZW94, ZLY06, ZLP19, ZC19b, ZLSY20, Zho92, ZP20b, ABT18a, AR10, AD23, APS20, AS00, ASM21, And08, Art02, AHT13, AH13, ASK15, BCD18, BBC16, Bar00, BDL08, BM13, Bet16, BP03, Bon01, Bon05, Bou08, Bra05, BS02, Bro07, BFK12, BFK13, BFK19, CS08, Cai11, CSD12, CCGA13, Cha99, CLY17, CGP22, CP17, Cha16b, CCY16, CH21, CF14, CF98, CG11, CH11, DG98, DB03, DGH06, DC01, DRY21, DE07, DLRY08, DdRSK21, DFF23, DL15, DH13, DHT14, DPT12, DP10, DA14, DLPP14, DR11, Erc11, Fan05, FHK20, FF13, FB13, FLL13, FK99, FP12, FL00, FG04, FM04, GG08, GF15, GA16, GM15, GA04]. time [GA08, GM23, GB06, GAP09, GKL11, GLN15, GSO^{+}17, HPY02, Hal13, HNT23, HM13, HP23, Has13, HCT04, Hon97, HL11b, HR15, Hoy24, HW99, HDB98, HD99, HC00, HB01, Hur01, HK08, Huz07, ILT14, JMP12, JFML13, JW16, Jin18, Jir16, JT03, JT11, KM03, KP15, KWPV12, KYP20, Ken12, KH04b, KC10, KL10, KL11, Kom99, KGY18, KL23, KXS^{+}12, LC03, LLS02, LM00, LZ18, LD04, LV00, LLS08, Len16, LLRR^{+}21, LNVK02, Li14, LLY14, LRS21, Lie12, LP14, LL97, LPZ15, LTT18, LZZ22, LLS97, LOS12, Lub99, Mar99, MTJ14, MGRM10, MY04, MJ12, MT15, ME98, MIN^{+}16, MS00b, MRT07, Muk99, NHCLP08, NM11, Nie05, Nis09, PZ04, PD12, Par13, PT04, PT23, Pen07, PR03, PRC03, PPS20, PR10, PS00, Pro23, PSSS09, QR98, RCLM^{+}11, Rao10b]. time [Rao18, RY23, RB22, RS17, RNI13, RT22, SK23, Sbr11, Seo24, SL97, SB19, SO97b, Sme15, SW21, SA07, SCW19, ST05, SO12, Sto19, SR07, Stu01, Swe22, Sza22, Sza23, TKOP20, TNH23, TW02, TM98, TAM11, TC13, TJL23, Tri11, TC05a, TT97b, Unn04, UD09, Vel99, VW15, VY16, Wal00, WWG09, WL11, WP21, WD10, WC14, WX18, XZ22, XLT23, YLC21, Yau12, ZT18, ZKP22, Zha24, ZJN24, ZP20a, Zho12, Zho13, ZJ06, YB06, Aue22, Bos09, Nun20, Omb13, Pou16, Qui15, Wil16]. Time-Aggregation [Eng84]. Time-correlation [Li98]. Time-Dependent [GOV19, Pri96, Li14]. time-domain [KM03]. time-invariant [ZT18]. Time-Lag [DF80]. Time-Reversibility [BD92, Psa08, Pro23]. Time-scale [JM04]. Time-Series [Eva80, MMT05, Nie15, SPA20, Smi08, WL05, ZLY06, CH14, VP12, AR10, Cai11, DE07, GG08, GF15, GB06, HL11b, Ken12, KC10, LD04, LLY14, LPZ15, Nie05, VW15, WD10, Zho12, Zho13]. time-symmetric [LZ18]. Time-Transformed [CT08]. Time-Varying [Ano21a, AD04, BBKL17, BDH^{+}18, Pet19, BP11, PH02, CGP22, CP17, DHT14, FG04, Lie12, LP14, RCLM^{+}11, YLC21, ZT18]. Times [Kur21, FKK12]. TMA [LQ20]. TOC [Ano16g, Ano16h]. Toeplitz [LP04, Tak24]. Tohru [Omb13]. Tools [Neš16]. Total [Bos16, Cao19, Cha16a, Hal14, Kar16, Kil18, Lat17, Lu18, McC15, McL17, Mcn15, Nea13, Neš16, Omb13, Pou17, Qui15, Rao16a, Rao16b, Rao17, Ter14, Tur18, Wan21]. Totals [SW86a]. Tracking [BW00, LK98, Whi05, DZ17, Whi02]. Tractable [JC17]. Trade [GJ02]. Transaction [CHS17]. Transaction-Level [CHS17]. Transfer [Bha86, Gri91, Kav93, LHR82, Lii85, MW97, CGM08]. Transform [DGJ06, HC20, ST97, DR11, KM99, Rig92]. Transformation [GS13]. Transformations [ADD19, Cor95, GH91, PR09, FM98, JM04]. Transformed [Bla14, CT08, PF21]. Transforms [Sto85, Yaj89, Wal00]. Transition [BBKL17, Ber01, LS06, XZL20, dJ86, BF10, HK22, Kil11, ZL24a]. Transitions [SLN99, LNV98, Sol04]. Transmission [CK15]. Treating [AK10]. Trend [AP95, DH98, Gil99, HH05, KLN04a, MKN22, NVS06, SL00, Sib01, TSL08, Ano99d, CHLT15, CP16a, EP17, Fos13, HLT10, KP10, KPRN03, KLN04b, LL12, MT15, NR11, Rao12b, kTR98, TT97a, VN17, YLC12, ZSW22]. Trending [HMW19, TY10, WNS22]. Trends [Gra88, Has00, Has01, HD96, Mar20, NV96, CCGA13, CG11, HMV08, KT16, McC13, Ott21, TJL23, WX18]. triangular [PPS14, Wan05]. Trigonometric [BSG18, KH94]. trimmed [Hil13]. Trimming [AV05]. Trindade [Pou17]. TSMARS [LR02]. Tucker [Aue22]. Tunnicliffe [McL17]. Turbulence* [LHR82]. Tweedie [Fei20]. Two [Ala89, BR06, CGN15, CD86, EK13, HT99, LT83, PR88, SS90, Sch96, Vog99, Wri95, ZLSY20, DL15, Efr19, HOS15, JW16, KPY22, LL18, PY22, PRC03, RCLM^{+}11, SH09]. Two-Channel [SS90]. two-dimensional [RCLM^{+}11]. Two-Sided [Sch96]. Two-Stage [CGN15, Wri95, Efr19, HOS15]. Two-Step [ZLSY20, EK13, KPY22]. two-steps [PY22]. Type [AN92, Fin16, Ger18, GL96, HH93a, Hua90a, Kil16, NV96, SR88, Wan93a, PLNL22, TT97a, ZS17, NR11, Wie13].

U.S. [ZC19a]. Uhlenbeck [OV04, WHY22]. Uhlig [PI22]. UK [Rao14, Rao17]. Unbiased [AF91, Zie99, LR21, Lug06, VY16, Wan16]. Unbiasedness [Fin85]. unbounded [LP04, Tak24]. Uncertainty [LS06, Kil98, NAJ12, PH02]. unconditional [LD04, SF98]. Uncorrelated [MIN^{+}16, FR07, Kat12]. Uncorrelatedness [WS20]. Underdispersed [MA20]. unequal [DL15, Jin18, LL18]. unevenly [GM23]. Unified [Eng84, HP14, Kim16, KJ85, PW84, Lie05, LP19]. Uniform [BDL08, GP06, Mas96]. unifying [GPRV00]. Unimodality [MA93]. Uniqueness [BDH^{+}18, GZ15]. Unit [AT87a, AT87b, AES06, Bea18, BK03, BF96, Bre94, CISG16, CP03, CL95b, Fos13, GRT17, GL96, Hal95, HR02, HL06, HCH00, HR04, HK90, HMW19, HH16, Ioa10, Kap05, Kar16, KLN04a, LNV98, MR12, Ott21, OT98, PF95, Pau84, Pon06, RA92, SS16, SH87, San18, SS96b, SF98, SS01, ST99, SLN99, Tay98, Tay03, YR95, AG08, CT08, CHLT15, Cha15b, DK13, FM98, GPRV00, HL02, KT16, KP08, LLS02, Lar98, LD04, LT03, LLT14, LLZZ22, LG11, Lub99, NK24, NR11, PSU08, PRW04, Psa01, Rod13, Sen07, SL97, SS98, Sol04, Sto19, Swe03, kTR98, Tay05, Vog99, Wes13, WNS22, Xia01, Yab12, YLC12, ZS17]. Unit-Root [AES06, HL06, KLN04a, Ioa10, Kap05, LLZZ22, PSU08]. units [Sme15]. unity [BP18, Sta20]. Univariate [FN97, HH93c, KOV94, KMS15, Lim87, SPA20, BFK19, PS07, VY16]. University [Cha16a, Hal14, Kar16, Leo13, Nea13, Neš16]. Unknown [Hid92, Kab94, Kak96, MM93, BEL06, BH03, DCCL03, GLL06, GA08, HLX10, ILT14, LM00, Lug06, Sel10, Spe10, VY16, Wri98]. Unobservable [Ott88]. Unobserved [Hot89, Pou16, Sin93, Cha16a]. Unrestricted [CRT15]. Unstable [Bea18]. unusual [Pen07]. US$39.99 [Kar16]. US$95.00 [Kar16]. USA [Bos16, Cao19, Hal14]. Use [KT01, PT81b, BFZ02, Tay05, TW02]. Useful [Tua84b]. Uses [HP92]. Using [AiL15, BiS17, dBC07, CL95b, CT87b, Cub95, Del96, Gae00, Gil99, GL94, HH93a, Huz81, JC17, Joy92, Kav89, KF08, Lu18, MFM05, ML83, MS07, Mil19, MB97, Nun20, Oza82, PBT00, PV98, PW05, Rei94, Rig92, SS96b, SS82, Stu01, TA88, Ton82, Tsa07, YR92, Yua00b, Ano99d, BWH19, CHLT15, CK22, Cra03, DdRSK21, DFF23, GH11, HOS15, HL18, HB01, LR02, LL18, MS23, MIN^{+}16, MMNT20, Mil06, NHCLP08, Olm23, PSSS09, Rao11, RSW08, TT97a, Tri12, UT12, ZC19b].

Valid [Kak99b]. Validation [Fas00, Kav89, De 01, Sko01, Vel00]. Validatory [HB90]. Validity [Tan84, FP18]. Value [BT06, GLP10, DP20, EFT16, Mar12, MW22]. Value-at-Risk [GLP10, DP20]. Valued [ABT18b, AOA87, DK17, FLO06, GMLS15, Gor18, JGY91, Lat17, Mic20, NR07, Sch16, Sto87, AD23, AS24, BMH08, EMNR09, HA21, JJL12, KY09, KT11, KL09, Lat98, LJ23, NLR16, SA07, XZ22, Zha24, ZBD06, Zhu11]. Values [CL95b, Pou89, Swi90, AK10, Bon05]. Vanishing [Hog19]. var [IC05, Bra13, CRT15, Cav14a, Cav14b, Hoy24, JN19, TSL08, WR08]. Variability [De 98b, Kra16]. Variable [Che95, FRP99, Hal92, Hal95, ST04, ZKG22, BBK23, CGP22, LLY14, Mar07b, THN24]. Variables [Abr87, AD84, ADD19, Cha95, New80, Pou88, Wal87b, DE07, KL10]. Variance [Bha93, CJ82, IP20, Kni87, KT94c, KT01, MMT98, MN95, Nas93, SF93, SLL^{+}20, TP85, TZ02, TT82, AMS^{+}17, Bha97, BM10, CT06a, CZ12, CW99, DHT14, HLZ23, Hil13, Joh03, KS18, KP08, KW12, LV00, Li14, LM08, MR12, NKC15, PS00, Sen07, SLL97, TT97a, TNZ04, VW15, ZL12a]. variance-ratio [NKC15, TNZ04]. variance-type [TT97a]. Variances [BG95, Pfe94, CF98, TP03]. Variate [SS96a, TCCG19, Cai11, EK13, KL11, Nie11, TP03, Tri12, ZSW22]. variates [DdRSK21]. Variation [BW00, HH81, LW19]. variogram [RT17b]. Various [Lim87]. VARMA [Arb08, AV08, Kat12, Maï12, MS07, ZC19a]. VARs [GH19a, HN21, Sta20]. VARX [PPT93]. Varying [Ano21a, AD04, BBKL17, BDH^{+}18, Dun81, GKY18, Pet19, Wal87a, ZLSY20, BP11, CGP22, CP17, DHT14, FG04, Lie12, LP14, MS00a, MS01, Ott21, PH02, RCLM^{+}11, YLC21, ZT18]. Vector [AK90b, BMY99, BR06, CV06, Dav91, HT93, KP89, KP93, LH96, Lüt85, MK15, PS92, Pap94, Pap05, PS95, RA92, She88, ZT94, Zho92, AN08, BBK23, BB14, DdM13, FB21, JR22, Joh03, JS24, Kak99a, KD03, KM21, KP21, LK21, LZ24, MMM22, Mau02, MR18, SL00, SAZ13, Swe22, UD09, ZJN24]. vectors [MS00a]. Version [Rav89]. Versus [AP95, HS98, MS08a, PY22]. very [AG08]. Via [BT06, Jan82, MT94a, MT94b, Neu96, Wah89, FHK20, Hog19, KS08b, KGY18, LW19, Mar99, MW05, TS14, VP12]. vibrometry [LL06]. Vicinity [FM96]. view [HK17]. Viewpoint [RBY92]. Ville [BDH^{+}18]. Vine [BS15b]. virus [VPWD11]. VMA [Cav14b]. Vol [Ano21a, Fei20]. Volatility [Ano21a, BCCR19, Bea18, Ber07, BK19, CPR18, CK22, GS21, Hal14, HL17, Jen04, KS08b, KS23a, KS05, LKB15, LDH19, LW19, Pet19, SHL22, SLL^{+}20, SY20, BL13, CT08, CHLT15, Com04, EFT16, HL18, HWBD11, LLL22, McC13, PA23, SBS23, Tri12, Wan16, YHN99, YFL^{+}14, YSX^{+}22, Zaf08]. Volume [Ano97b, Ano98b, Ano00b, Ano01d, Ano03, Ano07, NNC91, Ano99c, Ano02, Ano04, Ano05b, Ano06a]. vs [PS07].

Walk [CL95b, Hal92, BH01, Ing01]. Walker [CADF11, Hyn93, SR88]. walks [Ste05]. Walsh [HC20, Sto85, Sto87, Sto90]. Walsh-Fourier [Sto85, Sto87, Sto90]. Walter [Lu18]. warming [EP17]. Warning [BHL90]. Wasserstein [AK24, ZKP22]. Watson [Ali83]. Wavelet [CN17, DL15, Neu96, ST97, VN00, WC10, AG16, DP10, FNV08, Gao97, GA16, KM99, LL18, MKN22, MRT07, Nan14, RCLM^{+}11, RT09, ST05]. Wavelet-Based [DL15, VN00, RCLM^{+}11]. wavelet-Fisz [FNV08]. Wavelets [Gil99, Pri96, AVF98, AG24, Ano99d]. Way [ZW94]. Wayne [Kil18]. Weak [FRS11, HMS13, Hua12, FZ23, Hid07, Maï12, San17]. Weakly [Sto85, XX18, BDL08, Bra11, Bra22, MMM22, Psa01, PV15, Sta20]. Weighted [BCK23, FRS11, HH20, Iri02, ST91, YL20, ZC19a, ZW12, BM09, CP17, DGK14, GF15, HI15, LZ20, LPS23]. Weighted-Covariance [ZC19a]. Weighting [SSW15, ZLP19]. Weights [LDH19, DP20]. Which [HP92]. White [And93b, FR97, RB92, BCD18]. Whitney [Mau11]. Whittle [AG16, AG24, BDL08, Iac10, JC17, Kul85, LRS21, Nie11]. Wiener [Góm07, RY23, Wan09]. Wiener-Kolmogorov [Góm07]. Wigner [BDH^{+}18]. Wilcoxon [Bet16, Ger18]. Wilcoxon-Type [Ger18]. Wild [CF24, DLLN15]. wildfires [NSK^{+}11]. Wiley [Bos16, Rao16b, Rao17, Tur18, Wil16]. William [Tur18]. Wilson [McL17]. Wind [TCCG19]. Window [Bha83, Büh96, CAP94, KP13]. Wishart [AS21, Tri12]. Within [Kra16]. Within-Group [Kra16]. Without [Aok91, Lim92]. Wold [KM21]. Woodward [Kil18]. worth [PS00].

X [Jan05b, BEvdW12, Pfe94, Pou17]. X-11 [Jan05b, Pfe94]. X-ray [BEvdW12]. xiii [Nea13].

yellow [VPWD11]. yield [HKVW22, SPH18]. York [Cao19, Hal14, Nun20, Yao20]. Yuanhua [Ter14]. Yule [CADF11, Hyn93, SR88].

Z [Mau11]. Zero [BS15a, FM96, GT19b, HK90, VY90, CWDL97, JJL12, Mar99, PP16a]. Zero-Crossing [HK90, CWDL97]. zero-crossings [Mar99]. Zero-Modified [BS15a]. Zeros [BS15a]. Zhang [Wan21]. Zou [Wan21]. Zucchini [Lu18]. Zumbach [McC15].

References

Arvanitis:2020:LTG

[AA20] Stelios Arvanitis and Sofia Anyfantaki. On the limit theory of the Gaussian SQMLE in the EGARCH(1,1) model. Journal of Time Series Analysis, 41(2):341–350, March 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Aknouche:2022:PAC

[AAD22] Abdelhakim Aknouche, Bader Almohaimeed, and Stefanos Dimitrakopoulos. Periodic autoregressive conditional duration. Journal of Time Series Analysis, 43(1):5–29, January 2022. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Andel:1986:NTM

[AB86] Jiři Anděl and Tomáŝ Bartoň. A note on the threshold AR(1) model with Cauchy innovations. Journal of Time Series Analysis, 7 (1):1–5, January 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Abraham:1999:IGA

[AB99] B. Abraham and N. Balakrishna. Inverse Gaussian autoregressive models. Journal of Time Series Analysis, 20(6):605–618, November 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Aknouche:2009:QML

[AB09] Abdelhakim Aknouche and Abdelouahab Bibi. Quasi-maximum likelihood estimation of periodic GARCH and periodic ARMA-GARCH processes. Journal of Time Series Analysis, 30(1):19–46, January 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Abril:1987:ADS

[Abr87] Juan Carlos Abril. The approximate densities of some quadratic forms of stationary random variables. Journal of Time Series Analysis, 8(3):249–259, May 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Akashi:2018:RRS

[ABT18a] Fumiya Akashi, Shuyang Bai, and Murad S. Taqqu. Robust regression on stationary time series: a self-normalized resampling approach. Journal of Time Series Analysis, 39(3):417–432, May 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Aknouche:2018:NBQ

[ABT18b] Abdelhakim Aknouche, Sara Bendjeddou, and Nassim Touche. Negative binomial quasi-likelihood inference for general integervalued time series models. Journal of Time Series Analysis, 39(2): 192–211, March 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Abraham:1993:EMA

[AC93] Bovas Abraham and Alice Chuang. Expectation-maximization algorithms and the estimation of time series models in the presence of outliers. Journal of Time Series Analysis, 14(3):221–234, May 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anderson:1996:HOM

[AC96] Oliver D. Anderson and Zhao-Guo Chen. Higher order moments of sample autocovariances and sample autocorrelations from an independent time series. Journal of Time Series Analysis, 17(4): 323–331, July 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Audrino:2018:OPB

[AC18] Francesco Audrino and Lorenzo Camponovo. Oracle properties, bias correction, and bootstrap inference for adaptive lasso for time series M-estimators. Journal of Time Series Analysis, 39(2):111– 128, March 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Alpay:2001:EPD

[ACL01] D. Alpay, A. Chevreuil, and Ph. Loubaton. An extension problem for discrete-time periodically correlated stochastic processes. Journal of Time Series Analysis, 22(1):1–11, January 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Andersen:2023:ACP

[ACN23] Torben Andersen, Kim Christensen, and Ingmar Nolte. Announcement: Call for papers for special issue in honour of Stephen J. Taylor. Journal of Time Series Analysis, 44(4):336, July 2023. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anderson:1984:IDE

[AD84] B. D. O. Anderson and M. Deistler. Identifiability in dynamic errors-in-variables models. Journal of Time Series Analysis, 5(1): 1–13, January 1984. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Allen:1999:NBE

[AD99] Michael Allen and Somnath Datta. A note on bootstrapping Mestimators in ARMA models. Journal of Time Series Analysis, 20 (4):365–379, July 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Arvanitis:2004:TDM

[AD04] Stelios Arvanitis and Antonis Demos. Time dependence and moments of a family of time-varying parameter GARCH in mean models. Journal of Time Series Analysis, 25(1):1–25, January 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Aknouche:2023:ACP

[AD23] Abdelhakim Aknouche and Stefanos Dimitrakopoulos. Autoregressive conditional proportion: a multiplicative-error model for (0, 1)- valued time series. Journal of Time Series Analysis, 44(4):393–417, July 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anderson:2019:SGC

[ADD19] Brian D. O. Anderson, Manfred Deistler, and Jean-Marie Dufour. On the sensitivity of Granger causality to errors-in-variables, linear transformations and subsampling. Journal of Time Series Analysis, 40(1):102–123, January 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Akashi:2018:CPD

[ADL18] Fumiya Akashi, Holger Dette, and Yan Liu. Change-point detection in autoregressive models with no moment assumptions. Journal of Time Series Analysis, 39(5):763–786, September 2018. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Askanazi:2018:CIF

[ADSS18] Ross Askanazi, Francis X. Diebold, Frank Schorfheide, and Minchul Shin. On the comparison of interval forecasts. Journal of Time Series Analysis, 39(6):953–965, November 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Allal:2006:ODE

[AE06] Jelloul Allal and Saïd El Melhaoui. Optimal detection of exponential component in autoregressive models. Journal of Time Series Analysis, 27(6):793–810, November 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Aparicio:2006:RUR

[AES06] Felipe Aparicio, Alvaro Escribano, and Ana E. Sipols. Range unitroot (RUR) tests: Robust against nonlinearities, error distributions, structural breaks and outliers. Journal of Time Series Analysis, 27(4):545–576, July 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Azzalini:1991:ENU

[AF91] A. Azzalini and A. C. Frigo. An explicit nearly unbiased estimate of the AR(1) parameter for repeated measurements. Journal of Time Series Analysis, 12(4):273–281, July 1991. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ahmad:2016:PQC

[AF16] Ali Ahmad and Christian Francq. Poisson QMLE of count time series models. Journal of Time Series Analysis, 37(3):291–314, May 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Aknouche:2022:SEM

[AF22] Abdelhakim Aknouche and Christian Francq. Stationarity and ergodicity of Markov switching positive conditional mean models. Journal of Time Series Analysis, 43(3):436–459, May 2022. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Adams:1995:PEP

[AG95] G. J. Adams and G. C. Goodwin. Parameter estimation for periodic ARMA models. Journal of Time Series Analysis, 16(2): 127–145, March 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Andrews:2008:ASV

[AG08] Donald W. K. Andrews and Patrik Guggenberger. Asymptotics for stationary very nearly unit root processes. Journal of Time Series Analysis, 29(1):203–212, January 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Achard:2016:MWW

[AG16] Sophie Achard and Irène Gannaz. Multivariate wavelet Whittle estimation in long-range dependence. Journal of Time Series Analysis, 37(4):476–512, July 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Achard:2024:LWE

[AG24] Sophie Achard and Irène Gannaz. Local Whittle estimation with (quasi-)analytic wavelets. Journal of Time Series Analysis, 45(3): 421–443, May 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Allende:1992:RGM

[AH92] Hector Allende and Siegfried Heiler. Recursive generalized m estimates for autoregressive moving-average models. Journal of Time Series Analysis, 13(1):1–18, January 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Aue:2013:SBT

[AH13] Alexander Aue and Lajos Horváth. Structural breaks in time series. Journal of Time Series Analysis, 34(1):1–16, January 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Astill:2018:RTM

[AHL^{+}18] Sam Astill, David I. Harvey, Stephen J. Leybourne, Robert Sollis, and A. M. Robert Taylor. Real-time monitoring for explosive financial bubbles. Journal of Time Series Analysis, 39(6):863–891, November 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Aue:2017:FGA

[AHP17] Alexander Aue, Lajos Horváth, and Daniel F. Pellatt. Functional generalized autoregressive conditional heteroskedasticity. Journal of Time Series Analysis, 38(1):3–21, January 2017. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Aue:2006:ERC

[AHS06] Alexander Aue, Lajos Horváth, and Josef Steinebach. Estimation in random coefficient autoregressive models. Journal of Time Series Analysis, 27(1):61–76, January 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Astill:2013:BTA

[AHT13] Sam Astill, David I. Harvey, and A. M. Robert Taylor. A bootstrap test for additive outliers in non-stationary time series. Journal of Time Series Analysis, 34(4):454–465, July 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Arrieta-ibarra:2015:TPF

[AiL15] Imanol Arrieta-ibarra and Ignacio N. Lobato. Testing for predictability in financial returns using statistical learning procedures. Journal of Time Series Analysis, 36(5):672–686, September 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ansley:1990:FSS

[AK90a] Craig F. Ansley and Robert Kohn. Filtering and smoothing in state space models with partially diffuse initial conditions. Journal of Time Series Analysis, 11(4):275–293, July 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ansley:1990:NSR

[AK90b] Craig F. Ansley and Robert Kohn. A note on square root filtering for vector autoregressive moving-average models. Journal of Time Series Analysis, 11(3):181–183, May 1990. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Andersson:2010:TMV

[AK10] Jonas Andersson and Dimitris Karlis. Treating missing values in INAR(1) models: an application to syndromic surveillance data. Journal of Time Series Analysis, 31(1):12–19, January 2010. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anastasiou:2024:WDB

[AK24] Andreas Anastasiou and Tobias Kley. Wasserstein distance bounds on the normal approximation of empirical autocovariances and cross-covariances under non-stationarity and stationarity. Journal of Time Series Analysis, 45(3):361–375, May 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Akaike:1980:SAB

[Aka80] Hirotugu Akaike. Seasonal adjustment by a Bayesian modeling. Journal of Time Series Analysis, 1(1):1–13, January 1980. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Alagon:1989:SDT

[Ala89] Javier Alagón. Spectral discrimination for two groups of time series. Journal of Time Series Analysis, 10(3):203–214, May 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ali:1983:NAD

[Ali83] Mukhtar M. Ali. A note on approximating the distribution of the Durbin-Watson statistic. Journal of Time Series Analysis, 4(4): 217–220, July 1983. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anderson:1980:SLF

[AM80] T. W. Anderson and Raúl P. Mentz. On the structure of the likelihood function of autoregressive and moving average models. Journal of Time Series Analysis, 1(2):83–94, March 1980. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anderson:2005:PEP

[AM05] Paul L. Anderson and Mark M. Meerschaert. Parameter estimation for periodically stationary time series. Journal of Time Series Analysis, 26(4):489–518, July 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Avarucci:2007:PCB

[AM07] Marco Avarucci and Domenico Marinucci. Polynomial cointegration between stationary processes with long memory. Journal of Time Series Analysis, 28(6):923–942, November 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Arvanitis:2018:MEA

[AM18] Stelios Arvanitis and Tassos Magdalinos. Mildly explosive autoregression under stationary conditional heteroskedasticity. Journal of Time Series Analysis, 39(6):892–908, November 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Albert:2017:MSD

[AMS^{+}17] Stefan Albert, Michael Messer, Julia Schiemann, Jochen Roeper, and Gaby Schneider. Multi-scale detection of variance changes in renewal processes in the presence of rate change points. Journal of Time Series Analysis, 38(6):1028–1052, November 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anderson:2013:FPI

[AMZ13] Paul L. Anderson, Mark M. Meerschaert, and Kai Zhang. Forecasting with prediction intervals for periodic autoregressive moving average models. Journal of Time Series Analysis, 34(2):187–193, March 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Agiakloglou:1992:EED

[AN92] C. Agiakloglou and P. Newbold. Empirical evidence on Dickey– Fuller-type tests. Journal of Time Series Analysis, 13(6):471–483, November 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Agiakloglou:1994:LMT

[AN94] Christos Agiakloglou and Paul Newbold. Lagrange multiplier tests for fractional difference. Journal of Time Series Analysis, 15(3): 253–262, May 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ahlgren:2008:TAS

[AN08] Niklas Ahlgren and Jukka Nyblom. Tests against stationary and explosive alternatives in vector autoregressive models. Journal of Time Series Analysis, 29(3):421–443, May 2008. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Andel:1987:LPG

[And87] Jirí Andel. On linear processes with given moments. Journal of Time Series Analysis, 8(4):373–378, July 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Andel:1989:NNA

[And89] Jirí Andel. Non-negative autoregressive processes. Journal of Time Series Analysis, 10(1):1–11, January 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anderson:1992:PAP

[And92] O. D. Anderson. Partial autocorrelation properties for nonstationary autoregressive moving-average models. Journal of Time Series Analysis, 13(6):485–500, November 1992. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Andel:1993:TSM

[And93a] Jirí Andel. A time series model with suddenly changing parameters. Journal of Time Series Analysis, 14(2):111–123, March 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anderson:1993:EGL

[And93b] Oliver D. Anderson. Exact general-lag serial correlation moments and approximate low-lag partial correlation moments for Gaussian white noise. Journal of Time Series Analysis, 14(6):551–574, November 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anderson:1997:GFT

[And97] T. W. Anderson. Goodness-of-fit tests for autoregressive processes. Journal of Time Series Analysis, 18(4):321–339, July 1997. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Andrews:2008:RBE

[And08] Beth Andrews. Rank-based estimation for autoregressive moving average time series models. Journal of Time Series Analysis, 29(1): 51–73, January 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:1982:BR

[Ano82] Anonymous. Book reviews. Journal of Time Series Analysis, 3(4): 283–285, July 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:1986:A

[Ano86] Anonymous. Announcement. Journal of Time Series Analysis, 7 (3):i, May 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:1987:C

[Ano87] Anonymous. Correction. Journal of Time Series Analysis, 8(3):i, May 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:1992:C

[Ano92] Anonymous. Correction. Journal of Time Series Analysis, 13(3): 281–282, May 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:1994:AAS

[Ano94] Anonymous. The Australian Academy of Science establishes a Hannan Medal for Distinguished Research in the Mathematical Sciences. Journal of Time Series Analysis, 15(6):649, November 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:1995:BR

[Ano95] Anonymous. Book review. Journal of Time Series Analysis, 16(3): 355–358, May 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:1997:BR

[Ano97a] Anonymous. Book reviews. Journal of Time Series Analysis, 18 (5):529–534, September 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:1997:JTS

[Ano97b] Anonymous. Journal of Time Series Analysis: Index to volume 18 1997. Journal of Time Series Analysis, 18(6):663–664, November 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:1997:C

[Ano97c] Anonymous. Corrigendum. Journal of Time Series Analysis, 18 (3):320, May 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:1998:BR

[Ano98a] Anonymous. Book review. Journal of Time Series Analysis, 19 (5):627–628, September 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:1998:IV

[Ano98b] Anonymous. Index to volume 19, 1998. Journal of Time Series Analysis, 19(6):755–756, November 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:1999:BRa

[Ano99a] Anonymous. Book review. Journal of Time Series Analysis, 20(3): 361–363, May 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:1999:BRb

[Ano99b] Anonymous. Book review. Journal of Time Series Analysis, 20 (6):715–716, November 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:1999:JTS

[Ano99c] Anonymous. Journal of Time Series Analysis: index to volume 20 1999. Journal of Time Series Analysis, 20(6):717–718, November 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:1999:CTO

[Ano99d] Anonymous. Corrigendum: testing for the onset of trend, using wavelets. Journal of Time Series Analysis, 20(6):i, November 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2000:BR

[Ano00a] Anonymous. Book review. Journal of Time Series Analysis, 21 (1):111–112, January 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2000:IV

[Ano00b] Anonymous. Index to volume 21, 2000. Journal of Time Series Analysis, 21(6):739–740, November 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2001:BRa

[Ano01a] Anonymous. Book review. Journal of Time Series Analysis, 22 (1):125–126, January 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2001:BRb

[Ano01b] Anonymous. Book review. Journal of Time Series Analysis, 22(2): 251–252, March 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2001:BRc

[Ano01c] Anonymous. Book reviews. Journal of Time Series Analysis, 22(3): 375–377, May 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2001:IV

[Ano01d] Anonymous. Index to volume: 22 2001. Journal of Time Series Analysis, 22(6):755–756, November 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2002:JTS

[Ano02] Anonymous. Journal of Time Series Analysis: index to volume 23 2002. Journal of Time Series Analysis, 23(6):753–754, November 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2003:JTS

[Ano03] Anonymous. Journal of Time Series Analysis index to volume 24 2003. Journal of Time Series Analysis, 24(6):755–756, November 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2004:JTS

[Ano04] Anonymous. Journal of Time Series Analysis: index to volume 25 2004. Journal of Time Series Analysis, 25(6):943–945, November 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2005:BR

[Ano05a] Anonymous. Book reviews 2. Journal of Time Series Analysis, 26 (1):152–153, January 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2005:JTS

[Ano05b] Anonymous. Journal of Time Series Analysis: index to volume 26 2005. Journal of Time Series Analysis, 26(6):945–946, November 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2005:E

[Ano05c] Anonymous. Erratum. Journal of Time Series Analysis, 26(1):155– 156, January 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2005:EBR

[Ano05d] Anonymous. Erratum: Book review. Journal of Time Series Analysis, 26(5):787, September 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2005:OEA

[Ano05e] Anonymous. Online early announcement. Journal of Time Series Analysis, 26(6):943, November 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2006:JTS

[Ano06a] Anonymous. Journal of Time Series Analysis: index to volume 27 2006. Journal of Time Series Analysis, 27(6):945–946, November 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2006:C

[Ano06b] Anonymous. Corrigendum. Journal of Time Series Analysis, 27(6): i–ii, November 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2007:JTS

[Ano07] Anonymous. Journal of Time Series Analysis index to Volume 28 2007. Journal of Time Series Analysis, 28(6):943–944, November 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2009:CP

[Ano09a] Anonymous. Call for papers. Journal of Time Series Analysis, 30 (4):466, July 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2009:C

[Ano09b] Anonymous. Corrigendum. Journal of Time Series Analysis, 30 (2):259, March 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2010:C

[Ano10] Anonymous. Corrigendum. Journal of Time Series Analysis, 31 (3):227, May 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2016:IIIa

[Ano16a] Anonymous. Issue information — info page. Journal of Time Series Analysis, 37(1):2, January 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2016:IIIb

[Ano16b] Anonymous. Issue information — info page. Journal of Time Series Analysis, 37(2):146, March 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2016:IIIc

[Ano16c] Anonymous. Issue information — info page. Journal of Time Series Analysis, 37(3):289–290, May 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2016:IIId

[Ano16d] Anonymous. Issue information — info page. Journal of Time Series Analysis, 37(4):433–434, July 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2016:IIIe

[Ano16e] Anonymous. Issue information — info page. Journal of Time Series Analysis, 37(5):577–578, September 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2016:IIIf

[Ano16f] Anonymous. Issue information — info page. Journal of Time Series Analysis, 37(6):721–722, November 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2016:IITa

[Ano16g] Anonymous. Issue information — TOC. Journal of Time Series Analysis, 37(1):1, January 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2016:IITb

[Ano16h] Anonymous. Issue information — TOC. Journal of Time Series Analysis, 37(2):145, March 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2017:IIa

[Ano17a] Anonymous. Issue information. Journal of Time Series Analysis, 38(1):1–2, January 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2017:IIb

[Ano17b] Anonymous. Issue information. Journal of Time Series Analysis, 38(2):147–148, March 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2017:IIc

[Ano17c] Anonymous. Issue information. Journal of Time Series Analysis, 38(3):393–394, May 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2017:IId

[Ano17d] Anonymous. Issue information. Journal of Time Series Analysis, 38(4):511–512, July 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2017:IIe

[Ano17e] Anonymous. Issue information. Journal of Time Series Analysis, 38(5):637–638, September 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2017:IIf

[Ano17f] Anonymous. Issue information. Journal of Time Series Analysis, 38(6):807–808, November 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2018:IIa

[Ano18a] Anonymous. Issue information. Journal of Time Series Analysis, 39(1):1–2, January 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2018:IIb

[Ano18b] Anonymous. Issue information. Journal of Time Series Analysis, 39(2):109–110, March 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2018:IIc

[Ano18c] Anonymous. Issue information. Journal of Time Series Analysis, 39(3):239–240, May 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2018:IId

[Ano18d] Anonymous. Issue information. Journal of Time Series Analysis, 39(4):469–470, July 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2018:IIe

[Ano18e] Anonymous. Issue information. Journal of Time Series Analysis, 39(5):637–638, September 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2018:IIf

[Ano18f] Anonymous. Issue information. Journal of Time Series Analysis, 39(6):811–812, November 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2019:IIa

[Ano19a] Anonymous. Issue information. Journal of Time Series Analysis, 40(1):1–2, January 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2019:IIb

[Ano19b] Anonymous. Issue information. Journal of Time Series Analysis, 40(2):161–162, March 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2019:IIc

[Ano19c] Anonymous. Issue information. Journal of Time Series Analysis, 40(3):265–266, May 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2019:IId

[Ano19d] Anonymous. Issue information. Journal of Time Series Analysis, 40(4):383–384, July 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2019:IIe

[Ano19e] Anonymous. Issue information. Journal of Time Series Analysis, 40(5):629–630, September 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2019:IIf

[Ano19f] Anonymous. Issue information. Journal of Time Series Analysis, 40(6):867–868, November 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2020:IIa

[Ano20a] Anonymous. Issue information. Journal of Time Series Analysis, 41(1):1–2, January 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2020:IIb

[Ano20b] Anonymous. Issue information. Journal of Time Series Analysis, 41(2):175–176, March 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2020:IIc

[Ano20c] Anonymous. Issue information. Journal of Time Series Analysis, 41(3):365–366, May 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2020:IId

[Ano20d] Anonymous. Issue information. Journal of Time Series Analysis, 41(4):487–488, July 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2020:IIe

[Ano20e] Anonymous. Issue information. Journal of Time Series Analysis, 41(5):603–604, September 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2020:IIf

[Ano20f] Anonymous. Issue information. Journal of Time Series Analysis, 41(6):731–732, November 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2020:PYN

[Ano20g] Anonymous. Publish your next paper open access in Journal of Time Series Analysis. Journal of Time Series Analysis, 41(4):491, July 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2021:CQB

[Ano21a] Anonymous. Correction to: Quasi-Bayesian Estimation of Time-Varying Volatility in DSGE Models by Katerina Petrova J. Time Series Anal, Vol. 40, No. 1 (2019). Journal of Time Series Analysis, 42(2):267, March 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic). See [Pet19].

Anonymous:2021:IIa

[Ano21b] Anonymous. Issue information. Journal of Time Series Analysis, 42(1):1–2, January 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2021:IIb

[Ano21c] Anonymous. Issue information. Journal of Time Series Analysis, 42(2):137–138, March 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2021:IIc

[Ano21d] Anonymous. Issue information. Journal of Time Series Analysis, 42(3):269–270, May 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2021:IId

[Ano21e] Anonymous. Issue information. Journal of Time Series Analysis, 42(4):375–376, July 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2021:IIe

[Ano21f] Anonymous. Issue information. Journal of Time Series Analysis, 42(5-6):493–494, September 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2022:IIa

[Ano22a] Anonymous. Issue information. Journal of Time Series Analysis, 43(1):1–2, January 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2022:IIb

[Ano22b] Anonymous. Issue information. Journal of Time Series Analysis, 43(2):155–156, March 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2022:IIc

[Ano22c] Anonymous. Issue information. Journal of Time Series Analysis, 43(3):343–344, May 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2022:IId

[Ano22d] Anonymous. Issue information. Journal of Time Series Analysis, 43(4):509–510, July 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2022:IIe

[Ano22e] Anonymous. Issue information. Journal of Time Series Analysis, 43(5):667–668, September 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2022:IIf

[Ano22f] Anonymous. Issue information. Journal of Time Series Analysis, 43(6):853–854, November 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2022:NAE

[Ano22g] Anonymous. New associate editors. Journal of Time Series Analysis, 43(6):855, November 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2023:EA

[Ano23a] Anonymous. Editorial announcement. Journal of Time Series Analysis, 44(3):261, May 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2023:IIa

[Ano23b] Anonymous. Issue information. Journal of Time Series Analysis, 44(1):1–2, January 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2023:IIb

[Ano23c] Anonymous. Issue information. Journal of Time Series Analysis, 44(2):149–150, March 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2023:IIc

[Ano23d] Anonymous. Issue information. Journal of Time Series Analysis, 44(3):259–260, May 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2023:IId

[Ano23e] Anonymous. Issue information. Journal of Time Series Analysis, 44(4):333–334, July 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2023:IIe

[Ano23f] Anonymous. Issue information. Journal of Time Series Analysis, 44(5-6):437–438, September 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2024:IIa

[Ano24a] Anonymous. Issue information. Journal of Time Series Analysis, 45(1):1–2, January 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2024:IIb

[Ano24b] Anonymous. Issue information. Journal of Time Series Analysis, 45(2):161–162, March 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anonymous:2024:IIc

[Ano24c] Anonymous. Issue information. Journal of Time Series Analysis, 45(3):331–332, May 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Agiakloglou:1993:BEF

[ANW93] Christos Agiakloglou, Paul Newbold, and Mark Wohar. Bias in an estimator of the fractional difference parameter. Journal of Time Series Analysis, 14(3):235–246, May 1993. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Arteche:2009:BBB

[AO09] Josu Arteche and Jesus Orbe. Bootstrap-based bandwidth choice for log-periodogram regression. Journal of Time Series Analysis, 30(6):591–617, November 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Al-Osh:1987:FOI

[AOA87] M. A. Al-Osh and A. A. Alzaid. First-order integer-valued autoregressive (INAR(1)) process. Journal of Time Series Analysis, 8(3): 261–275, May 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Aoki:1991:SST

[Aok91] Masanao Aoki. A state space time series modelling method without individual detrending. Journal of Time Series Analysis, 12(1):1– 26, January 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Arellano:1995:TTS

[AP95] Consuelo Arellano and Sastry G. Pantula. Testing for trend stationarity versus difference stationarity. Journal of Time Series Analysis, 16(2):147–164, March 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ashley:1986:DTN

[APH86] Richard A. Ashley, Douglas M. Patterson, and Melvin J. Hinich. A diagnostic test for nonlinear serial dependence in time series fitting errors. Journal of Time Series Analysis, 7(3):165–178, May 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Aktekin:2020:FMN

[APS20] Tevfik Aktekin, Nicholas G. Polson, and Refik Soyer. A family of multivariate non-Gaussian time series models. Journal of Time Series Analysis, 41(5):691–721, September 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Al-Qassam:1989:FEA

[AQL89] M. S. Al-Qassam and J. A. Lane. Forecasting exponential autoregressive models of order 1. Journal of Time Series Analysis, 10(2): 95–113, March 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Arteche:2000:SIS

[AR00] Josu Arteche and Peter M. Robinson. Semiparametric inference in seasonal and cyclical long memory processes. Journal of Time Series Analysis, 21(1):1–25, January 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Antunes:2006:HTS

[AR06] Ana Mónica C. Antunes and Tata Subba Rao. On hypotheses testing for the selection of spatio-temporal models. Journal of Time Series Analysis, 27(5):767–791, September 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Aknouche:2010:IID

[AR10] Abdelhakim Aknouche and Nadia Rabehi. On an independent and identically distributed mixture bilinear time-series model. Journal of Time Series Analysis, 31(2):113–131, March 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Arbues:2008:EPT

[Arb08] Ignacio Arbués. An extended portmanteau test for VARMA models with mixing nonlinear constraints. Journal of Time Series Analysis, 29(5):741–761, September 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Akamanam:1986:EBT

[ARS86] S. I. Akamanam, M. Bhaskara Rao, and K. Subramanyam. On the ergodicity of bilinear time series models. Journal of Time Series Analysis, 7(3):157–163, May 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Arteche:2002:SRT

[Art02] Josu Arteche. Semiparametric robust tests on seasonal or cyclical long memory time series. Journal of Time Series Analysis, 23(3): 251–285, May 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Arvanitis:2014:SEI

[Arv14] Stelios Arvanitis. A simple example of an indirect estimator with discontinuous limit theory in the MA(1) model. Journal of Time Series Analysis, 35(6):536–557, November 2014. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anderson:2000:SIG

[AS00] T. W. Anderson and M. A. Stephens. Sign invariance in goodnessof-fit tests for time series. Journal of Time Series Analysis, 21 (5):489–496, September 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Asai:2021:QML

[AS21] Manabu Asai and Mike K. P. So. Quasi-maximum likelihood estimation of conditional autoregressive Wishart models. Journal of Time Series Analysis, 42(3):271–294, May 2021. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Aknouche:2024:MTB

[AS24] Abdelhakim Aknouche and Manuel G. Scotto. A multiplicative thinning-based integer-valued GARCH model. Journal of Time Series Analysis, 45(1):4–26, January 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Al-Sulami:2019:SDD

[ASJLZ19] Dawlah Al-Sulami, Zhenyu Jiang, Zudi Lu, and Jun Zhu. On a semiparametric data-driven nonlinear model with penalized spatiotemporal lag interactions. Journal of Time Series Analysis, 40(3): 327–342, May 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Azimmohseni:2015:SRD

[ASK15] M. Azimmohseni, A. R. Soltani, and M. Khalafi. Simulation of real discrete time Gaussian multivariate stationary processes with given spectral densities. Journal of Time Series Analysis, 36(6): 783–796, November 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anderson:2021:PTS

[ASM21] Paul L. Anderson, Farzad Sabzikar, and Mark M. Meerschaert. Parsimonious time series modeling for high frequency climate data. Journal of Time Series Analysis, 42(4):442–470, July 2021. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anderson:1986:WDN

[AT86] T. W. Anderson and Akimichi Takemura. Why do noninvertible estimated moving averages occur?*. Journal of Time Series Analysis, 7(4):235–254, July 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ahtola:1987:DLS

[AT87a] Juha Ahtola and George C. Tiao. Distributions of least squares estimators of autoregressive parameters for a process with complex roots on the unit circle. Journal of Time Series Analysis, 8(1):1– 14, January 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ahtola:1987:NAI

[AT87b] Juha Ahtola and George C. Tiao. A note on asymptotic inference in autoregressive models with roots on the unit circle. Journal of Time Series Analysis, 8(1):15–19, January 1987. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Abadir:1999:DCI

[AT99] Karim M. Abadir and A. M. Robert Taylor. On the definitions of (co-)integration. Journal of Time Series Analysis, 20(2):129–137, March 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Antunes:2003:BAE

[ATT03] M. Antunes, M. A. Amaral Turkman, and K. F. Turkman. A Bayesian approach to event prediction. Journal of Time Series Analysis, 24(6):631–646, November 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Audrino:2005:LLN

[Aud05] Francesco Audrino. Local likelihood for non-parametric ARCH(1) models. Journal of Time Series Analysis, 26(2):251–278, March 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Aue:2022:BRT

[Aue22] Alexander Aue. Book review: Time Series: a First Course with Bootstrap Starter, by McElroy, Tucker S. and Politis, Dimitris N.. Published by CRC Press, 2020. 586 pp. ISBN: 978-1-4398-7651-0. Journal of Time Series Analysis, 43(2):341–342, March 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anderson:1993:ARP

[AV93] P. L. Anderson and A. V. Vecchia. Asymptotic results for periodic autoregressive moving-average processes. Journal of Time Series Analysis, 14(1):1–18, January 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Arteche:2005:TTS

[AV05] J. Arteche and C. Velasco. Trimming and tapering semi-parametric estimates in asymmetric long memory time series. Journal of Time Series Analysis, 26(4):581–611, July 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Athanasopoulos:2008:CVM

[AV08] George Athanasopoulos and Farshid Vahid. A complete VARMA modelling methodology based on scalar components. Journal of Time Series Analysis, 29(3):533–554, May 2008. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Abry:1998:LRD

[AVF98] Patrice Abry, Darryl Veitch, and Patrick Flandrin. Long-range dependence: revisiting aggregation with wavelets. Journal of Time Series Analysis, 19(3):253–266, May 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Antoniano-Villalobos:2016:NMS

[AVW16] Isadora Antoniano-Villalobos and Stephen G. Walker. A nonparametric model for stationary time series. Journal of Time Series Analysis, 37(1):126–142, January 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Abraham:1988:STD

[AY88] Bovas Abraham and Nihal Yatawara. A score test for detection of time series outliers. Journal of Time Series Analysis, 9(2):109–119, March 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Anderson:1996:AAT

[AY96] T. W. Anderson and Linfeng You. Adequacy of asymptotic theory for goodness-of-fit criteria for spectral distributions. Journal of Time Series Analysis, 17(6):533–552, November 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Azzalini:1981:ROL

[Azz81] A. Azzalini. Replicated observations of low order autoregressive time series. Journal of Time Series Analysis, 2(2):63–70, March 1981. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Azzalini:1982:AFP

[Azz82] A. Azzalini. Approximate filtering of parameter driven processes. Journal of Time Series Analysis, 3(4):219–223, July 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bai:1993:PSR

[Bai93] Jushan Bai. On the partial sums of residuals in autoregressive and moving average models. Journal of Time Series Analysis, 14(3): 247–260, May 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bai:1994:LSE

[Bai94] Jushan Bai. Least squares estimation of a shift in linear processes. Journal of Time Series Analysis, 15(5):453–472, September 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Barone:1987:MGI

[Bar87] Piero Barone. A method for generating independent realizations of a multivariate normal stationary and invertible ARMA(p, q) process. Journal of Time Series Analysis, 8(2):125–130, March 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bardet:2000:TPS

[Bar00] Jean-Marc Bardet. Testing for the presence of self-similarity of Gaussian time series having stationary increments. Journal of Time Series Analysis, 21(5):497–515, September 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bardet:2024:NEL

[Bar24] Jean-Marc Bardet. A new estimator for LARCH processes. Journal of Time Series Analysis, 45(1):103–132, January 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Basawa:1998:BR

[Bas98] Iswar Basawa. Book review. Journal of Time Series Analysis, 19(2):251–252, March 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Battaglia:1983:IAM

[Bat83] Francesco Battaglia. Inverse autocovariances and a measure of linear determinism for a stationary process. Journal of Time Series Analysis, 4(2):79–87, March 1983. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Battaglia:1988:EIC

[Bat88] Francesco Battaglia. On the estimation of the inverse correlation function. Journal of Time Series Analysis, 9(1):1–10, January 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bauer:2005:CCS

[Bau05] Dietmar Bauer. Comparing the CCA subspace method to pseudo maximum likelihood methods in the case of no exogenous inputs. Journal of Time Series Analysis, 26(5):631–668, September 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Beltrato:1987:DBK

[BB87] Kaizô I. Beltrato and Peter Bloomfield. Determining the bandwidth of a kernel spectrum estimate. Journal of Time Series Analysis, 8(1):21–38, January 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Baragona:1995:LII

[BB95] Roberto Baragona and Francesco Battaglia. Linear interpolators and the inverse correlation function of non-stationary time series. Journal of Time Series Analysis, 16(6):531–538, November 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Brockwell:1999:CNE

[BB99] A. E. Brockwell and P. J. Brockwell. A class of non-embeddable ARMA processes. Journal of Time Series Analysis, 20(5):483– 486, September 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bardet:2007:IMF

[BB07a] Jean-Marc Bardet and Pierre Bertrand. Identification of the multiscale fractional Brownian motion with biomechanical applications. Journal of Time Series Analysis, 28(1):1–52, January 2007. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bianco:2007:REU

[BB07b] Ana Bianco and Graciela Boente. Robust estimators under semiparametric partly linear autoregression: Asymptotic behaviour and bandwidth selection. Journal of Time Series Analysis, 28(2): 274–306, March 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bondon:2012:LSE

[BB12] Pascal Bondon and Natalia Bahamonde. Least squares estimation of ARCH models with missing observations. Journal of Time Series Analysis, 33(6):880–891, November 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bhattacharjee:2014:EAM

[BB14] Monika Bhattacharjee and Arup Bose. Estimation of autocovariance matrices for infinite dimensional vector linear process. Journal of Time Series Analysis, 35(3):262–281, May 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Baragona:2016:ELO

[BBC16] Roberto Baragona, Francesco Battaglia, and Domenico Cucina. Empirical likelihood for outlier detection and estimation in autoregressive time series. Journal of Time Series Analysis, 37(3): 315–336, May 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bertsche:2023:DGV

[BBK23] Dominik Bertsche, Ralf Brüggemann, and Christian Kascha. Directed graphs and variable selection in large vector autoregressive models. Journal of Time Series Analysis, 44(2):223–246, March 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bazzi:2017:TVT

[BBKL17] Marco Bazzi, Francisco Blasques, Siem Jan Koopman, and Andre Lucas. Time-varying transition probabilities for Markov regime switching models. Journal of Time Series Analysis, 38(3):458–478, May 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Barnard:1995:ASC

[BC95] Roger W. Barnard and Kamal C. Chanda. An application of the Schur–Cohn algorithm to time series analysis. Journal of Time Series Analysis, 16(5):445–449, September 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Baragona:1997:OCA

[BC97] R. Baragona and F. Carlucci. An optimality criterion for aggregating a set of time series in a composite index. Journal of Time Series Analysis, 18(1):1–9, January 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Baillie:2001:EGM

[BC01] Richard T. Baillie and Huimin Chung. Estimation of GARCH models from the autocorrelations of the squares of a process. Journal of Time Series Analysis, 22(6):631–650, November 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bertelli:2002:NCA

[BC02] Stefano Bertelli and Massimiliano Caporin. A note on calculating autocovariances of long-memory processes. Journal of Time Series Analysis, 23(5):503–508, September 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bowden:2012:SSR

[BC12] Ross S. Bowden and Brenton R. Clarke. A single series representation of multiple independent ARMA processes. Journal of Time Series Analysis, 33(2):304–311, March 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Baillie:2019:LMR

[BCCR19] Richard T. Baillie, Fabio Calonaci, Dooyeon Cho, and Seunghwa Rho. Long memory, realized volatility and heterogeneous autoregressive models. Journal of Time Series Analysis, 40(4):609–628, July 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bagchi:2018:STW

[BCD18] Pramita Bagchi, Vaidotas Characiejus, and Holger Dette. A simple test for white noise in functional time series. Journal of Time Series Analysis, 39(1):54–74, January 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Berentsen:2017:SPL

[BCFFT17] Geir Drage Berentsen, Ricardo Cao, Mario Francisco-Fernández, and Dag TjØstheim. Some properties of local Gaussian correlation and other nonlinear dependence measures. Journal of Time Series Analysis, 38(2):352–380, March 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bhattacharjee:2023:WPC

[BCK23] Monika Bhattacharjee, Nilanjan Chakraborty, and Hira L. Koul. Weighted l_{1}-penalized corrected quantile regression for highdimensional temporally dependent measurement errors. Journal of Time Series Analysis, 44(5-6):442–473, September 2023. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bertail:2015:BRS

[BCT15] Patrice Bertail, Stéphan Clémençon, and Jessica Tressou. Bootstrapping robust statistics for Markovian data applications to regenerative R-statistics and L-statistics. Journal of Time Series Analysis, 36(3):462–480, May 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Breidt:1992:TRI

[BD92] F. J. Breidt and R. A. Davis. Time-reversibility, identifiability and independence of innovations for stationary time series. Journal of Time Series Analysis, 13(5):377–390, September 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Breidt:1995:IBP

[BDD95] F. Jay Breidt, Richard A. Davis, and William T. M. Dunsmuir. Improved bootstrap prediction intervals for autoregressions. Journal of Time Series Analysis, 16(2):177–200, March 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Birr:2018:WVS

[BDH^{+}18] Stefan Birr, Holger Dette, Marc Hallin, Tobias Kley, and Stanislav Volgushev. On Wigner–Ville spectra and the uniqueness of timevarying copula-based spectral densities. Journal of Time Series Analysis, 39(3):242–250, May 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bardet:2008:ULT

[BDL08] Jean-Marc Bardet, Paul Doukhan, and José Rafael León. Uniform limit theorems for the integrated periodogram of weakly dependent time series and their applications to Whittle’s estimate. Journal of Time Series Analysis, 29(5):906–945, September 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Banerjee:1998:ECM

[BDM98] Anindya Banerjee, Juan Dolado, and Ricardo Mestre. Errorcorrection mechanism tests for cointegration in a single-equation framework. Journal of Time Series Analysis, 19(3):267–283, May 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bradley:2021:PMR

[BDP21] Richard C. Bradley, Richard A. Davis, and Dimitris N. Politis. Preface to the Murray Rosenblatt memorial special issue of JTSA. Journal of Time Series Analysis, 42(5-6):495–498, September 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Beran:2012:RTI

[BDS12] Jan Beran, Bikramjit Das, and Dieter Schell. On robust tail index estimation for linear long-memory processes. Journal of Time Series Analysis, 33(3):406–423, May 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Beare:2018:URT

[Bea18] Brendan K. Beare. Unit root testing with unstable volatility. Journal of Time Series Analysis, 39(6):816–835, November 2018. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Becker:2006:STP

[BEL06] Ralf Becker, Walter Enders, and Junsoo Lee. A stationarity test in the presence of an unknown number of smooth breaks. Journal of Time Series Analysis, 27(3):381–409, May 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Berchtold:2001:EMT

[Ber01] Andre Berchtold. Estimation in the mixture transition distribution model. Journal of Time Series Analysis, 22(4):379–397, July 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Beran:2007:EUL

[Ber07] Jan Beran. On M-estimation under long-range dependence in volatility. Journal of Time Series Analysis, 28(1):138–153, January 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Betken:2016:TCP

[Bet16] Annika Betken. Testing for change-points in long-range dependent time series by means of a self-normalized Wilcoxon test. Journal of Time Series Analysis, 37(6):785–809, November 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Breidt:2012:ASR

[BEvdW12] F. Jay Breidt, Andreea Erciulescu, and Mark van der Woerd. Autocovariance structures for radial averages in small-angle X-ray scattering experiments. Journal of Time Series Analysis, 33(5): 704–717, September 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Boswijk:1996:URP

[BF96] H. Peter Boswijk and Philip Hans Franses. Unit roots in periodic autoregressions. Journal of Time Series Analysis, 17(3):221–245, May 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Berlinet:1997:BFN

[BF97] Alain Berlinet and Christian Francq. On Bartlett’s formula for non-linear processes. Journal of Time Series Analysis, 18(6):535– 552, November 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bartolucci:2010:NMT

[BF10] Francesco Bartolucci and Alessio Farcomeni. A note on the mixture transition distribution and hidden Markov models. Journal of Time Series Analysis, 31(2):132–138, March 2010. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Brockwell:2012:HFS

[BFK12] Peter J. Brockwell, Vincenzo Ferrazzano, and Claudia Klüppelberg. High-frequency sampling of a continuous-time ARMA process. Journal of Time Series Analysis, 33(1):152–160, January 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Brockwell:2013:HFS

[BFK13] Peter J. Brockwell, Vincenzo Ferrazzano, and Claudia Klüppelberg. High-frequency sampling and kernel estimation for continuous-time moving average processes. Journal of Time Series Analysis, 34(3): 385–404, May 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bucher:2019:CCS

[BFK19] Axel Bücher, Jean-David Fermanian, and Ivan Kojadinovic. Combining cumulative sum change-point detection tests for assessing the stationarity of univariate time series. Journal of Time Series Analysis, 40(1):124–150, January 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Broze:2002:EUH

[BFZ02] Laurence Broze, Christian Francq, and Jean-Michel Zakoïan. Efficient use of higher-lag autocorrelations for estimating autoregressive processes. Journal of Time Series Analysis, 23(3):287–312, May 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Beran:1995:TEV

[BG95] Jan Beran and Theo Gasser. Testing equality of variances for paired time series. Journal of Time Series Analysis, 16(2):165– 176, March 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Beran:2000:EDF

[BG00] Jan Beran and Sucharita Ghosh. Estimation of the dominating frequency for stationary and nonstationary fractional autoregressive models. Journal of Time Series Analysis, 21(5):517–533, September 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Beran:2020:EMD

[BG20] Jan Beran and Sucharita Ghosh. Estimating the mean direction of strongly dependent circular time series. Journal of Time Series Analysis, 41(2):210–228, March 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Babayan:2021:ERR

[BGT21] Nikolay M. Babayan, Mamikon S. Ginovyan, and Murad S. Taqqu. Extensions of Rosenblatt’s results on the asymptotic behavior of the prediction error for deterministic stationary sequences. Journal of Time Series Analysis, 42(5-6):622–652, September 2021. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bell:1991:IKF

[BH91] William Bell and Steven Hillmer. Initializing the Kalman filter for nonstationary time series models. Journal of Time Series Analysis, 12(4):283–300, July 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bera:1992:TCH

[BH92] A. K. Bera and M. L. Higgins. A test for conditional heteroskedasticity in time series models. Journal of Time Series Analysis, 13 (6):501–519, November 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bentarzi:1994:IPM

[BH94] Mohamed Bentarzi and Marc Hallin. On the invertibility of periodic moving-average models. Journal of Time Series Analysis, 15(3):263–268, May 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Busetti:2001:TPR

[BH01] Fabio Busetti and Andrew Harvey. Testing for the presence of a random walk in series with structural breaks. Journal of Time Series Analysis, 22(2):127–150, March 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Busetti:2003:FCS

[BH03] Fabio Busetti and Andrew Harvey. Further comments on stationarity tests in series with structural breaks at unknown points. Journal of Time Series Analysis, 24(2):137–140, March 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Burridge:2008:CEO

[BH08] Peter Burridge and Daniela Hristova. Consistent estimation and order selection for nonstationary autoregressive processes with stable innovations. Journal of Time Series Analysis, 29(4):695–718, July 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Busetti:2010:TSS

[BH10] Fabio Busetti and Andrew Harvey. Tests of strict stationarity based on quantile indicators. Journal of Time Series Analysis, 31(6):435– 450, November 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bayer:2013:CNC

[BH13] Christian Bayer and Christoph Hanck. Combining non-cointegration tests. Journal of Time Series Analysis, 34(1):83–95, January 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bhattacharyya:1982:LPD

[Bha82] M. N. Bhattacharyya. Lydia Pinkham data remodelled. Journal of Time Series Analysis, 3(2):81–102, March 1982. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bhansali:1983:EOM

[Bha83] R. J. Bhansali. Estimation of the order of a moving average model from autoregressive and window estimates of the inverse correlation function. Journal of Time Series Analysis, 4(3):137–162, May 1983. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bhansali:1986:CAT

[Bha86] R. J. Bhansali. The criterion autoregressive transfer function of Parzen. Journal of Time Series Analysis, 7(2):79–104, March 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bhansali:1989:EMA

[Bha89] R. J. Bhansali. Estimation of the moving-average representation of a stationary process by autoregressive model fitting. Journal of Time Series Analysis, 10(3):215–232, May 1989. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bhansali:1993:EPE

[Bha93] R. J. Bhansali. Estimation of the prediction error variance and an R^{2} measure by autoregressive model fitting. Journal of Time Series Analysis, 14(2):125–146, March 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bhansali:1997:RAS

[Bha97] R. J. Bhansali. Robustness of the autoregressive spectral estimate for linear processes with infinite variance. Journal of Time Series Analysis, 18(3):213–229, May 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Beckman:1990:ACF

[BHL90] Stig-Inge Beckman, Jan Holst, and Georg Lindgren. Alarm characteristics for a flood warning system with deterministic components. Journal of Time Series Analysis, 11(1):1–18, January 1990. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bloomfield:1994:PCS

[BHL94] Peter Bloomfield, Harry L. Hurd, and Robert B. Lund. Periodic correlation in stratospheric ozone data. Journal of Time Series Analysis, 15(2):127–150, March 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Berkes:2009:ENR

[BHL09] István Berkes, Lajos Horváth, and Shiqing Ling. Estimation in nonstationary random coefficient autoregressive models. Journal of Time Series Analysis, 30(4):395–416, July 2009. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Berkes:2011:TSC

[BHLS11] István Berkes, Lajos Horváth, Shiqing Ling, and Johannes Schauer. Testing for structural change of AR model to threshold AR model. Journal of Time Series Analysis, 32(5):547–565, September 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Boshnakov:2009:GTS

[BI09] Georgi N. Boshnakov and Bisher M. Iqelan. Generation of time series models with given spectral properties. Journal of Time Series Analysis, 30(3):349–368, May 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Boshnakov:2012:MEM

[BI12] Georgi N. Boshnakov and Bisher M. Iqelan. Maximum entropy models for general lag patterns. Journal of Time Series Analysis, 33(1):112–120, January 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Banerjee:2017:TPC

[BiS17] Anindya Banerjee and Josep Lluís Carrion i Silvestre. Testing for panel cointegration using common correlated effects estimators. Journal of Time Series Analysis, 38(4):610–636, July 2017. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bandyopadhyay:2017:SDT

[BJR17] Soutir Bandyopadhyay, Carsten Jentsch, and Suhasini Subba Rao. A spectral domain test for stationarity of spatio-temporal data. Journal of Time Series Analysis, 38(2):326–351, March 2017. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Blake:2003:PST

[BK03] Andrew P. Blake and George Kapetanios. Pure significance tests of the unit root hypothesis against nonlinear alternatives. Journal of Time Series Analysis, 24(3):253–267, May 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Blake:2007:TNN

[BK07] Andrew P. Blake and George Kapetanios. Testing for neglected nonlinearity in cointegrating relationships*. Journal of Time Series Analysis, 28(6):807–826, November 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Betken:2019:TCL

[BK19] Annika Betken and Rafał Kulik. Testing for change in long-memory stochastic volatility time series. Journal of Time Series Analysis, 40(5):707–738, September 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Braumann:2021:SIA

[BKM21] Alexander Braumann, Jens-Peter Kreiss, and Marco Meyer. Simultaneous inference for autocovariances based on autoregressive sieve bootstrap. Journal of Time Series Analysis, 42(5-6):534– 553, September 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Barnett:1997:RBE

[BKS97] Glen Barnett, Robert Kohn, and Simon Sheather. Robust Bayesian estimation of autoregressive–moving-average models. Journal of Time Series Analysis, 18(1):11–28, January 1997. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Basawa:2001:LSP

[BL01] I. V. Basawa and Robert Lund. Large sample properties of parameter estimates for periodic ARMA models. Journal of Time Series Analysis, 22(6):651–663, November 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Brockwell:2013:ICP

[BL13] Peter Brockwell and Alexander Lindner. Integration of CARMA processes and spot volatility modelling. Journal of Time Series Analysis, 34(2):156–167, March 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Brockwell:2019:SEI

[BL19] Peter J. Brockwell and Alexander Lindner. Sampling, embedding and inference for CARMA processes. Journal of Time Series Analysis, 40(2):163–181, March 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Brockwell:2021:ANC

[BL21] Peter J. Brockwell and Alexander Lindner. Aspects of non-causal and non-invertible CARMA processes. Journal of Time Series Analysis, 42(5-6):777–790, September 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Blasques:2014:TPN

[Bla14] Francisco Blasques. Transformed polynomials for nonlinear autoregressive models of the conditional mean. Journal of Time Series Analysis, 35(3):218–238, May 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Biscay:2005:ENA

[BLL05] R. J. Biscay, Marc Lavielle, and Carenne Ludeña. Estimation of nonparametric autoregressive time series models under dynamical constraints. Journal of Time Series Analysis, 26(3):371–397, May 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Boshnakov:2009:MEP

[BLL09] Georgi N. Boshnakov and Sophie Lambert-Lacroix. Maximum entropy for periodically correlated processes from nonconsecutive autocovariance coefficients. Journal of Time Series Analysis, 30(5): 467–486, September 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Brockwell:1992:EST

[BLT92] Peter J. Brockwell, Jian Liu, and Richard L. Tweedie. On the existence of stationary threshold autoregressive moving-average processes. Journal of Time Series Analysis, 13(2):95–107, March 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Birch:1981:CIR

[BM81] Jeffrey B. Birch and R. Douglas Martin. Confidence intervals for robust estimates of the first order autoregressive parameter. Journal of Time Series Analysis, 2(4):205–220, July 1981. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Batts:1989:PPT

[BM89] John T. Batts and Robert F. McNown. The predictive performance of three autoregressive moving-average models: A Monte Carlo investigation. Journal of Time Series Analysis, 10(4):301–314, July 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Billard:1991:EPE

[BM91] L. Billard and Fouad Y. Mohamed. Estimation of the parameters of an Ear(p) process. Journal of Time Series Analysis, 12(3):179–192, May 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bose:2003:EAP

[BM03] Arup Bose and Kanchan Mukherjee. Estimating the ARCH parameters by solving linear equations. Journal of Time Series Analysis, 24(2):127–136, March 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bell:2004:CAS

[BM04] William R. Bell and Donald E. K. Martin. Computation of asymmetric signal extraction filters and mean squared error for ARIMA component models. Journal of Time Series Analysis, 25(4):603– 623, July 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bose:2009:BWL

[BM09] Arup Bose and Kanchan Mukherjee. Bootstrapping a weighted linear estimator of the ARCH parameters. Journal of Time Series Analysis, 30(3):315–331, May 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Borkowski:2010:PSE

[BM10] Piotr Borkowski and Jan Mielniczuk. Postmodel selection estimators of variance function for nonlinear autoregression. Journal of Time Series Analysis, 31(1):50–63, January 2010. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bartlett:2013:ENN

[BM13] A. Bartlett and W. P. McCormick. Estimation for non-negative time series with heavy-tail innovations. Journal of Time Series Analysis, 34(1):96–115, January 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bu:2008:MLE

[BMH08] Ruijun Bu, Brendan McCabe, and Kaddour Hadri. Maximum likelihood estimation of higher-order integer-valued autoregressive processes. Journal of Time Series Analysis, 29(6):973–994, November 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bu:2009:C

[BMH09] Ruijun Bu, Brendan McCabe, and Kaddour Hadri. Corrigendum. Journal of Time Series Analysis, 30(2):260–261, March 2009. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ben:1999:REV

[BMY99] Marta Garcia Ben, Elena J. Martinez, and Victor J. Yohai. Robust estimation in vector autoregressive moving-average models. Journal of Time Series Analysis, 20(4):381–399, July 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Burman:1992:DDE

[BN92] P. Burman and D. Nolan. Data-dependent estimation of prediction functions. Journal of Time Series Analysis, 13(3):189–207, May 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Battaglia:2005:ODE

[BO05] Francesco Battaglia and Lia Orfei. Outlier detection and estimation in nonlinear time series. Journal of Time Series Analysis, 26 (1):107–121, January 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bollerslev:1988:CSG

[Bol88] Tim Bollerslev. On the correlation structure for the generalized autoregressive conditional heteroskedastic process. Journal of Time Series Analysis, 9(2):121–131, March 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bondon:2001:RRM

[Bon01] Pascal Bondon. Recursive relations for multistep prediction of a stationary time series. Journal of Time Series Analysis, 22(4): 399–410, July 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bondon:2005:IMV

[Bon05] Pascal Bondon. Influence of missing values on the prediction of a stationary time series. Journal of Time Series Analysis, 26(4): 519–525, July 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Boshnakov:1996:RCP

[Bos96] Georgi N. Boshnakov. Recursive computation of the parameters of periodic autoregressive moving-average processes. Journal of Time Series Analysis, 17(4):333–349, July 1996. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Boshnakov:2009:BRT

[Bos09] Georgi N. Boshnakov. Book review: Time series analysis with applications in R series: Springer texts in statistics, 2nd edition. Journal of Time Series Analysis, 30(6):708–709, November 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Boshnakov:2010:BRI

[Bos10] Georgi N. Boshnakov. Book review: Introductory Time Series with R. Journal of Time Series Analysis, 31(5):406, September 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Boshnakov:2016:BRI

[Bos16] Georgi N. Boshnakov. Book review: Introduction to Time Series Analysis and Forecasting, 2nd Edition, Wiley Series in Probability and Statistics, by Douglas C. Montgomery, Cheryl L. Jennings and Murat Kulahci (eds). Published by John Wiley and Sons, Hoboken, NJ, USA, 2015. Total number of pages: 672 Hardcover: ISBN: 978-1-118-74511-3, e-book: ISBN: 978-1-118-74515-1, etext: ISBN: 978-1-118-74495-6. Journal of Time Series Analysis, 37(6):864, November 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Boutahar:2008:IPC

[Bou08] Mohamed Boutahar. Identification of persistent cycles in non-Gaussian long-memory time series. Journal of Time Series Analysis, 29(4):653–672, July 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Blanke:2003:OSD

[BP03] D. Blanke and B. Pumo. Optimal sampling for density estimation in continuous time. Journal of Time Series Analysis, 24(1):1–23, January 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bondon:2007:CAP

[BP07] Pascal Bondon and Wilfredo Palma. A class of antipersistent processes. Journal of Time Series Analysis, 28(2):261–273, March 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Battaglia:2011:TVM

[BP11] Francesco Battaglia and Mattheos K. Protopapas. Time-varying multi-regime models fitting by genetic algorithms. Journal of Time Series Analysis, 32(3):237–252, May 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Baek:2012:STS

[BP12] Changryong Baek and Vladas Pipiras. Statistical tests for a single change in mean against long-range dependence. Journal of Time Series Analysis, 33(1):131–151, January 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bykhovskaya:2018:BLT

[BP18] Anna Bykhovskaya and Peter C. B. Phillips. Boundary limit theory for functional local to unity regression. Journal of Time Series Analysis, 39(4):523–562, July 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bagnato:2012:AGD

[BPN12] Luca Bagnato, Antonio Punzo, and Orietta Nicolis. The autodependogram: a graphical device to investigate serial dependences. Journal of Time Series Analysis, 33(2):233–254, March 2012. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Brailsford:2002:SFF

[BPT02] T. J. Brailsford, Jack H. W. Penm, and R. D. Terrell. Selecting the forgetting factor in subset autoregressive modelling. Journal of Time Series Analysis, 23(6):629–649, November 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bouhaddioui:2006:GPT

[BR06] Chafik Bouhaddioui and Roch Roy. A generalized portmanteau test for independence of two infinite-order vector autoregressive series. Journal of Time Series Analysis, 27(4):505–544, July 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bravo:2005:BEE

[Bra05] Francesco Bravo. Blockwise empirical entropy tests for time series regressions. Journal of Time Series Analysis, 26(2):185–210, March 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bravo:2011:IGM

[Bra11] Francesco Bravo. Improved generalized method of moments estimators for weakly dependent observations. Journal of Time Series Analysis, 32(6):680–698, November 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bramati:2013:COT

[Bra13] Maria Caterina Bramati. A class of optimal tests for contemporaneous non-causality in VAR models. Journal of Time Series Analysis, 34(3):330–344, May 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bravo:2016:LIT

[Bra16] Francesco Bravo. Local information theoretic methods for smooth coefficients dynamic panel data models. Journal of Time Series Analysis, 37(5):690–708, September 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bradley:2021:SBF

[Bra21] Richard C. Bradley. On some basic features of strictly stationary, reversible Markov chains. Journal of Time Series Analysis, 42(5-6):499–533, September 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bravo:2022:MSM

[Bra22] Francesco Bravo. Misspecified semiparametric model selection with weakly dependent observations. Journal of Time Series Analysis, 43(4):558–586, July 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Breitung:1994:SST

[Bre94] Jorg Breitung. Some simple tests of the moving-average unit root hypothesis. Journal of Time Series Analysis, 15(4):351–370, July 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Brillinger:1980:CLS

[Bri80] David R. Brillinger. The comparison of least squares and thirdorder periodogram procedures in the estimation of bifrequency. Journal of Time Series Analysis, 1(2):95–102, March 1980. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Brillinger:2012:NBE

[Bri12] David R. Brillinger. The Nicholson blowfly experiments: some history and EDA. Journal of Time Series Analysis, 33(5):718– 723, September 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Brockwell:1995:NED

[Bro95] Peter J. Brockwell. A note on the embedding of discrete-time ARMA processes. Journal of Time Series Analysis, 16(5):451– 460, September 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Brockwell:2007:LBA

[Bro07] A. E. Brockwell. Likelihood-based analysis of a class of generalized long-memory time series models. Journal of Time Series Analysis, 28(3):386–407, May 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Burman:1998:SMS

[BS98] Prabir Burman and Robert Shumway. Semiparametric modeling of seasonal time series. Journal of Time Series Analysis, 19(2): 127–145, March 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Breitung:2002:TAS

[BS02] JÖRg Breitung and Norman R. Swanson. Temporal aggregation and spurious instantaneous causality in multiple time series models. Journal of Time Series Analysis, 23(6):651–665, November 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bandt:2007:OPT

[BS07] Chstoph Bandt and Faten Shiha. Order patterns in time series. Journal of Time Series Analysis, 28(5):646–665, September 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Barreto-Souza:2015:ZMG

[BS15a] Wagner Barreto-Souza. Zero-modified geometric INAR(1) process for modelling count time series with deflation or inflation of zeros. Journal of Time Series Analysis, 36(6):839–852, November 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Beare:2015:VCS

[BS15b] Brendan K. Beare and Juwon Seo. Vine copula specifications for stationary multivariate Markov chains. Journal of Time Series Analysis, 36(2):228–246, March 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Beran:2018:LTR

[BSG18] Jan Beran, Britta Steffens, and Sucharita Ghosh. On local trigonometric regression under dependence. Journal of Time Series Analysis, 39(4):592–617, July 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Beare:2017:CLP

[BSS17] Brendan K. Beare, Juwon Seo, and Won-Ki Seo. Cointegrated linear processes in Hilbert space. Journal of Time Series Analysis, 38(6):1010–1027, November 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Beran:1994:ELM

[BT94] Jan Beran and Norma Terrin. Estimation of the long-memory parameter, based on a multivariate central limit theorem. Journal of Time Series Analysis, 15(3):269–278, May 1994. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Burridge:2006:AOD

[BT06] Peter Burridge and A. M. Robert Taylor. Additive outlier detection via extreme-value theory. Journal of Time Series Analysis, 27(5): 685–701, September 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bai:2013:MLT

[BT13] Shuyang Bai and Murad S. Taqqu. Multivariate limit theorems in the context of long-range dependence. Journal of Time Series Analysis, 34(6):717–743, November 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Buhlmann:1996:LAL

[Büh96] Peter Bühlmann. Locally adaptive lag-window spectral estimation. Journal of Time Series Analysis, 17(3):247–270, May 1996. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Belcher:2000:TSE

[BW00] John Belcher and Granville Tunnicliffe Wilson. Time scale estimation by tracking parameter variation. Journal of Time Series Analysis, 21(3):237–248, May 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bruder:2018:BBJ

[BW18] Stefan Bruder and Michael Wolf. Balanced bootstrap joint confidence bands for structural impulse response functions. Journal of Time Series Analysis, 39(5):641–664, September 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Bradley:2019:STM

[BWH19] Jonathan R. Bradley, Christopher K. Wikle, and Scott H. Holan. Spatio-temporal models for big multinomial data using the conditional multivariate logit-beta distribution. Journal of Time Series Analysis, 40(3):363–382, May 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chen:1999:PTA

[CA99] Z. G. Chen and O. D. Anderson. Polyvariograms and their asymptotes. Journal of Time Series Analysis, 20(5):387–512, September 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chen:2011:SYW

[CADF11] Weitian Chen, Brian D. O. Anderson, Manfred Deistler, and Alexander Filler. Solutions of Yule–Walker equations for singular AR processes. Journal of Time Series Analysis, 32(5):531–538, September 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cai:2011:MVT

[Cai11] Yuzhi Cai. Multi-variate time-series simulation. Journal of Time Series Analysis, 32(5):566–579, September 2011. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cameron:1987:ANP

[Cam87] M. A. Cameron. An automatic non-parametric spectrum estimator. Journal of Time Series Analysis, 8(4):379–387, July 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Campbell:2004:BST

[Cam04] Edward P. Campbell. Bayesian selection of threshold autoregressive models. Journal of Time Series Analysis, 25(4):467–482, July 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cao:2019:DDA

[Cao19] Jiguo Cao. Book review: Dynamic Data Analysis, by James Ramsay and Giles Hooker. Published by Springer, New York, USA, 2017. Total number of pages: 230. ISSN: 0172-7397. Journal of Time Series Analysis, 40(1):158–159, January 2019. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chen:1994:LWE

[CAP94] Gemai Chen, Bovas Abraham, and Shelton Peiris. Lag window estimation of the degree of differencing in fractionally integrated time series models. Journal of Time Series Analysis, 15(5):473– 487, September 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cartwright:1985:FTS

[Car85] Phillip A. Cartwright. Forecasting time series: A comparative analysis of alternative classes of time series models. Journal of Time Series Analysis, 6(4):203–211, July 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cavicchioli:2014:ALF

[Cav14a] Maddalena Cavicchioli. Analysis of the likelihood function for Markov-switching VAR(CH) models. Journal of Time Series Analysis, 35(6):624–639, November 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cavicchioli:2014:DNR

[Cav14b] Maddalena Cavicchioli. Determining the number of regimes in Markov switching VAR and VMA models. Journal of Time Series Analysis, 35(2):173–186, March 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Caporale:2013:MLR

[CCGA13] Guglielmo Maria Caporale, Juncal Cuñado, and Luis A. Gil-Alana. Modelling long-run trends and cycles in financial time series data. Journal of Time Series Analysis, 34(3):405–421, May 2013. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chen:2016:BCE

[CCY16] Kun Chen, Ngai Hang Chan, and Chun Yip Yau. Bartlett correction of empirical likelihood for non-Gaussian short-memory time series. Journal of Time Series Analysis, 37(5):624–649, September 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Coates:1986:TCT

[CD86] D. S. Coates and P. J. Diggle. Tests for comparing two estimated spectral densities. Journal of Time Series Analysis, 7(1):7–20, January 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chang:1994:ROT

[CD94] Ming Chun Chang and David A. Dickey. Recognizing overdifferenced time series. Journal of Time Series Analysis, 15(1):1–18, January 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chen:2009:RLR

[CD09] Willa W. Chen and Rohit S. Deo. The restricted likelihood ratio test at the boundary in autoregressive series. Journal of Time Series Analysis, 30(6):618–630, November 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chen:2012:RLR

[CD12] Willa W. Chen and Rohit S. Deo. The restricted likelihood ratio test for autoregressive processes. Journal of Time Series Analysis, 33(2):325–339, March 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cooil:1998:DET

[CF98] Bruce Cooil and Luke Froeb. A difference estimator for testing equality of variances for paired time series. Journal of Time Series Analysis, 19(3):285–290, May 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Christou:2014:QLI

[CF14] Vasiliki Christou and Konstantinos Fokianos. Quasi-likelihood inference for negative binomial time series models. Journal of Time Series Analysis, 35(1):55–78, January 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cho:2024:MCP

[CF24] Haeran Cho and Piotr Fryzlewicz. Multiple change point detection under serial dependence: Wild contrast maximisation and gappy Schwarz algorithm. Journal of Time Series Analysis, 45(3):479– 494, May 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Clarke:1982:CPS

[CG82] B. R. Clarke and E. J. Godolphin. Comparative power studies for goodness of fit tests of time series models. Journal of Time Series Analysis, 3(3):141–151, May 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Canepa:2007:IQL

[CG07] A. Canepa and L. G. Godfrey. Improvement of the quasi-likelihood ratio test in ARMA models: some results for bootstrap methods. Journal of Time Series Analysis, 28(3):434–453, May 2007. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Craigmile:2011:STM

[CG11] Peter F. Craigmile and Peter Guttorp. Space-time modelling of trends in temperature series. Journal of Time Series Analysis, 32 (4):378–395, July 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chan:2019:EFO

[CG19] Kung-Sik Chan and Greta Goracci. On the ergodicity of firstorder threshold autoregressive moving-average processes. Journal of Time Series Analysis, 40(2):256–264, March 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chiogna:2008:AIS

[CGM08] Monica Chiogna, Carlo Gaetan, and Guido Masarotto. Automatic identification of seasonal transfer function models by means of iterative stepwise and genetic algorithms. Journal of Time Series Analysis, 29(1):37–50, January 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chronopoulos:2015:DBC

[CGN15] Dimitris K. Chronopoulos, Claudia Girardone, and John C. Nankervis. Double bootstrap confidence intervals in the two-stage DEA approach. Journal of Time Series Analysis, 36(5):653–662, September 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chan:2022:SVS

[CGP22] Ngai Hang Chan, Linhao Gao, and Wilfredo Palma. Simultaneous variable selection and structural identification for time-varying coefficient models. Journal of Time Series Analysis, 43(4):511–531, July 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cressie:2011:ESI

[CH11] Noel Cressie and Scott H. Holan. Editorial: Special issue on time series in the environmental sciences. Journal of Time Series Analysis, 32(4):337–338, July 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Caivano:2014:TSM

[CH14] Michele Caivano and Andrew Harvey. Time-series models with an EGB2 conditional distribution. Journal of Time Series Analysis, 35(6):558–571, November 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cheng:2015:MNS

[CH15] Raymond Cheng and Charles B. Harris. Mixed-norm spaces and prediction of SαS moving averages. Journal of Time Series Analysis, 36(6):853–875, November 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chen:2021:REL

[CH21] Kun Chen and Rui Huang. Robust empirical likelihood for time series. Journal of Time Series Analysis, 42(1):4–18, January 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chanda:1987:AED

[Cha87] Kamal C. Chanda. Asymptotic expansions for the distributions of serial correlations. Journal of Time Series Analysis, 8(3):283–291, May 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chanda:1991:SCL

[Cha91] Kamal C. Chanda. Stationarity and central limit theorem associated with bilinear time series models. Journal of Time Series Analysis, 12(4):301–313, July 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chanda:1995:LSA

[Cha95] Kamal C. Chanda. Large sample analysis of autoregressive movingaverage models with errors in variables. Journal of Time Series Analysis, 16(1):1–15, January 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chambers:1999:NMS

[Cha99] Marcus J. Chambers. A note on modelling seasonal processes in continuous time. Journal of Time Series Analysis, 20(2):139–143, March 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chanda:2005:LSP

[Cha05a] Kamal C. Chanda. Large sample properties of spectral estimators for a class of stationary nonlinear processes. Journal of Time Series Analysis, 26(1):1–16, January 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chandler:2005:BR

[Cha05b] Richard E. Chandler. Book reviews. Journal of Time Series Analysis, 26(5):783–784, September 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chambers:2015:CSL

[Cha15a] Marcus J. Chambers. The calculation of some limiting distributions arising in near-integrated models with GLS detrending. Journal of Time Series Analysis, 36(4):562–586, July 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chambers:2015:TUR

[Cha15b] Marcus J. Chambers. Testing for a unit root in a near-integrated model with skip-sampled data. Journal of Time Series Analysis, 36 (5):630–649, September 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chambers:2016:BRU

[Cha16a] Marcus J. Chambers. Book review: Unobserved Components and Time Series Econometrics, edited by Siem Jan Koopman and Neil Shephard. Published by Oxford University Press, Oxford, 2015. Total number of pages: 400. ISBN: 978-0-19-968366-6. Journal of Time Series Analysis, 37(6):862–863, November 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chavez:2016:CMM

[Cha16b] Gordon Chavez. Conditional and marginal mutual information in Gaussian and hyperbolic decay time series. Journal of Time Series Analysis, 37(6):851–861, November 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chambers:2019:FDE

[Cha19] Marcus J. Chambers. Frequency domain estimation of continuous time cointegrated models with mixed frequency and mixed sample data. Journal of Time Series Analysis, 40(6):887–913, November 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cheng:1991:PEE

[Che91] Qiansheng Cheng. Parameter estimation in exponential models. Journal of Time Series Analysis, 12(1):27–40, January 1991. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cheung:1993:TFI

[Che93] Yin-Wong Cheung. Tests for fractional integration: A Monte Carlo investigation. Journal of Time Series Analysis, 14(4):331–345, July 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chen:1995:TVS

[Che95] Rong Chen. Threshold variable selection in open-loop threshold autoregressive models. Journal of Time Series Analysis, 16(5): 461–481, September 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chen:2006:ALF

[Che06] Wen-Den Chen. An approximate likelihood function for panel data with a mixed ARMA(p, q) remainder disturbance model. Journal of Time Series Analysis, 27(6):911–921, November 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chen:2009:BRA

[Che09] Willa W. Chen. Book review: Analysis of Integrated and Cointegrated Time Series with R, 2nd edition. Journal of Time Series Analysis, 30(5):575, September 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chiu:1991:LEP

[Chi91] Shean-Tsong Chiu. A linear estimation procedure for the parameters of autoregressive moving-average processes. Journal of Time Series Analysis, 12(4):315–327, July 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cavaliere:2015:TUR

[CHLT15] Giuseppe Cavaliere, David I. Harvey, Stephen J. Leybourne, and A. M. Robert Taylor. Testing for unit roots under multiple possible trend breaks and non-stationary volatility using bootstrap minimum Dickey–Fuller statistics. Journal of Time Series Analysis, 36(5):603–629, September 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Choi:1991:ADG

[Cho91] Byoung Seon Choi. On the asymptotic distribution of the generalized partial autocorrelation function in autoregressive movingaverage processes. Journal of Time Series Analysis, 12(3):193–205, May 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cao:2017:DTL

[CHS17] Wen Cao, Clifford Hurvich, and Philippe Soulier. Drift in transaction-level asset price models. Journal of Time Series Analysis, 38(5):769–790, September 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chung:1996:GFI

[Chu96] Ching-Fan Chung. A generalized fractionally integrated autoregressive moving-average process. Journal of Time Series Analysis, 17(2):111–140, March 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chu:2012:LTD

[Chu12] Ba Chu. Limit theorems for the discount sums of moving averages. Journal of Time Series Analysis, 33(1):1–12, January 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Carrion-I-Silvestre:2016:BBU

[CISG16] Josep Lluís Carrion-I-Silvestre and María Dolores Gadea. Bounds, breaks and unit root tests. Journal of Time Series Analysis, 37(2): 165–181, March 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cantarelis:1982:LVE

[CJ82] N. Cantarelis and F. R. Johnston. On-line variance estimation for the steady state Bayesian forecasting model. Journal of Time Series Analysis, 3(4):225–234, July 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chen:2013:GIG

[CK13] Jhih-Gang Chen and Biing-Shen Kuo. Gaussian inference in general AR(1) models based on difference. Journal of Time Series Analysis, 34(4):447–453, July 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Conrad:2015:TMG

[CK15] Christian Conrad and Menelaos Karanasos. On the transmission of memory in GARCH-in-Mean models. Journal of Time Series Analysis, 36(5):706–720, September 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chun:2022:SHA

[CK22] Dohyun Chun and Donggyu Kim. State heterogeneity analysis of financial volatility using high-frequency financial data. Journal of Time Series Analysis, 43(1):105–124, January 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Constantinou:2018:TSF

[CKR18] Panayiotis Constantinou, Piotr Kokoszka, and Matthew Reimherr. Testing separability of functional time series. Journal of Time Series Analysis, 39(5):731–747, September 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chen:1995:BIT

[CL95a] Cathy W. S. Chen and Jack C. Lee. Bayesian inference of threshold autoregressive models. Journal of Time Series Analysis, 16(5):483– 492, September 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cheung:1995:EFS

[CL95b] Yin-Wong Cheung and Kon S. Lai. Estimating finite sample critical values for unit root tests using pure random walk processes: A note. Journal of Time Series Analysis, 16(5):493–498, September 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cappuccio:1997:SRB

[CL97] Nunzio Cappuccio and Diego Lubian. Spurious regressions between I(1) processes with long memory errors. Journal of Time Series Analysis, 18(4):341–354, July 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Comte:2000:SON

[CL00] Fabienne Comte and Offer Lieberman. Second-order noncausality in multivariate GARCH processes. Journal of Time Series Analysis, 21(5):535–557, September 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chen:2001:FCA

[CL01] Rong Chen and Lon-Mu Liu. Functional coefficient autoregressive models: estimation and tests of hypotheses. Journal of Time Series Analysis, 22(2):151–173, March 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cappuccio:2006:LAD

[CL06] Nunzio Cappuccio and Diego Lubian. Local asymptotic distributions of stationarity tests. Journal of Time Series Analysis, 27(3): 323–345, May 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cleur:2001:MLE

[Cle01] Eugene M. Cleur. Maximum likelihood estimates of a class of onedimensional stochastic differential equation models from discrete data. Journal of Time Series Analysis, 22(5):505–515, September 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cline:2007:ELE

[Cli07] Daren B. H. Cline. Evaluating the Lyapounov exponent and existence of moments for threshold AR–ARCH models. Journal of Time Series Analysis, 28(2):241–260, March 2007. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chen:2014:NSQ

[CLL14] Min Chen, Dong Li, and Shiqing Ling. Non-stationarity and quasimaximum likelihood estimation on a double autoregressive model. Journal of Time Series Analysis, 35(3):189–202, May 2014. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chan:2017:FMH

[CLY17] Ngai Hang Chan, Ye Lu, and Chun Yip Yau. Factor modelling for high-dimensional time series: inference and model selection. Journal of Time Series Analysis, 38(2):285–307, March 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Camba-Mendez:2005:ERS

[CMK05] Gonzalo Camba-Mendez and George Kapetanios. Estimating the rank of the spectral density matrix. Journal of Time Series Analysis, 26(1):37–48, January 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Coursey:1986:POE

[CN86] Don Coursey and Hans Nyquist. A procedure for obtaining Mestimates in regression models with serially dependent errors. Journal of Time Series Analysis, 7(4):255–267, July 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cardinali:2017:LSW

[CN17] Alessandro Cardinali and Guy P. Nason. Locally stationary wavelet packet processes: Basis selection and model fitting. Journal of Time Series Analysis, 38(2):151–174, March 2017. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cavaliere:2017:CBT

[CNR17] Giuseppe Cavaliere, Heino Bohn Nielsen, and Anders Rahbek. On the consistency of bootstrap testing for a parameter on the boundary of the parameter space. Journal of Time Series Analysis, 38(4): 513–534, July 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Comte:1996:SEL

[Com96] F. Comte. Simulation and estimation of long memory continuous time models. Journal of Time Series Analysis, 17(1):19–36, January 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Comte:2004:KDS

[Com04] Fabienne Comte. Kernel deconvolution of stochastic volatility models. Journal of Time Series Analysis, 25(4):563–582, July 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Corradi:1995:NTI

[Cor95] Valentina Corradi. Nonlinear transformations of integrated time series: A reconsideration. Journal of Time Series Analysis, 16 (6):539–549, November 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cox:1991:LRD

[Cox91] D. R. Cox. Long-range dependence, non-linearity and time irreversibility. Journal of Time Series Analysis, 12(4):329–335, July 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cox:1994:BRD

[Cox94] D. R. Cox. Book review: Developments in Time Series Analysis, T. Subba Rao, Editor. Journal of Time Series Analysis, 15(2): 251–252, March 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chang:2003:SBT

[CP03] Yoosoon Chang and Joon Y. Park. A sieve bootstrap for the test of a unit root. Journal of Time Series Analysis, 24(4):379–400, July 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chang:2016:ISB

[CP16a] Seong Yeon Chang and Pierre Perron. Inference on a structural break in trend with fractionally integrated errors. Journal of Time Series Analysis, 37(4):555–574, July 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Choi:2016:QAC

[CP16b] Seokwoo Jake Choi and Stephen Portnoy. Quantile autoregression for censored data. Journal of Time Series Analysis, 37(5):603– 623, September 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chandler:2017:REP

[CP17] Gabe Chandler and Wolfgang Polonik. Residual empirical processes and weighted sums for time-varying processes with applications to testing for homoscedasticity. Journal of Time Series Analysis, 38(1):72–98, January 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Carnero:2007:EOI

[CPR07] M. Angeles Carnero, Daniel Peña, and Esther Ruiz. Effects of outliers on the identification and estimation of GARCH models.

Journal of Time Series Analysis, 28(4):471–497, July 2007. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cavaliere:2015:RDB

[CPR15] Giuseppe Cavaliere, Dimitris N. Politis, and Anders Rahbek. Recent developments in bootstrap methods for dependent data. Journal of Time Series Analysis, 36(3):269–271, May 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cavaliere:2018:FVB

[CPR18] Giuseppe Cavaliere, Rasmus Søndergaard Pedersen, and Anders Rahbek. The fixed volatility bootstrap for a class of Arch(q) models. Journal of Time Series Analysis, 39(6):920–941, November 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chanda:1990:GLP

[CR90] K. C. Chanda and F. H. Ruymgaart. General linear processes: A property of the empirical process applied to density and mode estimation. Journal of Time Series Analysis, 11(3):185–199, May 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chen:1999:BAG

[CR99] Hui Chen and J. P. Romano. Bootstrap-assisted goodness-of-fit tests in the frequency domain. Journal of Time Series Analysis, 20(6):619–654, November 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Craigmile:2003:SCS

[Cra03] Peter F. Craigmile. Simulating a class of stationary Gaussian processes using the Davies-Harte algorithm, with application to long memory processes. Journal of Time Series Analysis, 24(5):505– 511, September 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cavaliere:2015:BDC

[CRT15] Giuseppe Cavaliere, Anders Rahbek, and A. M. Robert Taylor. Bootstrap determination of the co-integration rank in VAR models with unrestricted deterministic components. Journal of Time Series Analysis, 36(3):272–289, May 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Corradi:1984:NCB

[CS84] C. Corradi and C. Scarani. A note on the computation of the Bayesian decomposition of a time series. Journal of Time Series Analysis, 5(4):205–212, July 1984. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Corradi:1987:ICE

[CS87] Corrado Corradi and Claudia Scarani. Improving the computational efficiency of the Bayesian decomposition of a time series: a comment. Journal of Time Series Analysis, 8(2):131–133, March 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cai:2008:QSE

[CS08] Yuzhi Cai and Julian Stander. Quantile self-exciting threshold autoregressive time series models. Journal of Time Series Analysis, 29(1):186–202, January 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Christodoulakis:2011:SCH

[CS11] George A. Christodoulakis and Stephen E. Satchell. Stability conditions for heteroscedastic factor models with conditionally autoregressive betas. Journal of Time Series Analysis, 32(5):482–497, September 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Carcea:2015:GAF

[CS15] Marcel Carcea and Robert Serfling. A Gini autocovariance function for time series modelling. Journal of Time Series Analysis, 36(6): 817–838, November 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cai:2012:NBA

[CSD12] Yuzhi Cai, Julian Stander, and Neville Davies. A new Bayesian approach to quantile autoregressive time series model estimation and forecasting. Journal of Time Series Analysis, 33(4):684–698, July 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chan:1986:ETA

[CT86] K. S. Chan and H. Tong. On estimating thresholds in autoregressive models. Journal of Time Series Analysis, 7(3):179–190, May 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chan:1987:NED

[CT87a] K. S. Chan and H. Tong. A note on embedding a discrete parameter ARMA model in a continuous parameter ARMA model. Journal of Time Series Analysis, 8(3):277–281, May 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cipra:1987:EMA

[CT87b]  Tomás Cipra and Pavel Tlustý.  Estimation in multiple

autoregressive-moving average models using periodicity. Journal of Time Series Analysis, 8(3):293–300, May 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chan:1992:NTS

[CT92] Ngai Hang Chan and Lanh Tat Tran. Nonparametric tests for serial dependence. Journal of Time Series Analysis, 13(1):19–28, January 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chan:1996:AIN

[CT96] Ngai Hang Chan and Ruey S. Tsay. Asymptotic inference for non-invertible moving-average time series. Journal of Time Series Analysis, 17(1):1–17, January 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Choy:2001:SRM

[CT01] Kokyo Choy and Masanobu Taniguchi. Stochastic regression model with dependent disturbances. Journal of Time Series Analysis, 22(2):175–196, March 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cavaliere:2006:TNC

[CT06a] Giuseppe Cavaliere and A. M. Robert Taylor. Testing the null of cointegration in the presence of variance breaks. Journal of Time Series Analysis, 27(4):613–636, July 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chandra:2006:MDE

[CT06b] S. Ajay Chandra and Masanobu Taniguchi. Minimum α-divergence estimation for ARCH models. Journal of Time Series Analysis, 27(1):19–39, January 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cavaliere:2008:TTU

[CT08] Giuseppe Cavaliere and A. M. Robert Taylor. Time-transformed unit root tests for models with non-stationary volatility. Journal of Time Series Analysis, 29(2):300–330, March 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Coke:2010:REM

[CT10] Geoffrey Coke and Min Tsao. Random effects mixture models for clustering electrical load series. Journal of Time Series Analysis, 31(6):451–464, November 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chambers:2020:DPC

[CT20] Marcus J. Chambers and A. M. Robert Taylor. Deterministic parameter change models in continuous and discrete time. Journal of Time Series Analysis, 41(1):134–145, January 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cubadda:1995:NTS

[Cub95] Gianluca Cubadda. A note on testing for seasonal cointegration using principal components in the frequency domain. Journal of Time Series Analysis, 16(5):499–508, September 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Corander:2006:BAM

[CV06] Jukka Corander and Mattias Villani. A Bayesian approach to modelling graphical vector autoregressions. Journal of Time Series Analysis, 27(1):141–156, January 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cooper:1982:IMT

[CW82] D. M. Cooper and E. F. Wood. Identifying multivariate time series models. Journal of Time Series Analysis, 3(3):153–164, May 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Corradi:1999:STV

[CW99] Valentina Corradi and Halbert White. Specification tests for the variance of a diffusion. Journal of Time Series Analysis, 20(3): 253–270, May 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chen:2000:HFE

[CWD00] Zhao-Guo Chen, Ka Ho Wu, and Rainer Dahlhaus. Hidden frequency estimation with data tapers. Journal of Time Series Analysis, 21(2):113–142, March 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Cheng:1997:ZCR

[CWDL97] Ximing Cheng, Yougui Wu, Jinguan Du, and Huowang Liu. The zero-crossing rate of p th-order autoregressive processes. Journal of Time Series Analysis, 18(4):355–374, July 1997. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chan:2012:NSA

[CZ12] Ngai Hang Chan and Rongmao Zhang. Non-stationary autoregressive processes with infinite variance. Journal of Time Series Analysis, 33(6):916–934, November 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chambers:2019:EMM

[CZ19] Marcus J. Chambers and Peter A. Zadrozny. Econometric modelling with mixed frequency and temporally aggregated data. Journal of Time Series Analysis, 40(6):869–871, November 2019. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dienes:2014:LMP

[DA14] Christopher Dienes and Alexander Aue. On-line monitoring of pollution concentrations with autoregressive moving average time series. Journal of Time Series Analysis, 35(3):239–261, May 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dahlhaus:1983:SAT

[Dah83] Rainer Dahlhaus. Spectral analysis with tapered data. Journal of Time Series Analysis, 4(3):163–175, May 1983. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dahlhaus:1985:ADB

[Dah85] Rainer Dahlhaus. On the asymptotic distribution of Bartlett’s U_{p}- statistic. Journal of Time Series Analysis, 6(4):213–227, July 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Datta:1997:NDE

[Dat97] Somnath Datta. A note on l 1 density estimation for linear processes. Journal of Time Series Analysis, 18(4):375–383, July 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Davidson:1991:CPV

[Dav91] James Davidson. The cointegration properties of vector autoregression models. Journal of Time Series Analysis, 12(1):41–62, January 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dai:1998:STB

[DB98] Yuqing Dai and L. Billard. A space-time bilinear model and its identification. Journal of Time Series Analysis, 19(6):657–679, November 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dai:2003:MLE

[DB03] Yuqing Dai and L. Billard. Maximum likelihood estimation in space time bilinear models. Journal of Time Series Analysis, 24(1): 25–44, January 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Castro:2007:UHP

[dBC07] Tomas del Barrio Castro. Using the HEGY procedure when not all roots are present. Journal of Time Series Analysis, 28(6):910– 922, November 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Castro:2022:CPI

[dBCCO22] Tomás del Barrio Castro, Gianluca Cubadda, and Denise R. Osborn. On cointegration for processes integrated at different frequencies. Journal of Time Series Analysis, 43(3):412–435, May 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Castro:2012:NPT

[dBCO12] Tomás del Barrio Castro and Denise R. Osborn. Non-parametric testing for seasonally and periodically integrated processes. Journal of Time Series Analysis, 33(3):424–437, May 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Castro:2019:TAS

[dBCRT19] Tomás del Barrio Castro, Paulo M. M. Rodrigues, and A. M. Robert Taylor. Temporal aggregation of seasonally near-integrated processes. Journal of Time Series Analysis, 40(6):872–886, November 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Daniels:2001:HAC

[DC01] Michael J. Daniels and Noel Cressie. A hierarchical approach to covariance function estimation for time series. Journal of Time Series Analysis, 22(3):253–266, May 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dai:2023:TCP

[DC23] Shan Dai and Ngai Hang Chan. Testing of constant parameters for semi-parametric functional coefficient models with integrated covariates. Journal of Time Series Analysis, 44(5-6):474– 486, September 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

DeJong:2003:SUI

[DCCL03] Piet De Jong and Singfat Chu-Chun-Lin. Smoothing with an unknown initial condition. Journal of Time Series Analysis, 24(2): 141–148, March 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dehay:2015:BBP

[DD15] Dominique Dehay and Anna E. Dudek. Block bootstrap for Poisson-sampled almost periodic processes. Journal of Time Series Analysis, 36(3):327–351, May 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ducharme:2004:GFT

[DdM04] Gilles R. Ducharme and Pierre Lafaye de Micheaux. Goodness-of-fit tests of normality for the innovations in ARMA models. Journal of Time Series Analysis, 25(3):373–395, May 2004. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Duchesne:2013:DRA

[DdM13] Pierre Duchesne and Pierre Lafaye de Micheaux. Distributions for residual autocovariances in parsimonious periodic vector autoregressive models with applications. Journal of Time Series Analysis, 34(4):496–507, July 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Davis:2021:IIT

[DdRSK21] Richard A. Davis, Thiago do Rêgo Sousa, and Claudia Klüppelberg. Indirect inference for time series using the empirical characteristic function and control variates. Journal of Time Series Analysis, 42(5-6):653–684, September 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

DeJong:1991:SAS

[De 91] Piet De Jong. Stable algorithms for the state space model. Journal of Time Series Analysis, 12(2):143–157, March 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

DeGooijer:1998:TMA

[De 98a] Jan De Gooijer. On threshold moving-average models. Journal of Time Series Analysis, 19(1):1–18, January 1998. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

DeLuna:1998:IAF

[De 98b] Xavier De Luna. An improvement of Akaike’s FPE criterion to reduce its variability. Journal of Time Series Analysis, 19(4):457– 471, July 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

DeGooijer:2001:CVC

[De 01] Jan G. De Gooijer. Cross-validation criteria for SETAR model selection. Journal of Time Series Analysis, 22(3):267–281, May 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Davis:2007:MTS

[DE07] Ginger M. Davis and Katherine B. Ensor. Multivariate time-series analysis with categorical and continuous variables in an lstr model. Journal of Time Series Analysis, 28(6):867–885, November 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Debowski:2011:PHD

[Deb11] Lukasz Debowski. On processes with hyperbolically decaying autocorrelations. Journal of Time Series Analysis, 32(5):580–584, September 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Degerine:1987:MLE

[Deg87] Serge Degerine. Maximum likelihood estimation of autocovariance matrices from replicated short time series. Journal of Time Series Analysis, 8(2):135–146, March 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Delgado:1996:TSI

[Del96] Miguel A. Delgado. Testing serial independence using the sample distribution function. Journal of Time Series Analysis, 17(3):271– 285, May 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Deo:1997:ATC

[Deo97] R. S. Deo. Asymptotic theory for certain regression models with long memory errors. Journal of Time Series Analysis, 18(4):385– 393, July 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Deaton:1980:GDT

[DF80] Michael L. Deaton and Robert V. Foutz. Group delay and the timelag relationship between stochastic processes. Journal of Time Series Analysis, 1(2):111–118, March 1980. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

De:2011:DSB

[DF11] Swarup De and Álvaro E. Faria. Dynamic spatial Bayesian models for radioactivity deposition. Journal of Time Series Analysis, 32 (6):607–617, November 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Davis:2023:CMT

[DFF23] Richard A. Davis, Leon Fernandes, and Konstantinos Fokianos. Clustering multivariate time series using energy distance. Journal of Time Series Analysis, 44(5-6):487–504, September 2023. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Doukhan:2021:MNP

[DFR21] Paul Doukhan, Konstantinos Fokianos, and Joseph Rynkiewicz. Mixtures of nonlinear Poisson autoregressions. Journal of Time Series Analysis, 42(1):107–135, January 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dahlhaus:1998:OSL

[DG98] Rainer Dahlhaus and Liudas Giraitis. On the optimal segment length for parameter estimates for locally stationary time series. Journal of Time Series Analysis, 19(6):629–655, November 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Das:2020:SMD

[DG20] Soumya Das and Marc G. Genton. On the stationary marginal distributions of subclasses of multivariate SETAR processes of order one. Journal of Time Series Analysis, 41(3):406–420, May 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dalla:2006:CEM

[DGH06] Violetta Dalla, Liudas Giraitis, and Javier Hidalgo. Consistent estimation of the memory parameter for nonlinear time series. Journal of Time Series Analysis, 27(2):211–251, March 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Darolles:2006:SLT

[DGJ06] Serge Darolles, Christian Gourieroux, and Joann Jasiak. Structural Laplace transform and compound autoregressive models. Journal of Time Series Analysis, 27(4):477–503, July 2006. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dalla:2014:SWS

[DGK14] Violetta Dalla, Liudas Giraitis, and Hira L. Koul. Studentizing weighted sums of linear processes. Journal of Time Series Analysis, 35(2):151–172, March 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Djogbenou:2015:BIR

[DGP15] Antoine Djogbenou, Sílvia Gonçalves, and Benoit Perron. Bootstrap inference in regressions with estimated factors and serial correlation. Journal of Time Series Analysis, 36(3):481–502, May 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Durham:2019:BIA

[DGPHS19] Garland Durham, John Geweke, Susan Porter-Hudak, and Fallaw Sowell. Bayesian inference for ARFIMA models. Journal of Time Series Analysis, 40(4):388–410, July 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Deo:1998:LTF

[DH98] Rohit S. Deo and Clifford M. Hurvich. Linear trend with fractionally integrated errors. Journal of Time Series Analysis, 19(4): 379–397, July 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dehay:2013:EDF

[DH13] D. Dehay and H. L. Hurd. Empirical determination of the frequencies of an almost periodic time series. Journal of Time Series Analysis, 34(2):262–279, March 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dunsmuir:2017:MEP

[DH17] William Dunsmuir and Jieyi He. Marginal estimation of parameter driven binomial time series models. Journal of Time Series Analysis, 38(1):120–144, January 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Demetrescu:2022:ASE

[DH22] Matei Demetrescu and Mehdi Hosseinkouchack. Autoregressive spectral estimates under ignored changes in the mean. Journal of Time Series Analysis, 43(2):329–340, March 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dovonon:2012:IAL

[DHJ12] Prosper Dovonon, Alastair R. Hall, and Kalidas Jana. Inference about long run canonical correlations. Journal of Time Series Analysis, 33(4):665–683, July 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Demetrescu:2014:IBC

[DHT14] Matei Demetrescu, Christoph Hanck, and Adina I. Tarcolea. IVbased cointegration testing in dependent panels with time-varying variance. Journal of Time Series Analysis, 35(5):393–406, September 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dickinson:1982:SSS

[Dic82] Bradley W. Dickinson. Sufficient statistics for stationary discretetime Gaussian random processes. Journal of Time Series Analysis, 3(3):165–168, May 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dittmann:2000:RBT

[Dit00] Ingolf Dittmann. Residual-based tests for fractional cointegration: a Monte Carlo study. Journal of Time Series Analysis, 21(6):615– 647, November 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dittmann:2004:ECM

[Dit04] Ingolf Dittmann. Error correction models for fractionally cointegrated time series. Journal of Time Series Analysis, 25(1):27–32, January 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

deJong:1986:STS

[dJ86] Piet de Jong. State transition specification in state-space models. Journal of Time Series Analysis, 7(3):213–216, May 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

deJong:1994:SNS

[dJCCL94] Pidt de Jong and Singfat Chu-Chun-Lin. Stationary and nonstationary state space models. Journal of Time Series Analysis, 15(2):151–166, March 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Darroch:1986:SFM

[DJM86] John Darroch, Miloslav Jirina, and John McDonald. The sum of finite moving average processes. Journal of Time Series Analysis, 7(1):21–25, January 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Demetrescu:2013:PUR

[DK13] Matei Demetrescu and Robinson Kruse. The power of unit root tests against nonlinear local alternatives. Journal of Time Series Analysis, 34(1):40–61, January 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Diop:2017:TPC

[DK17] Mamadou Lamine Diop and William Kengne. Testing parameter change in general integer-valued time series. Journal of Time Series Analysis, 38(6):880–894, November 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dette:2011:TNP

[DKV11] Holger Dette, Tatjana Kinsvater, and Mathias Vetter. Testing non-parametric hypotheses for stationary processes by estimating minimal distances. Journal of Time Series Analysis, 32(5):447– 461, September 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Decowski:2015:WBT

[DL15] Jonathan Decowski and Linyuan Li. Wavelet-based tests for comparing two time series with unequal lengths. Journal of Time Series Analysis, 36(2):189–208, March 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Doukhan:2015:DWB

[DLLN15] Paul Doukhan, Gabriel Lang, Anne Leucht, and Michael H. Neumann. Dependent wild bootstrap for the empirical process. Journal of Time Series Analysis, 36(3):290–314, May 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dudek:2014:GBB

[DLPP14] Anna E. Dudek, Jacek Leśkow, Efstathios Paparoditis, and Dimitris N. Politis. A generalized block bootstrap for seasonal time series. Journal of Time Series Analysis, 35(2):89–114, March 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Davis:2008:BDC

[DLRY08] Richard A. Davis, Thomas C. M. Lee, and Gabriel A. Rodriguez-Yam. Break detection for a class of nonlinear time series models. Journal of Time Series Analysis, 29(5):834–867, September 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dehay:1996:RSE

[DM96] D. Dehay and V. Monsan. Random sampling estimation for almost periodically correlated processes. Journal of Time Series Analysis, 17(5):425–445, September 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Didier:2012:SCM

[DMHF12] Gustavo Didier, Scott A. McKinley, David B. Hill, and John Fricks. Statistical challenges in microrheology. Journal of Time Series Analysis, 33(5):724–743, September 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dargahi-Noubary:1995:SMI

[DN95] G. R. Dargahi-Noubary. Stochastic modeling and identification of seismic records based on established deterministic formulations. Journal of Time Series Analysis, 16(2):201–220, March 1995. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dargahi-Noubary:1999:LDG

[DN99] G. R. Dargahi-Noubary. A linear discriminant for Gaussian time series. Journal of Time Series Analysis, 20(2):145–153, March 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dargahi-Noubary:1981:SRD

[DNL81] G. R. Dargahi-Noubary and P. J. Laycock. Spectral ratio discriminants and information theory. Journal of Time Series Analysis, 2 (2):71–86, March 1981. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

DaSilva:2004:DEH

[DO04] Maria Eduarda Da Silva and Vera Lúcia Oliveira. Difference equations for the higher-order moments and cumulants of the INAR(1) model. Journal of Time Series Analysis, 25(3):317–333, May 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dodson:2005:BR

[Dod05] C. T. J. Dodson. Book reviews. Journal of Time Series Analysis, 26(5):780–782, September 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Didier:2010:AWD

[DP10] Gustavo Didier and Vladas Pipiras. Adaptive wavelet decompositions of stationary time series. Journal of Time Series Analysis, 31(3):182–209, May 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Du:2020:BPV

[DP20] Zaichao Du and Pei Pei. Backtesting portfolio value-at-risk with estimated portfolio weights. Journal of Time Series Analysis, 41 (5):605–619, September 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

DiMarzio:2012:NPS

[DPT12] Macro Di Marzio, Agnese Panzera, and Charles C. Taylor. Nonparametric smoothing and prediction for nonlinear circular time series. Journal of Time Series Analysis, 33(4):620–630, July 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dette:2023:DRC

[DQ23] Holger Dette and Pascal Quanz. Detecting relevant changes in the spatiotemporal mean function. Journal of Time Series Analysis, 44(5-6):505–532, September 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dwivedi:2011:TSO

[DR11] Yogesh Dwivedi and Suhasini Subba Rao. A test for second-order stationarity of a time series based on the discrete Fourier transform. Journal of Time Series Analysis, 32(1):68–91, January 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Drouiche:2007:TSF

[Dro07] K. Drouiche. A test for spectrum flatness. Journal of Time Series Analysis, 28(6):793–806, November 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Douc:2021:NSC

[DRS21] Randal Douc, François Roueff, and Tepmony Sim. Necessary and sufficient conditions for the identifiability of observation-driven models. Journal of Time Series Analysis, 42(2):140–160, March 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Das:2021:SMS

[DRY21] Sourav Das, Suhasini Subba Rao, and Junho Yang. Spectral methods for small sample time series: a complete periodogram approach. Journal of Time Series Analysis, 42(5-6):597–621, September 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dunsmuir:1991:SCA

[DS91] William T. M. Dunsmuir and Nancy M. Spencer. Strong consistency and asymptotic normality of /1 estimates of the autoregressive moving-average model. Journal of Time Series Analysis, 12(2):95–104, March 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dette:2004:SCS

[DS04] Holger Dette and Ingrid Spreckelsen. Some comments on specification tests in nonparametric absolutely regular processes. Journal of Time Series Analysis, 25(2):159–172, March 2004. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Davies:1980:AMA

[DSW80] Neville Davies, Trevor Spedding, and William Watson. Autoregressive moving average processes with non-normal residuals. Journal of Time Series Analysis, 1(2):103–109, March 1980. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Duchesne:2005:RPS

[Duc05] Pierre Duchesne. Robust and powerful serial correlation tests with new robust estimates in ARX models. Journal of Time Series Analysis, 26(1):49–81, January 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dufour:1981:RTS

[Duf81] Jean-Marie Dufour. Rank tests for serial dependence. Journal of Time Series Analysis, 2(3):117–128, May 1981. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dunsmuir:1981:EPV

[Dun81] W. Dunsmuir. Estimation of periodically varying means and standard deviations in time series data. Journal of Time Series Analysis, 2(3):129–153, May 1981. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Duong:1984:COA

[Duo84] Quang Phuc Duong. On the choice of the order of autoregressive models: a ranking and selection approach. Journal of Time Series Analysis, 5(3):145–157, May 1984. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Drost:2008:LAN

[DVW08] Feike C. Drost, Ramon Van Den Akker, and Bas J. M. Werker. Local asymptotic normality and efficient estimation for INAR(p) models. Journal of Time Series Analysis, 29(5):783–801, September 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Didier:2017:ADP

[DZ17] Gustavo Didier and Kui Zhang. The asymptotic distribution of the pathwise mean squared displacement in single particle tracking experiments. Journal of Time Series Analysis, 38(3):395–416, May 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Davis:2018:SPE

[DZ18] Richard A. Davis and Jing Zhang. Semi-parametric estimation for non-Gaussian non-minimum phase ARMA models. Journal of Time Series Analysis, 39(3):251–272, May 2018. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Dark:2010:IDM

[DZQ10] Jonathan Dark, Xibin Zhang, and Nan Qu. Influence diagnostics for multivariate GARCH processes. Journal of Time Series Analysis, 31(4):278–291, July 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Eichler:2017:GMM

[EDD17] Michael Eichler, Rainer Dahlhaus, and Johannes Dueck. Graphical modeling for multivariate Hawkes processes with nonparametric link functions. Journal of Time Series Analysis, 38(2):225–242, March 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

ElGhini:2006:ARE

[EF06] Ahmed El Ghini and Christian Francq. Asymptotic relative efficiency of goodness-of-fit tests based on inverse and ordinary autocorrelations. Journal of Time Series Analysis, 27(6):843–855, November 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Elsaied:2014:RFI

[EF14] Hanan Elsaied and Roland Fried. Robust fitting of INARCH models. Journal of Time Series Analysis, 35(6):517–535, November 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Efromovich:2014:ENP

[Efr14] Sam Efromovich. Efficient non-parametric estimation of the spectral density in the presence of missing observations. Journal of Time Series Analysis, 35(5):407–427, September 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Efromovich:2019:TSE

[Efr19] Sam Efromovich. On two-stage estimation of the spectral density with assigned risk in presence of missing data. Journal of Time Series Analysis, 40(2):203–224, March 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Efromovich:2020:MRN

[Efr20] Sam Efromovich. Missing not at random and the nonparametric estimation of the spectral density. Journal of Time Series Analysis, 41(5):652–675, September 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

ElGhourabi:2016:CEV

[EFT16] Mohamed El Ghourabi, Christian Francq, and Fedya Telmoudi. Consistent estimation of the value at risk when the error distribution of the volatility model is misspecified. Journal of Time Series Analysis, 37(1):46–76, January 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Esmaeili:2013:TSE

[EK13] Habib Esmaeili and Claudia Klüppelberg. Two-step estimation of a multi-variate Lévy process. Journal of Time Series Analysis, 34 (6):668–690, November 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Eltinge:1994:CTC

[Elt94] John L. Eltinge. Comparison of time and cross-sectional aggregation under a time series random component model. Journal of Time Series Analysis, 15(2):167–181, March 1994. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Escribano:2002:NEC

[EM02] Alvaro Escribano and Santiago Mira. Nonlinear error correction models. Journal of Time Series Analysis, 23(5):509–522, September 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Elek:2008:LTC

[EM08] Péter Elek and László Márkus. A light-tailed conditionally heteroscedastic model with applications to river flows. Journal of Time Series Analysis, 29(1):14–36, January 2008. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Enciso-Mora:2009:EOS

[EMNR09] Víctor Enciso-Mora, Peter Neal, and T. Subba Rao. Efficient order selection algorithms for integer-valued ARMA processes. Journal of Time Series Analysis, 30(1):1–18, January 2009. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Engel:1984:UAS

[Eng84] E. M. R. A. Engel. A unified approach to the study of sums, products, time-aggregation and other functions of ARMA processes. Journal of Time Series Analysis, 5(3):159–171, May 1984. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Escribano:1994:CCF

[EP94] Alvaro Escribano and Daniel Peña. Cointegration and common factors. Journal of Time Series Analysis, 15(6):577–586, November 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Estrada:2017:EAW

[EP17] Francisco Estrada and Pierre Perron. Extracting and analyzing the warming trend in global and hemispheric temperatures. Journal of Time Series Analysis, 38(5):711–732, September 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ercolani:2008:BR

[Erc08] Joanne S. Ercolani. Book review. Journal of Time Series Analysis, 29(4):738–740, July 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ercolani:2011:APF

[Erc11] Joanne S. Ercolani. On the asymptotic properties of a feasible estimator of the continuous time long memory parameter. Journal of Time Series Analysis, 32(5):512–517, September 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ergemen:2019:PHT

[EV19] Yunus Emre Ergemen and Carlos Velasco. Persistence heterogeneity testing in panels with interactive fixed effects. Journal of Time Series Analysis, 40(4):573–589, July 2019. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Evans:1980:TST

[Eva80] Paul Evans. A time-series test of the natural-rate hypothesis. Journal of Time Series Analysis, 1(2):119–133, March 1980. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fotopoulos:2003:RBD

[FA03] Stergios B. Fotopoulos and Sung K. Ahn. Rank Based Dickey– Fuller Test Statistics. Journal of Time Series Analysis, 24(6):647– 662, November 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fang:2005:EEG

[Fan05] Yue Fang. The effect of the estimation on goodness-of-fit tests in time series models. Journal of Time Series Analysis, 26(4): 527–541, July 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fasso:2000:RAD

[Fas00] Alessandro Fasso. Residual autocorrelation distribution in the validation data set. Journal of Time Series Analysis, 21(2):143–153, March 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Feng:2013:OCR

[FB13] Yuanhua Feng and Jan Beran. Optimal convergence rates in nonparametric regression with fractional time series errors. Journal of Time Series Analysis, 34(1):30–39, January 2013. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Funovits:2021:ISS

[FB21] Bernd Funovits and Alexander Braumann. Identifiability of structural singular vector autoregressive models. Journal of Time Series Analysis, 42(4):431–441, July 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fearnhead:2005:BR

[Fea05] Paul Fearnhead. Book review. Journal of Time Series Analysis, 26(3):487–488, May 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Feigin:2020:CRC

[Fei20] Paul D. Feigin. Correction to: Random Coefficient Autoregressive Processes: a Markov Chain Analysis of Stationarity and Finiteness of Moments by Paul D. Feigin and Richard L. Tweedie J. Time Series Anal., Vol. 6, No. 1 (1985). Journal of Time Series Analysis, 41(6):899–900, November 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic). See [FT85].

Fernandez:1990:ETM

[Fer90] F. Javier Fernández. Estimation and testing of a multivariate exponential smoothing model. Journal of Time Series Analysis, 11(2): 89–105, March 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fokianos:2010:IIP

[FF10] Konstantinos Fokianos and Roland Fried. Interventions in IN-GARCH processes. Journal of Time Series Analysis, 31(3):210– 225, May 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fasen:2013:SEH

[FF13] Vicky Fasen and Florian Fuchs. Spectral estimates for highfrequency sampled continuous-time autoregressive moving average processes. Journal of Time Series Analysis, 34(5):532–551, September 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Figuerola-Ferretti:2015:TME

[FFGM15] Isabel Figuerola-Ferretti, Christopher L. Gilbert, and J. Roderick McCrorie. Testing for mild explosivity and bubbles in LME nonferrous metals prices. Journal of Time Series Analysis, 36(5):763– 782, September 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Francq:2004:LSP

[FG04] Christian Francq and Antony Gautier. Large sample properties of parameter least squares estimates for time-varying ARMA models.

Journal of Time Series Analysis, 25(5):765–783, September 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fasen-Hartmann:2020:RES

[FHK20] Vicky Fasen-Hartmann and Sebastian Kimmig. Robust estimation of stationary continuous-time ARMA models via indirect inference. Journal of Time Series Analysis, 41(5):620–651, September 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Feuerverger:1994:EFI

[FHW94] Andrey Feuerverger, Peter Hall, and Andrew T. A. Wood. Estimation of fractal index and fractal dimension of a Gaussian process by counting the number of level crossings. Journal of Time Series Analysis, 15(6):587–606, November 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Findley:1984:SAA

[Fin84] David F. Findley. On some ambiguities associated with the fitting of ARMA models to time series. Journal of Time Series Analysis, 5 (4):213–225, July 1984. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Findley:1985:UPA

[Fin85] D. F. Findley. On the unbiasedness property of AIC for exact or approximating linear stochastic time series models. Journal of Time Series Analysis, 6(4):229–252, July 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fink:2016:CDM

[Fin16] Holger Fink. Conditional distributions of Mandelbrot–van Ness fractional lévy processes and continuous-time ARMA–GARCHtype models with long memory. Journal of Time Series Analysis, 37(1):30–45, January 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fahrmeir:1987:RMN

[FK87] Ludwig Fahrmeir and Heinz Kaufmann. Regression models for nonstationary categorical time series. Journal of Time Series Analysis, 8(2):147–160, March 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fokianos:1999:SAA

[FK99] Konstantinos Fokianos and Benjamin Kedem. A stochastic approximation algorithm for the adaptive control of time series following generalized linear models. Journal of Time Series Analysis, 20(3): 289–308, May 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fokianos:2004:PLI

[FK04] Konstantinos Fokianos and Benjamin Kedem. Partial likelihood inference for time series following generalized linear models. Journal of Time Series Analysis, 25(2):173–197, March 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fink:2013:BRC

[FK13] Thorsten Fink and Jens-Peter Kreiss. Bootstrap for random coefficient autoregressive models. Journal of Time Series Analysis, 34 (6):646–667, November 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Francke:2010:LFS

[FKD10] Marc K. Francke, Siem Jan Koopman, and Aart F. De Vos. Likelihood functions for state space models with diffuse initial conditions. Journal of Time Series Analysis, 31(6):407–414, November 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Franke:2012:CTS

[FKK12] Jürgen Franke, Claudia Kirch, and Joseph Tadjuidje Kamgaing. Changepoints in times series of counts. Journal of Time Series Analysis, 33(5):757–770, September 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Franke:2002:PNA

[FKMN02] J. Franke, J.-P. Kreiss, E. Mammen, and M. H. Neumann. Properties of the nonparametric autoregressive bootstrap. Journal of Time Series Analysis, 23(5):555–585, September 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Foutz:2000:AFS

[FL00] Robert V. Foutz and Hoonja Lee. Adaptive Fourier series and the analysis of periodicities in time series data. Journal of Time Series Analysis, 21(6):649–662, November 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fong:2004:SRC

[FL04] P. W. Fong and W. K. Li. Some results on cointegration with random coefficients in the error correction form: estimation and testing. Journal of Time Series Analysis, 25(3):419–441, May 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Figueroa-Lopez:2013:NRR

[FLL13] José E. Figueroa-López and Michael Levine. Nonparametric regression with rescaled time series errors. Journal of Time Series Analysis, 34(3):345–361, May 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ferland:2006:IVG

[FLO06] René Ferland, Alain Latour, and Driss Oraichi. Integer-valued GARCH process. Journal of Time Series Analysis, 27(6):923–942, November 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Florens:1985:CDM

[FM85] J.-P. Florens and M. Mouchart. Conditioning in dynamic models. Journal of Time Series Analysis, 6(1):15–34, January 1985. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fernandez-Macho:1996:SML

[FM96] F. Javier Fernández-Macho. Spectral maximum likelihood estimation of a signal-to-noise ratio lying in the vicinity of zero. Journal of Time Series Analysis, 17(5):447–459, September 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Franses:1998:TUR

[FM98] Philip Hans Franses and Michael McAleer. Testing for unit roots and non-linear transformations. Journal of Time Series Analysis, 19(2):147–164, March 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Freeland:2004:ALC

[FM04] R. K. Freeland and B. P. M. McCabe. Analysis of low count time series data by Poisson autoregression. Journal of Time Series Analysis, 25(5):701–722, September 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fay:2002:NFP

[FMS02] Gilles Fay, Eric Moulines, and Philippe Soulier. Nonlinear functionals of the periodogram. Journal of Time Series Analysis, 23 (5):523–553, September 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Flores:1997:GTU

[FN97] Rafael Flores and Alfonso Novales. A general test for univariate seasonality. Journal of Time Series Analysis, 18(1):29–48, January 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fryzlewicz:2008:WFA

[FNV08] Piotr Fryzlewicz, Guy P. Nason, and Rainer Von Sachs. A wavelet-Fisz approach to spectrum estimation. Journal of Time Series Analysis, 29(5):868–880, September 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fokianos:2010:AML

[Fok10] Konstantinos Fokianos. Antedependence models for longitudinal data. Journal of Time Series Analysis, 31(6):494, November 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fokianos:2011:MOC

[Fok11] Konstantinos Fokianos. Modeling ordered choices, a primer. Journal of Time Series Analysis, 32(5):585, September 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fossati:2013:URT

[Fos13] Sebastian Fossati. Unit root testing with stationary covariates and a structural break in the trend function. Journal of Time Series Analysis, 34(3):368–384, May 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fokianos:2012:BAT

[FP12] Konstantinos Fokianos and Vasilis J. Promponas. Biological applications of time series frequency domain clustering. Journal of Time Series Analysis, 33(5):744–756, September 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fragkeskou:2016:IFO

[FP16] Maria Fragkeskou and Efstathios Paparoditis*. Inference for the fourth-order innovation cumulant in linear time series. Journal of Time Series Analysis, 37(2):240–266, March 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fragkeskou:2018:ERV

[FP18] Maria Fragkeskou and Efstathios Paparoditis. Extending the range of validity of the autoregressive (sieve) bootstrap. Journal of Time Series Analysis, 39(3):356–379, May 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fortune:2021:LLT

[FPS21] Timothy Fortune, Magda Peligrad, and Hailin Sang. A local limit theorem for linear random fields. Journal of Time Series Analysis, 42(5-6):696–710, September 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fotopoulos:1983:CPE

[FR83] S. B. Fotopoulos and W. D. Ray. Components of prediction errors for a stationary process with estimated parameters. Journal of Time Series Analysis, 4(1):1–8, January 1983. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Francq:1997:WND

[FR97] Christian Francq and Michel Roussignol. On white noises driven by hidden Markov chains. Journal of Time Series Analysis, 18 (6):553–578, November 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Francq:2007:MPT

[FR07] Christian Francq and Hamdi Raïssi. Multivariate portmanteau test for autoregressive models with uncorrelated but nonindependent errors. Journal of Time Series Analysis, 28(3):454–470, May 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Franke:1984:RPI

[Fra84] Jurgen Franke. On the robust prediction and interpolation of time series in the presence of correlated noise. Journal of Time Series Analysis, 5(4):227–244, July 1984. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Franses:2005:EAS

[Fra05] Philip Hans Franses. The econometric analysis of seasonal time series. Journal of Time Series Analysis, 26(2):319–321, March 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ferreira:1999:VBK

[FRP99] Eva Ferreira and Juan Manuel Rodriguez-Poo. Variable bandwidth kernel estimators of the spectral density. Journal of Time Series Analysis, 20(3):271–287, May 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fermin:2017:RMA

[FRR17] Lisandro Javier Fermin, Ricardo Rios, and Luis Angel Rodriguez. A Robbins–Monro algorithm for non-parametric estimation of NAR process with Markov switching: Consistency. Journal of Time Series Analysis, 38(6):809–837, November 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Francq:2011:APW

[FRS11] Christian Francq, Roch Roy, and Abdessamad Saidi. Asymptotic properties of weighted least squares estimation in weak PARMA models. Journal of Time Series Analysis, 32(6):699–723, November 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Francq:2001:CHD

[FRZ01] Christian Francq, Michel Roussignol, and Jean-Michel Zakoian. Conditional heteroskedasticity driven by hidden Markov chains. Journal of Time Series Analysis, 22(2):197–220, March 2001. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Fruhwirth-Schnatter:1994:DAD

[FS94] Sylvia Frühwirth-Schnatter. Data augmentation and dynamic linear models. Journal of Time Series Analysis, 15(2):183–202, March 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Feigin:1985:RCA

[FT85] Paul D. Feigin and Richard L. Tweedie. Random coefficient autoregressive processes: a Markov chain analysis of stationarity and finiteness of moments. Journal of Time Series Analysis, 6(1):1– 14, January 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic). See correction [Fei20].

Francq:2009:BFG

[FZ09] Christian Francq and Jean-Michel Zakoïan. Bartlett’s formula for a general class of nonlinear processes. Journal of Time Series Analysis, 30(4):449–465, July 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Francq:2023:OEF

[FZ23] Christian Francq and Jean-Michel Zakoïan. Optimal estimating function for weak location-scale dynamic models. Journal of Time Series Analysis, 44(5-6):533–555, September 2023. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gil-Alana:2001:TSC

[GA01] L. A. Gil-Alana. Testing stochastic cycles in macroeconomic time series. Journal of Time Series Analysis, 22(4):411–430, July 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gil-Alana:2004:JTF

[GA04] Luis A. Gil-Alana. A joint test of fractional integration and structural breaks at a known period of time. Journal of Time Series Analysis, 25(5):691–700, September 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gil-Alana:2008:FIS

[GA08] Luis A. Gil-Alana. Fractional integration and structural breaks at unknown periods of time. Journal of Time Series Analysis, 29 (1):163–185, January 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ghanbarzadeh:2016:WCC

[GA16] Mitra Ghanbarzadeh and Mina Aminghafari. A wavelet characterization of continuous-time periodically correlated processes with application to simulation. Journal of Time Series Analysis, 37 (6):741–762, November 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gabr:1988:TOM

[Gab88] M. M. Gabr. On the third-order moment structure and bispectral analysis of some bilinear time series. Journal of Time Series Analysis, 9(1):11–20, January 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gaetan:2000:SAM

[Gae00] Carlo Gaetan. Subset ARMA model identification using genetic algorithms. Journal of Time Series Analysis, 21(5):559–570, September 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gao:2001:PES

[GAHT01] Jiti Gao, Vo Anh, Chris Heyde, and Quang Tieng. Parameter estimation of stochastic processes with long-range dependence and intermittency. Journal of Time Series Analysis, 22(5):517–535, September 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gao:1997:CTW

[Gao97] Hong-Ye Gao. Choice of thresholds for wavelet shrinkage estimate of the spectrum. Journal of Time Series Analysis, 18(3):231–251, May 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gomez:2009:NSS

[GAP09] Víctor Gómez and Félix Aparicio-Pérez. A new state-space methodology to disaggregate multivariate time series. Journal of Time Series Analysis, 30(1):97–124, January 2009. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Guo:1998:SIR

[GB98] Jiin-Huarng Guo and L. Billard. Some inference results for causal autoregressive processes on a plane. Journal of Time Series Analysis, 19(6):681–691, November 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gomez:1999:BND

[GB99] Victor Gomez and Jorg Breitung. The Beveridge–Nelson decomposition: a different perspective with new results. Journal of Time Series Analysis, 20(5):527–535, September 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Godolphin:2006:EIM

[GB06] E. J. Godolphin and S. R. Bane. On the evaluation of the information matrix for multiplicative seasonal time-series models. Journal of Time Series Analysis, 27(2):167–190, March 2006. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gao:2017:EDS

[GBY17] Wei Gao, Wicher Bergsma, and Qiwei Yao. Estimation for dynamic and static panel probit models with large individual effects. Journal of Time Series Analysis, 38(2):266–284, March 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gerlach:1999:DTS

[GCK99] Richard Gerlach, Chris Carter, and Robert Kohn. Diagnostics for time series analysis. Journal of Time Series Analysis, 20(3): 309–330, May 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gamerman:2013:NGF

[GdSF13] Dani Gamerman, Thiago Rezende dos Santos, and Glaura C. Franco. A non-Gaussian family of state-space models with exact marginal likelihood. Journal of Time Series Analysis, 34(6):625– 645, November 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gerstenberger:2018:RWT

[Ger18] Carina Gerstenberger. Robust Wilcoxon-type estimation of change-point location under short-range dependence. Journal of Time Series Analysis, 39(1):90–104, January 2018. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gerstenberger:2021:RDB

[Ger21] Carina Gerstenberger. Robust discrimination between long-range dependence and a change in mean. Journal of Time Series Analysis, 42(1):34–62, January 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gallagher:2015:WPT

[GF15] Colin M. Gallagher and Thomas J. Fisher. On weighted portmanteau tests for time-series goodness-of-fit. Journal of Time Series Analysis, 36(1):67–83, January 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Godolphin:2007:NIM

[GG07] E. J. Godolphin and J. D. Godolphin. A note on the information matrix for multiplicative seasonal autoregressive moving-average models. Journal of Time Series Analysis, 28(5):783–791, September 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gagliardini:2008:DTS

[GG08] P. Gagliardini and C. Gourieroux. Duration time-series models with proportional hazard. Journal of Time Series Analysis, 29(1): 74–124, January 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Grzesiek:2020:MCD

[GGSW20] Aleksandra Grzesiek, Prashant Giri, S. Sundar, and Agnieszka WyLomańska. Measures of cross-dependence for bidimensional periodic AR(1) model with α-stable distribution. Journal of Time Series Analysis, 41(6):785–807, November 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Granger:1991:NTI

[GH91] C. W. J. Granger and Jeff Hallman. Nonlinear transformations of integrated time series. Journal of Time Series Analysis, 12(3): 207–224, May 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ghosh:2003:DBP

[GH03] Malay Ghosh and Jungeun Heo. Default Bayesian priors for regression models with first-order autoregressive residuals. Journal of Time Series Analysis, 24(3):269–282, May 2003. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Garcia-Hiernaux:2011:FLD

[GH11] Alfredo García-Hiernaux. Forecasting linear dynamical systems using subspace methods. Journal of Time Series Analysis, 32(5): 462–468, September 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gotz:2019:GCT

[GH19a] Thomas B. Götz and Alain W. Hecq. Granger causality testing in mixed-frequency VARs with possibly (co)integrated processes. Journal of Time Series Analysis, 40(6):914–935, November 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gupta:2019:OSI

[GH19b] Abhimanyu Gupta and Javier Hidalgo. Order selection and inference with long memory dependent data. Journal of Time Series Analysis, 40(4):425–446, July 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gorecki:2018:TNF

[GHHK18] Tomasz Górecki, Siegfried Hörmann, Lajos Horváth, and Piotr Kokoszka. Testing normality of functional time series. Journal of Time Series Analysis, 39(4):471–487, July 2018. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gilbert:1999:AOT

[Gil99] S. D. Gilbert. A testing for the onset of trend, using wavelets. Journal of Time Series Analysis, 20(5):513–526, September 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Girardin:2007:RES

[Gir07] Valerie Girardin. Relative entropy and spectral constraints: some invariance properties of the ARMA class. Journal of Time Series Analysis, 28(6):844–866, November 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Giurcanu:2017:OET

[Giu17] Mihai C. Giurcanu. Oracle M-estimation for time series models. Journal of Time Series Analysis, 38(3):479–504, May 2017. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Granger:1980:ILM

[GJ80] C. W. J. Granger and Roselyne Joyeux. An introduction to longmemory time series models and fractional differencing. Journal of Time Series Analysis, 1(1):15–29, January 1980. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gourieroux:2001:SSM

[GJ01] Christian Gourieroux and Joann Jasiak. State-space models with finite dimensional dependence. Journal of Time Series Analysis, 22 (6):665–678, November 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gourieroux:2002:NAA

[GJ02] Christian Gouriéroux and Joann Jasiak. Nonlinear autocorrelograms: an application to inter-trade durations. Journal of Time Series Analysis, 23(2):127–154, March 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Granger:2006:DMO

[GJ06] Clive W. J. Granger and Yongil Jeon. Dynamics of model overfitting measured in terms of autoregressive roots. Journal of Time Series Analysis, 27(3):347–365, May 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gourieroux:2016:FPS

[GJ16] Christian Gourieroux and Joann Jasiak. Filtering, prediction and simulation methods for noncausal processes. Journal of Time Series Analysis, 37(3):405–430, May 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gourieroux:2023:DDI

[GJ23] Christian Gourieroux and Joann Jasiak. Dynamic deconvolution and identification of independent autoregressive sources. Journal of Time Series Analysis, 44(2):151–180, March 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gosmann:2021:NAO

[GKD21] Josua Gösmann, Tobias Kley, and Holger Dette. A new approach for open-end sequential change point monitoring. Journal of Time Series Analysis, 42(1):63–84, January 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gong:2011:PRA

[GKL11] Zhiyun Gong, Peter Kiessler, and Robert Lund. A predictionresidual approach for identifying rare events in periodic time series. Journal of Time Series Analysis, 32(4):407–419, July 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Giraitis:2018:IMH

[GKY18] Liudas Giraitis, George Kapetanios, and Tony Yates. Inference on multivariate heteroscedastic time varying random coefficient models. Journal of Time Series Analysis, 39(2):129–149, March 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Granger:1994:UMI

[GL94] Clive Granger and Jin-Lung Lin. Using the mutual information coefficient to identify lags in nonlinear models. Journal of Time Series Analysis, 15(4):371–384, July 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gonzalo:1996:RPT

[GL96] Jesus Gonzalo and Tae-Hwy Lee. Relative power of t type tests for stationary and unit root processes. Journal of Time Series Analysis, 17(1):37–47, January 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gourieroux:2019:NBA

[GL19] Christian Gouriéroux and Yang Lu. Negative binomial autoregressive process with stochastic intensity. Journal of Time Series Analysis, 40(2):225–247, March 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gong:2020:TSN

[GL20] Huan Gong and Dong Li. On the three-step non-Gaussian quasimaximum likelihood estimation of heavy-tailed double autoregressive models. Journal of Time Series Analysis, 41(6):883–891, November 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gaitan:2021:EPA

[GL21] Rodrigo Saul Gaitan and Keh-Shin Lii. On the estimation of periodicity or almost periodicity in inhomogeneous gamma pointprocess data. Journal of Time Series Analysis, 42(5-6):711–736, September 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gassiat:2006:ESE

[GLL06] Elisabeth Gassiat and Céline Lévy-Leduc. Efficient semiparametric estimation of the periods in a superposition of periodic functions with unknown shape. Journal of Time Series Analysis, 27(6):877– 910, November 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Goncalves:2016:DEP

[GLML16] Esmeralda Gonçalves, Joana Leite, and NazarÉ Mendes-Lopes. On the distribution estimation of power threshold GARCH processes.

Journal of Time Series Analysis, 37(5):579–602, September 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gregory:2015:SBB

[GLN15] Karl B. Gregory, Soumendra N. Lahiri, and Daniel J. Nordman. A smooth block bootstrap for statistical functionals and time series. Journal of Time Series Analysis, 36(3):442–461, May 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gong:2010:ELI

[GLP10] Yun Gong, Zhouping Li, and Liang Peng. Empirical likelihood intervals for conditional value-at-risk in ARCH/GARCH models. Journal of Time Series Analysis, 31(2):65–75, March 2010. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gingras:1985:SDE

[GM85] D. F. Gingras and E. Masry. Spectral density estimation from nonlinearly observed data. Journal of Time Series Analysis, 6(2): 63–80, March 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ghysels:2015:TCT

[GM15] Eric Ghysels and J. Isaac Miller. Testing for cointegration with temporally aggregated and mixed-frequency time series. Journal of Time Series Analysis, 36(6):797–816, November 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Giraitis:2023:EUS

[GM23] Liudas Giraitis and Fulvia Marotta. Estimation on unevenly spaced time series. Journal of Time Series Analysis, 44(5-6):556– 577, September 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Goncalves:2015:IDD

[GMLS15] E. Gonçalves, N. Mendes-Lopes, and F. Silva. Infinitely divisible distributions in integer-valued GARCH models. Journal of Time Series Analysis, 36(4):503–527, July 2015. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Grose:2015:BCP

[GMP15] Simone D. Grose, Gael M. Martin, and Donald S. Poskitt. Bias correction of persistence measures in fractionally integrated models. Journal of Time Series Analysis, 36(5):721–740, September 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Granger:2004:DMP

[GMR04] C. W. Granger, E. Maasoumi, and J. Racine. A dependence metric for possibly nonlinear processes. Journal of Time Series Analysis, 25(5):649–669, September 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Galvao:2011:TQA

[GMRO11] Antonio F. Galvao, Jr., Gabriel Montes-Rojas, and Jose Olmo. Threshold quantile autoregressive models. Journal of Time Series Analysis, 32(3):253–267, May 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gomez:2007:WKF

[Góm07] Víctor Gómez. Wiener-Kolmogorov filtering and smoothing for multivariate series with state-space structure. Journal of Time Series Analysis, 28(3):361–385, May 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gorrostieta:2012:EDB

[GOP^{+}12] Cristina Gorrostieta, Hernando Ombao, Raquel Prado, Shaun Patel, and Emad Eskandar. Exploring dependence between brain signals in a monkey during learning. Journal of Time Series Analysis, 33(5):771–778, September 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gordon:1981:ALS

[Gor81] William B. Gordon. Accuracy of linear spectral estimates of bandlimited signals. Journal of Time Series Analysis, 2(3):173–184, May 1981. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gorgi:2018:IVA

[Gor18] Paolo Gorgi. Integer-valued autoregressive models with survival probability driven by a stochastic recurrence equation. Journal of Time Series Analysis, 39(2):150–171, March 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gorrostieta:2019:TDD

[GOV19] Cristina Gorrostieta, Hernando Ombao, and Rainer Von Sachs. Time-dependent dual-frequency coherence in multivariate nonstationary time series. Journal of Time Series Analysis, 40(1): 3–22, January 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gonzalo:2002:LLE

[GP02] Jesús Gonzalo and Jean-Yves Pitarakis. Lag length estimation in large dimensional systems. Journal of Time Series Analysis, 23(4): 401–423, July 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Giraitis:2006:ULT

[GP06] Liudas Giraitis and Peter C. B. Phillips. Uniform limit theory for stationary autoregression. Journal of Time Series Analysis, 27(1): 51–60, January 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Geweke:1983:EAL

[GPH83] John Geweke and Susan Porter-Hudak. The estimation and application of long memory time series models. Journal of Time Series Analysis, 4(4):221–238, July 1983. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gonen:2000:LDU

[GPRV00] Mithat Gonen, Madan L. Puri, Frits H. Ruymgaart, and Martien C. A. Van Zuijlen. The limiting density of unit root test statistics: a unifying technique. Journal of Time Series Analysis, 21(3):249– 260, May 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Grant:2017:PSD

[GQ17] Andrew J. Grant and Barry G. Quinn. Parametric spectral discrimination. Journal of Time Series Analysis, 38(6):838–864, November 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gabr:1981:EPS

[GR81] M. M. Gabr and T. Subba Rao. The estimation and prediction of subset bilinear time series models with applications. Journal of Time Series Analysis, 2(3):155–171, May 1981. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Granger:1982:ATS

[Gra82] C. W. J. Granger. Acronyms in time series analysis (ATSA). Journal of Time Series Analysis, 3(2):103–107, March 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Granger:1988:MGT

[Gra88] C. W. J. Granger. Models that generate trends. Journal of Time Series Analysis, 9(4):329–343, July 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Grahn:1995:CLS

[Gra95] T. Grahn. A conditional least squares approach to bilinear time series estimation. Journal of Time Series Analysis, 16(5):509–529, September 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Grillenzoni:1991:IRE

[Gri91] Carlo Grillenzoni. Iterative and recursive estimation of transfer functions. Journal of Time Series Analysis, 12(2):105–127, March 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Giraitis:1997:ROS

[GRS97] Liudas Giraitis, Peter M. Robinson, and Alexander Samarov. Rate optimal semiparametric estimation of the memory parameter of the Gaussian time series with long-range dependence. Journal of Time Series Analysis, 18(1):49–60, January 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Georgiev:2017:URT

[GRT17] Iliyan Georgiev, Paulo M. M. Rodrigues, and A. M. Robert Taylor. Unit root tests and heavy-tailed innovations. Journal of Time Series Analysis, 38(5):733–768, September 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Guinness:2013:TAI

[GS13] Joseph Guinness and Michael L. Stein. Transformation to approximate independence for locally stationary Gaussian processes. Journal of Time Series Analysis, 34(5):574–590, September 2013. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Girardin:2020:FGB

[GS20] Valerie Girardin and Rachid Senoussi. Filling the gap between continuous and discrete time dynamics of autoregressive processes. Journal of Time Series Analysis, 41(4):590–602, July 2020. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gong:2021:NEF

[GS21] Chen Gong and David S. Stoffer. A note on efficient fitting of stochastic volatility models. Journal of Time Series Analysis, 42 (2):186–200, March 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Guillaumin:2017:ANS

[GSO^{+}17] Arthur P. Guillaumin, Adam M. Sykulski, Sofia C. Olhede, Jeffrey J. Early, and Jonathan M. Lilly. Analysis of non-stationary modulated time series with applications to oceanographic surface flow measurements. Journal of Time Series Analysis, 38(5):668– 710, September 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Grublyte:2017:QQA

[GSS17] Ieva Grublyte, Donatas Surgailis, and Andrius Skarnulis. QMLE for quadratic ARCH model with long memory. Journal of Time Series Analysis, 38(4):535–551, July 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Grzesiek:2020:STD

[GSTW20] Aleksandra Grzesiek, Grzegorz Sikora, Marek Teuerle, and Agnieszka Wyłomańska. Spatio-temporal dependence measures for bivariate AR(1) models with α-stable noise. Journal of Time Series Analysis, 41(3):454–475, May 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Geweke:1993:BTA

[GT93] John Geweke and Nobuhiko Terui. Bayesian threshold autoregressive models for nonlinear time series. Journal of Time Series Analysis, 14(5):441–454, September 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gao:2019:SFA

[GT19a] Zhaoxing Gao and Ruey S. Tsay. A structural-factor approach to modeling high-dimensional time series and space-time data. Journal of Time Series Analysis, 40(3):343–362, May 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Goto:2019:RZC

[GT19b] Yuichi Goto and Masanobu Taniguchi. Robustness of zero crossing estimator. Journal of Time Series Analysis, 40(5):815–830, September 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Guegan:1987:DRB

[Gue87] Dominique. Guegan. Different representations for bilinear models. Journal of Time Series Analysis, 8(4):389–408, July 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Guo:2003:DSS

[Guo03] Wensheng Guo. Dynamic state-space models. Journal of Time Series Analysis, 24(2):149–158, March 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Giummole:2010:IPL

[GV10] Federica Giummolè and Paolo Vidoni. Improved prediction limits for a general class of Gaussian models. Journal of Time Series Analysis, 31(6):483–493, November 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gray:1988:CNP

[GZ88] H. L. Gray and Nien Fan Zhang. On a class of nonstationary processes. Journal of Time Series Analysis, 9(2):133–154, March 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gourieroux:2015:UMA

[GZ15] Christian Gouriéroux and Jean-Michel Zakoïan. On uniqueness of moving average representations of heavy-tailed stationary processes. Journal of Time Series Analysis, 36(6):876–887, November 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Guyton:1986:RPP

[GZF86] Deborah A. Guyton, Nien-Fan Zhang, and Robert V. Foutz. A random parameter process for modeling and forecasting time series. Journal of Time Series Analysis, 7(2):105–115, March 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gray:1989:GFP

[GZW89] Henry L. Gray, Nien-Fan Zhang, and Wayne A. Woodward. On generalized fractional processes. Journal of Time Series Analysis, 10(3):233–257, May 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Gray:1994:GFP

[GZW94] Henry L. Gray, Nien-Fan Zhang, and Wayne A. Woodward. On generalized fractional processes — a correction. Journal of Time Series Analysis, 15(5):561–562, September 1994. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Huang:1993:ENS

[HA93] D. Huang and V. V. Anh. Estimation of the non-stationary factor in aruma models. Journal of Time Series Analysis, 14(1):27–46, January 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hu:2021:IVA

[HA21] Xiaofei Hu and Beth Andrews. Integer-valued asymmetric GARCH modeling. Journal of Time Series Analysis, 42(5-6):737–751, September 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Haddad:1995:RPI

[Had95] John N. Haddad. The recursive property of the inverse of the covariance matrix of a moving-average process of general order. Journal of Time Series Analysis, 16(6):551–554, November 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Haddad:2004:CFC

[Had04] John N. Haddad. On the closed form of the covariance matrix and its inverse of the causal ARMA process. Journal of Time Series Analysis, 25(4):443–448, July 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hall:1992:JHT

[Hal92] Alastair Hall. Joint hypothesis tests for a random walk based on instrumental variable estimators. Journal of Time Series Analysis, 13(1):29–45, January 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hall:1994:OIM

[Hal94a] Alastair Hall. Order identification in misspecified autoregressive time series models. Journal of Time Series Analysis, 15(3):279– 283, May 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hallin:1994:PNA

[Hal94b] Marc Hallin. On the Pitman non-admissibility of correlogrambased methods. Journal of Time Series Analysis, 15(6):607–611, November 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hall:1995:RAU

[Hal95] Alastair Hall. Residual autocovariances and unit root tests based on instrumental variable estimators from time series regression models. Journal of Time Series Analysis, 16(6):555–569, November 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hall:2012:BRO

[Hal12] Alastair R. Hall. Book review: The Oxford Handbook of Economic Forecasts. Journal of Time Series Analysis, 33(3):530–531, May 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hall:2013:ETS

[Hal13] Alastair R. Hall. Economic time series: modeling and seasonality. Journal of Time Series Analysis, 34(2):282–283, March 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hall:2014:BRD

[Hal14] Alastair R. Hall. Book review: Dynamic Models for Volatility and Heavy Tails: with Applications to Financial and Economic Time Series, by A. C. Harvey. Published by Cambridge University Press, 2013 New York, Usa. Total Number of Pages: 261. Price: $36.99. ISBN: 978-1-107-63002-4. Journal of Time Series Analysis, 35(2): 187–188, March 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Harvey:1981:FSP

[Har81] A. C. Harvey. Finite sample prediction and overdifferencing. Journal of Time Series Analysis, 2(4):221–232, July 1981. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hasan:1982:NTS

[Has82] T. Hasan. Nonlinear time series regression for a class of amplitude modulated consinusoids. Journal of Time Series Analysis, 3(2): 109–122, March 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hassler:1993:PR

[Has93a] Uwe Hassler. The periodogram regression. Journal of Time Series Analysis, 14(5):549, September 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hassler:1993:RSE

[Has93b] Uwe Hassler. Regression of spectral estimators with fractionally integrated time series. Journal of Time Series Analysis, 14(4): 369–380, July 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hassler:1994:MSL

[Has94] Uwe Hassler. (mis)specification of long memory in seasonal time series. Journal of Time Series Analysis, 15(1):19–30, January 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hasseler:2000:SRL

[Has00] Uwe Hasseler. Simple regressions with linear time trends. Journal of Time Series Analysis, 21(1):27–32, January 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hassler:2001:ELT

[Has01] Uwe Hassler. The effect of linear time trends on the KPSS test for cointegration. Journal of Time Series Analysis, 22(3):283–292, May 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hassler:2013:ETA

[Has13] Uwe Hassler. Effect of temporal aggregation on multiple time series in the frequency domain. Journal of Time Series Analysis, 34(5):

562–573, September 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hurvich:1990:CVC

[HB90] Clifford M. Hurvich and Kaizô I. Beltrato. Cross-validatory choice of a spectrum estimate and its connections with AIC. Journal of Time Series Analysis, 11(2):121–137, March 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hurvich:1993:ALF

[HB93] Clifford M. Hurvich and Kaizo I. Beltrao. Asymptotics for the low-frequency ordinates of the periodogram of a long-memory time series. Journal of Time Series Analysis, 14(5):455–472, September 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hurvich:1994:APA

[HB94a] Clifford M. Hurvich and Kaizo I. Beltrao. Acknowledgement of priority for “asymptotics for the low-frequency ordinates of the periodogram of a long-memory time series”. Journal of Time Series Analysis, 15(1):64, January 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hurvich:1994:ASE

[HB94b] Clifford M. Hurvich and Kaizo I. Beltrao. Automatic semiparametric estimation of the memory parameter of a long-memory time series. Journal of Time Series Analysis, 15(3):285–302, May 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hurvich:2001:BSE

[HB01] Clifford M. Hurvich and Julia Brodsky. Broadband semiparametric estimation of the memory parameter of a long-memory time series using fractional exponential models. Journal of Time Series Analysis, 22(2):221–249, March 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hwang:2005:ERC

[HB05] S. Y. Hwang and I. V. Basawa. Explosive random-coefficient AR(1) processes and related asymptotics for least-squares estimation. Journal of Time Series Analysis, 26(6):807–824, November 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hurvich:2000:ETP

[HC00] Clifford M. Hurvich and Willa W. Chen. An efficient taper for potentially overdifferenced long-memory time series. Journal of Time Series Analysis, 21(2):155–180, March 2000. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hu:2004:PBM

[HC04] Yu-Pin Hu and Rouh-Jane Chou. On the Peña–Box model. Journal of Time Series Analysis, 25(6):811–830, November 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Huang:2020:WFT

[HC20] Zhelin Huang and Ngai Hang Chan. Walsh Fourier transform of locally stationary time series. Journal of Time Series Analysis, 41(2):312–340, March 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hasegawa:2000:BUR

[HCH00] Hikaru Hasegawa, Anoop Chaturvedi, and Tran Van Hoa. Bayesian unit root test in nonnormal AR(1) model. Journal of Time Series Analysis, 21(3):261–280, May 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hjellvik:2004:NET

[HCT04] Vidar Hjellvik, Rong Chen, and Dag Tjøstheim. Nonparametric estimation and testing in panels of intercorrelated time series. Journal of Time Series Analysis, 25(6):831–872, November 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Heyde:1996:RST

[HD96] C. C. Heyde and W. Dai. On the robustness to small trends of estimation based on the smoothed periodogram. Journal of Time Series Analysis, 17(2):141–150, March 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hurvich:1999:PSN

[HD99] Clifford M. Hurvich and Rohit S. Deo. Plug-in selection of the number of frequencies in regression estimates of the memory parameter of a long-memory time series. Journal of Time Series Analysis, 20(3):331–341, May 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hurvich:1998:MSE

[HDB98] Clifford M. Hurvich, Rohit Deo, and Julia Brodsky. The mean squared error of Geweke and Porter-Hudak’s estimator of the memory parameter of a long-memory time series. Journal of Time Series Analysis, 19(1):19–46, January 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Henry:2001:APS

[Hen01a] Marc Henry. Averaged periodogram spectral estimation with longmemory conditional heteroscedasticity. Journal of Time Series Analysis, 22(4):431–459, July 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Henry:2001:RAB

[Hen01b] Marc Henry. Robust automatic bandwidth for long memory. Journal of Time Series Analysis, 22(3):293–316, May 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hurd:1991:GMD

[HG91] Harry L. Hurd and Neil L. Gerr. Graphical methods for determining the presence of periodic correlation. Journal of Time Series Analysis, 12(4):337–350, July 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Herzberg:1981:IAP

[HH81] Agnes M. Herzberg and J. S. Hickie. An investigation of Andrews’ plots to show time variation of model parameters. Journal of Time Series Analysis, 2(4):233–262, July 1981. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hall:1993:PEW

[HH93a] Peter Hall and Jeffrey D. Hart. On the probability of error when using a general Akaike-type criterion to estimate autoregression order. Journal of Time Series Analysis, 14(4):347–368, July 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hannan:1993:LFE

[HH93b] E. J. Hannan and D. Huang. On-line frequency estimation. Journal of Time Series Analysis, 14(2):147–161, March 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hodges:1993:CME

[HH93c] P. E. Hodges and D. F. Hale. A computational method for estimating densities of non-Gaussian nonstationary univariate time series. Journal of Time Series Analysis, 14(2):163–178, March 1993. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ho:2005:PTR

[HH05] Hwai-Chung Ho and Nan-Jung Hsu. Polynomial trend regression with long-memory errors. Journal of Time Series Analysis, 26(3): 323–354, May 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Horvath:2012:CPD

[HH12] Lajos Horváth and Marie Husková. Change-point detection in panel data. Journal of Time Series Analysis, 33(4):631–648, July 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hosseinkouchack:2016:PUR

[HH16] Mehdi Hosseinkouchack and Uwe Hassler. Powerful unit root tests free of nuisance parameters. Journal of Time Series Analysis, 37 (4):533–554, July 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hassler:2020:HWP

[HH20] Uwe Hassler and Mehdi Hosseinkouchack. Harmonically weighted processes. Journal of Time Series Analysis, 41(1):41–66, January 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hsiao:2018:IEF

[HHI18] Wei-Cheng Hsiao, Hao-Yun Huang, and Ching-Kang Ing. Interval estimation for a first-order positive autoregressive process. Journal of Time Series Analysis, 39(3):447–467, May 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Haggan:1984:SAS

[HHP84] V. Haggan, S. M. Heravi, and M. B. Priestley. A study of the application of state-dependent models in non-linear time series analysis. Journal of Time Series Analysis, 5(2):69–102, March 1984. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hualde:2015:SFA

[HI15] Javier Hualde and Fabrizio Iacone. Small-b and fixed-b asymptotics for weighted covariance estimation in fractional cointegration. Journal of Time Series Analysis, 36(4):528–540, July 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hualde:2019:FBI

[HI19] Javier Hualde and Fabrizio Iacone. Fixed bandwidth inference for fractional cointegration. Journal of Time Series Analysis, 40(4): 544–572, July 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hidalgo:1992:ASE

[Hid92] F. Javier Hidalgo. Adaptive semiparametric estimation in the presence of autocorrelation of unknown form. Journal of Time Series Analysis, 13(1):47–78, January 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hidalgo:1997:NPE

[Hid97] Javier Hidalgo. Non-parametric estimation with strongly dependent multivariate time series. Journal of Time Series Analysis, 18(2):95–122, March 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hidalgo:2007:NTW

[Hid07] Javier Hidalgo. A nonparametric test for weak dependence against strong cycles and its bootstrap analogue. Journal of Time Series Analysis, 28(3):307–349, May 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hill:2013:LTT

[Hil13] Jonathan B. Hill. Least tail-trimmed squares for infinite variance autoregressions. Journal of Time Series Analysis, 34(2):168–186, March 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hinich:1982:TGL

[Hin82] Melvin J. Hinich. Testing for Gaussianity and linearity of a stationary time series. Journal of Time Series Analysis, 3(3):169–176, May 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hallin:1987:LQS

[HIP87] Marc. Hallin, Jean-François Ingenbleek, and Madan L. Puri. Linear and quadratic serial rank tests for randomness against serial dependence. Journal of Time Series Analysis, 8(4):409–424, July 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hirukawa:2006:MNP

[Hir06] Masayuki Hirukawa. A modified nonparametric prewhitened covariance estimator. Journal of Time Series Analysis, 27(3):441– 476, May 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hannan:1986:RAM

[HK86] E. J. Hannan and L. Kavalieris. Regression, autoregression models. Journal of Time Series Analysis, 7(1):27–49, January 1986. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

He:1990:ZCR

[HK90] Shuyuan He and Benjamin Kedem. The zero-crossing rate of autoregressive processes and its link to unit roots. Journal of Time Series Analysis, 11(3):201–213, May 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Huskova:2008:BCI

[HK08] Marie Husková and Claudia Kirch. Bootstrapping confidence intervals for the change-point of time series. Journal of Time Series Analysis, 29(6):947–972, November 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Harris:2014:PAT

[HK14] David Harris and Hsein Kew. Portmanteau autocorrelation tests under q-dependence and heteroskedasticity. Journal of Time Series Analysis, 35(3):203–217, May 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hafner:2017:MFA

[HK17] Franziska Häfner and Claudia Kirch. Moving Fourier analysis for locally stationary processes with the bootstrap in view. Journal of Time Series Analysis, 38(6):895–922, November 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Holleland:2020:SST

[HK20] Sondre Hølleland and Hans Arnfinn Karlsen. A stationary spatiotemporal GARCH model. Journal of Time Series Analysis, 41(2): 177–209, March 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Heiner:2022:ADM

[HK22] Matthew Heiner and Athanasios Kottas. Autoregressive density modeling with the Gaussian process mixture transition distribution. Journal of Time Series Analysis, 43(2):157–177, March 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hormann:2015:EFL

[HKK15] Siegfried Hörmann, Lukasz Kidziński, and Piotr Kokoszka. Estimation in functional lagged regression. Journal of Time Series Analysis, 36(4):541–561, July 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hallin:2023:SIJ

[HKK23] Marc Hallin, Yoshihide Kakizawa, and Hira Koul. Special issue of the Journal of Time Series Analysis in honor of Professor Masanobu Taniguchi. Journal of Time Series Analysis, 44(5-6): 440–441, September 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Holmes:2024:MPO

[HKV24] Mark Holmes, Ivan Kojadinovic, and Alex Verhoijsen. Multipurpose open-end monitoring procedures for multivariate observations based on the empirical distribution function. Journal of Time Series Analysis, 45(1):27–56, January 2024. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Horvath:2022:IFF

[HKVW22] Lajos Horváth, Piotr Kokoszka, Jeremy VanderDoes, and Shixuan Wang. Inference in functional factor models with applications to yield curves. Journal of Time Series Analysis, 43(6):872–894, November 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Haldrup:2002:RUR

[HL02] Niels Haldrup and Peter Lildholdt. On the robustness of unit root tests in the presence of double unit roots. Journal of Time Series Analysis, 23(2):155–171, March 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Harvey:2006:PUR

[HL06] David I. Harvey and Stephen J. Leybourne. Power of a unit-root test and the initial condition. Journal of Time Series Analysis, 27 (5):739–752, September 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Herwartz:2011:GLS

[HL11a] Helmut Herwartz and Helmut Lütkepohl. Generalized least squares estimation for cointegration parameters under conditional heteroskedasticity. Journal of Time Series Analysis, 32(3):281–291, May 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hong:2011:DMA

[HL11b] Yongmiao Hong and Yoon-Jin Lee. Detecting misspecifications in autoregressive conditional duration models and non-negative timeseries processes. Journal of Time Series Analysis, 32(1):1–32, January 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Harvey:2017:VMG

[HL17] Andrew Harvey and Rutger-Jan Lange. Volatility modeling with a generalized t distribution. Journal of Time Series Analysis, 38(2): 175–190, March 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Harvey:2018:MIB

[HL18] Andrew Harvey and Rutger-Jan Lange. Modeling the interactions between volatility and returns using EGARCH-M. Journal of Time Series Analysis, 39(6):909–919, November 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Holst:1994:RES

[HLHT94] Ulla Holst, Georg Lindgren, Jan Holst, and Mikael Thuvesholmen. Recursive estimation in switching autoregressions with a Markov regime. Journal of Time Series Analysis, 15(5):489–506, September 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hurn:2003:ESM

[HLM03] A. S. Hurn, K. A. Lindsay, and V. L. Martin. On the efficacy of simulated maximum likelihood for estimating the parameters of stochastic differential Equations*. Journal of Time Series Analysis, 24(1):45–63, January 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Harvey:2010:IIC

[HLT10] David I. Harvey, Stephen J. Leybourne, and A. M. Robert Taylor. The impact of the initial condition on robust tests for a linear trend. Journal of Time Series Analysis, 31(4):292–302, July 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Harvey:2010:TND

[HLX10] David I. Harvey, Stephen J. Leybourne, and Lisa Xiao. Testing for nonlinear deterministic components when the order of integration is unknown. Journal of Time Series Analysis, 31(5):379–391, September 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Harvey:2023:EVF

[HLZ23] David I. Harvey, Stephen J. Leybourne, and Yang Zu. Estimation of the variance function in structural break autoregressive models with non-stationary and explosive segments. Journal of Time Series Analysis, 44(2):181–205, March 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Harvey:2003:NBH

[HM03] David I. Harvey and Terence C. Mills. A note on Busetti–Harvey tests for stationarity in series with structural breaks. Journal of Time Series Analysis, 24(2):159–164, March 2003. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Han:2013:TPC

[HM13] Lu Han and Brendan McCabe. Testing for parameter constancy in non-Gaussian time series. Journal of Time Series Analysis, 34(1): 17–29, January 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Heinen:2013:WIE

[HMS13] Florian Heinen, Stefanie Michael, and Philipp Sibbertsen. Weak identification in the ESTAR model and a new model. Journal of Time Series Analysis, 34(2):238–261, March 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hassler:2008:FCP

[HMV08] Uwe Hassler, Francesc Marmol, and Carlos Velasco. Fractional cointegration in the presence of linear trends. Journal of Time Series Analysis, 29(6):1088–1103, November 2008. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Herwartz:2019:HRU

[HMW19] Helmut Herwartz, Simone Maxand, and Yabibal M. Walle. Heteroskedasticity-robust unit root testing for trending panels. Journal of Time Series Analysis, 40(5):649–664, September 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hopwood:1980:TSA

[HN80] W. S. Hopwood and P. Newbold. Time series analysis in accounting: a survey and analysis of recent issues. Journal of Time Series Analysis, 1(2):135–144, March 1980. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hotta:1993:EAP

[HN93] L. K. Hotta and J. Cardosc Neto. The effect of aggregation on prediction in autoregressive integrated moving-average models. Journal of Time Series Analysis, 14(3):261–269, May 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hormann:2021:PSV

[HN21] Siegfried Hörmann and Gilles Nisol. Prediction of singular VARs and an application to generalized dynamic factor models. Journal of Time Series Analysis, 42(3):295–313, May 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hallin:2023:FMH

[HNT23] Marc Hallin, Gilles Nisol, and Shahin Tavakoli. Factor models for high-dimensional functional time series I: Representation results. Journal of Time Series Analysis, 44(5-6):578–600, September 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Haggan:1984:SSA

[HO84] V. Haggan and O. B. Oyetunji. On the selection of subset autoregressive time series models. Journal of Time Series Analysis, 5(2):

103–113, March 1984. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hognas:1986:CSN

[Hög86] Göran Högnäs. Comparison of some non-linear autoregressive processes. Journal of Time Series Analysis, 7(3):205–211, May 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hoga:2018:DTR

[Hog18] Yannick Hoga. Detecting tail risk differences in multivariate time series. Journal of Time Series Analysis, 39(5):665–689, September 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hoga:2019:ELB

[Hog19] Yannick Hoga. Extending the limits of backtesting via the ‘vanishing p’-approach. Journal of Time Series Analysis, 40(5):858–866, September 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hokstad:1983:MDC

[Hok83] Per Hokstad. A method for diagnostic checking of time series models. Journal of Time Series Analysis, 4(3):177–183, May 1983. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hong:1997:OST

[Hon97] Yongmiao Hong. One-sided testing for conditional heteroskedasticity in time series models. Journal of Time Series Analysis, 18(3): 253–277, May 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hosoya:2001:ETS

[Hos01] Yuzo Hosoya. Elimination of third-series effect and defining partial measures of causality. Journal of Time Series Analysis, 22(5):537– 554, September 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hosoya:2005:FIP

[Hos05] Yuzo Hosoya. Fractional invariance principle. Journal of Time Series Analysis, 26(3):463–486, May 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hall:2015:SBI

[HOS15] Alastair R. Hall, Denise R. Osborn, and Nikolaos Sakkas. Structural break inference using information criteria in models estimated by two-stage least squares. Journal of Time Series Analysis, 36 (5):741–762, September 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hotta:1989:IUC

[Hot89] Luiz Koodi Hotta. Identification of unobserved components models. Journal of Time Series Analysis, 10(3):259–270, May 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hoyos:2020:MFS

[Hoy20] Milena Hoyos. Mixed first- and second-order cointegrated continuous time models with mixed stock and flow data. Journal of Time Series Analysis, 41(2):249–267, March 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hoyos:2024:FOC

[Hoy24] Milena Hoyos. A first order continuous time VAR with random coefficients. Journal of Time Series Analysis, 45(1):57–77, January 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hinich:1992:NDT

[HP92] Melvin J. Hinich and Douglas M. Patterson. A new diagnostic test of model inadequacy which uses the martingale difference criterion. Journal of Time Series Analysis, 13(3):233–252, May 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hill:2014:UIE

[HP14] Jonathan Hill and Liang Peng. Unified interval estimation for random coefficient autoregressive models. Journal of Time Series Analysis, 35(3):282–297, May 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hafner:2017:ATM

[HP17] Christian M. Hafner and Arie Preminger. On asymptotic theory for ARCH(∞) models. Journal of Time Series Analysis, 38(6):865– 879, November 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Harvey:2023:RSM

[HP23] Andrew Harvey and Dario Palumbo. Regime switching models for circular and linear time series. Journal of Time Series Analysis, 44(4):374–392, July 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Horvath:2017:DMM

[HPW17] Lajos Horváth, William Pouliot, and Shixuan Wang. Detecting atmost-m changes in linear regression models. Journal of Time Series Analysis, 38(4):552–590, July 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hall:2002:PNE

[HPY02] Peter Hall, Liang Peng, and Qiwei Yao. Prediction and nonparametric estimation for time series with heavy tails. Journal of Time Series Analysis, 23(3):313–331, May 2002. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hannan:1989:RCA

[HQ89] E. J. Hannan and B. G. Quinn. The resolution of closely adjacent spectral lines. Journal of Time Series Analysis, 10(1):13–31, January 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Harvey:1988:EEN

[HR88] A. C. Harvey and P. M. Robinson. Efficient estimation of nonstationary time series regression. Journal of Time Series Analysis, 9(3):201–214, May 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hosking:1993:ASS

[HR93] J. R. M. Hosking and Nalini Ravishanker. Approximate simultaneous significance intervals for residual autocorrelations of autoregressive moving-average time series models. Journal of Time Series Analysis, 14(1):19–26, January 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hurvich:1995:EMP

[HR95] Clifford M. Hurvich and Bonnie K. Ray. Estimation of the memory parameter for nonstationary or noninvertible fractionally integrated processes. Journal of Time Series Analysis, 16(1):17–41, January 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hansen:2002:ACU

[HR02] Henrik Hansen and Anders Rahbek. Approximate conditional unit root inference. Journal of Time Series Analysis, 23(1):1–28, January 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hassler:2004:SUR

[HR04] Uwe Hassler and Paulo M. M. Rodrigues. Seasonal unit root tests under structural breaks*. Journal of Time Series Analysis, 25(1): 33–53, January 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Horvath:2015:TEM

[HR15] Lajos Horváth and Gregory Rice. Testing equality of means when the observations are from functional time series. Journal of Time Series Analysis, 36(1):84–108, January 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Harvey:1998:TDV

[HS98] Andrew Harvey and Mariane Streibel. Tests for deterministic versus indeterministic cycles. Journal of Time Series Analysis, 19 (5):505–529, September 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hidalgo:2004:ELE

[HS04] Javier Hidalgo and Philippe Soulier. Estimation of the location and exponent of the spectral singularity of a long memory process. Journal of Time Series Analysis, 25(1):55–81, January 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hallin:2005:TNC

[HS05] Marc Hallin and Abdessamad Saidi. Testing non-correlation and non-causality between multivariate ARMA time series. Journal of Time Series Analysis, 26(1):83–105, January 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hwang:2011:SBN

[HS11] Eunju Hwang and Dong Wan Shin. Stationary bootstrapping for non-parametric estimator of nonlinear autoregressive model. Journal of Time Series Analysis, 32(3):292–303, May 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hsu:2023:TSC

[HST23] Nan-Jung Hsu, Lai Heng Sim, and Ruey S. Tsay. Testing for symmetric correlation matrices with applications to factor models. Journal of Time Series Analysis, 44(5-6):622–643, September 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hardle:1986:STS

[HT86] Wolfgang Härdle and Pham-Dinh Tuan. Some theory on Msmoothing of time series. Journal of Time Series Analysis, 7(3): 191–204, May 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hannan:1988:TDE

[HT88] E. J. Hannan and P. J. Thomson. Time delay estimation. Journal of Time Series Analysis, 9(1):21–33, January 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hurvich:1993:CAI

[HT93] Clifford M. Hurvich and Chih-Ling Tsai. A corrected Akaike information criterion for vector autoregressive model selection. Journal of Time Series Analysis, 14(3):271–279, May 1993. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

He:1999:PAF

[HT99] Changli He and Timo Teräsvirta. Properties of the autocorrelation function of squared observations for second-order GARCH processes under two sets of parameter constraints. Journal of Time Series Analysis, 20(1):23–30, January 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hosoya:2010:NMF

[HT10] Yuzo Hosoya and Taro Takimoto. A numerical method for factorizing the rational spectral density matrix. Journal of Time Series Analysis, 31(4):229–240, July 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Huang:1990:LTR

[Hua90a] Dawei Huang. Levinson-type recursive algorithms for least-squares autoregression. Journal of Time Series Analysis, 11(4):295–315, July 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Huang:1990:SOG

[Hua90b] Dawei Huang. Selecting order for general autoregressive models by minimum description length. Journal of Time Series Analysis, 11(2):107–119, March 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Huang:1996:LHF

[Hua96] Dawei Huang. On low and high frequency estimation. Journal of Time Series Analysis, 17(4):351–365, July 1996. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hualde:2012:WCM

[Hua12] Javier Hualde. Weak convergence to a modified fractional Brownian motion. Journal of Time Series Analysis, 33(3):519–529, May 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hurvich:2001:MSB

[Hur01] Clifford M. Hurvich. Model selection for broadband semiparametric estimation of long memory in time series. Journal of Time Series Analysis, 22(6):679–709, November 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Huzii:1981:ECA

[Huz81] Mituaki Huzii. Estimation of coefficients of an autoregressive process by using a higher order moment. Journal of Time Series Analysis, 2(2):87–93, March 1981. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Huzii:1988:SPC

[Huz88] Mituaki Huzii. Some properties of conditional quasi-likelihood functions for time series model fitting. Journal of Time Series Analysis, 9(4):345–353, July 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Huzii:2007:EGD

[Huz07] Mituaki Huzii. Embedding a Gaussian discrete-time autoregressive moving average process in a Gaussian continuous-time autoregressive moving average process. Journal of Time Series Analysis, 28 (4):498–520, July 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hardle:1992:KRS

[HV92] Wolfgang Härdle and Philippe Vieu. Kernel regression smoothing of time series. Journal of Time Series Analysis, 13(3):209–232, May 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hotta:1999:ADS

[HV99] Luiz K. Hotta and Klaus L. Vasconcellos. Aggregation and disaggregation of structural time series models. Journal of Time Series Analysis, 20(2):155–171, March 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hashimzade:2008:FAA

[HV08] Nigar Hashimzade and Timothy J. Vogelsang. Fixed-b asymptotic approximation of the sampling behaviour of nonparametric spectral density estimators. Journal of Time Series Analysis, 29 (1):142–162, January 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Huang:2015:SNC

[HVS15] Yinxiao Huang, Stanislav Volgushev, and Xiaofeng Shao. On selfnormalization for censored dependent data. Journal of Time Series Analysis, 36(1):109–124, January 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

He:1989:EDP

[HW89] S. W. He and J. G. Wang. On embedding a discrete-parameter ARMA model in a continuous-parameter ARMA model. Journal of Time Series Analysis, 10(4):315–323, July 1989. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hall:1995:SDT

[HW95a] Peter Hall and Rodney C. L. Wolff. On the strength of dependence of a time series generated by a chaotic map. Journal of Time Series Analysis, 16(6):571–583, November 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hamilton:1995:CRP

[HW95b] David Hamilton and Ka Ho Wu. Confidence regions for parameters in the AR(1) model. Journal of Time Series Analysis, 16(3):249– 265, May 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Huerta:1999:BIP

[HW99] Gabriel Huerta and Mike West. Bayesian inference on periodicities and component spectral structure in time series. Journal of Time Series Analysis, 20(4):401–416, July 1999. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Haywood:2009:TIM

[HW09] John Haywood and Granville Tunnicliffe Wilson. A test for improved multi-step forecasting. Journal of Time Series Analysis, 30 (6):682–707, November 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Huang:2011:CSV

[HWBD11] Wenying Huang, Ke Wang, F. Jay Breidt, and Richard A. Davis. A class of stochastic volatility models for environmental applications. Journal of Time Series Analysis, 32(4):364–377, July 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hyndman:1993:YWE

[Hyn93] Rob J. Hyndman. Yule–Walker estimates for continuous-time autoregressive models. Journal of Time Series Analysis, 14(3):281– 296, May 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hong-zhi:1992:NNA

[Hz92] An Hong-zhi. Non-negative autoregressive models. Journal of Time Series Analysis, 13(4):283–295, July 1992. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hong-Zhi:1993:ERA

[HZF93] An Hong-Zhi and Huang Fuchun. Estimation for regressive and autoregressive models with non-negative residual errors. Journal of Time Series Analysis, 14(2):179–191, March 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Hong-Zhi:1983:NAE

[HZZGH83] An Hong-Zhi, Chen Zhao-Guo, and E. J. Hannan. A note on ARMA estimation. Journal of Time Series Analysis, 4(1):9–17, January 1983. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Iacone:2010:LWE

[Iac10] Fabrizio Iacone. Local Whittle estimation of the memory parameter in presence of deterministic components. Journal of Time Series Analysis, 31(1):37–49, January 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ioannidis:2005:ESV

[IC05] E. E. Ioannidis and G. A. Chronis. Extreme spectra of var models and orders of near-cointegration. Journal of Time Series Analysis, 26(3):399–421, May 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Icaza:1999:SSE

[IJ99] Gloria Icaza and Richard Jones. A state-space EM algorithm for longitudinal data. Journal of Time Series Analysis, 20(5):537– 550, September 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Iacone:2019:SDC

[IL19] Fabrizio Iacone and Stepána Lazarová. Semiparametric detection of changes in long range dependence. Journal of Time Series Analysis, 40(5):693–706, September 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Iacone:2014:FTB

[ILT14] Fabrizio Iacone, Stephen J. Leybourne, and A. M. Robert Taylor. A fixed-b test for a break in level at an unknown time under fractional integration. Journal of Time Series Analysis, 35(1):40–54, January 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Iacus:2018:DTA

[IMR18] Stefano M. Iacus, Lorenzo Mercuri, and Edit Rroji. Discrete-time approximation of a Cogarch(p, q) model and its estimation. Journal of Time Series Analysis, 39(5):787–809, September 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ing:2001:NMS

[Ing01] C. K. Ing. A note on mean-squared prediction errors of the least squares predictors in random walk models. Journal of Time Series Analysis, 22(6):711–724, November 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ioannidis:2010:URT

[Ioa10] Evangelos E. Ioannidis. Unit-root testing: on the asymptotic equivalence of Dickey–Fuller with the log–log slope of a fitted autoregressive spectrum. Journal of Time Series Analysis, 31(3):153–166, May 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ioannidis:2011:AIC

[Ioa11] Evangelos E. Ioannidis. Akaike’s information criterion correction for the least-squares autoregressive spectral estimator. Journal of Time Series Analysis, 32(6):618–630, November 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ioannidis:2022:NNP

[Ioa22] Evangelos E. Ioannidis. A new non-parametric cross-spectrum estimator. Journal of Time Series Analysis, 43(5):808–827, September 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Iglesias:2008:FST

[IP08] Emma M. Iglesias and Garry D. A. Phillips. Finite sample theory of QMLE in ARCH models with dynamics in the mean equation. Journal of Time Series Analysis, 29(4):719–737, July 2008. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Iglesias:2020:FRP

[IP20] Emma M. Iglesias and Garry D. A. Phillips. Further results on pseudo-maximum likelihood estimation and testing in the constant elasticity of variance continuous time model. Journal of Time Series Analysis, 41(2):357–364, March 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Irizarry:2002:WEH

[Iri02] Rafael A. Irizarry. Weighted estimation of harmonic components in a musical sound signal. Journal of Time Series Analysis, 23(1): 29–48, January 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ishiguro:1984:CEI

[Ish84] Makio Ishiguro. Computationally efficient implementation of a Bayesian seasonal adjustment procedure. Journal of Time Series Analysis, 5(4):245–253, July 1984. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ing:2003:ECM

[IY03] Ching-Kang Ing and Shu-Hui Yu. On estimating conditional meansquared prediction error in autoregressive models. Journal of Time Series Analysis, 24(4):401–422, July 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jenkins:1981:SAM

[JA81] Gwilym M. Jenkins and Athar S. Alavi. Some aspects of modelling and forecasting multivariate time series. Journal of Time Series Analysis, 2(1):1–47, January 1981. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Janacek:1982:DDD

[Jan82] G. J. Janacek. Determining the degree of differencing for time series via the log spectrum. Journal of Time Series Analysis, 3(3): 177–183, May 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Janacek:2005:BRM

[Jan05a] G. Janacek. Book review: Measuring Business Cycles in Economic Time Series. Journal of Time Series Analysis, 26(4):627–628, July 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Janacek:2005:BRS

[Jan05b] G. Janacek. Book review: Seasonal adjustment with the X-11 method. Journal of Time Series Analysis, 26(4):626–627, July 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Janacek:2010:TSA

[Jan10] G. Janacek. Time series analysis forecasting and control. Journal of Time Series Analysis, 31(4):303, July 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jasiak:2003:FOA

[Jas03] Joann Jasiak. First-order autoregressive processes with heterogeneous persistence. Journal of Time Series Analysis, 24(3):283–309, May 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jesus:2017:IWL

[JC17] Joao Jesus and Richard E. Chandler. Inference with the Whittle likelihood: A tractable approach using estimating functions. Journal of Time Series Analysis, 38(2):204–224, March 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jensen:2004:SBI

[Jen04] Mark J. Jensen. Semiparametric Bayesian inference of longmemory stochastic volatility models. Journal of Time Series Analysis, 25(6):895–922, November 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jentsch:2012:NFD

[Jen12] Carsten Jentsch. A new frequency domain approach of testing for covariance stationarity and for periodic stationarity in multivariate linear processes. Journal of Time Series Analysis, 33(2):177–192, March 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jandhyala:2013:ISM

[JFML13] Venkata Jandhyala, Stergios Fotopoulos, Ian MacNeill, and Pengyu Liu. Inference for single and multiple change-points in time series. Journal of Time Series Analysis, 34(4):423–446, July 2013. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jin-Guan:1991:IVA

[JGY91] Du Jin-Guan and Li Yuan. The integer-valued autoregressive (INAR(p)) model. Journal of Time Series Analysis, 12(2):129– 142, March 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jin:2018:FDT

[Jin18] Lei Jin. A frequency-domain test to check equality in spectral densities of multiple time series with unequal lengths. Journal of Time Series Analysis, 39(4):618–633, July 2018. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jirak:2016:ORC

[Jir16] Moritz Jirak. Optimal rate of convergence for empirical quantiles and distribution functions for time series. Journal of Time Series Analysis, 37(6):825–836, November 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jazi:2012:FOI

[JJL12] Mansour Aghababaei Jazi, Geoff Jones, and Chin-Diew Lai. Firstorder integer valued AR processes with zero inflated Poisson innovations. Journal of Time Series Analysis, 33(6):954–963, November 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jacobs:1983:SDA

[JL83] P. A. Jacobs and P. A. W. Lewis. Stationary discrete autoregressive-moving average time series generated by mixtures. Journal of Time Series Analysis, 4(1):19–36, January 1983. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jentsch:2020:ECF

[JLMB20] Carsten Jentsch, Anne Leucht, Marco Meyer, and Carina Beering. Empirical characteristic functions-based estimation and distance correlation for locally stationary processes. Journal of Time Series Analysis, 41(1):110–133, January 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jorda:2004:TST

[JM04] Òscar Jordà and Massimiliano Marcellino. Time-scale transformations of discrete time processes. Journal of Time Series Analysis, 25(6):873–894, November 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jach:2012:SIM

[JMP12] Agnieszka Jach, Tucker McElroy, and Dimitris N. Politis. Subsampling inference for the mean of heavy-tailed long-memory time series. Journal of Time Series Analysis, 33(1):96–111, January 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic). See correction [JMP16].

Jach:2016:CSI

[JMP16] Agnieszka Jach, Tucker S. McElroy, and Dimitris N. Politis. Corrigendum to ‘Subsampling Inference for the Mean of Heavy-Tailed Long-Memory Time Series’ by A. Jach, T. S. McElroy and D. N. Politis. Journal of Time Series Analysis, 37(5):713–720, September 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic). See [JMP12].

Jensen:2014:FFD

[JN14] Andreas Noack Jensen and Morten Ørregaard Nielsen. A fast fractional difference algorithm. Journal of Time Series Analysis, 35 (5):428–436, September 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Johansen:2018:TCF

[JN18] Søren Johansen and Morten Ørregaard Nielsen. Testing the CVAR in the fractional CVAR model. Journal of Time Series Analysis, 39 (6):836–849, November 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Johansen:2019:NCF

[JN19] Søren Johansen and Morten Ørregaard Nielsen. Nonstationary cointegration in the fractionally cointegrated VAR model. Journal of Time Series Analysis, 40(4):519–543, July 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jimenez:2006:AIM

[JO06] J. C. Jimenez and T. Ozaki. An approximate innovation method for the estimation of diffusion processes from discrete data. Journal of Time Series Analysis, 27(1):77–97, January 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Johansen:2003:AVE

[Joh03] Søren Johansen. The asymptotic variance of the estimated roots in a cointegrated vector autoregressive model. Journal of Time Series Analysis, 24(6):663–678, November 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Joyeux:1987:SCP

[Joy87] Roselyne. Joyeux. Slowly changing processes and harmonizability. Journal of Time Series Analysis, 8(4):425–431, July 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Joyeux:1992:TSC

[Joy92] Roselyne Joyeux. Tests for seasonal cointegration using principal components. Journal of Time Series Analysis, 13(2):109–118, March 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jian:1999:CEN

[JP99] Huang Jian and Yudi Pawitan. Consistent estimation for non-Gaussian non-causal autoregessive processes. Journal of Time Series Analysis, 20(4):417–423, July 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jentsch:2015:BBT

[JPP15] Carsten Jentsch, Dimitris N. Politis, and Efstathios Paparoditis. Block bootstrap theory for multivariate integrated and cointegrated processes. Journal of Time Series Analysis, 36(3):416–441, May 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jentsch:2022:GBV

[JR22] Carsten Jentsch and Lena Reichmann. Generalized binary vector autoregressive processes. Journal of Time Series Analysis, 43(2): 285–311, March 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Janacek:1990:CMN

[JS90] G. J. Janacek and A. L. Swift. A class of models for non-normal time series. Journal of Time Series Analysis, 11(1):19–31, January 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Johansen:2024:ACE

[JS24] Søren Johansen and Anders Rygh Swensen. Adjustment coefficients and exact rational expectations in cointegrated vector autoregressive models. Journal of Time Series Analysis, 45(2):248– 268, March 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jung:2003:TSD

[JT03] Robert C. Jung and A. R. Tremayne. Testing for serial dependence in time series models of counts. Journal of Time Series Analysis, 24(1):65–84, January 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jung:2011:CCM

[JT11] Robert C. Jung and A. R. Tremayne. Convolution-closed models for count time series with applications. Journal of Time Series Analysis, 32(3):268–280, May 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Janas:1995:CNL

[JvS95] Daniel Janas and Rainer von Sachs. Consistency for non-linear functions of the periodogram of tapered data. Journal of Time Series Analysis, 16(6):585–606, November 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jin:2016:NTC

[JW16] Lei Jin and Suojin Wang. A new test for checking the equality of the correlation structures of two time series. Journal of Time Series Analysis, 37(3):355–368, May 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Jammalamadaka:1999:INO

[JWW99] S. Rao Jammalamadaka, Chengou Wu, and Weiqung Wang. The influence of numerical and observational errors on the likelihood of an ARMA series. Journal of Time Series Analysis, 20(2):223–235, March 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kurozumi:2007:EEI

[KA07] Eiji Kurozumi and Yoichi Arai. Efficient estimation and inference in cointegrating regressions with structural change. Journal of Time Series Analysis, 28(4):545–575, July 2007. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kurozumi:2008:TNH

[KA08] Eiji Kurozumi and Yoichi Arai. Test for the null hypothesis of cointegration with reduced size distortion. Journal of Time Series Analysis, 29(3):476–500, May 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kabaila:1983:AEE

[Kab83] Paul Kabaila. On the asymptotic efficiency of estimators of the parameters of an ARMA process. Journal of Time Series Analysis, 4(1):37–47, January 1983. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kabaila:1987:RLB

[Kab87] Paul Kabaila. On Rissanen’s lower bound on the accumulated mean-square prediction error. Journal of Time Series Analysis, 8(3):301–309, May 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kabaila:1993:BPI

[Kab93] Paul Kabaila. On bootstrap predictive inference for autoregressive processes. Journal of Time Series Analysis, 14(5):473–484, September 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kabaila:1994:DSA

[Kab94] Paul Kabaila. The detection of a single additive outlier of unknown position. Journal of Time Series Analysis, 15(5):507–522, September 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kabaila:1999:RPP

[Kab99] Paul Kabaila. The relevance property for prediction intervals. Journal of Time Series Analysis, 20(6):655–662, November 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kakizawa:1996:TOA

[Kak96] Yoshihide Kakizawa. Third-order asymptotic properties of estimators in Gaussian ARMA processes with unknown mean. Journal of Time Series Analysis, 17(4):367–377, July 1996. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kakizawa:1999:NAE

[Kak99a] Yoshihide Kakizawa. Note on the asymptotic efficiency of sample covariances in Gaussian vector stationary processes. Journal of Time Series Analysis, 20(5):551–558, September 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kakizawa:1999:VEE

[Kak99b] Yoshihide Kakizawa. Valid Edgeworth expansions of some estimators and bootstrap confidence intervals in first-order autoregression. Journal of Time Series Analysis, 20(3):343–359, May 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kakizawa:2006:BPE

[Kak06] Yoshihide Kakizawa. Bernstein polynomial estimation of a spectral density. Journal of Time Series Analysis, 27(2):253–287, March 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kakizawa:2013:FDG

[Kak13] Yoshihide Kakizawa. Frequency domain generalized empirical likelihood method. Journal of Time Series Analysis, 34(6):691–716, November 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kang:1981:NSC

[Kan81] Heejoon Kang. Necessary and sufficient conditions for causality testing in multivariate ARMA models. Journal of Time Series Analysis, 2(2):95–101, March 1981. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kanto:1987:FIA

[Kan87] Antti J. Kanto. A formula for the inverse autocorrelation function of an autoregressive process. Journal of Time Series Analysis, 8(3): 311–312, May 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kapetanios:2001:MST

[Kap01] George Kapetanios. Model selection in threshold models. Journal of Time Series Analysis, 22(6):733–754, November 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kapetanios:2005:URT

[Kap05] George Kapetanios. Unit-root testing against the alternative hypothesis of up to m structural breaks. Journal of Time Series Analysis, 26(1):123–133, January 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Karanasos:2001:PAM

[Kar01] Menelaos Karanasos. Prediction in ARMA models with GARCH in mean effects. Journal of Time Series Analysis, 22(5):555–576, September 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Karavias:2016:BRA

[Kar16] Y. Karavias. Book review: Almost All About Unit Roots: Foundations, Developments, and Applications, by In Choi. Published by Cambridge University Press, Cambridge, 2015. Total number of pages: 295. ISBN: 978-1-107-48250-0 (paperback), price: 24.99£; (US$39.99) ISBN: 978-1-107-09733-9 (hardback), price:

60.00£(US$95.00). Journal of Time Series Analysis, 37(1):143– 144, January 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kassam:1982:RHT

[Kas82] Saleem A. Kassam. Robust hypothesis testing and robust time series interpolation and regression. Journal of Time Series Analysis, 3(3):185–194, May 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Katayama:2008:IPS

[Kat08] Naoya Katayama. An improvement of the portmanteau statistic. Journal of Time Series Analysis, 29(2):359–370, March 2008. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Katayama:2009:MPT

[Kat09] Naoya Katayama. On multiple portmanteau tests*. Journal of Time Series Analysis, 30(5):487–504, September 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Katayama:2012:CSP

[Kat12] Naoya Katayama. Chi-squared portmanteau tests for structural VARMA models with uncorrelated errors. Journal of Time Series Analysis, 33(6):863–872, November 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kavalieris:1989:EOA

[Kav89] L. Kavalieris. The estimation of the order of an autoregression using recursive residuals and cross-validation. Journal of Time Series Analysis, 10(3):271–281, May 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kavalieris:1993:TFE

[Kav93] L. Kavalieris. Transfer function estimation. Journal of Time Series Analysis, 14(5):485–496, September 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:1990:EFO

[KBB90] Won Kyung Kim, L. Billard, and I. V. Basawa. Estimation for the first-order diagonal bilinear time series model. Journal of Time Series Analysis, 11(3):215–229, May 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:1996:BSK

[KC96] Tae Yoon Kim and Dennis D. Cox. Bandwidth selection in kernel smoothing of time series. Journal of Time Series Analysis, 17(1): 49–63, January 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:2010:BRS

[KC10] Jaehee Kim and Sooyoung Cheon. A Bayesian regime-switching time-series model. Journal of Time Series Analysis, 31(5):365– 378, September 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Katzfuss:2011:STS

[KC11] Matthias Katzfuss and Noel Cressie. Spatio-temporal smoothing and EM estimation for massive remote-sensing data sets. Journal of Time Series Analysis, 32(4):430–446, July 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Karmakar:2022:LTP

[KCW22] Sayar Karmakar, Marek Chudý, and Wei Biao Wu. Long-term prediction intervals with many covariates. Journal of Time Series Analysis, 43(4):587–609, July 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Koopman:2000:FFS

[KD00] S. J. Koopman and J. Durbin. Fast filtering and smoothing for multivariate state space models. Journal of Time Series Analysis, 21(3):281–296, May 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Koopman:2003:FSS

[KD03] S. J. Koopman and J. Durbin. Filtering and smoothing of state vector for diffuse state-space models. Journal of Time Series Analysis, 24(1):85–98, January 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kashyap:1988:ELM

[KE88] R. L. Kashyap and Kie-Bum Eom. Estimation in long-memory time series model. Journal of Time Series Analysis, 9(1):35–41, January 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kedem:1987:DPH

[Ked87] Benjamin Kedem. Detection of periodicities by higher-order crossings. Journal of Time Series Analysis, 8(1):39–50, January 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Keenan:1997:CLT

[Kee97] Daniel M. Keenan. A central limit theorem for m ( n ) autocovariances. Journal of Time Series Analysis, 18(1):61–78, January 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Keich:2003:STD

[Kei03] U. Keich. Stationary tangent: the discrete and non-smooth case. Journal of Time Series Analysis, 24(2):173–192, March 2003. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kengne:2012:TPC

[Ken12] William Charky Kengne. Testing for parameter constancy in general causal time-series models. Journal of Time Series Analysis, 33(3):503–518, May 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kreiss:1992:BSA

[KF92] Jens-Peter Kreiss and Jürgen Franke. Bootstrapping stationary autoregressive moving-average models. Journal of Time Series Analysis, 13(4):297–317, July 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Koreisha:2008:ULS

[KF08] Sergio G. Koreisha and Yue Fang. Using least squares to generate forecasts in regressions with serial correlation. Journal of Time Series Analysis, 29(3):555–580, May 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kaufmann:2002:BAS

[KFS02] Sylvia Kaufmann and Sylvia Frühwirth-Schnatter. Bayesian analysis of switching ARCH models. Journal of Time Series Analysis, 23(4):425–458, July 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kong:2018:PIA

[KGY18] Juanjuan Kong, Lijie Gu, and Lijian Yang. Prediction interval for autoregressive time series via oracally efficient estimation of multistep-ahead innovation distribution function. Journal of Time Series Analysis, 39(5):690–708, September 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kavalieris:1994:DNT

[KH94] L. Kavalieris and E. J. Hannan. Determining the number of terms in a trigonometric regression. Journal of Time Series Analysis, 15 (6):613–625, November 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:1998:TCM

[KH98a] Jaehee H. Kim and Jeffrey D. Hart. Tests for change in a mean function when the data are dependent. Journal of Time Series Analysis, 19(4):399–424, July 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kuan:1998:CPE

[KH98b] Chung-Ming Kuan and Chih-Chiang Hsu. Change-point estimation of fractionally integrated processes. Journal of Time Series Analysis, 19(6):693–708, November 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kabaila:1999:APE

[KH99] Paul Kabaila and Zhisong He. On assessing prediction error in autoregressive models. Journal of Time Series Analysis, 20(6): 663–670, November 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kabaila:2001:PIC

[KH01] Paul Kabaila and Zhisong He. On prediction intervals for conditionally heteroscedastic processes. Journal of Time Series Analysis, 22(6):725–731, November 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kabaila:2004:API

[KH04a] Paul Kabaila and Zhisong He. The adjustment of prediction intervals to account for errors in parameter estimation. Journal of Time Series Analysis, 25(3):351–358, May 2004. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:2004:KMS

[KH04b] Tae Yoon Kim and Sun Young Hwang. Kernel matching scheme for block bootstrap of time series data. Journal of Time Series Analysis, 25(2):199–216, March 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kavalieris:2003:GLS

[KHS03] L. Kavalieris, E. J. Hannan, and M. Salau. Generalized least squares estimation of ARMA models. Journal of Time Series Analysis, 24(2):165–172, March 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kilian:1998:ALO

[Kil98] Lutz Kilian. Accounting for lag order uncertainty in autoregressions: the endogenous lag order bootstrap algorithm. Journal of Time Series Analysis, 19(5):531–548, September 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kilic:2011:TCI

[Kil11] Rehim Kiliç. Testing for co-integration and nonlinear adjustment in a smooth transition error correction model. Journal of Time Series Analysis, 32(6):647–660, November 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kilic:2016:TLS

[Kil16] Rehim Kiliç. Tests for linearity in star models: Supwald and LM-type tests. Journal of Time Series Analysis, 37(5):660–674, September 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Killick:2018:BRA

[Kil18] Rebecca Killick. Book review: Applied Time Series Analysis With R, Second Edition by Wayne A. Woodward, Henry L. Gray, and Alan C. Elliott (eds). Published by CRC Press, 2017. Total number of pages: 618. ISBN: 978-1-4987-3422-6. Journal of Time Series Analysis, 39(1):107, January 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:1991:CEF

[Kim91] Peter T. Kim. Consistent estimation of the fourth-order cumulant spectral density. Journal of Time Series Analysis, 12(1):63–71, January 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:2015:HTA

[Kim15] Seonjin Kim. Hypothesis testing for ARCH models: a multiple quantile regressions approach. Journal of Time Series Analysis, 36(1):26–38, January 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:2016:SIU

[Kim16] Donggyu Kim. Statistical inference for unified GARCH–Itô models with high-frequency financial data. Journal of Time Series Analysis, 37(4):513–532, July 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kitagawa:1981:NTS

[Kit81] Genshiro Kitagawa. A nonstationary time series model and its fitting by a recursive filter. Journal of Time Series Analysis, 2(2): 103–116, March 1981. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Koslov:1985:UAC

[KJ85] Judith W. Koslov and Richard H. Jones. A unified approach to confidence bounds for the autoregressive spectral estimator. Journal of Time Series Analysis, 6(3):141–151, May 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kirch:2012:TPS

[KK12] Claudia Kirch and Joseph Tadjuidje Kamgaing. Testing for parameter stability in nonlinear autoregressive models. Journal of Time Series Analysis, 33(3):365–385, May 2012. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:2020:CHE

[KK20] Mihyun Kim and Piotr Kokoszka. Consistency of the Hill estimator for time series observed with measurement errors. Journal of Time Series Analysis, 41(3):421–435, May 2020. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kidzinski:2018:PCA

[KKJ18] Lukasz Kidziński, Piotr Kokoszka, and Neda Mohammadi Jouzdani. Principal components analysis of periodically correlated functional time series. Journal of Time Series Analysis, 39(4):502–522, July 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kulperger:1998:TIT

[KL98] R. J. Kulperger and R. A. Lockhart. Tests of independence in time series. Journal of Time Series Analysis, 19(2):165–185, March 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kang:2009:PCT

[KL09] Jiwon Kang and Sangyeol Lee. Parameter change test for random coefficient integer-valued autoregressive processes with application to polio data analysis. Journal of Time Series Analysis, 30(2): 239–258, March 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:2010:CLT

[KL10] Tae Yoon Kim and Zhi-Ming Luo. Central limit theorems for nonparametric estimators with real-time random variables. Journal of Time Series Analysis, 31(5):337–347, September 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:2011:REC

[KL11] Byungsoo Kim and Sangyeol Lee. Robust estimation for the covariance matrix of multi-variate time series. Journal of Time Series Analysis, 32(5):469–481, September 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:2013:REC

[KL13] Byungsoo Kim and Sangyeol Lee. Robust estimation for copula parameter in SCOMDY models. Journal of Time Series Analysis, 34(3):302–314, May 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kong:2023:SCT

[KL23] Jiajie Kong and Robert Lund. Seasonal count time series. Journal of Time Series Analysis, 44(1):93–124, January 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Klemela:2008:DEL

[Kle08] Jussi Klemelä. Density estimation with locally identically distributed data and with locally stationary data. Journal of Time Series Analysis, 29(1):125–141, January 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kustosz:2016:TBS

[KLM16] Christoph P. Kustosz, Anne Leucht, and Christine H. Müller. Tests based on simplicial depth for AR(1) models with explosion. Journal of Time Series Analysis, 37(6):763–784, November 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:2004:BDF

[KLN04a] Tae-Hwan Kim, Stephen Leybourne, and Paul Newbold. Behaviour of Dickey–Fuller Unit-Root Tests Under Trend Misspecification. Journal of Time Series Analysis, 25(5):755–764, September 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:2004:AMS

[KLN04b] Tae-Hwan Kim, Stephen J. Leybourne, and Paul Newbold. Asymptotic mean-squared forecast error when an autoregression with linear trend is fitted to data generated by an I(0) or I(1) process. Journal of Time Series Analysis, 25(4):583–602, July 2004. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:2018:NPS

[KLN18] Young Min Kim, Soumendra N. Lahiri, and Daniel J. Nordman. Non-parametric spectral density estimation under long-range dependence. Journal of Time Series Analysis, 39(3):380–401, May 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Klein:1990:FIM

[KM90] André Klein and Guy Mélard. Fisher’s information matrix for seasonal autoregressive-moving average models. Journal of Time Series Analysis, 11(3):231–237, May 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kato:1999:SDW

[KM99] Takeshi Kato and Elias Masry. On the spectral density of the wavelet transform of fractional Brownian motion. Journal of Time Series Analysis, 20(5):559–563, September 1999. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kato:2003:TDS

[KM03] Takeshi Kato and Elias Masry. A time-domain semi-parametric estimate for strongly dependent continuous-time stationary processes. Journal of Time Series Analysis, 24(6):679–703, November 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Klein:2004:ACA

[KM04] André Klein and Guy Mélard. An algorithm for computing the asymptotic Fisher information matrix for seasonal SISO models. Journal of Time Series Analysis, 25(5):627–648, September 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kapetanios:2009:PEM

[KM09] George Kapetanios and Massimiliano Marcellino. A parametric estimation method for dynamic factor models of large dimensions. Journal of Time Series Analysis, 30(2):208–238, March 2009. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Krampe:2021:EWM

[KM21] Jonas Krampe and Timothy L. McMurry. Estimating Wold matrices and vector moving average processes. Journal of Time Series Analysis, 42(2):201–221, March 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kalliovirta:2015:GMA

[KMS15] Leena Kalliovirta, Mika Meitz, and Pentti Saikkonen. A Gaussian mixture autoregressive model for univariate time series. Journal of Time Series Analysis, 36(2):247–266, March 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kuan:2017:QRQ

[KMX17] Chung-Ming Kuan, Christos Michalopoulos, and Zhijie Xiao. Quantile regression on quantile ranges — a threshold approach. Journal of Time Series Analysis, 38(1):99–119, January 2017. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Knight:1987:RCC

[Kni87] Keith Knight. Rate of convergence of centred estimates of autoregressive parameters for infinite variance autoregressions. Journal of Time Series Analysis, 8(1):51–60, January 1987. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kellard:2015:IJJ

[KOC15] Neil Kellard, Denise Osborn, and Jerry Coakley. Introduction to the JTSA John Nankervis Memorial Issue. Journal of Time Series Analysis, 36(5):601–602, September 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kamgaing:2009:APD

[KOD09] Joseph Tadjuidje Kamgaing, Hernando Ombao, and Richard A. Davis. Autoregressive processes with data-driven regime switching. Journal of Time Series Analysis, 30(5):505–533, September 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kokoszka:2012:NPE

[Kok12] Piotr S. Kokoszka. Non-parametric econometrics. Journal of Time Series Analysis, 33(1):175, January 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kokoszka:2013:BR

[Kok13] Plotr S. Kokoszka. Book review. Journal of Time Series Analysis, 34(1):138, January 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Komaki:1999:EMP

[Kom99] Fumiyasu Komaki. An estimating method for parametric spectral densities of Gaussian time series. Journal of Time Series Analysis, 20(1):31–50, January 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kadi:1994:RAU

[KOV94] A. Kadi, G. Oppenheim, and M. C. Viano. Random aggregation of univariate and multivariate linear processes. Journal of Time Series Analysis, 15(1):31–43, January 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:2022:CQA

[KOW22] Donggyu Kim, Minseog Oh, and Yazhen Wang. Conditional quantile analysis for realized GARCH models. Journal of Time Series Analysis, 43(4):640–665, July 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kunihama:2012:EEP

[KOZ12] Tsuyoshi Kunihama, Yasuhiro Omori, and Zhengjun Zhang. Efficient estimation and particle filter for max-stable processes. Journal of Time Series Analysis, 33(1):61–80, January 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Koreisha:1989:FLE

[KP89] Sergio Koreisha and Tarmo Pukkila. Fast linear estimation methods for vector autoregressive moving-average models. Journal of Time Series Analysis, 10(4):325–339, July 1989. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Koreisha:1990:GLS

[KP90] Sergio Koreisha and Tarmo Pukkila. A generalized least-squares approach for estimation of autoregressive moving-average models. Journal of Time Series Analysis, 11(2):139–151, March 1990. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Koreisha:1993:DOV

[KP93] Sergio G. Koreisha and Tarmo Pukkila. Determining the order of a vector autoregression when the number of component series is large. Journal of Time Series Analysis, 14(1):47–69, January 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Koreisha:1995:ISA

[KP95] Sergio G. Koreisha and Tarmo Pukkila. The identification of seasonal autoregressive models. Journal of Time Series Analysis, 16(3):267–290, May 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kourogenis:2008:TUR

[KP08] Nikolaos Kourogenis and Nikitas Pittis. Testing for a unit root under errors with just barely infinite variance. Journal of Time Series Analysis, 29(6):1066–1087, November 2008. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kejriwal:2010:SPD

[KP10] Mohitosh Kejriwal and Pierre Perron. A sequential procedure to determine the number of breaks in trend with an integrated or stationary noise component. Journal of Time Series Analysis, 31 (5):305–328, September 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Khan:2013:MTW

[KP13] Md Atikur Rahman Khan and D. S. Poskitt. Moment tests for window length selection in singular spectrum analysis of short- and long-memory processes. Journal of Time Series Analysis, 34(2): 141–155, March 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kechagias:2015:DRM

[KP15] Stefanos Kechagias and Vladas Pipiras. Definitions and representations of multivariate long-range dependent time series. Journal of Time Series Analysis, 36(1):1–25, January 2015. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kechagias:2020:MBL

[KP20] Stefanos Kechagias and Vladas Pipiras. Modeling bivariate longrange dependence with general phase. Journal of Time Series Analysis, 41(2):268–292, March 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Krampe:2021:SCE

[KP21] Jonas Krampe and Efstathios Paparoditis. Sparsity concepts and estimation procedures for high-dimensional vector autoregressive models. Journal of Time Series Analysis, 42(5-6):554–579, September 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:2003:TLT

[KPRN03] Tae-Hwan Kim, Stephan Pfaffenzeller, Tony Rayner, and Paul Newbold. Testing for linear trend with application to relative primary commodity prices. Journal of Time Series Analysis, 24(5): 539–551, September 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Karanasos:2004:APL

[KPS04] Menelaos Karanasos, Zacharias Psaradakis, and Martin Sola. On the autocorrelation properties of long-memory GARCH processes. Journal of Time Series Analysis, 25(2):265–282, March 2004. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kapetanios:2019:GFD

[KPT19] George Kapetanios, Fotis Papailias, and A. M. Robert Taylor. A generalised fractional differencing bootstrap for long memory processes. Journal of Time Series Analysis, 40(4):467–492, July 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic). See corrigendum [KPT21].

Kapetanios:2021:CGF

[KPT21] George Kapetanios, Fotis Papailias, and A. M. Robert Taylor. Corrigendum to “A Generalised Fractional Differencing Bootstrap for Long Memory Processes” Journal of Time Series Analysis 40: 467– 492 (2019) DOI: 10.1111/jtsa.12460. Journal of Time Series Analysis, 42(4):492, July 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic). See [KPT19].

Kejriwal:2022:TSP

[KPY22] Mohitosh Kejriwal, Pierre Perron, and Xuewen Yu. A two-step procedure for testing partial parameter stability in cointegrated regression models. Journal of Time Series Analysis, 43(2):219– 237, March 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Krishnamurthy:1998:CEL

[KR98] Vikram Krishnamurthy and Tobias Ryden. Consistent estimation of linear and non-linear autoregressive models with Markov regime. Journal of Time Series Analysis, 19(3):291–307, May 1998. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kokoszka:2013:DOF

[KR13] Piotr Kokoszka and Matthew Reimherr. Determining the order of the functional autoregressive model. Journal of Time Series Analysis, 34(1):116–129, January 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Krafty:2016:DAT

[Kra16] Robert T. Krafty. Discriminant analysis of time series in the presence of within-group spectral variability. Journal of Time Series Analysis, 37(4):435–450, July 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kristensen:2009:SEB

[Kri09] Dennis Kristensen. On stationarity and ergodicity of the bilinear model with applications to GARCH models. Journal of Time Series Analysis, 30(1):125–144, January 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Krizmanic:2022:MLP

[Kri22] Danijel Krizmanić. Maxima of linear processes with heavy-tailed innovations and random coefficients. Journal of Time Series Analysis, 43(2):238–262, March 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Krogstad:1982:CP

[Kro82] Harald E. Krogstad. On the covariance of the periodogram. Journal of Time Series Analysis, 3(3):195–207, May 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ku:1994:NPS

[KS94] Simon Ku and Eugene Seneta. The number of peaks in a stationary sample and orthant probabilities. Journal of Time Series Analysis, 15(4):385–403, July 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kobayashi:2005:TEA

[KS05] Masahito Kobayashi and Xiuhong Shi. Testing for EGARCH against stochastic volatility models. Journal of Time Series Analysis, 26(1):135–150, January 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kabaila:2008:IPL

[KS08a] Paul Kabaila and Khreshna Syuhada. Improved prediction limits for AR(p) and ARCH(p) processes. Journal of Time Series Analysis, 29(2):213–223, March 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:2008:FSV

[KS08b] Jeongeun Kim and David S. Stoffer. Fitting stochastic volatility models in the presence of irregular sampling via particle methods and the EM algorithm. Journal of Time Series Analysis, 29(5): 811–833, September 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Koul:2018:ADS

[KS18] Hira L. Koul and Donatas Surgailis. Asymptotic distributions of some scale estimators in nonlinear models with long memory errors having infinite variance. Journal of Time Series Analysis, 39(3): 273–298, May 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Koul:2019:ADB

[KS19] Hira L. Koul and Donatas Surgailis. Asymptotic distribution of the bias corrected least squares estimators in measurement error linear regression models under long memory. Journal of Time Series Analysis, 40(4):493–518, July 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kim:2023:VMS

[KS23a] Donggyu Kim and Minseok Shin. Volatility models for stylized facts of high-frequency financial data. Journal of Time Series Analysis, 44(3):262–279, May 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kurozumi:2023:ABB

[KS23b] Eiji Kurozumi and Anton Skrobotov. On the asymptotic behavior of bubble date estimators. Journal of Time Series Analysis, 44(4): 359–373, July 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kooperberg:1995:LEP

[KST95a] Charles Kooperberg, Charles J. Stone, and Young K. Truong. Logspline estimation of a possibly mixed spectral distribution. Journal of Time Series Analysis, 16(4):359–388, July 1995. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kooperberg:1995:RCL

[KST95b] Charles Kooperberg, Charles J. Stone, and Young K. Truong. Rate of convergence for logspline spectral density estimation. Journal of Time Series Analysis, 16(4):389–401, July 1995. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kakizawa:1994:AES

[KT94a] Yoshihide Kakizawa and Masanobu Taniguchi. Asymptotic efficiency of the sample covariances in a Gaussian stationary process. Journal of Time Series Analysis, 15(3):303–311, May 1994. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kedem:1994:IFA

[KT94b] Benjamin Kedem and James Troendle. An iterative filtering algorithm for non-Fourier frequency estimation. Journal of Time Series Analysis, 15(1):45–63, January 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kokoszka:1994:IVS

[KT94c] Piotr S. Kokoszka and Murad S. Taqqu. Infinite variance stable ARMA processes. Journal of Time Series Analysis, 15(2):203–220, March 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kokoszka:2001:COU

[KT01] Piotr S. Kokoszka and Murad S. Taqqu. Can one use the Durbin– Levinson algorithm to generate infinite variance fractional ARIMA time series? Journal of Time Series Analysis, 22(3):317–337, May 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kurozumi:2010:RSD

[KT10] Eiji Kurozumi and Shinya Tanaka. Reducing the size distortion of the KPSS test. Journal of Time Series Analysis, 31(6):415– 426, November 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kachour:2011:OSI

[KT11] M. Kachour and L. Truquet. A p-order signed integer-valued autoregressive (SINAR(p)) model. Journal of Time Series Analysis, 32(3):223–236, May 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Karavias:2016:LPF

[KT16] Yiannis Karavias and Elias Tzavalis. Local power of fixed-t panel unit root tests with serially correlated errors and incidental trends. Journal of Time Series Analysis, 37(2):222–239, March 2016. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kohda:2000:CPC

[KTL00] Tohru Kohda, Akio Tsuneda, and Anthony J. Lawrance. Correlational properties of Chebyshev chaotic sequences. Journal of Time Series Analysis, 21(2):181–191, March 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tam:1998:SMA

[kTR98] Wing kuen Tam and Gregory Reinsel. Seasonal moving-average unit root tests in the presence of a linear trend. Journal of Time Series Analysis, 19(5):609–625, September 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kulperger:1985:OPW

[Kul85] Reg Kulperger. On an optimality property of Whittle’s Gaussian estimate of the parameter of the spectrum of a time series. Journal of Time Series Analysis, 6(4):253–259, July 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kumar:1986:ISB

[Kum86] Kuldeep Kumar. On the identification of some bilinear time series models. Journal of Time Series Analysis, 7(2):117–122, March 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kunst:1997:TCN

[Kun97] Robert M. Kunst. Testing for cyclical non-stationarity in autoregressive processes. Journal of Time Series Analysis, 18(2):123–135, March 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kurita:2011:LPL

[Kur11] Takamitsu Kurita. Local power of likelihood-based tests for cointegrating rank: Comparative analysis of full and partial systems. Journal of Time Series Analysis, 32(6):672–679, November 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kurozumi:2017:MPC

[Kur17] Eiji Kurozumi. Monitoring parameter constancy with endogenous regressors. Journal of Time Series Analysis, 38(5):791–805, September 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kurozumi:2018:CSD

[Kur18] Eiji Kurozumi. Confidence sets for the date of a structural change at the end of a sample. Journal of Time Series Analysis, 39(6):850– 862, November 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kurozumi:2021:ABD

[Kur21] Eiji Kurozumi. Asymptotic behavior of delay times of bubble monitoring tests. Journal of Time Series Analysis, 42(3):314–337, May 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kokoszka:2004:SMH

[KW04] Piotr Kokoszka and Michael Wolf. Subsampling the mean of heavytailed dependent observations. Journal of Time Series Analysis, 25(2):217–234, March 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kulik:2012:CVE

[KW12] Rafał Kulik and Cornelia Wichelhaus. Conditional variance estimation in regression models with long memory. Journal of Time Series Analysis, 33(3):468–483, May 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Keenan:2012:MNT

[KWPV12] Daniel M. Keenan, Xin Wang, Steven M. Pincus, and Johannes D. Veldhuis. Modelling the nonlinear time dynamics of multidimensional hormonal systems. Journal of Time Series Analysis, 33(5): 779–796, September 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Krafty:2012:ESS

[KXS^{+}12] Robert T. Krafty, Shuangyan Xiong, David S. Stoffer, Daniel J. Buysse, and Martica Hall. Enveloping spectral surfaces: covariate dependent spectral analysis of categorical time series. Journal of Time Series Analysis, 33(5):797–806, September 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kachour:2009:FOR

[KY09] M. Kachour and J. F. Yao. First-order rounded integer-valued autoregressive (RINAR(1)) process. Journal of Time Series Analysis, 30(4):417–448, July 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Kejriwal:2020:BPD

[KYP20] Mohitosh Kejriwal, Xuewen Yu, and Pierre Perron. Bootstrap procedures for detecting multiple persistence shifts in heteroskedastic time series. Journal of Time Series Analysis, 41(5):676–690, September 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Larsson:1998:BCU

[Lar98] Rolf Larsson. Bartlett corrections for unit root test statistics. Journal of Time Series Analysis, 19(4):425–438, July 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Latour:1998:ESS

[Lat98] Alain Latour. Existence and stochastic structure of a non-negative integer-valued autoregressive process. Journal of Time Series Analysis, 19(4):439–455, July 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Latour:2017:BRH

[Lat17] Alain Latour. Book review: Handbook of Discrete-Valued Time Series, edited by R. A. Davis, S. H. Holan, R. Lund, R. and Ravishanker. Published by Hall/CRC, Boca Raton, Florida, 2015. Total number of pages: 464. ISBN: 978-1-4665-7773-2. Journal of Time Series Analysis, 38(3):508–509, May 2017. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Layton:1984:FND

[Lay84] Allan P. Layton. A further note on the detection of Granger instantaneous causality. Journal of Time Series Analysis, 5(1):15–18, January 1984. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Laycock:1998:BR

[Lay98] Patric Laycock. Book review. Journal of Time Series Analysis, 19(3):377–378, May 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lund:2000:RPL

[LB00] Robert Lund and I. V. Basawa. Recursive prediction and likelihood evaluation for periodic ARMA models. Journal of Time Series Analysis, 21(1):75–93, January 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2011:AAR

[LB11] Weiming Li and Z. D. Bai. Analysis of accumulated rounding errors in autoregressive processes. Journal of Time Series Analysis, 32 (5):518–530, September 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lund:2009:TES

[LBV09] Robert Lund, Hany Bassily, and Brani Vidakovic. Testing equality of stationary autocovariances. Journal of Time Series Analysis, 30(3):332–348, May 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lai:2003:DEL

[LC03] Dejian Lai and Guanrong Chen. Distribution of the estimated Lyapunov exponents from noisy chaotic time series. Journal of Time Series Analysis, 24(6):705–720, November 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lee:2004:LDU

[LD04] Taiyeong Lee and David A. Dickey. Limiting distributions of unconditional maximum likelihood unit root test statistics in seasonal time-series models. Journal of Time Series Analysis, 25(4): 551–561, July 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Liu:2019:SCO

[LDH19] Jun Liu, Rohit Deo, and Clifford Hurvich. The slow convergence of ordinary least squares estimators of α, β and portfolio weights under long-memory stochastic volatility. Journal of Time Series Analysis, 40(4):590–608, July 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lahiri:2019:ELL

[LDN19] Soumendra N. Lahiri, Ujjwal Das, and Daniel J. Nordman. Empirical likelihood for a long range dependent process subordinated to a Gaussian process. Journal of Time Series Analysis, 40(4): 447–466, July 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ledolter:1990:ODT

[Led90] Johannes Ledolter. Outlier diagnostics in time series analysis. Journal of Time Series Analysis, 11(4):317–324, July 1990. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lee:2016:PSP

[Lee16] Dong Jin Lee. Parametric and semi-parametric efficient tests for parameter instability. Journal of Time Series Analysis, 37(4):451– 475, July 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lenart:2016:GRS

[Len16] Lukasz Lenart. Generalized resampling scheme with application to spectral density matrix in almost periodically correlated class of time series. Journal of Time Series Analysis, 37(3):369–404, May 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Leonenko:2013:BRD

[Leo13] Nikolai Leonenko. Book review: Domenico Marinucci and Giovanni Peccati, Random Fields on the Sphere: Representation, Limit Theorems and Cosmological Applications, London Mathematical Society Lecture Notes Series 389. Published by the Cambridge University Press, Cambridge, 2011. Number of Pages: 341. Price £40.00, ISBN 978-0-521-17561-6. Journal of Time Series Analysis, 34(5): 602–603, September 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Levy:2002:CFD

[Lev02] D. Levy. Cointegration in frequency domain. Journal of Time Series Analysis, 23(3):333–339, May 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lu:2011:TUR

[LG11] Zhiping Lu and Dominique Guegan. Testing unit roots and long range dependence of foreign exchange. Journal of Time Series Analysis, 32(6):631–638, November 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:1983:ERC

[LH83] W. K. Li and Y. V. Hui. Estimation of random coefficient autoregressive process: an empirical Bayes approach. Journal of Time Series Analysis, 4(2):89–94, March 1983. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lindoff:1996:BCR

[LH96] B. Lindoff and J. Holst. Bias and covariance of the recursive least squares estimator with exponential forgetting in vector autoregressions. Journal of Time Series Analysis, 17(6):553–570, November 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lii:1982:ETD

[LHR82] K. S. Lii, K. N. Helland, and M. Rosenblatt. Estimating threedimensional energy transfer in isotropic turbulence*. Journal of Time Series Analysis, 3(1):1–28, January 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:1984:ASI

[Li84] W. K. Li. On the autocorrelation structure and identification of some bilinear time series. Journal of Time Series Analysis, 5(3): 173–181, May 1984. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:1993:EBD

[Li93] Ta-Hsin Li. Estimation and blind deconvolution of autoregressive systems with nonstationary binary inputs. Journal of Time Series Analysis, 14(6):575–588, November 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:1998:TCA

[Li98] Ta-Hsin Li. Time-correlation analysis of nonstationary time series. Journal of Time Series Analysis, 19(1):47–67, January 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2006:SNM

[Li06] Lei M. Li. Some notes on mutual information between past and future. Journal of Time Series Analysis, 27(2):309–322, March 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2012:NMA

[Li12a] Dong Li. A note on moving-average models with feedback. Journal of Time Series Analysis, 33(6):873–879, November 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2012:RSA

[Li12b] Ta-Hsin Li. On robust spectral analysis by least absolute deviations. Journal of Time Series Analysis, 33(2):298–303, March 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2014:QPT

[Li14] Ta-Hsin Li. Quantile periodogram and time-dependent variance. Journal of Time Series Analysis, 35(4):322–340, July 2014. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2020:LME

[Li20] Qian Li. Location multiplicative error models with quasi maximum likelihood estimation. Journal of Time Series Analysis, 41(3):387– 405, May 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2024:CBS

[Li24] Yifan Li. Correcting the bias of the sample cross-covariance estimator. Journal of Time Series Analysis, 45(2):214–247, March 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Liebscher:2005:TUA

[Lie05] Eckhard Liebscher. Towards a unified approach for proving geometric ergodicity and mixing properties of nonlinear autoregressive processes. Journal of Time Series Analysis, 26(5):669–689, September 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lieberman:2012:SBA

[Lie12] Offer Lieberman. A similarity-based approach to time-varying coefficient non-stationary autoregression. Journal of Time Series Analysis, 33(3):484–502, May 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lii:1985:TFM

[Lii85] Keh-Shin Lii. Transfer function model order and parameter estimation. Journal of Time Series Analysis, 6(3):153–169, May 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lim:1987:CSV

[Lim87] K. S. Lim. A comparative study of various univariate time series models for Canadian lynx data. Journal of Time Series Analysis, 8(2):161–176, March 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lim:1992:STI

[Lim92] K. S. Lim. On the stability of a threshold AR(1) without intercepts. Journal of Time Series Analysis, 13(2):119–132, March 1992. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Liu:1989:EGM

[Liu89a] Jian Liu. On the existence of a general multiple bilinear time series. Journal of Time Series Analysis, 10(4):341–355, July 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Liu:1989:SCE

[Liu89b] Jian Liu. A simple condition for the existence of some stationary bilinear time series. Journal of Time Series Analysis, 10(1):33– 39, January 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Liu:1992:SRK

[Liu92] Jian Liu. Spectral radius, Kronecker products and stationarity. Journal of Time Series Analysis, 13(4):319–325, July 1992. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Liu:2012:FMS

[Liu12] Ji-Chun Liu. A family of Markov-switching GARCH processes. Journal of Time Series Analysis, 33(6):892–902, November 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lee:2023:BRC

[LJ23] Sangyeol Lee and Minyoung Jo. Bivariate random coefficient integer-valued autoregressive models: Parameter estimation and change point test. Journal of Time Series Analysis, 44(5-6):644– 666, September 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ljung:1988:LMT

[Lju88] Greta M. Ljung. On the Lagrange multiplier test for autoregressive moving-average models. Journal of Time Series Analysis, 9(4): 355–359, July 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:1998:TAF

[LK98] Ta-Hsin Li and Benjamin Kedem. Tracking abrupt frequency changes. Journal of Time Series Analysis, 19(1):69–82, January 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2021:JDS

[LK21] Nan Li and Simon S. Kwok. Jointly determining the state dimension and lag order for Markov-switching vector autoregressive models. Journal of Time Series Analysis, 42(4):471–491, July 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lee:2015:TVS

[LKB15] Taewook Lee, Moosup Kim, and Changryong Baek. Tests for volatility shifts in GARCH against long-range dependence. Journal of Time Series Analysis, 36(2):127–153, March 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Leybourne:2005:ESM

[LKN05] Stephen Leybourne, Tae-Hwan Kim, and Paul Newbold. Examination of some more powerful modifications of the Dickey–Fuller test. Journal of Time Series Analysis, 26(3):355–369, May 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lawrance:1992:RRA

[LL92] A. J. Lawrance and P. A. W. Lewis. Reversed residuals in autoregressive time series analysis. Journal of Time Series Analysis, 13(3):253–266, May 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Leibowitz:1995:APS

[LL95] Daniela Leibowitz and Elia M. Leibowitz. An algorithm for a period search in a sparsely covered time series at a fixed phase. Journal of Time Series Analysis, 16(2):221–236, March 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ling:1997:DCN

[LL97] Shiqing Ling and W. K. Li. Diagnostic checking of nonlinear multivariate time series with multivariate arch errors. Journal of Time Series Analysis, 18(5):447–464, September 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lambert-Lacroix:2005:EAC

[LL05] Sophie Lambert-Lacroix. Extension of autocovariance coefficients sequence for periodically correlated processes. Journal of Time Series Analysis, 26(3):423–435, May 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Levy-Leduc:2006:EFE

[LL06] Céline Lévy-Leduc. Efficient frequency estimation from a particular almost periodic function with application to laser vibrometry. Journal of Time Series Analysis, 27(5):637–669, September 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2010:ATT

[LL10] Jing Li and Junsoo Lee. ADL tests for threshold cointegration. Journal of Time Series Analysis, 31(4):241–254, July 2010. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lee:2012:RER

[LL12] Jaechoul Lee and Robert Lund. A refined efficiency rate for ordinary least squares and generalized least squares estimators for a linear trend with autoregressive errors. Journal of Time Series Analysis, 33(2):312–324, March 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2018:TET

[LL18] Linyuan Li and Kewei Lu. Tests for the equality of two processes’ spectral densities with unequal lengths using wavelet methods. Journal of Time Series Analysis, 39(1):4–27, January 2018. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Levy-Leduc:2011:RES

[LLBM^{+}11] Céline Lévy-Leduc, Hélène Boistard, Eric Moulines, Murad S. Taqqu, and Valderio A. Reisen. Robust estimation of the scale and of the autocovariance function of Gaussian short- and longrange dependent processes. Journal of Time Series Analysis, 32(2): 135–156, March 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lin:2009:LLE

[LLG09] Zhengyan Lin, Degui Li, and Jiti Gao. Local linear M-estimation in non-parametric spatial regression. Journal of Time Series Analysis, 30(3):286–314, May 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2013:MMG

[LLL13] Muyi Li, Wai Keung Li, and Guodong Li. On mixture memory GARCH models. Journal of Time Series Analysis, 34(6):606–624, November 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Liu:2022:TVJ

[LLL22] Guangying Liu, Meiyao Liu, and Jinguan Lin. Testing the volatility jumps based on the high frequency data. Journal of Time Series Analysis, 43(5):669–694, September 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Levy-Leduc:2008:FEB

[LLMR08] C. Lévy-Leduc, E. Moulines, and F. Roueff. Frequency estimation based on the cumulated Lomb-Scargle periodogram. Journal of Time Series Analysis, 29(6):1104–1131, November 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Laib:2008:REP

[LLOS08] Naâmane Laïb, Mohamed Lemdani, and Elias Ould-Saïd. On residual empirical processes of GARCH-SM models: application to conditional symmetry tests. Journal of Time Series Analysis, 29(5): 762–782, September 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Leonardi:2021:IBI

[LLRR^{+}21] Florencia Leonardi, Matías Lopez-Rosenfeld, Daniela Rodriguez, Magno T. F. Severino, and Mariela Sued. Independent block identification in multivariate time series. Journal of Time Series Analysis, 42(1):19–33, January 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lopes:1997:SDC

[LLS97] Artur Lopes, Selvia Lopes, and Rafael R. Souza. On the spectral density of a class of chaotic time series. Journal of Time Series Analysis, 18(5):465–474, September 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lanne:2002:CUR

[LLS02] Markku Lanne, Helmut Lütkepohl, and Pentti Saikkonen. Comparison of unit root tests for time series with level shifts. Journal of Time Series Analysis, 23(6):667–685, November 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lenart:2008:STA

[LLS08] Łukasz Lenart, Jacek Leśkow, and Rafał Synowiecki. Subsampling in testing autocovariance for periodically correlated time series. Journal of Time Series Analysis, 29(6):995–1018, November 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2014:HBA

[LLT14] Guodong Li, Chenlei Leng, and Chih-Ling Tsai. A hybrid bootstrap approach to unit root tests. Journal of Time Series Analysis, 35(4):299–321, July 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2014:SVS

[LLY14] Degao Li, Guodong Li, and Jinhong You. Significant variable selection and autoregressive order determination for time-series partially linear models. Journal of Time Series Analysis, 35(5):478– 490, September 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Liao:2022:RTU

[LLZZ22] Guili Liao, Qimeng Liu, Rongmao Zhang, and Shifang Zhang. Rank test of unit-root hypothesis with AR-GARCH errors. Journal of Time Series Analysis, 43(5):695–719, September 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:1988:AMN

[LM88] W. K. Li and A. I. McLeod. Arma modelling with non-Gaussian innovations. Journal of Time Series Analysis, 9(2):155–168, March 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:1994:SRA

[LM94a] W. K. Li and T. K. Mak. On the squared residual autocorrelations in non-linear time series with conditional heteroskedasticity. Journal of Time Series Analysis, 15(6):627–636, November 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lye:1994:NLT

[LM94b] Jenny N. Lye and Vance L. Martin. Non-linear time series modelling and distributional flexibility. Journal of Time Series Analysis, 15(1):65–84, January 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lii:1995:SRS

[LM95] Keh-Shin Lii and Elias Masry. On the selection of random sampling schemes for the spectral estimation of continuous time processes. Journal of Time Series Analysis, 16(3):291–311, May 1995. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lavielle:2000:LSE

[LM00] Marc Lavielle and Eric Moulines. Least-squares estimation of an unknown number of shifts in a time series. Journal of Time Series Analysis, 21(1):33–59, January 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lagos:2004:ILR

[LM04] Bernardo M. Lagos and Pedro A. Morettin. Improvement of the likelihood ratio test statistic in ARMA models. Journal of Time Series Analysis, 25(1):83–101, January 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lin:2008:PTA

[LM08] J.-W. Lin and A. I. McLeod. Portmanteau tests for ARMA models with infinite variance. Journal of Time Series Analysis, 29(3): 600–617, May 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Leybourne:1999:SPP

[LN99] Stephen Leybourne and Paul Newbold. On the size properties of Phillips–Perron tests. Journal of Time Series Analysis, 20(1):51– 61, January 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Leybourne:1998:URS

[LNV98] Stephen Leybourne, Paul Newbold, and Dimitrios Vougas. Unit roots and smooth transitions. Journal of Time Series Analysis, 19(1):83–97, January 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Leybourne:2002:DTC

[LNVK02] Stephen J. Leybourne, Paul Newbold, Dimitrios Vougas, and Tae-Hwan Kim. A direct test for cointegration between a pair of time series. Journal of Time Series Analysis, 23(2):173–191, March 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lim:2016:CQP

[LO16] Yaeji Lim and Hee-Seok Oh. Composite quantile periodogram for spectral analysis. Journal of Time Series Analysis, 37(2):195–221, March 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lobato:1997:CAC

[Lob97] Ignacio N. Lobato. Consistency of the averaged cross-periodogram in long memory series. Journal of Time Series Analysis, 18(2): 137–155, March 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Loges:2004:SMD

[Log04] Wilfried Loges. The stationary marginal distribution of a threshold AR(1) process. Journal of Time Series Analysis, 25(1):103–125, January 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Loredo-Osti:2012:ERD

[LOS12] J. C. Loredo-Osti and Brajendra C. Sutradhar. Estimation of regression and dynamic dependence parameters for non-stationary multinomial time series. Journal of Time Series Analysis, 33(3): 458–467, May 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Louni:2008:ODA

[Lou08] Hamid Louni. Outlier detection in ARMA models. Journal of Time Series Analysis, 29(6):1057–1065, November 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lieberman:2004:EBA

[LP04] Offer Lieberman and Peter C. B. Phillips. Error bounds and asymptotic expansions for Toeplitz product functionals of unbounded spectra. Journal of Time Series Analysis, 25(5):733–753, September 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic). See corrigendum [Tak24].

Landajo:2010:STU

[LP10a] Manuel Landajo and María José Presno. Stationarity testing under nonlinear models. Some asymptotic results. Journal of Time Series Analysis, 31(5):392–405, September 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Luati:2010:HSE

[LP10b] Alessandra Luati and Tommaso Proietti. Hyper-spherical and elliptical stochastic cycles. Journal of Time Series Analysis, 31(3): 169–181, May 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lieberman:2014:NRL

[LP14] Offer Lieberman and Peter C. B. Phillips. Norming rates and limit theory for some time-varying coefficient autoregressions. Journal of Time Series Analysis, 35(6):592–623, November 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Liu:2019:ATU

[LP19] Xiaohui Liu and Liang Peng. Asymptotic theory and unified confidence region for an autoregressive model. Journal of Time Series Analysis, 40(1):43–65, January 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Leipus:2020:ELM

[LPPS20] Remigijus Leipus, Anne Philippe, Vytaute Pilipauskaite, and Donatas Surgailis. Estimating long memory in panel randomcoefficient AR(1) data. Journal of Time Series Analysis, 41(4): 520–535, July 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lahiri:2018:E

[LPR18] Soumendra N. Lahiri, Dimitris N. Politis, and Peter M. Robinson. Editorial. Journal of Time Series Analysis, 39(3):241, May 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lin:1999:RMT

[LPS99] T. C. Lin, M. Pourahmadi, and A. Schick. Regression models with time series errors. Journal of Time Series Analysis, 20(4):425–433, July 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2023:GEC

[LPS23] Xiaoyan Li, Jiazhu Pan, and Anchao Song. Geometric ergodicity and conditional self-weighted M-estimator of a GRCAR(p) model with heavy-tailed errors. Journal of Time Series Analysis, 44(4): 418–436, July 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ling:2015:ISB

[LPZ15] Shiqing Ling, Liang Peng, and Fukang Zhu. Inference for a special bilinear time-series model. Journal of Time Series Analysis, 36(1): 61–66, January 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2020:MDT

[LQ20] Dong Li and Jiaming Qiu. The marginal density of a TMA(1) process. Journal of Time Series Analysis, 41(3):476–484, May 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lewis:1988:PEM

[LR88] Richard A. Lewis and Gregory C. Reinsel. Prediction error of multivariate time series with mis-specified models. Journal of Time Series Analysis, 9(1):43–57, January 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lewis:2002:NMP

[LR02] Peter A. W. Lewis and Bonnie K. Ray. Nonlinear modelling of periodic threshold autoregressions using TSMARS. Journal of Time Series Analysis, 23(4):459–471, July 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2021:SNU

[LR21] Yifan Li and Yao Rao. A simple nearly unbiased estimator of cross-covariances. Journal of Time Series Analysis, 42(2):240–266, March 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2021:LWE

[LRS21] Degui Li, Peter M. Robinson, and Han Lin Shang. Local Whittle estimation of long-range dependence for functional time series. Journal of Time Series Analysis, 42(5-6):685–695, September 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lanne:2003:RSD

[LS03a] Markku Lanne and Pentti Saikkonen. Reducing size distortions of parametric stationarity tests. Journal of Time Series Analysis, 24 (4):423–439, July 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2003:TSC

[LS03b] Dingding Li and Thanasis Stengos. Testing serial correlation in semiparametric time series models. Journal of Time Series Analysis, 24(3):311–335, May 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lopes:2006:BMU

[LS06] Hedibert F. Lopes and Esther Salazar. Bayesian model uncertainty in smooth transition autoregressions. Journal of Time Series Analysis, 27(1):99–117, January 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lin:2016:BDS

[LSSC16] Ming Lin, Eric A. Suess, Robert H. Shumway, and Rong Chen. Bayesian deconvolution of signals observed on arrays. Journal of Time Series Analysis, 37(6):837–850, November 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lim:1983:SAD

[LT83] K. S. Lim and H. Tong. A statistical approach to difference-delay equation modelling in ecology — two case studies. Journal of Time Series Analysis, 4(4):239–267, July 1983. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lii:1992:DSN

[LT92] K.-S. Lii and T.-H. Tsou. Detecting sinusoids in non-Gaussian noise. Journal of Time Series Analysis, 13(5):391–409, September 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lii:1995:BAR

[LT95] Keh-Shin Lii and Tai-Houn Tsou. Bispectral analysis of randomly sampled data. Journal of Time Series Analysis, 16(1):43–66, January 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Leybourne:2003:SUR

[LT03] Stephen Leybourne and A. M. Robert Taylor. Seasonal unit root tests based on forward and reverse estimation. Journal of Time Series Analysis, 24(4):441–460, July 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lacal:2017:LGA

[LT17] Virginia Lacal and Dag TjØstheim. Local Gaussian autocorrelation and tests for serial independence. Journal of Time Series Analysis, 38(1):51–71, January 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Leybourne:2018:SIJ

[LT18] Stephen Leybourne and Robert Taylor. Special issue of the Journal of Time Series Analysis in honour of Professor Paul Newbold: Guest Editors’ introduction. Journal of Time Series Analysis, 39 (6):814–815, November 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Leybourne:2007:CSB

[LTK07] Stephen Leybourne, Robert Taylor, and Tae-Hwan Kim. CUSUM of squares-based tests for a change in persistence. Journal of Time Series Analysis, 28(3):408–433, May 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Liu:2018:ATT

[LTT18] Yan Liu, Yurie Tamura, and Masanobu Taniguchi. Asymptotic theory of test statistic for sphericity of high-dimensional time series. Journal of Time Series Analysis, 39(3):402–416, May 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2020:NFA

[LTZ20] Degui Li, Jiraroj Tosasukul, and Wenyang Zhang. Nonlinear factor-augmented predictive regression models with functional coefficients. Journal of Time Series Analysis, 41(3):367–386, May 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lu:2018:BRH

[Lu18] Zudi Lu. Book review: Hidden Markov Models for Time Series: An Introduction Using R, 2nd Edition, by Walter Zucchini, Iain L. Macdonald, and Roland Langrock. Monographs on Statistics and Applied Probability 150, Published by CRC Press, 2016. Total number of pages: 28 + 370. ISBN: 978-1-4822-5383-2 (Hardback). Journal of Time Series Analysis, 39(1):105–106, January 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lubian:1999:LME

[Lub99] Diego Lubian. Long-memory errors in time series regressions with a unit root. Journal of Time Series Analysis, 20(5):565–577, September 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Luger:2006:MUE

[Lug06] Richard Luger. Median-unbiased estimation and exact inference methods for first-order autoregressive models with conditional heteroscedasticity of unknown form. Journal of Time Series Analysis, 27(1):119–128, January 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lutkepohl:1982:DMT

[Lüt82] Helmut Lütkepohl. Differencing multiple time series: Another look at Canadian money and income data. Journal of Time Series Analysis, 3(4):235–243, July 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lutkepohl:1985:CCE

[Lüt85] Helmut Lütkepohl. Comparison of criteria for estimating the order of a vector autoregressive process. Journal of Time Series Analysis, 6(1):35–52, January 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Leipus:2000:MLM

[LV00] Remigijus Leipus and Marie-Claude Viano. Modelling longmemory time series with finite or infinite variance: a general approach. Journal of Time Series Analysis, 21(1):61–74, January 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lam:1991:PSA

[LW91] Raymond L. H. Lam and Donald G. Watts. Profile summaries for ARIMA time series model parameters. Journal of Time Series Analysis, 12(3):225–235, May 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Liu:2019:VEJ

[LW19] Weiyi Liu and Mingjin Wang. Volatility estimation and jump testing via realized information variation. Journal of Time Series Analysis, 40(5):753–787, September 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Libert:1993:ISS

[LWL93] Gaëtan Libert, Liang Wang, and Bao Liu. An innovation state space approach for time series forecasting. Journal of Time Series Analysis, 14(6):589–601, November 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:1996:MSO

[LX96] Lei Li and Zhongjie Xie. Model selection and order determination for time series by information between the past and the future. Journal of Time Series Analysis, 17(1):65–84, January 1996. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2001:BTS

[LX01] Hongyi Li and Zhijie Xiao. Bootstrapping time series regressions with integrated processes. Journal of Time Series Analysis, 22(4): 461–480, July 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Liu:2020:RLI

[LXT20] Yan Liu, Yujie Xue, and Masanobu Taniguchi. Robust linear interpolation and extrapolation of stationary time series in L^{p}. Journal of Time Series Analysis, 41(2):229–248, March 2020. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lavitas:2018:TSS

[LZ18]  Liliya Lavitas and Ting Zhang.  A time-symmetric self-

normalization approach for inference of time series. Journal of Time Series Analysis, 39(5):748–762, September 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Li:2020:IAE

[LZ20] Dong Li and Ke Zhu. Inference for asymmetric exponentially weighted moving average models. Journal of Time Series Analysis, 41(1):154–162, January 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lin:2024:VLD

[LZ24] Yuchang Lin and Qianqian Zhu. On vector linear double autoregression. Journal of Time Series Analysis, 45(3):376–397, May 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Liu:2022:MND

[LZZ22] Mengya Liu, Fukang Zhu, and Ke Zhu. Modeling normalcydominant ordinal time series: an application to air quality level. Journal of Time Series Analysis, 43(3):460–478, May 2022. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Minozzo:1993:UEL

[MA93] M. Minozzo and A. Azzalini. On the unimodality of the exact likelihood function for normal AR(2) series. Journal of Time Series Analysis, 14(5):497–509, September 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ma:2002:EML

[Ma02] Chunsheng Ma. Exact maximum likelihood estimation of an ARMA(1, 1) model with incomplete data. Journal of Time Series Analysis, 23(1):49–56, January 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Melo:2020:CMP

[MA20] Moizes Melo and Airlane Alencar. Conway–Maxwell–Poisson autoregressive moving average model for equidispersed, underdispersed, and overdispersed count data. Journal of Time Series Analysis, 41(6):830–857, November 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mainassara:2024:PTP

[MA24] Y. Boubacar Maïnassara and A. Ilmi Amir. Portmanteau tests for periodic ARMA models with dependent errors. Journal of Time Series Analysis, 45(2):164–188, March 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mainassara:2012:SWV

[Maï12] Y. Boubacar Maïnassara. Selection of weak VARMA models by modified Akaike’s information criteria. Journal of Time Series Analysis, 33(1):121–130, January 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Marshall:1992:SSM

[Mar92a] Pablo Marshall. State space models with diffuse initial conditions. Journal of Time Series Analysis, 13(5):411–414, September 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Martin:1992:TTS

[Mar92b] Vance L. Martin. Threshold time series models as multimodal distribution jump processes. Journal of Time Series Analysis, 13(1): 79–94, January 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Marmol:1995:SRB

[Mar95] Francesc Marmol. Spurious regressions between i( d ) processes. Journal of Time Series Analysis, 16(3):313–321, May 1995. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Martin:1999:DPA

[Mar99] Donald E. K. Martin. Detection of periodic autocorrelation in time series data via zero-crossings. Journal of Time Series Analysis, 20 (4):435–452, July 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Marinucci:2000:SRC

[Mar00] D. Marinucci. Spectral regression for cointegrated time series with long-memory innovations. Journal of Time Series Analysis, 21 (6):685–705, November 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Marcellino:2007:PBD

[Mar07a] Massimiliano Marcellino. Pooling-based data interpolation and backdating. Journal of Time Series Analysis, 28(1):53–71, January 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Marsh:2007:COT

[Mar07b] Patrick Marsh. Constructing optimal tests on a lagged dependent variable. Journal of Time Series Analysis, 28(5):723–743, September 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Martin:2012:EVA

[Mar12] Rodney A. Martin. Extreme value analysis of optimal level-crossing prediction for linear Gaussian processes. Journal of Time Series Analysis, 33(4):583–607, July 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Marsh:2020:PPE

[Mar20] Patrick Marsh. Properties of the power envelope for tests against both stationary and explosive alternatives: The effect of trends. Journal of Time Series Analysis, 41(1):146–153, January 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Masry:1996:MLP

[Mas96] Elias Masry. Multivariate local polynomial regression for time series:uniform strong consistency and rates. Journal of Time Series Analysis, 17(6):571–599, November 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mauricio:2002:AEL

[Mau02] JosÉ Alberto Mauricio. An algorithm for the exact likelihood of a stationary vector autoregressive-moving average model. Journal of Time Series Analysis, 23(4):473–486, July 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mauget:2011:TSA

[Mau11] Steve Mauget. Time series analysis based on running Mann-Whitney Z Statistics. Journal of Time Series Analysis, 32(1): 47–53, January 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mayer:2022:EIA

[May22] Alexander Mayer. Estimation and inference in adaptive learning models with slowly decreasing gains. Journal of Time Series Analysis, 43(5):720–749, September 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mazzoni:2012:FCD

[Maz12] Thomas Mazzoni. Fast continuous-discrete DAF-filters. Journal of Time Series Analysis, 33(2):193–210, March 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mitchell:1997:ECM

[MB97] Heather Mitchell and Peter Brockwell. Estimation of the coefficients of a multivariate linear filter using the innovations algorithm. Journal of Time Series Analysis, 18(2):157–179, March 1997. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McCloskey:2013:ELM

[McC13] Adam McCloskey. Estimation of the long-memory stochastic volatility model parameters that is robust to level shifts and deterministic trends. Journal of Time Series Analysis, 34(3):285–301, May 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McCabe:2015:BRD

[McC15] Brendan McCabe. Book review: Discrete Time Series, Processes, and Applications in Finance, by Gilles Zumbach. Springer Finance Series. Published by Springer, Heidelberg, Berlin, 2013. Total Number of Pages: 315. ISBN: 978-3-642-31741-5. Journal of Time Series Analysis, 36(1):125, January 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McElroy:2018:RCB

[McE18] Tucker McElroy. Recursive computation for block-nested covariance matrices. Journal of Time Series Analysis, 39(3):299–312, May 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McLeod:1993:NAM

[McL93] A. I. McLeod. A note on ARMA model parameter redundancy. Journal of Time Series Analysis, 14(2):207–208, March 1993. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McLeod:1994:DCP

[McL94] A. I. McLeod. Diagnostic checking of periodic autoregression models with application. Journal of Time Series Analysis, 15(2):221– 233, March 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McLeod:1995:DCP

[McL95] A. I. McLeod. Diagnostic checking of periodic autoregression models with application. Journal of Time Series Analysis, 16(6):647– 648, November 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McLeod:1998:HDT

[McL98] A. I. McLeod. Hyperbolic decay time series. Journal of Time Series Analysis, 19(4):473–483, July 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McLeod:2017:BRM

[McL17] A. I. McLeod. Book review: Models for Dependent Time Series, by Granville Tunnicliffe Wilson, Marco Reale and John Haywood Published by CRC Press, 2016. Total number of pages: 323. ISBN 978-1-58488-650-1. Journal of Time Series Analysis, 38(3):505– 507, May 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mcneil:2015:BRD

[Mcn15] Alexander J. Mcneil. Book review: Dependence Modeling with Copulas, by Harry Joe. Monographs on Statistics and Applied probability 134, Published by CRC Press, 2015. Total number of pages: 18 + 462. ISBN: 978-1-4665-8322-1 (Hardback). Journal of Time Series Analysis, 36(4):599–600, July 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Melard:1989:CES

[MdS89] Guy Mélard and Annie Herteleer de Schutter. Contributions to evolutionary spectral theory. Journal of Time Series Analysis, 10(1):41–63, January 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McGee:1998:THC

[ME98] Monnie McGee and Katherine Ensor. Tests for harmonic components in the spectra of categorical time series. Journal of Time Series Analysis, 19(3):309–323, May 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Melard:1985:EES

[Mél85] Guy Mélard. Examples of the evolutionary spectrum theory. Journal of Time Series Analysis, 6(2):81–90, March 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Martin:2005:IBI

[MFM05] Gael M. Martin, Catherine S. Forbes, and Vance L. Martin. Implicit Bayesian inference using option prices. Journal of Time Series Analysis, 26(3):437–462, May 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Meyn:1993:GED

[MG93] Sean P. Meyn and Lei Guo. Geometric ergodicity of a doubly stochastic time series model. Journal of Time Series Analysis, 14 (1):93–108, January 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ma:2000:HRE

[MG00] Yanyuan Ma and Marc G. Genton. Highly robust estimation of the autocovariance function. Journal of Time Series Analysis, 21 (6):663–684, November 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Matilla-Garcia:2010:STT

[MGRM10] Mariano Matilla-García, José Miguel Rodríguez, and Manuel Ruiz Marín. A symbolic test for testing independence between time series. Journal of Time Series Analysis, 31(2):76–85, March 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Michel:2020:LDN

[Mic20] Jon Michel. The limiting distribution of a non-stationary integer valued GARCH(1,1) process. Journal of Time Series Analysis, 41(2):351–356, March 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Milhoj:1984:MEM

[Mil84] Anders Milhøj. Multiplicative exponential models for stationary time series. Journal of Time Series Analysis, 5(1):19–35, January 1984. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Miller:1995:EML

[Mil95] James W. Miller. Exact maximum likelihood estimation in autoregressive processes. Journal of Time Series Analysis, 16(6):607–615, November 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mills:2004:BR

[Mil04] Terence C. Mills. Book review. Journal of Time Series Analysis, 25(2):315–316, March 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mills:2005:BR

[Mil05] Terence C. Mills. Book reviews. Journal of Time Series Analysis, 26(5):782–783, September 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mills:2006:IEU

[Mil06] Terence C. Mills. Introductory econometrics: using Monte Carlo simulation with Microsoft Excel(R). Journal of Time Series Analysis, 27(6):943–944, November 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Miller:2010:CRM

[Mil10] J. Isaac Miller. Cointegrating regressions with messy regressors and an application to mixed-frequency series. Journal of Time Series Analysis, 31(4):255–277, July 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Miller:2019:TCR

[Mil19] J. Isaac Miller. Testing cointegrating relationships using irregular and non-contemporaneous series with an application to paleoclimate data. Journal of Time Series Analysis, 40(6):936–950, November 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Miettinen:2016:SUS

[MIN^{+}16] Jari Miettinen, Katrin Illner, Klaus Nordhausen, Hannu Oja, Sara Taskinen, and Fabian J. Theis. Separation of uncorrelated stationary time series using autocovariance matrices. Journal of Time Series Analysis, 37(3):337–354, May 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McElroy:2012:SIA

[MJ12] Tucker McElroy and Agnieszka Jach. Subsampling inference for the autocovariances and autocorrelations of long-memory heavy-tailed linear time series. Journal of Time Series Analysis, 33(6):935– 953, November 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Martin:1993:EPP

[MK93] Donald E. K. Martin and Benjamin Kedem. Estimation of the period of periodically correlated sequences. Journal of Time Series Analysis, 14(2):193–205, March 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Meyer:2015:VAS

[MK15] Marco Meyer and Jens-Peter Kreiss. On the vector autoregressive sieve bootstrap. Journal of Time Series Analysis, 36(3):377–397, May 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McGonigle:2022:TLS

[MKN22] Euan T. McGonigle, Rebecca Killick, and Matthew A. Nunes. Trend locally stationary wavelet processes. Journal of Time Series Analysis, 43(6):895–917, November 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mainassara:2022:PTC

[MKS22] Yacouba Boubacar Maïnassara, Othman Kadmiri, and Bruno Saussereau. Portmanteau test for a class of multivariate asymmetric power GARCH model. Journal of Time Series Analysis, 43 (6):964–1002, November 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McLeod:1983:DCA

[ML83] A. I. McLeod and W. K. Li. Diagnostic checking ARMA time series models using squared-residual autocorrelations. Journal of Time Series Analysis, 4(4):269–273, July 1983. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McCabe:1997:PAT

[MLS97] B. P. M. McCabe, S. J. Leybourne, and Y. Shin. A parametric approach to testing the null of cointegration. Journal of Time Series Analysis, 18(4):395–413, July 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mikulski:1991:OML

[MM91] Piotr W. Mikulski and Michael J. Monsour. Optimality of the maximum likelihood estimator in first-order autoregressive processes. Journal of Time Series Analysis, 12(3):237–253, May 1991. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McCormick:1993:ENA

[MM93] William P. McCormick and George Mathew. Estimation for nonnegative autoregressive processes with an unknown location parameter. Journal of Time Series Analysis, 14(1):71–92, January 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mahdi:2012:IMP

[MM12] Esam Mahdi and A. Ian McLeod. Improved multivariate portmanteau test. Journal of Time Series Analysis, 33(2):211–222, March 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McAleer:1988:TST

[MMH88] Michael McAleer, C. R. McKenzie, and A. D. Hall. Testing separate time series models. Journal of Time Series Analysis, 9(2): 169–189, March 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Masini:2022:REH

[MMM22] Ricardo P. Masini, Marcelo C. Medeiros, and Eduardo F. Mendes. Regularized estimation of high-dimensional vector autoregressions with weakly dependent innovations. Journal of Time Series Analysis, 43(4):532–557, July 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Miettinen:2020:ECH

[MMNT20] Jari Miettinen, Markus Matilainen, Klaus Nordhausen, and Sara Taskinen. Extracting conditionally heteroskedastic components using independent component analysis. Journal of Time Series Analysis, 41(2):293–311, March 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Meintanis:2024:GFT

[MMOV24] Simos Meintanis, Bojana Milosević, Marko Obradović, and Mirjana Veljović. Goodness-of-fit tests for the multivariate Student-t distribution based on i.i.d. data, and for GARCH observations. Journal of Time Series Analysis, 45(2):298–319, March 2024. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mentz:1998:RVE

[MMT98] Raul P. Mentz, Pedro A. Morettin, and Clélia Toloi. On residual variance estimation in autoregressive models. Journal of Time Series Analysis, 19(2):187–208, March 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McCabe:2005:APD

[MMT05] B. P. M. McCabe, G. M. Martin, and A. R. Tremayne. Assessing persistence in discrete nonstationary time-series models. Journal of Time Series Analysis, 26(2):305–317, March 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Miller:1995:GVR

[MN95] John P. Miller and Paul Newbold. A generalized variance ratio test of ARIMA(p, 1, q) model specification. Journal of Time Series Analysis, 16(4):403–413, July 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Miguel:2002:AFI

[MO02] Jesús Miguel and Pilar Olave. Adjusting forecast intervals in ARCH-M models. Journal of Time Series Analysis, 23(5):587– 598, September 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mokkadem:1987:MAN

[Mok87] Abdelkader Mokkadem. Sur un modéle autorégressif non linéaire, ergodicité et ergodicité géométrique. (French) [On a non-linear autoregressive model, ergodicity and geometric ergodicity]. Journal of Time Series Analysis, 8(2):195–204, March 1987. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Monti:1998:PEP

[Mon98] Anna Clara Monti. A proposal for estimation of the parameters of multivariate moving-average models. Journal of Time Series Analysis, 19(2):209–219, March 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Morettin:1983:NCL

[Mor83] Pedro A. Morettin. A note on a central limit theorem for stationary processes. Journal of Time Series Analysis, 4(1):49–52, January 1983. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Makogin:2021:LRD

[MORS21] Vitalii Makogin, Marco Oesting, Albert Rapp, and Evgeny Spodarev. Long range dependence for stable random processes. Journal of Time Series Analysis, 42(2):161–185, March 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Morris:1984:KFA

[MP84] N. D. Morris and D. Pfeffermann. A Kalman filter approach to the forecasting of monthly time series affected by Morris festivals. Journal of Time Series Analysis, 5(4):255–268, July 1984. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Maravall:1987:PSA

[MP87] Agustin Maravall and David A. Pierce. A prototypical seasonal adjustment model. Journal of Time Series Analysis, 8(2):177–193, March 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mackisack:1990:SPA

[MP90] M. S. Mackisack and D. S. Poskitt. Some properties of autoregressive estimates for processes with mixed spectra. Journal of Time Series Analysis, 11(4):325–337, July 1990. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McMurry:2010:BTE

[MP10] Timothy L. McMurry and Dimitris N. Politis. Banded and tapered estimates for autocovariance matrices and the linear process bootstrap. Journal of Time Series Analysis, 31(6):471–482, November 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Martins:2016:ITF

[MP16] Luis Filipe Martins and Pierre Perron. Improved tests for forecast comparisons in the presence of instabilities. Journal of Time Series Analysis, 37(5):650–659, September 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McMurry:2018:EMP

[MP18] Timothy L. McMurry and Dimitris N. Politis. Estimating MA parameters through factorization of the autocovariance matrix and an MA-sieve bootstrap. Journal of Time Series Analysis, 39(3):

433–446, May 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Melard:1991:CEA

[MPR91] Guy Melard, Marianne Paesmans, and Roch Roy. Consistent estimation of the asymptotic covariance structure of multivariate serial correlations. Journal of Time Series Analysis, 12(4):351–361, July 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Moreno:2012:URB

[MR12] Marta Moreno and Juan Romo. Unit root bootstrap tests under infinite variance. Journal of Time Series Analysis, 33(1):32–47, January 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McElroy:2018:IKL

[MR18] Tucker McElroy and Anindya Roy. The inverse Kullback–Leibler method for fitting vector moving averages. Journal of Time Series Analysis, 39(2):172–191, March 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Montes-Rojas:2019:MQI

[MR19] Gabriel Montes-Rojas. Multivariate quantile impulse response functions. Journal of Time Series Analysis, 40(5):739–752, September 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Moulines:2007:SDW

[MRT07] E. Moulines, F. Roueff, and M. S. Taqqu. On the spectral density of the wavelet coefficients of long-memory time series with application to the log-regression estimation of the memory parameter. Journal of Time Series Analysis, 28(2):155–187, March 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Marriott:1992:RAN

[MS92] J. M. Marriott and A. F. M. Smith. Reparametrization aspects of numerical Bayesian methodology for autoregressive movingaverage models. Journal of Time Series Analysis, 13(4):327–343, July 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Meerschaert:2000:MAR

[MS00a] Mark M. Meerschaert and Hans-Peter Scheffler. Moving averages of random vectors with regularly varying tails. Journal of Time Series Analysis, 21(3):297–328, May 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Moulines:2000:DDO

[MS00b] Eric Moulines and Philippe Soulier. Data driven order selection for projection estimator of the spectral density of time series with long range dependence. Journal of Time Series Analysis, 21(2): 193–218, March 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Meerschaert:2001:SCC

[MS01] Mark M. Meerschaert and Hans-Peter Scheffler. Sample crosscorrelations for moving averages with regularly varying tails. Journal of Time Series Analysis, 22(4):481–492, July 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Metaxoglou:2007:MLE

[MS07] Konstantinos Metaxoglou and Aaron Smith. Maximum likelihood estimation of VARMA models using a state-space EM algorithm. Journal of Time Series Analysis, 28(5):666–685, September 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mallick:2008:GVC

[MS08a] Taslim S. Mallick and Brajendra C. Sutradhar. GQL versus conditional GQL inferences for non-stationary time series of counts with overdispersion. Journal of Time Series Analysis, 29(2):402– 420, March 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Meitz:2008:SNA

[MS08b] Mika Meitz and Pentti Saikkonen. Stability of nonlinear AR-GARCH models. Journal of Time Series Analysis, 29(3):453–475, May 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Merkatas:2023:SIU

[MS23] Christos Merkatas and Simo Särkkä. System identification using autoregressive Bayesian neural networks with nonparametric noise models. Journal of Time Series Analysis, 44(3):319–330, May 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Moeanaddin:1990:NED

[MT90] R. Moeanaddin and Howell Tong. Numerical evaluation of distributions in non-linear autoregression. Journal of Time Series Analysis, 11(1):33–48, January 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McCulloch:1994:BAA

[MT94a] Robert E. McCulloch and Ruey S. Tsay. Bayesian analysis of autoregressive time series via the Gibbs sampler. Journal of Time Series Analysis, 15(2):235–250, March 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McCulloch:1994:SAE

[MT94b] Robert E. McCulloch and Ruey S. Tsay. Statistical analysis of economic time series via Markov switching models. Journal of Time Series Analysis, 15(5):523–539, September 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McElroy:2015:SEN

[MT15] Tucker McElroy and Thomas Trimbur. Signal extraction for nonstationary multivariate time series with illustrations for trend inflation. Journal of Time Series Analysis, 36(2):209–227, March 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Martin:2014:EMM

[MTJ14] Vance L. Martin, Andrew R. Tremayne, and Robert C. Jung. Efficient method of moments estimators for integer time series models. Journal of Time Series Analysis, 35(6):491–516, November 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Morton:2004:CMH

[MTW04] Alex S. Morton and Granville Tunnicliffe-Wilson. A class of modified high-order autoregressive models with improved resolution of low-frequency cycles. Journal of Time Series Analysis, 25(2):235– 250, March 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mukherjee:1999:AQR

[Muk99] Kanchan Mukherjee. Asymptotics of quantiles and rank scores in nonlinear time series. Journal of Time Series Analysis, 20(2):

173–192, March 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Murthy:1985:FOA

[Mur85] D. N. P. Murthy. First order auto-regressive model parameter estimation with periodic observations. Journal of Time Series Analysis, 6(2):91–95, March 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Medeiros:2003:DCF

[MV03] Marcelo C. Medeiros and Alvaro Veiga. Diagnostic checking in a flexible nonlinear time series model. Journal of Time Series Analysis, 24(4):461–482, July 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Moore:1987:EBS

[MVS87] Mike I. Moore, Andy W. Visser, and Tim G. L. Shirtcliffe. Experiences with the Brillinger spectral estimator applied to simulated irregularly observed processes. Journal of Time Series Analysis, 8 (4):433–442, July 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Muller:1997:ILS

[MW97] Daniel Muller and William W. S. Wei. Iterative least squares estimation and identification of the transfer function model. Journal of Time Series Analysis, 18(6):579–592, November 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mena:2005:SAM

[MW05] Ramsés H. Mena and Stephen G. Walker. Stationary autoregressive models via a Bayesian nonparametric approach. Journal of Time Series Analysis, 26(6):789–805, November 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Miller:2016:IRB

[MW16] J. Isaac Miller and Xi Wang. Implementing residual-based KPSS tests for cointegration with data subject to temporal aggregation and mixed sampling frequencies. Journal of Time Series Analysis, 37(6):810–824, November 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mentemeier:2022:AIE

[MW22] Sebastian Mentemeier and Olivier Wintenberger. Asymptotic independence ex machina: Extreme value theory for the diagonal SRE model. Journal of Time Series Analysis, 43(5):750–780, September 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wong:1996:BAE

[mWK96] Chi ming Wong and Robert Kohn. A Bayesian approach to estimating and forecasting additive nonparametric autoregressive models. Journal of Time Series Analysis, 17(2):203–220, March 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Muller:1997:BMN

[MWM97] Peter Müller, Mike West, and Steven MacEachern. Bayesian models for non-linear autoregressions. Journal of Time Series Analysis, 18(6):593–614, November 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Muler:2002:REA

[MY02] Nora Muler and Victor J. Yohai. Robust estimates for arch processes. Journal of Time Series Analysis, 23(3):341–375, May 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Matsuda:2004:TSC

[MY04] Yasumasa Matsuda and Yoshihiro Yajima. On testing for separable correlations of multivariate time series. Journal of Time Series Analysis, 25(4):501–528, July 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

McLeod:2006:PAP

[MZ06] A. I. McLeod and Y. Zhang. Partial autocorrelation parameterization for subset autoregression. Journal of Time Series Analysis, 27(4):599–612, July 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Mielniczuk:2009:NPL

[MZW09] Jan Mielniczuk, Zhou Zhou, and Wei Biao Wu. On nonparametric prediction of linear processes. Journal of Time Series Analysis, 30 (6):652–673, November 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nagahara:2003:NGF

[Nag03] Yuichi Nagahara. Non-Gaussian filter and smoother based on the Pearson distribution system. Journal of Time Series Analysis, 24 (6):721–738, November 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nam:2012:QUC

[NAJ12] Christopher F. H. Nam, John A. D. Aston, and Adam M. Johansen. Quantifying the uncertainty in change points. Journal of Time Series Analysis, 33(5):807–823, September 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nanamiya:2014:MWC

[Nan14] Kei Nanamiya. Modelling for the wavelet coefficients of ARFIMA processes. Journal of Time Series Analysis, 35(4):341–356, July 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nassiuma:1993:NSA

[Nas93] Dankit Nassiuma. Non-stationary autoregressive moving-average processes with infinite variance. Journal of Time Series Analysis, 14(3):297–304, May 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nassiuma:1994:SSS

[Nas94] Dankit K. Nassiuma. Symmetric stable sequences with missing observations. Journal of Time Series Analysis, 15(3):313–323, May 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Newbold:1983:CPC

[NB83] P. Newbold and T. Bos. On q-conditioned partial correlations. Journal of Time Series Analysis, 4(1):53–55, January 1983. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nieto-Barajas:2016:BNP

[NBQ16] Luis E. Nieto-Barajas and Fernando A. Quintana. A Bayesian nonparametric dynamic AR model for multiple time series analysis. Journal of Time Series Analysis, 37(5):675–689, September 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Neal:2013:BRB

[Nea13] Peter Neal. Book review: Bayesian Theory and Applications, by Paul Damien, Petros Dellaportas, Nicholas G. Polson and David A. Stephens (eds). Published by Oxford University Press, 2013. Total Number of Pages: xiii + 702. ISBN 978-0-19-969560-7. Journal of Time Series Analysis, 34(6):744, November 2013. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

NesLehova:2016:BRQ

[Neš16] Johanna G. NešLehová. Book review: Quantitative Risk Management: Concepts, Techniques and Tools, by Alexander J. Mc-Neil, Rüdiger Frey and Paul Embrechts. Revised edition. Published by Princeton University Press, 2015. Total number of pages:

720. ISBN: 978-0-691-16627-8 (Hardback). Journal of Time Series Analysis, 37(3):431–432, May 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Neumann:1996:SDE

[Neu96] Michael H. Neumann. Spectral density estimation via nonlinear wavelet methods for stationary non-Gaussian time series. Journal of Time Series Analysis, 17(6):601–633, November 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Newbold:1980:NRB

[New80] Paul Newbold. A note on relations between seasonally adjusted variables. Journal of Time Series Analysis, 1(1):31–35, January 1980. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nielsen:2019:SIJ

[NH19] Morten Ørregaard Nielsen and Javier Hualde. Special issue of the Journal of Time Series Analysis in honour of the 35th anniversary of the publication of Geweke and Porter-Hudak (1983): Guest Editors’ introduction. Journal of Time Series Analysis, 40(4):386–387, July 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nakamori:2008:DQE

[NHCLP08] Seiichi Nakamori, Aurora Hermoso-Carazo, and Josefa Linares-Pérez. Design of quadratic estimators using covariance information in linear discrete-time stochastic systems. Journal of Time Series Analysis, 29(3):501–512, May 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nielsen:2005:SET

[Nie05] Morten Ørregaard Nielsen. Semiparametric estimation in timeseries regression with long-range dependence. Journal of Time Series Analysis, 26(2):279–304, March 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nielsen:2011:LWE

[Nie11] Frank S. Nielsen. Local Whittle estimation of multi-variate fractionally integrated processes. Journal of Time Series Analysis, 32(3):317–335, May 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nielsen:2015:ACS

[Nie15] Morten Ørregaard Nielsen. Asymptotics for the conditional-sum-of-squares estimator in multivariate fractional time-series models. Journal of Time Series Analysis, 36(2):154–188, March 2015. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nishiyama:2009:GFT

[Nis09] Yoichi Nishiyama. Goodness-of-fit test for a nonlinear time series. Journal of Time Series Analysis, 30(6):674–681, November 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nishi:2024:SLM

[NK24] Mikihito Nishi and Eiji Kurozumi. Stochastic local and moderate departures from a unit root and its application to unit root testing. Journal of Time Series Analysis, 45(1):133–157, January 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nankervis:2015:GVR

[NKC15] John C. Nankervis, Periklis Kougoulis, and Jerry Coakley. Generalized variance-ratio tests in the presence of statistical dependence. Journal of Time Series Analysis, 36(5):687–705, September 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Na:2012:CPD

[NLL12] Okyoung Na, Jiyeon Lee, and Sangyeol Lee. Change point detection in copula ARMA–GARCH models. Journal of Time Series Analysis, 33(4):554–569, July 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nastic:2016:REI

[NLR16] Aleksandar S. Nastić, Petra N. Laketa, and Miroslav M. Ristić. Random environment integer-valued autoregressive process. Journal of Time Series Analysis, 37(2):267–287, March 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Narukawa:2011:BSP

[NM11] Masaki Narukawa and Yasumasa Matsuda. Broadband semiparametric estimation of long-memory time series by fractional exponential models. Journal of Time Series Analysis, 32(2):175–193, March 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nielsen:2021:IBE

[NN21] Morten Ørregaard Nielsen and Antoine L. Noël. To infinity and beyond: Efficient computation of ARCH(∞) models. Journal of Time Series Analysis, 42(3):338–354, May 2021. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Newton:1991:FGI

[NNC91] H. Joseph Newton, Gerald R. North, and Thomas J. Crowley. Forecasting global ice volume. Journal of Time Series Analysis, 12(3): 255–265, May 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nijman:1990:PIA

[NP90] Theo Nijman and Franz Palm. Parameter identification in ARMA processes in the presence of regular but incomplete sampling. Journal of Time Series Analysis, 11(3):239–248, May 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ng:1996:EEE

[NP96] Serena Ng and Pierre Perron. The exact error in estimating the spectral density at the origin. Journal of Time Series Analysis, 17(4):379–408, July 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nicholls:1980:ERC

[NQ80] D. F. Nicholls and B. G. Quinn. The estimation of random coefficient autoregressive models. I. Journal of Time Series Analysis, 1(1):37–46, January 1980. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nsiri:1993:IML

[NR93] Saïd Nsiri and Roch Roy. On the invertibility of multivariate linear processes. Journal of Time Series Analysis, 14(3):305–316, May 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Neal:2007:MIV

[NR07] Peter Neal and T. Subba Rao. MCMC for integer-valued ARMA processes. Journal of Time Series Analysis, 28(1):92–110, January 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nunes:2011:LTT

[NR11] Luis C. Nunes and Paulo M. M. Rodrigues. On LM-type tests for seasonal unit roots in the presence of a break in trend. Journal of Time Series Analysis, 32(2):108–134, March 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nielsen:2003:LAF

[NS03] B. Nielsen and N. Shephard. Likelihood analysis of a first-order autoregressive model with exponential innovations. Journal of Time Series Analysis, 24(3):337–344, May 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Neumeyer:2013:NNP

[NS13] Natalie Neumeyer and Leonie Selk. A note on non-parametric testing for Gaussian innovations in AR–ARCH models. Journal of Time Series Analysis, 34(3):362–367, May 2013. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nichols:2011:APP

[NSK^{+}11] Kevin Nichols, Frederic Paik Schoenberg, Jon E. Keeley, Andrew Bray, and David Diez. The application of prototype point processes for the summary and description of California wildfires. Journal of Time Series Analysis, 32(4):420–429, July 2011. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nordman:2007:ELC

[NSL07] Daniel J. Nordman, Philipp Sibbertsen, and Soumendra N. Lahiri. Empirical likelihood confidence intervals for the mean of a longrange dependent process. Journal of Time Series Analysis, 28(4): 576–599, July 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nunes:2020:BRT

[Nun20] Matthew Nunes. Book review: Time Series: a Data Analysis Approach Using R By Robert H. Shumway and David S. Stoffer. Published by Taylor and Francis Group, LLC, Boca Raton, London, New York, 2019. ISBN: 978-0-367-22109-6 (Hardback). Journal of Time Series Analysis, 41(3):485–486, May 2020. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Newbold:1996:BNT

[NV96] Paul Newbold and Dimitrios Vougas. Beveridge–Nelson-type trends for I(2) and some seasonal models. Journal of Time Series Analysis, 17(2):151–169, March 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Noriega:2006:SRU

[NVS06] Antonio E. Noriega and Daniel Ventosa-Santaulària. Spurious regression under broken-trend stationarity. Journal of Time Series Analysis, 27(5):671–684, September 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

OBrien:1987:TNL

[O’B87] C. O’Brien. A test for non-linearity of prediction in time series. Journal of Time Series Analysis, 8(3):313–327, May 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ochi:1983:AED

[Och83] Yoshimichi Ochi. Asymptotic expansions for the distribution of an estimator in the first-order autoregressive process. Journal of Time Series Analysis, 4(1):57–67, January 1983. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Odolphin:2003:DTS

[OJ03] E. J. G. Odolphin and S. E. Johnson. Decomposition of time series dynamic linear models. Journal of Time Series Analysis, 24(5): 513–527, September 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ozaki:2000:RLF

[OJHO00] T. Ozaki, J. C. Jimenez, and V. Haggan-Ozaki. The role of the likelihood function in the estimation of chaos models. Journal of Time Series Analysis, 21(4):363–387, July 2000. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Oke:1998:SRS

[Oke98] Thimothy Oke. Some results on specification search and pre-testing in an ARMA(1, 1) process. Journal of Time Series Analysis, 19 (6):709–722, November 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Olmo:2023:NPR

[Olm23] Jose Olmo. A nonparametric predictive regression model using partitioning estimators based on Taylor expansions. Journal of Time Series Analysis, 44(3):294–318, May 2023. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ouakasse:2017:NRE

[OM17] Abdelhamid Ouakasse and Guy Mélard. A new recursive estimation method for single input single output models. Journal of Time Series Analysis, 38(3):417–457, May 2017. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ombao:2013:BRT

[Omb13] Hernando Ombao. Book review: Time series modeling of neuroscience data, by Tohru Ozaki, published by CRC Press, 2012. Total number of pages: 548. Price: US $71.46. ISBN 978-1-4200-9460-2. Journal of Time Series Analysis, 34(6):745–746, November 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Oya:1998:DFL

[OT98] Kosuke Oya and Hiro Toda. Dickey–Fuller, Lagrange multiplier and combined tests for a unit root in autoregressive time series. Journal of Time Series Analysis, 19(3):325–347, May 1998. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Otter:1988:SDM

[Ott88] Pieter W. Otter. Structural, dynamic modelling in unobservable spaces of covariance-stationary stochastic processes. Journal of Time Series Analysis, 9(1):59–72, January 1988. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Otto:2021:URT

[Ott21] Sven Otto. Unit root testing with slowly varying trends. Journal of Time Series Analysis, 42(1):85–106, January 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Oppenheim:2004:ARP

[OV04] Georges Oppenheim and Marie-Claude Viano. Aggregation of random parameters Ornstein–Uhlenbeck or AR processes: some convergence results. Journal of Time Series Analysis, 25(3):335–350, May 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ozaki:1982:SAP

[Oza82] T. Ozaki. The statistical analysis of perturbed limit cycle processes using nonlinear time series models. Journal of Time Series Analysis, 3(1):29–41, January 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Poignard:2023:HDS

[PA23] Benjamin Poignard and Manabu Asai. High-dimensional sparse multivariate stochastic volatility models. Journal of Time Series Analysis, 44(1):4–22, January 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Paparoditis:1994:VAG

[Pap94] Efstathios Paparoditis. On vector autocorrelations and generalized second-order functions for time series. Journal of Time Series Analysis, 15(3):325–334, May 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Paparoditis:2005:TFV

[Pap05] Efstathios Paparoditis. Testing the fit of a vector autoregressive moving average model. Journal of Time Series Analysis, 26(4): 543–568, July 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Paparoditis:2022:RBS

[Pap22] Efstathios Paparoditis. Review of the book Stochastic Models for Time Series by Paul Doukhan. Journal of Time Series Analysis, 43 (1):154, January 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Parnell:2013:CTS

[Par13] Andrew C. Parnell. Climate time series analysis: classical statistical and bootstrap methods. Journal of Time Series Analysis, 34 (2):281, March 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Paulsen:1984:ODM

[Pau84] Jostein Paulsen. Order determination of multivariate autoregressive time series with unit roots. Journal of Time Series Analysis, 5(2):115–127, March 1984. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Piancastelli:2023:FBI

[PBSO23] Luiza S. C. Piancastelli, Wagner Barreto-Souza, and Hernando Ombao. Flexible bivariate INGARCH process with a broad range of contemporaneous correlation. Journal of Time Series Analysis, 44(2):206–222, March 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Penm:2000:RAS

[PBT00] J. H. W. Penm, T. J. Brailsford, and R. D. Terrell. A robust algorithm in sequentially selecting subset time series systems using neural networks. Journal of Time Series Analysis, 21(4):389–412, July 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Palma:2005:EES

[PC05] Wilfredo Palma and Ngai Hang Chan. Efficient estimation of seasonal long-range-dependent processes. Journal of Time Series Analysis, 26(6):863–892, November 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pavlopoulos:2023:HSF

[PC23] Harry Pavlopoulos and George Chronis. On highly skewed fractional log-stable noise sequences and their application. Journal of Time Series Analysis, 44(4):337–358, July 2023. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Planas:2002:CRA

[PD02] Christophe Planas and Raoul Depoutot. Controlling revisions in ARIMA-model-based seasonal adjustment. Journal of Time Series Analysis, 23(2):193–213, March 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Paraschakis:2012:FPE

[PD12] Konstantinos Paraschakis and Rainer Dahlhaus. Frequency and phase estimation in time series with quasi periodic components. Journal of Time Series Analysis, 33(1):13–31, January 2012. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pemberton:1987:ELS

[Pem87] John. Pemberton. Exact least squares multi-step prediction from nonlinear autoregressive models. Journal of Time Series Analysis, 8(4):443–448, July 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pena:1984:AFS

[Peñ84] Daniel Peña. The autocorrelation function of seasonal ARMA models. Journal of Time Series Analysis, 5(4):269–272, July 1984. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Penzer:2007:SSM

[Pen07] Jeremy Penzer. State space models for time series with patches of unusual observations. Journal of Time Series Analysis, 28(5):629– 645, September 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Perera:2004:MQL

[Per04] S. Perera. Maximum quasi-likelihood estimation for the NEAR(2) model. Journal of Time Series Analysis, 25(5):723–732, September 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pesavento:2007:RBT

[Pes07] Elena Pesavento. Residuals-based tests for the null of nocointegration: an analytical comparison. Journal of Time Series Analysis, 28(1):111–137, January 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Petruccelli:1986:CLS

[Pet86] Joseph D. Petruccelli. On the consistency of least squares estimators for a threshold AR(1) model. Journal of Time Series Analysis, 7(4):269–278, July 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Petruccelli:1989:APN

[Pet89] Joseph D. Petruccelli. Autoregressive processes with normal stationary distributions. Journal of Time Series Analysis, 10(1):65– 70, January 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Petrova:2019:QBE

[Pet19] Katerina Petrova. Quasi-Bayesian estimation of time-varying volatility in DSGE models. Journal of Time Series Analysis, 40 (1):151–157, January 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic). See correction [Ano21a].

Park:1995:AEU

[PF95] Heon Jin Park and Wayne A. Fuller. Alternative estimators and unit root tests for the autoregressive process. Journal of Time Series Analysis, 16(4):415–429, July 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pizzinga:2021:EIP

[PF21] Adrian Pizzinga and Marcelo Fernandes. Extensions to the invariance property of maximum likelihood estimation for affinetransformed state-space models. Journal of Time Series Analysis, 42(3):355–371, May 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pfeffermann:1994:GME

[Pfe94] D. Pfeffermann. A general method for estimating the variances of X-11 seasonally adjusted estimators. Journal of Time Series Analysis, 15(1):85–116, January 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Prado:2002:TVA

[PH02] Raquel Prado and Gabriel Huerta. Time-varying autoregressions with model order uncertainty. Journal of Time Series Analysis, 23 (5):599–618, September 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pena:2022:RBU

[PI22] Víctor Peña and Kaoru Irie. On the relationship between Uhlig extended and beta-Bartlett processes. Journal of Time Series Analysis, 43(1):147–153, January 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Piccolo:1982:SSI

[Pic82] Domenico Piccolo. The size of the stationarity and invertibility region of an autoregressive-moving average process. Journal of Time Series Analysis, 3(4):245–247, July 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Piccolo:1990:DMC

[Pic90] Domenico Piccolo. A distance measure for classifying ARIMA models. Journal of Time Series Analysis, 11(2):153–164, March 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pedeli:2013:CLE

[PK13] Xanthi Pedeli and Dimitris Karlis. On composite likelihood estimation of a multivariate INAR(1) model. Journal of Time Series Analysis, 34(2):206–220, March 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pascalau:2022:JTC

[PLNL22] Razvan Pascalau, Junsoo Lee, Saban Nazlioglu, and Yan (Olivia) Lu. Johansen-type cointegration tests with a Fourier function. Journal of Time Series Analysis, 43(5):828–852, September 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pourahmadi:1992:CCC

[PM92] Mohsen Pourahmadi and A. G. Miamee. Computation of canonical correlation between past and future of a time series. Journal of Time Series Analysis, 13(4):345–351, July 1992. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Politis:1994:MEP

[Pol94] Dimitris N. Politis. On the maximum entropy property of nonlinear autoregressions. Journal of Time Series Analysis, 15(5):541–543, September 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pons:2006:TMS

[Pon06] Gabriel Pons. Testing monthly seasonal unit roots with monthly and quarterly information. Journal of Time Series Analysis, 27(2): 191–209, March 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pope:1990:BEM

[Pop90] Alun Lloyd Pope. Biases of estimators in multivariate non-Gaussian autoregressions. Journal of Time Series Analysis, 11(3): 249–258, May 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Porat:1987:SAP

[Por87] Boaz Porat. Some asymptotic properties of the sample covariances of Gaussian autoregressive moving-average processes. Journal of Time Series Analysis, 8(2):205–220, March 1987. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Poskitt:2005:NSE

[Pos05] D. S. Poskitt. A note on the specification and estimation of AR-MAX systems. Journal of Time Series Analysis, 26(2):157–183, March 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Poskitt:2008:PSB

[Pos08] D. S. Poskitt. Properties of the sieve bootstrap for fractionally integrated and non-invertible processes. Journal of Time Series Analysis, 29(2):224–250, March 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Poskitt:2020:SSA

[Pos20] Donald S. Poskitt. On singular spectrum analysis and stepwise time series reconstruction. Journal of Time Series Analysis, 41(1): 67–94, January 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Potscher:1990:EAM

[Pöt90] B. M. Pötscher. Estimation of autoregressive moving-average order given an infinite number of models and approximation of spectral densities. Journal of Time Series Analysis, 11(2):165–179, March 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pourahmadi:1986:SSD

[Pou86] Mohsen Pourahmadi. On stationarity of the solution of a doubly stochastic model. Journal of Time Series Analysis, 7(2):123–131, March 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pourahmadi:1988:SSA

[Pou88] Mohsen Pourahmadi. Stationarity of the solution of X_{t} = A_{t}X_{t−1} + ϵ_{t} and analysis of non-Gaussian dependent random variables. Journal of Time Series Analysis, 9(3):225–239, May 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pourahmadi:1989:EIM

[Pou89] Mohsen Pourahmadi. Estimation and interpolation of missing values of a stationary time series. Journal of Time Series Analysis, 10(2):149–169, March 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pourahmadi:2005:BR

[Pou05] Mohsen Pourahmadi. Book reviews. Journal of Time Series Analysis, 26(5):784–785, September 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pourahmadi:2016:BRT

[Pou16] Mohsen Pourahmadi. Book review: Time Series Modelling with Unobserved Components, by Matteo M. Pelagatti. Published by CRC Press, 2015, pages: 257. ISBN-13: 978-1-4822-2500-6. Journal of Time Series Analysis, 37(4):575–576, July 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pourahmadi:2017:BRS

[Pou17] Mohsen Pourahmadi. Book review: State-Space Methods for Time Series Analysis: Theory, Applications and Software, by Jose Casals, Alfredo Garcia-Hiernaux, Miguel Jerez, Sonia Sotoca, and A. Alexandre Trindade. Published by CRC Press, 2016. Total number of pages: 270. ISBN: 1-4822-1959-X. Journal of Time Series Analysis, 38(5):806, September 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Peiris:1988:PFD

[PP88] M. S. Peiris and B. J. C. Perera. On prediction with fractionally differenced ARIMA models. Journal of Time Series Analysis, 9(3): 215–220, May 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Poggi:1997:TLF

[PP97] Jean-Michel Poggi and Bruno Portier. A test of linearity for functional autoregressive models. Journal of Time Series Analysis, 18 (6):615–639, November 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Paparoditis:1999:LBP

[PP99] Efstathios Paparoditis and Dimitris N. Politis. The local bootstrap for periodogram statistics. Journal of Time Series Analysis, 20(2): 193–222, March 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Paparoditis:2012:NSD

[PP12] Efstathios Paparoditis and Dimitris N. Politis. Nonlinear spectral density estimation: thresholding the correlogram. Journal of Time Series Analysis, 33(3):386–397, May 2012. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Paparoditis:2016:NBN

[PP16a] Efstathios Paparoditis and Dimitris N. Politis. A note on the behaviour of nonparametric density and spectral density estimators at zero points of their support. Journal of Time Series Analysis, 37(2):182–194, March 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Puchstein:2016:TSM

[PP16b] Ruprecht Puchstein and Philip Preuß. Testing for stationarity in multivariate locally stationary processes. Journal of Time Series Analysis, 37(1):3–29, January 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Philippe:2014:CAT

[PPS14] Anne Philippe, Donata Puplinskaite, and Donatas Surgailis. Contemporaneous aggregation of triangular array of random-coefficient AR(1) processes. Journal of Time Series Analysis, 35(1):16–39, January 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pilavakis:2020:TEA

[PPS20] Dimitrios Pilavakis, Efstathios Paparoditis, and Theofanis Sapatinas. Testing equality of autocovariance operators for functional time series. Journal of Time Series Analysis, 41(4):571–589, July 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Penm:1993:RFS

[PPT93] Jack H. W. Penm, Jammie H. Penm, and R. D. Terrell. The recursive fitting of subset VARX models. Journal of Time Series Analysis, 14(6):603–619, November 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Potscher:1988:DBT

[PR88] B. M. Pötscher and E. Reschenhofer. Discriminating between two spectral densities in case of replicated observations. Journal of Time Series Analysis, 9(3):221–224, May 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Politis:1995:BCN

[PR95] Dimitris N. Politis and Joseph P. Romano. Bias-corrected nonparametric spectral estimation. Journal of Time Series Analysis, 16(1):67–103, January 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pai:1998:BAA

[PR98] Jeffrey S. Pai and Nalini Ravishanker. Bayesian analysis of autoregressive fractionally integrated moving-average processes. Journal of Time Series Analysis, 19(1):99–112, January 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Perron:2003:SAO

[PR03] Pierre Perron and Gabriel Rodríguez. Searching for additive outliers in nonstationary time series. Journal of Time Series Analysis, 24(2):193–220, March 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Proietti:2009:TSA

[PR09] Tommaso Proietti and Marco Riani. Transformations and seasonal adjustment. Journal of Time Series Analysis, 30(1):47–69, January 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ploberger:2010:TCM

[PR10] Werner Ploberger and Erhard Reschenhofer. Testing for cycles in multiple time series. Journal of Time Series Analysis, 31(6):427– 434, November 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Praetz:1982:MMC

[Pra82] Peter Praetz. The market model, CAPM and efficiency in the frequency domain. Journal of Time Series Analysis, 3(1):61–79, January 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pham:2003:TNC

[PRC03] Dinh Tuan Pham, Roch Roy, and Lyne Cédras. Tests for noncorrelation of two cointegrated ARMA time series. Journal of Time Series Analysis, 24(5):553–577, September 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Prewitt:1998:GFT

[Pre98] Kathryn Prewitt. Goodness-of-fit test in parametric time series models. Journal of Time Series Analysis, 19(5):549–574, September 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Priestley:1980:SDM

[Pri80] M. B. Priestley. State-dependent models: a general approach to non-linear time series analysis. Journal of Time Series Analysis, 1(1):47–71, January 1980. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Priestley:1994:PEJ

[Pri94] M. B. Priestley. Professor Edward James Hannan (1921–1994). Journal of Time Series Analysis, 15(2):234, March 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Priestley:1996:WTD

[Pri96] M. B. Priestley. Wavelets and time-dependent spectral analysis. Journal of Time Series Analysis, 17(1):85–103, January 1996. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Priestley:2004:BR

[Pri04] M. B. Priestley. Book reviews. Journal of Time Series Analysis, 25 (1):157, January 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Proietti:2003:LDS

[Pro03] Tommaso Proietti. Leave-k-out diagnostics in state-space models. Journal of Time Series Analysis, 24(2):221–236, March 2003. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Proietti:2023:PGT

[Pro23] Tommaso Proietti. Peaks, gaps, and time-reversibility of economic time series. Journal of Time Series Analysis, 44(1):43–68, January 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pascual:2004:BPI

[PRR04] Lorenzo Pascual, Juan Romo, and Esther Ruiz. Bootstrap predictive inference for ARIMA processes. Journal of Time Series Analysis, 25(4):449–465, July 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Politis:2004:IAP

[PRW04] Dimitris N. Politis, Joseph P. Romano, and Michael Wolf. Inference for autocorrelations in the possible presence of a unit root. Journal of Time Series Analysis, 25(2):251–263, March 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pawitan:1989:SED

[PS89] Yudianto Pawitan and R. H. Shumway. Spectral estimation and deconvolution for a linear time series model. Journal of Time Series Analysis, 10(2):115–129, March 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Paparoditis:1992:OIS

[PS92] Efstathios Paparoditis and Bernd Streitberg. Order identification statistics in stationary autoregressive moving-average models:vector autocorrelations and the bootstrap. Journal of Time Series Analysis, 13(5):415–434, September 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Poskitt:1995:RBG

[PS95] D. S. Poskitt and M. O. Salau. On the relationship between generalized least squares and Gaussian estimation of vector ARMA models. Journal of Time Series Analysis, 16(6):617–645, November 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pitt:1999:ACR

[PS99] Michael K. Pitt and Neil Shephard. Analytic convergence rates and parameterization issues for the Gibbs sampler applied to state space models. Journal of Time Series Analysis, 20(1):63–85, January 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pourahmadi:2000:PVI

[PS00] Mohsen Pourahmadi and E. S. Soofi. Prediction variance and information worth of observations in time series. Journal of Time Series Analysis, 21(4):413–434, July 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pawlak:2001:DPG

[PS01] M. Pawlak and W. Schmid. On the distributional properties of GARCH processes. Journal of Time Series Analysis, 22(3):339– 352, May 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Psaradakis:2003:DNR

[PS03] Zacharias Psaradakis and Nicola Spagnolo. On the determination of the number of regimes in Markov-switching autoregressive models. Journal of Time Series Analysis, 24(2):237–252, March 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Park:2006:SFS

[PS06a] Heungsun Park and Key-Il Shin. A shrinked forecast in stationary processes favouring percentage error. Journal of Time Series Analysis, 27(1):129–139, January 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Psaradakis:2006:JDS

[PS06b] Zacharias Psaradakis and Nicola Spagnolo. Joint determination of the state dimension and autoregressive order for models with Markov regime switching. Journal of Time Series Analysis, 27 (5):753–766, September 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pena:2007:MAM

[PS07] Daniel Peña and Ismael Sánchez. Measuring the advantages of multivariate vs. univariate forecasts. Journal of Time Series Analysis, 28(6):886–909, November 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Parente:2021:QML

[PS21] Paulo M. D. C. Parente and Richard J. Smith. Quasi-maximum likelihood and the kernel block bootstrap for nonlinear dynamic models. Journal of Time Series Analysis, 42(4):377–405, July 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Psaradakis:2001:BTA

[Psa01] Zacharias Psaradakis. Bootstrap tests for an autoregressive unit root in the presence of weakly dependent errors. Journal of Time Series Analysis, 22(5):577–594, September 2001. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Psaradakis:2008:ATR

[Psa08] Zacharias Psaradakis. Assessing time-reversibility under minimal assumptions. Journal of Time Series Analysis, 29(5):881–905, September 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Psaradakis:2009:SNT

[PSSS09] Zacharias Psaradakis, Martin Sola, Fabio Spagnolo, and Nicola Spagnolo. Selecting nonlinear time series models using information criteria. Journal of Time Series Analysis, 30(4):369–394, July 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Palm:2008:BUR

[PSU08] Franz C. Palm, Stephan Smeekes, and Jean-Pierre Urbain. Bootstrap unit-root tests: comparison and extensions. Journal of Time Series Analysis, 29(2):371–401, March 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pemberton:1981:NDN

[PT81a] J. Pemberton and H. Tong. A note on the distributions of nonlinear autoregressive stochastic models. Journal of Time Series Analysis, 2(1):49–52, January 1981. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Poskitt:1981:TSA

[PT81b] D. S. Poskitt and A. R. Tremayne. A time series application of the use of Monte Carlo methods to compare statistical tests. Journal of Time Series Analysis, 2(4):263–277, July 1981. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Penm:1982:RFS

[PT82] Jack H. W. Penm and R. D. Terrell. On the recursive fitting of subset autoregressions. Journal of Time Series Analysis, 3(1):43– 59, January 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Poskitt:1986:SAP

[PT86] D. S. Poskitt and A. R. Tremayne. Some aspects of the performance of diagnostic checks in bivariate time series models. Journal of Time Series Analysis, 7(3):217–233, May 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pipiras:2002:DFB

[PT02] Vladas Pipiras and Murad S. Taqqu. Deconvolution of fractional Brownian motion. Journal of Time Series Analysis, 23(4):487–501, July 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Peiris:2004:NFS

[PT04] S. Peiris and A. Thavaneswaran. A note on the filtering for some time series models. Journal of Time Series Analysis, 25(3):397– 407, May 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pfeffermann:2005:BAP

[PT05] Danny Pfeffermann and Richard Tiller. Bootstrap approximation to prediction MSE for state-space models with estimated parameters. Journal of Time Series Analysis, 26(6):893–916, November 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pena:2023:TAC

[PT23] Daniel Peña and Ruey S. Tsay. A testing approach to clustering scalar time series. Journal of Time Series Analysis, 44(5-6):667– 685, September 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Poggi:1998:EFI

[PV98] John-Michel Poggi and Marie-Claude Viano. An estimate of the fractal index using multiscale aggregates. Journal of Time Series Analysis, 19(2):221–233, March 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Psaradakis:2015:QBT

[PV15] Zacharias Psaradakis and Marián Vávra. A quantile-based test for symmetry of weakly dependent processes. Journal of Time Series Analysis, 36(4):587–598, July 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Piccolo:1984:UAA

[PW84] D. Piccolo and G. Tunnicliffe Wilson. A unified approach to ARMA model identification and preliminary estimation. Journal of Time Series Analysis, 5(3):183–204, May 1984. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pole:1989:RAD

[PW89] Andy Pole and Mike West. Reference analysis of the dynamic linear model. Journal of Time Series Analysis, 10(2):131–147, March 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Preminger:2005:UPL

[PW05] Arie Preminger and David Wettstein. Using the penalized likelihood method for model selection with nuisance parameters present only under the alternative: an application to switching regression models. Journal of Time Series Analysis, 26(5):715–741, September 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Phillips:2006:IAU

[PX06] Peter C. B. Phillips and Ke-Li Xu. Inference in autoregression under heteroskedasticity. Journal of Time Series Analysis, 27(2): 289–308, March 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Perron:2022:SCT

[PY22] Pierre Perron and Yohei Yamamoto. Structural change tests under heteroskedasticity: Joint estimation versus two-steps methods. Journal of Time Series Analysis, 43(3):389–411, May 2022. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Palma:2004:ACS

[PZ04] Wilfredo Palma and Mauricio Zevallos. Analysis of the correlation structure of square time series. Journal of Time Series Analysis, 25(4):529–550, July 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Perron:2017:TSM

[PZ17] Pierre Perron and Eduardo Zorita. Time series methods applied to climate change. Journal of Time Series Analysis, 38(5):639, September 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pang:2014:AIM

[PZC14] Tianxiao Pang, Danna Zhang, and Terence Tai-Leung Chong. asymptotic inferences for an AR(1) model with a change point: stationary and nearly non-stationary cases. Journal of Time Series Analysis, 35(2):133–150, March 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Qian:2014:FSS

[Qia14] Hang Qian. A flexible state space model and its applications. Journal of Time Series Analysis, 35(2):79–88, March 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Quinn:1981:ERC

[QN81] B. G. Quinn and D. F. Nicholls. The estimation of random coefficient autogressive models. II. Journal of Time Series Analysis, 2(3):185–203, May 1981. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Quinn:1982:TRA

[QN82] B. G. Quinn and D. F. Nicholls. Testing for the randomness of autoregressive coefficients. Journal of Time Series Analysis, 3(2): 123–135, March 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Qiou:1998:BIT

[QR98] Zuqiang Qiou and Nalini Ravishanker. Bayesian inference for time series with stable innovations. Journal of Time Series Analysis, 19(2):235–249, March 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Quenneville:2000:BPM

[QS00] Benoit Quenneville and Avinash C. Singh. Bayesian prediction mean squared error for state space models with estimated parameters. Journal of Time Series Analysis, 21(2):219–236, March 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Quinn:1982:NES

[Qui82] B. G. Quinn. A note on the existence of strictly stationary solutions to bilinear equations. Journal of Time Series Analysis, 3(4):249– 252, July 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Quinn:1988:NAO

[Qui88] B. G. Quinn. A note on AIC order determination for multivariate autoregressions. Journal of Time Series Analysis, 9(3):241–245, May 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Quinn:1989:ENT

[Qui89] B. G. Quinn. Estimating the number of terms in a sinusoidal regression. Journal of Time Series Analysis, 10(1):71–75, January 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Quinn:2000:KFE

[Qui00] B. G. Quinn. On Kay’s frequency estimator. Journal of Time Series Analysis, 21(6):707–712, November 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Quinn:2005:BR

[Qui05] Barry Quinn. Book reviews 1. Journal of Time Series Analysis, 26(1):151–152, January 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Quinn:2015:BRT

[Qui15] Barry G. Quinn. Book review: Time Series with Mixed Spectra, by Ta-Hsin Li. Published by CRC Press, 2014. Total number of pages: 680. ISBN: 978-1-58488-176-6 (hard cover), 978-1-42001-006-0 (e-book). Journal of Time Series Analysis, 36(2):267–268, March 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Reinsel:1992:VAM

[RA92] Gregory C. Reinsel and Sung K. Ahn. Vector autoregressive models with unit roots and reduced rank structure: Estimation. Likelihood ratio test, and forecasting. Journal of Time Series Analysis, 13(4): 353–375, July 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Raimondo:1996:TCP

[Rai96] M. Raimondo. Testing change-points in the explosive Gaussian autoregressive processes. Journal of Time Series Analysis, 17(5): 461–480, September 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2004:BRa

[Rao04a] B. L. S. Prakasa Rao. Book reviews. Journal of Time Series Analysis, 25(1):155–157, January 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2004:BRb

[Rao04b] T. Subba Rao. Book reviews. Journal of Time Series Analysis, 25(1):157–158, January 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2005:BRA

[Rao05] T. Subba Rao. Book review: Advanced Linear Modelling. Journal of Time Series Analysis, 26(4):628–629, July 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2008:SAS

[Rao08] Suhasini Subba Rao. Statistical analysis of a spatio-temporal model with location-dependent parameters and a test for spatial stationarity. Journal of Time Series Analysis, 29(4):673–694, July 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2010:BRH

[Rao10a] Suhasini Subba Rao. Book review: Handbook of Financial Time Series. Journal of Time Series Analysis, 31(1):64, January 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2010:NTS

[Rao10b] T. Subba Rao. Nonlinear time series: Semiparametric and nonparametric methods. Journal of Time Series Analysis, 31(3):226, May 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2010:TSA

[Rao10c] T. Subba Rao. Time series analysis. Journal of Time Series Analysis, 31(2):139, March 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2011:CPE

[Rao11] T. Subba Rao. Classification, parameter estimation and state estimation — an engineering approach using MATLAB. Journal of Time Series Analysis, 32(2):194, March 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2012:SST

[Rao12a] T. Subba Rao. Statistics for spatio-temporal data. Journal of Time Series Analysis, 33(4):699–700, July 2012. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2012:SMT

[Rao12b] Tata Subba Rao. Statistical methods for trend detection and analysis in the environmental sciences. Journal of Time Series Analysis, 33(1):176, January 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2013:SSS

[Rao13] T Subba Rao. Spatial statistics and spatio-temporal data. Journal of Time Series Analysis, 34(2):280, March 2013. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2014:BRR

[Rao14] T Subba Rao. Book review: Randall Douc, Eric Moulines and David S. Stoffer (2014) Nonlinear Time Series-Theory, Methods and Applications with R Examples. CRC Press, UK (A Chapman and Hall Book). Texts in Statistical Science. ISBN: 978-1-4665-0225-3 pages 531. Journal of Time Series Analysis, 35(6):640–641, November 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2016:BRI

[Rao16a] B. L. S. Prakasa Rao. Book review: An Introduction to Stochastic Orders, by Félix Belzunce, Carolina Martínez and Julio Mulero. Academic Press, Elsevier Ltd. 2016. Total number of pages: 157. ISBN: 978-0-12-803768-3 (Paperback). Journal of Time Series Analysis, 37(5):712, September 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2016:BRS

[Rao16b] T. Subba Rao. Book review: Statistics for Spatial Data, Revised Edition, by Noel Cressie. Published by Wiley Classics Library, John Wiley, 2015. Total number of pages: 928. ISBN: 978-1-119-11518-2. Journal of Time Series Analysis, 37(2):288, March 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2017:BRS

[Rao17] T. Subba Rao. Book review: Spatial and Spatio-Temporal Bayesian Models with R-INLA, by Marta Blangiardo and Michela Cameletti. Published by John Wiley and Sons, Chichester, UK, 2015. Total number of pages: 308. ISBN 978-1-118-32655-8. Journal of Time Series Analysis, 38(1):145–146, January 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2018:OSE

[Rao18] Suhasini Subba Rao. Orthogonal samples for estimators in time series. Journal of Time Series Analysis, 39(3):313–337, May 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Raveh:1989:NVS

[Rav89] Adi Raveh. A new version of structural persistence in prediction. Journal of Time Series Analysis, 10(1):77–93, January 1989. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ray:1988:AMS

[Ray88] D. Ray. Asymptotic mean square prediction error for a multivariate autoregressive model with random coefficients. Journal of Time Series Analysis, 9(1):73–80, January 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ray:1993:MLM

[Ray93] Bonnie K. Ray. Modeling long-memory processes for optimal longrange prediction. Journal of Time Series Analysis, 14(5):511–525, September 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Reschenhofer:1992:TWN

[RB92] E. Reschenhofer and I. M. Bomze. Testing for white noise against multimodal spectral alternatives. Journal of Time Series Analysis, 13(5):435–439, September 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Roy:2013:RCC

[RB13] Sugata Sen Roy and Sankha Bhattacharya. Rate of convergence in the central limit theorem for parameter estimation in a causal, invertible ARMA(p, q) model. Journal of Time Series Analysis, 34 (1):130–137, January 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ren:2022:AMM

[RB22] Benny Ren and Ian Barnett. Autoregressive mixture models for clustering time series. Journal of Time Series Analysis, 43(6):918– 937, November 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Reinsel:1992:MLE

[RBY92] Gregory C. Reinsel, Sabyasachi Basu, and Sook Fwe Yap. Maximum likelihood estimators in the multivariate autoregressive moving-average model from a generalized least squares viewpoint. Journal of Time Series Analysis, 13(2):133–145, March 1992. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ramirez-Cobo:2011:WBS

[RCLM^{+}11] Pepa Ramírez-Cobo, Kichun Sky Lee, Annalisa Molini, Amilcare Porporato, Gabriel Katul, and Brani Vidakovic. A wavelet-based spectral method for extracting self-similarity measures in timevarying two-dimensional rainfall maps. Journal of Time Series Analysis, 32(4):351–363, July 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2014:FDA

[RDB14] Tata Subba Rao, Sourav Das, and Georgi N. Boshnakov. A frequency domain approach for the estimation of parameters of spatiotemporal stationary random processes. Journal of Time Series Analysis, 35(4):357–377, July 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Reisen:1994:EFD

[Rei94] Valderio A. Reisen. Estimation of the fractional difference parameter in the ARIMA(p, d, q) model using the smoothed periodogram. Journal of Time Series Analysis, 15(3):335–350, May 1994. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:1980:TLS

[RG80] T. Subba Rao and M. M. Gabr. A test for linearity of stationary time series. Journal of Time Series Analysis, 1(2):145–158, March 1980. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:1989:ESI

[RG89] T. Subba Rao and M. M. Gabr. The estimation of spectrum, inverse spectrum and inverse autocovariances of a stationary time series. Journal of Time Series Analysis, 10(2):183–202, March 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Robbins:2011:MST

[RGLA11] Michael Robbins, Colin Gallagher, Robert Lund, and Alexander Aue. Mean shift testing in correlated data. Journal of Time Series Analysis, 32(5):498–511, September 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rigas:1992:SAS

[Rig92] A. G. Rigas. Spectral analysis of stationary point processes using the Fast Fourier Transform algorithm. Journal of Time Series Analysis, 13(5):441–450, September 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rigas:1996:SAS

[Rig96] A. G. Rigas. Spectral analysis of a stationary bivariate point process with applications to neurophysiological problems. Journal of Time Series Analysis, 17(2):171–187, March 1996. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Reisen:2010:PPS

[RMSF10] Valdério A. Reisen, Eric Moulines, Philippe Soulier, and Glaura C. Franco. On the properties of the periodogram of a stationary longmemory process over different epochs with applications. Journal of Time Series Analysis, 31(1):20–36, January 2010. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ravishanker:1990:DGA

[RMT90] Nalini Ravishanker, Edward L. Melnick, and Chih-Ling Tsai. Differential geometry of ARMA models. Journal of Time Series Analysis, 11(3):259–274, May 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ristic:2012:MM

[RN12] Miroslav M. Ristić and Aleksandar S. Nastić. A mixed INAR(p) model. Journal of Time Series Analysis, 33(6):903–915, November 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ristic:2013:GTS

[RNI13] Miroslav M. Ristić, Aleksandar S. Nastić, and Ana V. Miletić Ilić. A geometric time series model with dependent Bernoulli counting series. Journal of Time Series Analysis, 34(4):466–476, July 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Robinson:1983:NET

[Rob83] P. M. Robinson. Nonparametric estimators for time series. Journal of Time Series Analysis, 4(3):185–207, May 1983. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Robinson:1987:TSR

[Rob87] P. M. Robinson. Time series residuals with application to probability density estimation. Journal of Time Series Analysis, 8(3): 329–344, May 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Robinson:1994:EJH

[Rob94] P. M. Robinson. Edward J. Hannan, 1921–1994. Journal of Time Series Analysis, 15(6):563–576, November 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rodrigues:2013:RAU

[Rod13] Paulo M. M. Rodrigues. Recursive adjustment, unit root tests and structural breaks. Journal of Time Series Analysis, 34(1):62–82, January 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rozenholc:2001:NTC

[Roz01] Yves Rozenholc. Nonparametric tests of change-points with tapered data. Journal of Time Series Analysis, 22(1):13–43, January 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rubin:2020:FLR

[RP20] Tomás Rubín and Victor M. Panaretos. Functional lagged regression with sparse noisy observations. Journal of Time Series Analysis, 41(6):858–882, November 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rodriguez:2009:BPI

[RR09] Alejandro Rodriguez and Esther Ruiz. Bootstrap prediction intervals in state-space models. Journal of Time Series Analysis, 30(2): 167–178, March 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:1983:ESB

[RRW83] M. Bhaskara Rao, T. Subba Rao, and A. M. Walker. On the existence of some bilinear time series models. Journal of Time Series Analysis, 4(2):95–110, March 1983. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rice:2017:PBS

[RS17] Gregory Rice and Han Lin Shang. A plug-in bandwidth selection procedure for long-run covariance estimation with stationary functional time series. Journal of Time Series Analysis, 38(4):591–609, July 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rice:2019:ILC

[RS19] Gregory Rice and Marco Shum. Inference for the lagged crosscovariance operator between functional time series. Journal of Time Series Analysis, 40(5):665–692, September 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rodriguez:2015:PJD

[RSVM15] Frank Rodriguez, Soterios Soteri, and Leticia Veruete-McKay. Papers with John on the demand for mail. Journal of Time Series Analysis, 36(5):650–652, September 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rekkas:2008:IIF

[RSW08] M. Rekkas, Y. Sun, and A. Wong. Improved inference for first-order autocorrelation using likelihood analysis. Journal of Time Series Analysis, 29(3):513–532, May 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rabinovitch:1992:PNS

[RT92] A. Rabinovitch and R. Thieberger. ‘purifying’ noisy signals. Journal of Time Series Analysis, 13(3):267–280, May 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ray:2002:BMC

[RT02] Bonnie K. Ray and Ruey S. Tsay. Bayesian methods for changepoint detection in long-range dependent processes. Journal of Time Series Analysis, 23(6):687–705, November 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Roueff:2009:ANW

[RT09] F. Roueff and M. S. Taqqu. Asymptotic normality of wavelet estimators of the memory parameter for linear processes. Journal of Time Series Analysis, 30(5):534–558, September 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2017:NCF

[RT17a] T. Subba Rao and Gyorgy Terdik. A new covariance function and spatio-temporal prediction (kriging) for a stationary spatiotemporal random process. Journal of Time Series Analysis, 38 (6):936–959, November 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2017:FVF

[RT17b] Tata Subba Rao and Gyorgy Terdik. On the frequency variogram and on frequency domain methods for the analysis of spatiotemporal data. Journal of Time Series Analysis, 38(2):308–325, March 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Robinson:2017:AEM

[RT17c] P. M. Robinson and L. Taylor. Adaptive estimation in multiple time series with independent component errors. Journal of Time Series Analysis, 38(2):191–203, March 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Romano:2022:PTD

[RT22] Joseph P. Romano and Marius A. Tirlea. Permutation testing for dependence in time series. Journal of Time Series Analysis, 43 (5):781–807, September 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2014:OMB

[RTW14] T. Subba Rao and Granville Tunnicliffe-Wilson. Obituary: Maurice Bertram Priestley, MA, PhD, 1933–2013. Journal of Time Series Analysis, 35(1):1–3, January 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2017:ESI

[RW17] Tata Subba Rao and Granville Tunnicliffe Wilson. Editorial: Special issue to honor the memory of Maurice B. Priestley, 1933–2013. Journal of Time Series Analysis, 38(2):149–150, March 2017. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rice:2020:TCH

[RWZ20] Gregory Rice, Tony Wirjanto, and Yuqian Zhao. Tests for conditional heteroscedasticity of functional data. Journal of Time Series Analysis, 41(6):733–758, November 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Rao:2023:PPW

[RY23] Suhasini Subba Rao and Junho Yang. A prediction perspective on the Wiener–Hopf equations for time series. Journal of Time Series Analysis, 44(1):23–42, January 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Regnard:2010:SEC

[RZ10] Nazim Regnard and Jean-Michel Zakoïan. Structure and estimation of a class of nonstationary yet nonexplosive GARCH models. Journal of Time Series Analysis, 31(5):348–364, September 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Soltani:2007:SRV

[SA07] A. R. Soltani and M. Azimmohseni. Simulation of real-valued discrete-time periodically correlated Gaussian processes with prescribed spectral density matrices. Journal of Time Series Analysis, 28(2):225–240, March 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Saikkonen:1983:ARE

[Sai83] Pentti Saikkonen. Asymptotic relative efficiency of some tests of fit in time series models. Journal of Time Series Analysis, 4(1): 69–78, January 1983. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Saikkonen:1986:APS

[Sai86] Pentti Saikkonen. Asymptotic properties of some preliminary estimators for autoregressive moving average time series models. Journal of Time Series Analysis, 7(2):133–155, March 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sakaguchi:1991:RLB

[Sak91a] Fuminori Sakaguchi. A relation for ‘linearity’ of the bispectrum. Journal of Time Series Analysis, 12(3):267–272, May 1991. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sakai:1991:SDM

[Sak91b] Hideaki Sakai. On the spectral density matrix of a periodic ARMA process. Journal of Time Series Analysis, 12(1):73–82, January 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sakai:1993:DNT

[Sak93] Hideaki Sakai. The determination of the number of terms in a multichannel sinusoidal regression. Journal of Time Series Analysis, 14(6):621–628, November 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sandberg:2017:SMW

[San17] Rickard Sandberg. Sample moments and weak convergence to multivariate stochastic power integrals. Journal of Time Series Analysis, 38(6):1000–1009, November 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sandberg:2018:URT

[San18] Rickard Sandberg. Unit root testing in multiple smooth break models with nonlinear dynamics. Journal of Time Series Analysis, 39(6):942–952, November 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Savin:2015:PJ

[Sav15] Nathan E. (Gene) Savin. Papers with John. Journal of Time Series Analysis, 36(5):663–671, September 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Seong:2013:EVE

[SAZ13] Byeongchan Seong, Sung K. Ahn, and Peter A. Zadrozny. Estimation of vector error correction models with mixed-frequency data. Journal of Time Series Analysis, 34(2):194–205, March 2013. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shirota:2019:SIS

[SB19] Shinichiro Shirota and Sudipto Banerjee. Scalable inference for space–time Gaussian Cox processes. Journal of Time Series Analysis, 40(3):269–287, May 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Serban:2007:MDD

[SBLS07] Mihaela Serban, Anthony Brockwell, John Lehoczky, and Sanjay Srivastava. Modelling the dynamic dependence structure in multivariate financial time series. Journal of Time Series Analysis, 28 (5):763–782, September 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sbrana:2011:STS

[Sbr11] Giacomo Sbrana. Structural time series models and aggregation: some analytical results. Journal of Time Series Analysis, 32(3): 315–316, May 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Silva:2019:FRM

[SBS19] Rodrigo B. Silva and Wagner Barreto-Souza. Flexible and robust mixed Poisson INGARCH models. Journal of Time Series Analysis, 40(5):788–814, September 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sundararajan:2023:SSV

[SBS23] Raanju R. Sundararajan and Wagner Barreto-Souza. Studentt stochastic volatility model with composite likelihood EMalgorithm. Journal of Time Series Analysis, 44(1):125–147, January 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Chan:1988:ESE

[sC88] Kung sik Chan. On the existence of the stationary and ergodic near( p ) model. Journal of Time Series Analysis, 9(4):319–328, July 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sun:1997:SGS

[SC97] T. C. Sun and Milton Chaika. On simulation of a Gaussian stationary process. Journal of Time Series Analysis, 18(1):79–93, January 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Schmid:1996:OTT

[Sch96] W. Schmid. An outlier test for time series based on a two-sided predictor. Journal of Time Series Analysis, 17(5):497–510, September 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Schick:1998:AEA

[Sch98] Anton Schick. An adaptive estimator of the autocorrelation coefficient in regression models with autoregressive errors. Journal of Time Series Analysis, 19(5):575–589, September 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Schweer:2016:GFT

[Sch16] Sebastian Schweer. A goodness-of-fit test for integer-valued autoregressive processes. Journal of Time Series Analysis, 37(1): 77–98, January 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Schweikert:2022:OEE

[Sch22] Karsten Schweikert. Oracle efficient estimation of structural breaks in cointegrating regressions. Journal of Time Series Analysis, 43 (1):83–104, January 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Song:2019:BCE

[SCW19] Yuping Song, Ying Chen, and Zhouwei Wang. Bias correction estimation for a continuous-time asset return model with jumps. Journal of Time Series Analysis, 40(1):66–101, January 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sakarya:2022:SAH

[SdJ22] Neslihan Sakarya and Robert M. de Jong. The spectral analysis of the hodrick-prescott filter. Journal of Time Series Analysis, 43(3): 479–489, May 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Selukar:2010:ELE

[Sel10] Rajesh Selukar. Estimability of the linear effects in state space models with an unknown initial condition. Journal of Time Series Analysis, 31(3):167–168, May 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sen:2007:JHT

[Sen07] Amit Sen. Joint hypothesis tests for a unit root when there is a break in the innovation variance. Journal of Time Series Analysis, 28(5):686–700, September 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Seo:2024:FPC

[Seo24] Won-Ki Seo. Functional principal component analysis for cointegrated functional time series. Journal of Time Series Analysis, 45(2):320–330, March 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Smith:1993:VEQ

[SF93] B. Smith and C. Field. Variance estimation for quadratic statistics. Journal of Time Series Analysis, 14(4):381–395, July 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shin:1998:URT

[SF98] Dong Wan Shin and Wayne Fuller. Unit root tests based on unconditional maximum likelihood estimation for the autoregressive moving average. Journal of Time Series Analysis, 19(5):591–599, September 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Song:2005:PEE

[SF05] Peter X.-K. Song and Dingan Feng. On parameter estimation for exponential dispersion ARMA models. Journal of Time Series Analysis, 26(6):843–862, November 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Salazar:2011:TAL

[SF11] Esther Salazar and Marco A. R. Ferreira. Temporal aggregation of lognormal AR processes. Journal of Time Series Analysis, 32 (6):661–671, November 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Stockis:2010:GEC

[SFK10] Jean-Pierre Stockis, Jürgen Franke, and Joseph Tadjuidje Kamgaing. On geometric ergodicity of CHARME models. Journal of Time Series Analysis, 31(3):141–152, May 2010. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Samaranayake:1987:APS

[SH87] V. A. Samaranayake and David P. Hasza. The asymptotic properties of the sample autocorrelations for a multiple autoregressive process with one unit root. Journal of Time Series Analysis, 8(1): 79–93, January 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Samaranayake:1988:PPM

[SH88a] V. A. Samaranayake and David P. Hasza. Properties of predictors for multivariate autoregressive models with estimated parameters. Journal of Time Series Analysis, 9(4):361–383, July 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Stam:1988:MMF

[SH88b] Antonie Stam and Steven C. Hillmer. Marginals of multivariate first-order autoregressive time series models. Journal of Time Series Analysis, 9(1):89–97, January 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shephard:1990:PED

[SH90] N. G. Shephard and A. C. Harvey. On the probability of estimating a deterministic component in the local level model. Journal of Time Series Analysis, 11(4):339–347, July 1990. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sela:2009:CEM

[SH09] Rebecca J. Sela and Clifford M. Hurvich. Computationally efficient methods for two multivariate fractionally integrated models. Journal of Time Series Analysis, 30(6):631–651, November 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sela:2012:APE

[SH12] Rebecca J. Sela and Clifford M. Hurvich. The averaged periodogram estimator for a power law in coherency. Journal of Time Series Analysis, 33(2):340–363, March 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shao:2008:REP

[Sha08] Q. Shao. Robust estimation for periodic autoregressive time series. Journal of Time Series Analysis, 29(2):251–263, March 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shao:2011:STC

[Sha11] Xiaofeng Shao. A simple test of changes in mean in the possible presence of long-range dependence. Journal of Time Series Analysis, 32(6):598–606, November 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shea:1987:EMT

[She87] B. L. Shea. Estimation of multivariate time series. Journal of Time Series Analysis, 8(1):95–109, January 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shea:1988:NGI

[She88] B. L. Shea. A note on the generation of independent realizations of a vector autoregressive moving-average process. Journal of Time Series Analysis, 9(4):403–410, July 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shively:1988:ETS

[Shi88] Thomas S. Shively. An exact test for a stochastic coefficient in a time series regression model. Journal of Time Series Analysis, 9(1):81–88, January 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shin:1993:MLE

[Shi93] Dong Wan Shin. Maximum likelihood estimation for autoregressive processes disturbed by a moving average. Journal of Time Series Analysis, 14(6):629–643, November 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shin:1998:LDR

[Shi98] Dong Wan Shin. The limiting distribution of the residual processes in nonstationary autoregressive processes. Journal of Time Series Analysis, 19(6):723–736, November 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Song:2022:DSV

[SHL22] Yuping Song, Weijie Hou, and Zhengyan Lin. Double smoothed volatility estimation of potentially non-stationary jump-diffusion model of Shibor. Journal of Time Series Analysis, 43(1):53–82, January 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Svensson:1996:OPC

[SHLL96] A. Svensson, J. Holst, R. Lindquist, and G. Lindgren. Optimal prediction of catastrophes in autoregressive moving-average processes. Journal of Time Series Analysis, 17(5):511–531, September 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shumway:2005:BR

[Shu05] Robert H. Shumway. Book reviews. Journal of Time Series Analysis, 26(5):779–780, September 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sibbertsen:2001:ELR

[Sib01] Philipp Sibbertsen. S-estimation in the linear regression model with long-memory error terms under trend. Journal of Time Series Analysis, 22(3):353–363, May 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Simos:2008:EDM

[Sim08] Theodore Simos. The exact discrete model of a system of linear stochastic differential equations driven by fractional noise. Journal of Time Series Analysis, 29(6):1019–1031, November 2008. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Singer:1993:CTD

[Sin93] Hermann Singer. Continuous-time dynamical systems with sampled data, errors of measurement and unobserved components. Journal of Time Series Analysis, 14(5):527–545, September 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sato:1996:SPM

[SK96] Seisho Sato and Naoto Kunitomo. Some properties of the maximum likelihood estimator in the simultaneous switching autoregressive model. Journal of Time Series Analysis, 17(3):287–307, May 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sibbertsen:2009:TBP

[SK09] Philipp Sibbertsen and Robinson Kruse. Testing for a break in persistence under long-range dependencies. Journal of Time Series Analysis, 30(3):263–285, May 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sabzikar:2023:TFT

[SK23] Farzad Sabzikar and Piotr Kokoszka. Tempered functional time series. Journal of Time Series Analysis, 44(3):280–293, May 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Skold:2001:BCC

[Sko01] Martin Skold. A bias correction for cross-validation bandwidth selection when a kernel estimate is based on dependent data. Journal of Time Series Analysis, 22(4):493–503, July 2001. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Saikkonen:1996:TOD

[SL96a] Pentti Saikkonen and Ritva Luukkonen. Testing the order of differencing in time series regression. Journal of Time Series Analysis, 17(5):481–496, September 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shin:1996:DRA

[SL96b] Dong Wan Shin and Jong Hyup Lee. Distribution of residual autocorrelations in nonstationary autoregressive processes. Journal of Time Series Analysis, 17(1):105–109, January 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shin:1997:SMN

[SL97] Dong Wan Shin and Yoon Dong Lee. A study on misspecified nonstationary autoregressive time series with a unit root. Journal of Time Series Analysis, 18(5):475–484, September 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Saikkonen:2000:TAP

[SL00] Pentti Saikkonen and Helmut Lütkepohl. Trend adjustment prior to testing for the cointegrating rank of a vector autoregressive process. Journal of Time Series Analysis, 21(4):435–456, July 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shao:2004:CCA

[SL04a] Qin Shao and Robert Lund. Computation and characterization of autocorrelations and partial autocorrelations in periodic ARMA models. Journal of Time Series Analysis, 25(3):359–372, May 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shin:2004:ERI

[SL04b] Dong Wan Shin and Oesook Lee. M-estimation for regressions with integrated regressors and ARMA errors. Journal of Time Series Analysis, 25(2):283–299, March 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

So:1997:MMA

[SLL97] Mike K. P. So, W. K. Li, and K. Lam. Multivariate modelling of the autoregressive random variance process. Journal of Time Series Analysis, 18(4):429–446, July 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sun:2020:MVR

[SLL^{+}20] Yan Sun, Guanghua Lian, Zudi Lu, Jennifer Loveland, and Isaac Blackhurst. Modeling the variance of return intervals toward volatility prediction. Journal of Time Series Analysis, 41(4):492– 519, July 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sollis:1999:URA

[SLN99] Robert Sollis, Stephen Leybourne, and Paul Newbold. Unit roots and asymmetric smooth transitions. Journal of Time Series Analysis, 20(6):671–677, November 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Soltani:2006:MAR

[SM06] A. R. Soltani and M. Mohammadpour. Moving average representations for multivariate stationary processes. Journal of Time Series Analysis, 27(6):831–841, November 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Schmidt:2013:ESA

[SM13a] Daniel F. Schmidt and Enes Makalic. Estimation of stationary autoregressive models with the Bayesian LASSO. Journal of Time Series Analysis, 34(5):517–531, September 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sun:2013:SST

[SM13b] Jiajing Sun and Brendan P. McCabe. Score statistics for testing serial dependence in count data. Journal of Time Series Analysis, 34(3):315–329, May 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Smeekes:2015:BST

[Sme15] Stephan Smeekes. Bootstrap sequential tests to determine the order of integration of individual units in a time series panel. Journal of Time Series Analysis, 36(3):398–415, May 2015. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Smith:2008:EST

[Smi08] Daniel R. Smith. Evaluating specification tests for Markovswitching time-series models. Journal of Time Series Analysis, 29(4):629–652, July 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sakai:1997:BPA

[SO97a] Hideaki Sakai and Shyuichi Ohno. On backward periodic autoregressive processes. Journal of Time Series Analysis, 18(4):415–427, July 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shoji:1997:CSE

[SO97b] Isao Shoji and Tohru Ozaki. Comparative study of estimation methods for continuous time stochastic processes. Journal of Time Series Analysis, 18(5):485–506, September 1997. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Silva:2005:DEH

[SO05] Maria Eduarda Silva and Vera Lúcia Oliveira. Difference equations for the higher order moments and cumulants of the INAR(p) model.

Journal of Time Series Analysis, 26(1):17–36, January 2005. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Stoffer:2012:ESI

[SO12] David S. Stoffer and Hernando Ombao. Editorial: Special issue on time series analysis in the biological sciences. Journal of Time Series Analysis, 33(5):701–703, September 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sollis:2004:AAS

[Sol04] Robert Sollis. Asymmetric adjustment and smooth transitions: a combination of some unit root tests. Journal of Time Series Analysis, 25(3):409–417, May 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Souza:2007:TAB

[Sou07] Leonardo Rocha Souza. Temporal aggregation and bandwidth selection in estimating long memory. Journal of Time Series Analysis, 28(5):701–722, September 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sanchez:2001:PPO

[SP01] Ismael Sanchez and Daniel Pena. Properties of predictors in overdifferenced nearly nonstationary autoregression. Journal of Time Series Analysis, 22(1):45–66, January 2001. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shimotsu:2002:PLP

[SP02] Katsumi Shimotsu and Peter C. B. Phillips. Pooled log periodogram regression. Journal of Time Series Analysis, 23(1):57– 93, January 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sergides:2008:BLP

[SP08] Marios Sergides and Efstathios Paparoditis. Bootstrapping the local periodogram of locally stationary processes. Journal of Time Series Analysis, 29(2):264–299, March 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Solo:2012:TIB

[SP12] Victor Solo and Ahmed Pasha. A test for independence between a point process and an analogue signal. Journal of Time Series Analysis, 33(5):824–840, September 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sundararajan:2018:SSA

[SP18] Raanju Ragavendar Sundararajan and Mohsen Pourahmadi. Stationary subspace analysis of nonstationary processes. Journal of Time Series Analysis, 39(3):338–355, May 2018. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Spall:1993:CDN

[Spa93a] J. C. Spall. Correction to “The Distribution of Nonstationary Autoregressive Processes Under General Noise Conditions”. Journal of Time Series Analysis, 14(5):550, September 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic). See [Spa93b].

Spall:1993:DNA

[Spa93b] James C. Spall. The distribution of nonstationary autoregressive processes under general noise conditions. Journal of Time Series Analysis, 14(3):317–330, May 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic). See correction [Spa93a].

Sellers:2020:FUA

[SPA20] Kimberly F. Sellers, Stephen J. Peng, and Ali Arab. A flexible univariate autoregressive time-series model for dispersed count data. Journal of Time Series Analysis, 41(3):436–453, May 2020. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Spezia:2010:BAM

[Spe10] Luigi Spezia. Bayesian analysis of multivariate Gaussian hidden Markov models with an unknown number of regimes. Journal of Time Series Analysis, 31(1):1–11, January 2010. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shi:2018:CDC

[SPH18] Shuping Shi, Peter C. B. Phillips, and Stan Hurn. Change detection and the causal impact of the yield curve. Journal of Time Series Analysis, 39(6):966–987, November 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Silva:2019:BOD

[SPM19] Maria Eduarda Silva, Isabel Pereira, and Brendan McCabe. Bayesian outlier detection in non-Gaussian autoregressive time series. Journal of Time Series Analysis, 40(5):631–648, September 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sesay:1988:YWT

[SR88] S. A. O. Sesay and T. Subba Rao. Yule–Walker type difference equations for higher-order moments and cumulants for bilinear time series models. Journal of Time Series Analysis, 9(4):385–401, July 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sesay:1991:DEH

[SR91] S. A. O. Sesay and T. Subba Rao. Difference equations for higherorder moments and cumulants for the bilinear time series model Bl(p, 0, p, 1). Journal of Time Series Analysis, 12(2):159–177, March 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sesay:1992:FDE

[SR92] S. A. O. Sesay and T. Subba Rao. Frequency-domain estimation of bilinear time series models. Journal of Time Series Analysis, 13 (6):521–545, November 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Stramer:2007:BAN

[SR07] O. Stramer and G. O. Roberts. On Bayesian analysis of nonlinear continuous-time autoregression models. Journal of Time Series Analysis, 28(5):744–762, September 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sorbye:2017:PCP

[SR17] Sigrunn Holbek Sørbye and Håvard Rue. Penalised complexity priors for stationary autoregressive processes. Journal of Time Series Analysis, 38(6):923–935, November 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shou-Ren:1983:DSS

[SRHZT83] Wang Shou-Ren, An Hong-Zhi, and H. Tong. On the distribution of a simple stationary bilinear process. Journal of Time Series Analysis, 4(3):209–216, May 1983. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shumway:1982:ATS

[SS82] R. H. Shumway and D. S. Stoffer. An approach to time series smoothing and forecasting using the EM algorithm. Journal of Time Series Analysis, 3(4):253–264, July 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sakaguchi:1989:CLM

[SS89] Fuminori Sakaguchi and Hideaki Sakai. A composite linear model generating a stationary stochastic process with given third-order autocorrelation function. Journal of Time Series Analysis, 10(2): 171–181, March 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sakai:1990:SCB

[SS90] Hideaki Sakai and Fuminori Sakaguchi. Simultaneous confidence bands for the spectral estimate of two-channel autoregressive processes. Journal of Time Series Analysis, 11(1):49–56, January 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shin:1995:EMA

[SS95] Dong Wan Shin and Sahadeb Sarkar. Estimation of the multivariate autoregressive moving average having parameter restrictions and an application to rotational sampling. Journal of Time Series Analysis, 16(4):431–444, July 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shin:1996:EMV

[SS96a] Dong Wan Shin and Sahadeb Sarkar. Estimation of the multivariate autoregressive moving average having parameter restrictions and an application to rotational sampling. Journal of Time Series Analysis, 17(3):321, May 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shin:1996:TUR

[SS96b] Dong Wan Shin and Sahadeb Sarkar. Testing for a unit root in an AR(1) time series using irregularly observed data. Journal of Time Series Analysis, 17(3):309–321, May 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Snyder:1996:IKF

[SS96c] Ralph D. Snyder and Grant R. Saligari. Initialization of the Kalman filter with partially diffuse initial conditions. Journal of Time Series Analysis, 17(4):409–424, July 1996. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shin:1998:TUR

[SS98] Dong Wan Shin and Sahadeb Sarkar. Testing for a unit root in autoregressive moving-average models with missing data. Journal of Time Series Analysis, 19(5):601–608, September 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shin:2001:RMA

[SS01] Dong Wan Shin and Beong Soo So. Recursive mean adjustment for unit root tests. Journal of Time Series Analysis, 22(5):595– 612, September 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Saikkonen:2016:TUR

[SS16] Pentti Saikkonen and Rickard Sandberg. Testing for a unit root in noncausal autoregressive models. Journal of Time Series Analysis, 37(1):99–125, January 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sengupta:2015:DRW

[SSW15] Srijan Sengupta, Xiaofeng Shao, and Yingchuan Wang. The dependent random weighting. Journal of Time Series Analysis, 36(3): 315–326, May 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sang:2018:KEE

[SSX18] Hailin Sang, Yongli Sang, and Fangjun Xu. Kernel entropy estimation for linear processes. Journal of Time Series Analysis, 39(4): 563–591, July 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Stadtmuller:1985:ABD

[ST85] U. Stadtmüller and R. Trautner. Asymptotic behaviour of discrete linear processes. Journal of Time Series Analysis, 6(2):97–108, March 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Stensholt:1987:MBT

[ST87] Boonchai K. Stensholt and Dag Tjøstheim. Multiple bilinear time series models. Journal of Time Series Analysis, 8(2):221–233, March 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Samarov:1988:ESM

[ST88] Alexander Samarov and Murad S. Taqqu. On the efficiency of the sample mean in long-memory noise. Journal of Time Series Analysis, 9(2):191–200, March 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Swe:1991:HOA

[ST91] Myint Swe and Masanobu Taniguchi. Higher-order asymptotic properties of a weighted estimator for Gaussian ARMA processes. Journal of Time Series Analysis, 12(1):83–93, January 1991. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shibata:1997:CFE

[ST97] Ritei Shibata and Mutsumi Takagiwa. Consistency of frequency estimates based on the wavelet transform. Journal of Time Series Analysis, 18(6):641–662, November 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Smith:1999:LRT

[ST99] Richard J. Smith and A. M. Robert Taylor. Likelihood ratio tests for seasonal unit roots. Journal of Time Series Analysis, 20(4): 453–476, July 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sakiyama:2003:TCH

[ST03] Kenji Sakiyama and Masanobu Taniguchi. Testing composite hypotheses for locally stationary processes. Journal of Time Series Analysis, 24(4):483–504, July 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shi:2004:JRV

[ST04] Peide Shi and Chih-Ling Tsai. A joint regression variable and autoregressive order selection criterion. Journal of Time Series Analysis, 25(6):923–941, November 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Stoev:2005:ASS

[ST05] Stilian Stoev and Murad S. Taqqu. Asymptotic self-similarity and wavelet estimation for long-range dependent fractional autoregressive integrated moving average time series with stable innovations. Journal of Time Series Analysis, 26(2):211–249, March 2005. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Scotto:2003:ESS

[STA03] M. G. Scotto, K. F. Turkman, and C. W. Anderson. Extremes of some sub-sampled time series. Journal of Time Series Analysis, 24 (5):579–590, September 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Stauskas:2020:LTM

[Sta20] Ovidijus Stauskas. On the limit theory of mixed to unity VARs: Panel setting with weakly dependent errors. Journal of Time Series Analysis, 41(6):892–898, November 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Steland:2005:RWD

[Ste05] Ansgar Steland. Random walks with drift — a sequential approach. Journal of Time Series Analysis, 26(6):917–942, November 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Stoffer:1985:CLT

[Sto85] David S. Stoffer. Central limit theorems for finite Walsh-Fourier transforms of weakly stationary time series. Journal of Time Series Analysis, 6(4):261–267, July 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Stoffer:1987:WFA

[Sto87] David S. Stoffer. Walsh-Fourier analysis of discrete-valued time series. Journal of Time Series Analysis, 8(4):449–467, July 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Stoffer:1990:MWF

[Sto90] David S. Stoffer. Multivariate Walsh-Fourier analysis. Journal of Time Series Analysis, 11(1):57–73, January 1990. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Stoykov:2019:LSB

[Sto19] Marian Z. Stoykov. Least squares bias in time series with moderate deviations from a unit root. Journal of Time Series Analysis, 40(1): 23–42, January 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Stramer:1996:AMC

[Str96] O. Stramer. On the approximation of moments for continuous time threshold ARMA processes. Journal of Time Series Analysis, 17(2):189–202, March 1996. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Stulajter:2001:PTS

[Stu01] Frantisek Stulajter. Predictions in time series using multivariate regression models. Journal of Time Series Analysis, 22(3):365–373, May 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Smith:1997:CBM

[STY97] Jeremy Smith, Nick Taylor, and Sanjay Yadav. Comparing the bias and misspecification in ARFIMA models. Journal of Time Series Analysis, 18(5):507–527, September 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sun:2014:SPE

[Sun14] Yiguo Sun. Semi-parametric estimation of linear cointegrating models with nonlinear contemporaneous endogeneity. Journal of Time Series Analysis, 35(5):437–461, September 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Stram:1986:MND

[SW86a] Daniel O. Stram and William W. S. Wei. A methodological note on the disaggregation of time series totals. Journal of Time Series Analysis, 7(4):293–302, July 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Stram:1986:TAA

[SW86b] Daniel O. Stram and William W. S. Wei. Temporal aggregation in the ARIMA process. Journal of Time Series Analysis, 7(4): 279–292, July 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Smetanina:2021:ATQ

[SW21] Ekaterina Smetanina and Wei Biao Wu. Asymptotic theory for QMLE for the real-time GARCH(1,1) model. Journal of Time Series Analysis, 42(5-6):752–776, September 2021. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Swensen:2003:BUR

[Swe03] Anders Rygh Swensen. Bootstrapping unit root tests for integrated processes. Journal of Time Series Analysis, 24(1):99–126, January 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Swensen:2022:CNC

[Swe22] Anders Rygh Swensen. On causal and non-causal cointegrated vector autoregressive time series. Journal of Time Series Analysis, 43(2):178–196, March 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Swift:1990:OIV

[Swi90] A. L. Swift. Orders and initial values of non-stationary multivariate ARMA models. Journal of Time Series Analysis, 11(4):349–359, July 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Shao:2011:ACE

[SY11] Q. Shao and L. J. Yang. Autoregressive coefficient estimation in nonparametric analysis. Journal of Time Series Analysis, 32(6): 587–597, November 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Sun:2020:VAF

[SY20] Hao Sun and Bo Yu. Volatility asymmetry in functional threshold GARCH model. Journal of Time Series Analysis, 41(1):95–109, January 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Szabados:2022:RMS

[Sza22] Tamás Szabados. Regular multidimensional stationary time series. Journal of Time Series Analysis, 43(2):263–284, March 2022. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic). See corrigendum [Sza23].

Szabados:2023:CAR

[Sza23] Tamás Szabados. Corrigendum to the article “Regular multidimensional stationary time series”. Journal of Time Series Analysis, 44(3):331–332, May 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic). See [Sza22].

Thavaneswaran:1988:ENL

[TA88] A. Thavaneswaran and B. Abraham. Estimation for non-linear time series models using estimating equations. Journal of Time Series Analysis, 9(1):99–108, January 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Takabatake:2024:CEB

[Tak24] Tetsuya Takabatake. Corrigendum: Error bounds and asymptotic expansions for Toeplitz product functionals of unbounded spectra. Journal of Time Series Analysis, 45(1):158–160, January 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic). See [LP04].

Tamaki:2009:SOP

[Tam09] Kenichiro Tamaki. Second-order properties of locally stationary processes. Journal of Time Series Analysis, 30(1):145–166, January 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tesfaye:2011:ARF

[TAM11] Yonas Gebeyehu Tesfaye, Paul L. Anderson, and Mark M. Meerschaert. Asymptotic results for Fourier-PARMA time series. Journal of Time Series Analysis, 32(2):157–174, March 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taniguchi:1981:RRI

[Tan81] Masanobu Taniguchi. Robust regression and interpolation for time series. Journal of Time Series Analysis, 2(1):53–62, January 1981. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taniguchi:1984:VEE

[Tan84] Masanobu Taniguchi. Validity of Edgeworth expansions for statistics of time series. Journal of Time Series Analysis, 5(1):37–51, January 1984. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taniguchi:1987:TOA

[Tan87] Masanobu Taniguchi. Third order asymptotic properties of BLUE and LSE for a regression model with ARMA residual. Journal of Time Series Analysis, 8(1):111–114, January 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taufemback:2023:NPS

[Tau23] Cleiton Guollo Taufemback. Non-parametric short- and long-run Granger causality testing in the frequency domain. Journal of Time Series Analysis, 44(1):69–92, January 2023. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:1998:TUR

[Tay98] A. M. Robert Taylor. Testing for unit roots in monthly time series. Journal of Time Series Analysis, 19(3):349–368, May 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:2003:LOT

[Tay03] A. M. Robert Taylor. Locally optimal tests against unit roots in seasonal time series processes. Journal of Time Series Analysis, 24 (5):591–612, September 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:2005:USS

[Tay05] A. M. Robert Taylor. On the use of sub-sample unit root tests to detect changes in persistence. Journal of Time Series Analysis, 26 (5):759–778, September 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:2013:E

[Tay13a] Robert Taylor. Editorial. Journal of Time Series Analysis, 34(2): 139–140, March 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:2013:EA

[Tay13b] Robert Taylor. Editorial announcement. Journal of Time Series Analysis, 34(6):605, November 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:2018:EA

[Tay18a] Robert Taylor. Editorial announcement. Journal of Time Series Analysis, 39(6):813, November 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:2018:EJ

[Tay18b] Robert Taylor. Editorial, January 2018. Journal of Time Series Analysis, 39(1):3, January 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:2018:ES

[Tay18c] Robert Taylor. Editorial, September 2018. Journal of Time Series Analysis, 39(5):639, September 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:2019:EA

[Tay19] Robert Taylor. Editorial announcement. Journal of Time Series Analysis, 40(4):385, July 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:2020:EAJ

[Tay20] Robert Taylor. Editorial announcement: Journal of Time Series Analysis distinguished authors. Journal of Time Series Analysis, 41(4):489–490, July 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:2021:EA

[Tay21a] Robert Taylor. Editorial announcement. Journal of Time Series Analysis, 42(2):139, March 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:2021:EAJ

[Tay21b] Robert Taylor. Editorial announcement: Journal of Time Series Analysis distinguished authors 2020. Journal of Time Series Analysis, 42(1):3, January 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:2022:EAJ

[Tay22a] Robert Taylor. Editorial announcement: Journal of Time Series Analysis Distinguished Authors 2021. Journal of Time Series Analysis, 43(1):4, January 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:2022:EAP

[Tay22b] Robert Taylor. Editorial announcement: Professor Michael McAleer [obituary]. Journal of Time Series Analysis, 43(1):3, January 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:2023:EAa

[Tay23a] Robert Taylor. Editorial announcement. Journal of Time Series Analysis, 44(4):335, July 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:2023:EAb

[Tay23b] Robert Taylor. Editorial announcement. Journal of Time Series Analysis, 44(5-6):439, September 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:2023:EAJ

[Tay23c] Robert Taylor. Editorial announcement: Journal of Time Series Analysis Distinguished Authors 2022. Journal of Time Series Analysis, 44(1):3, January 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:2024:CPS

[Tay24a] Robert Taylor. Call for papers: Special issue on recent developments in time series methods for detecting bubbles and crashes. Journal of Time Series Analysis, 45(2):163, March 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taylor:2024:EAJ

[Tay24b] Robert Taylor. Editorial announcement: Journal of Time Series Analysis Distinguished Authors 2023. Journal of Time Series Analysis, 45(1):3, January 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tsai:2005:QML

[TC05a] Henghsiu Tsai and K. S. Chan. Quasi-maximum likelihood estimation for a class of continuous-time long-memory processes. Journal of Time Series Analysis, 26(5):691–713, September 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tsai:2005:TAS

[TC05b] Henghsiu Tsai and K. S. Chan. Temporal aggregation of stationary and nonstationary discrete-time processes. Journal of Time Series Analysis, 26(4):613–624, July 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tsai:2007:NNN

[TC07] Henghsiu Tsai and K. S. Chan. A note on non-negative ARMA processes. Journal of Time Series Analysis, 28(3):350–360, May 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Thornton:2013:CTA

[TC13] Michael A. Thornton and Marcus J. Chambers. Continuous-time autoregressive moving average processes in discrete time: representation and embeddability. Journal of Time Series Analysis, 34 (5):552–561, September 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tagle:2019:NGS

[TCCG19] Felipe Tagle, Stefano Castruccio, Paola Crippa, and Marc G. Genton. A non-Gaussian spatio-temporal model for daily wind speeds based on a multi-variate skew-t distribution. Journal of Time Series Analysis, 40(3):312–326, May 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Thomaidis:2011:DOS

[TD11] Nikos S. Thomaidis and George D. Dounias. On detecting the optimal structure of a neural network under strong statistical features in errors. Journal of Time Series Analysis, 32(3):204–222, May 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Terasvirta:1985:MMI

[Ter85] Timo Teräsvirta. Mink and muskrat interaction: A structural analysis. Journal of Time Series Analysis, 6(3):171–180, May 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Terdik:2005:BR

[Ter05] Gyorgy Terdik. Book reviews. Journal of Time Series Analysis, 26(5):786, September 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Terdik:2011:OSI

[Ter11] György Terdik. Optimal statistical inference in financial engineering. Journal of Time Series Analysis, 32(1):92, January 2011. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Terdik:2014:BRL

[Ter14] Gy Terdik. Book review: Long-memory Processes: Probabilistic Properties and Statistical Methods, by Jan Beran, Yuanhua Feng, Sucharita Ghosh, and Rafal Kulik. Published by Springer London, 2013. Total Number of Pages: 884. ISBN: 978-3-642-35511-0 (print), 978-3-642-35512-7 (online). Journal of Time Series Analysis, 35(4):390–392, July 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tewes:2018:BBE

[Tew18] Johannes Tewes. Block bootstrap for the empirical process of longrange dependent data. Journal of Time Series Analysis, 39(1): 28–53, January 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taniguchi:2012:GIC

[TH12] Masanobu Taniguchi and Junichi Hirukawa. Generalized information criterion. Journal of Time Series Analysis, 33(2):287–297, March 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Thanoon:1990:STA

[Tha90] B. Y. Thanoon. Subset threshold autoregression with applications. Journal of Time Series Analysis, 11(1):75–87, January 1990. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tseng:2024:GCT

[THN24] Neng-Fang Tseng, Ying-Chao Hung, and Junji Nakano. Granger causality tests based on reduced variable information. Journal of Time Series Analysis, 45(3):444–462, May 2024. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Thornton:2019:EDR

[Tho19] Michael A. Thornton. Exact discrete representations of linear continuous time models with mixed frequency data. Journal of Time Series Analysis, 40(6):951–967, November 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tian:1988:LPS

[Tia88] C. J. Tian. A limiting property of sample autocovariances of periodically correlated processes with application to period determination. Journal of Time Series Analysis, 9(4):411–417, July 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tigelaar:1985:INL

[Tig85] Harry H. Tigelaar. Identification of noisy linear systems with multiple ARMA inputs. Journal of Time Series Analysis, 6(2):109–115, March 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tjostheim:2023:SRT

[TJL23] Dag Tjøstheim, Martin Jullum, and Anders Løland. Some recent trends in embeddings of time series and dynamic networks. Journal of Time Series Analysis, 44(5-6):686–709, September 2023. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tjostheim:1986:SDS

[Tjø86] Dag Tjøstheim. Some doubly stochastic time series models. Journal of Time Series Analysis, 7(1):51–72, January 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taniguchi:1993:NPA

[TK93] Masanobu Taniguchi and Masao Kondo. Non-parametric approach in time series analysis. Journal of Time Series Analysis, 14(4): 397–408, July 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tanaka:2008:SPA

[TK08] Fuyuhiko Tanaka and Fumiyasu Komaki. A superharmonic prior for the autoregressive process of the second-order. Journal of Time Series Analysis, 29(3):444–452, May 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taniguchi:2020:MCD

[TKOP20] Masanobu Taniguchi, Shogo Kato, Hiroaki Ogata, and Arthur Pewsey. Models for circular data from time series spectra. Journal of Time Series Analysis, 41(6):808–829, November 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Terasvirta:1993:PNN

[TLG93] Timo Teräsvirta, Chien-Fu Lin, and Clive W. J. Granger. Power of the neural network linearity test. Journal of Time Series Analysis, 14(2):209–220, March 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Toloi:1993:SAA

[TM93] Clélia M. C. Toloi and Pedro A. Morettin. Spectral analysis for amplitude-modulated time series. Journal of Time Series Analysis, 14(4):409–432, July 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Terdik:1998:NTL

[TM98] Gy. Terdik and J. Math. A new test of linearity of time series based on the bispectrum. Journal of Time Series Analysis, 19(6):737– 753, November 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tavakoli:2023:FMH

[TNH23] Shahin Tavakoli, Gilles Nisol, and Marc Hallin. Factor models for high-dimensional functional time series II: Estimation and forecasting. Journal of Time Series Analysis, 44(5-6):601–621, September 2023. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tse:2004:SSO

[TNZ04] Y. K. Tse, K. W. Ng, and Xibin Zhang. A small-sample overlapping variance-ratio test. Journal of Time Series Analysis, 25(1):127– 135, January 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tomasek:1987:ASC

[Tom87] Ladislav. Tomásek. Asymptotic simultaneous confidence bands for autoregressive spectral density. Journal of Time Series Analysis, 8 (4):469–477, July 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tong:1981:NMB

[Ton81] H. Tong. A note on a Markov bilinear stochastic process in discrete time. Journal of Time Series Analysis, 2(4):279–284, July 1981. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tong:1982:NUT

[Ton82] H. Tong. A note on using threshold autoregressive models for multistep-ahead prediction of cyclical data. Journal of Time Series Analysis, 3(2):137–140, March 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tjostheim:1982:EIM

[TP82] Dag Tjøstheim and Jostein Paulsen. Empirical identification of multiple time series. Journal of Time Series Analysis, 3(4):265– 282, July 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tjostheim:1985:LSE

[TP85] Dag Tjøstheim and Jostein Paulsen. Least squares estimates and order determination procedures for autoregressive processes with a time dependent variance. Journal of Time Series Analysis, 6(2): 117–133, March 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tarami:2003:MVA

[TP03] B. Tarami and M. Pourahmadi. Multi-variate t autoregressions: innovations, prediction variances and exact likelihood equations. Journal of Time Series Analysis, 24(6):739–754, November 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Trimbur:2006:PHO

[Tri06] Thomas M. Trimbur. Properties of higher order stochastic cycles. Journal of Time Series Analysis, 27(1):1–17, January 2006. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Triantafyllopoulos:2011:RTC

[Tri11] K. Triantafyllopoulos. Real-time covariance estimation for the local level model. Journal of Time Series Analysis, 32(2):93–107, March 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Triantafyllopoulos:2012:MVS

[Tri12] K. Triantafyllopoulos. Multi-variate stochastic volatility modelling using Wishart autoregressive processes. Journal of Time Series Analysis, 33(1):48–60, January 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Trokic:2013:RFI

[Tro13] Mirza Trokić. Regulated fractionally integrated processes. Journal of Time Series Analysis, 34(5):591–601, September 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Truong:1994:STS

[TS94] Young K. Truong and Charles J. Stone. Semiparametric time series regression. Journal of Time Series Analysis, 15(4):405–428, July 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tang:2014:EEP

[TS14] L. Tang and Q. Shao. Efficient estimation for periodic autoregressive coefficients via residuals. Journal of Time Series Analysis, 35 (4):378–389, July 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tsay:1988:NLT

[Tsa88] Ruey S. Tsay. Non-linear time series analysis of blowfly population. Journal of Time Series Analysis, 9(3):247–263, May 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tsay:1989:IMT

[Tsa89] Ruey S. Tsay. Identifying multivariate time series models. Journal of Time Series Analysis, 10(4):357–372, July 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tsay:2007:UDB

[Tsa07] Wen-Jen Tsay. Using difference-based methods for inference in regression with fractionally integrated processes. Journal of Time Series Analysis, 28(6):827–843, November 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Trenkler:2008:TCR

[TSL08] Carsten Trenkler, Pentti Saikkonen, and Helmut Lütkepohl. Testing for the cointegrating rank of a VAR process with level shift and trend break. Journal of Time Series Analysis, 29(2):331–358, March 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tyssedal:1982:APT

[TT82] John S. Tyssedal and Dag Tjøstheim. Autoregressive processes with a time dependent variance. Journal of Time Series Analysis, 3(3):209–217, May 1982. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Teverovsky:1997:TLR

[TT97a] Vadim Teverovsky and Murad Taqqu. Testing for long-range dependence in the presence of shifting means or a slowly declining trend, using a variance-type estimator. Journal of Time Series Analysis, 18(3):279–304, May 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Turkman:1997:EBT

[TT97b] K. F. Turkman and M. A. Amaral Turkman. Extremes of bilinear time series models. Journal of Time Series Analysis, 18(3):305– 319, May 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tse:1999:NDM

[TT99] Y. K. Tse and A. K. C. Tsui. A note on diagnosing multivariate conditional heteroscedasticity models. Journal of Time Series Analysis, 20(6):679–691, November 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tuan:1984:EPA

[Tua84a] Pham Dinh Tuan. The estimation of parameters for autoregressive moving average models. Journal of Time Series Analysis, 5(1): 53–68, January 1984. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tuan:1984:NSS

[Tua84b] Pham Dinh Tuan. A note on some statistics useful in identifying the order of autoregressive moving average model. Journal of Time Series Analysis, 5(4):273–279, July 1984. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tuan:1986:FDA

[Tua86] Pham Dinh Tuan. A frequency domain approach to Lagrange multiplier test for autoregressive moving average models. Journal of Time Series Analysis, 7(1):73–78, January 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tuan:1987:EML

[Tua87] Pham Dinh Tuan. Exact maximum likelihood estimate and Lagrange multiplier test statistic for ARMA models. Journal of Time Series Analysis, 8(1):61–78, January 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tuan:1988:EAP

[Tua88] Pham Dinh Tuan. Estimation of autoregressive parameters and order selection for ARMA models. Journal of Time Series Analysis, 9(3):265–279, May 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tuan:1992:ADP

[Tua92] Dinh Pham Tuan. Approximate distribution of parameter estimators for first-order autoregressive models. Journal of Time Series Analysis, 13(2):147–170, March 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Turkman:1998:BR

[Tur98] K. F. Turkman. Book review. Journal of Time Series Analysis, 19(1):125–126, January 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Turkman:2011:BRI

[Tur11] Maria Antonia Amaral Turkman. Book review: Introduction to Time Series Modeling. Journal of Time Series Analysis, 32(3): 336, May 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Turkman:2012:BR

[Tur12] K. F. Turkman. Book review. Journal of Time Series Analysis, 33(6):964, November 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Turkman:2018:BRS

[Tur18] Maria Antónia Amaral Turkman. Book review: Statistical Intervals: A Guide for Practitioners and Researchers, Second Edition, by William Q. Meeker, Gerald J. Hahn, and Louis A. Escobar. Wiley Series in Probability and Statistics, Published by John Wiley and Sons, 2017. Total number of pages: 35 + 592. ISBN: 978-0-4716-8717-7. Journal of Time Series Analysis, 39(4):634–635, July 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Truong-Van:1983:GSA

[TV83] B. Truong-Van. Generalized seasonal ARIMA processes: Regularity/singularity criteria and linear prediction. Journal of Time Series Analysis, 4(2):111–126, March 1983. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Truong-van:2002:ALS

[TvV02] B. Truong-van and P. Varachaud. Asymptotic laws of successive least squares estimates for seasonal ARIMA models and application. Journal of Time Series Analysis, 23(6):707–731, November 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tunnicliffe-Wilson:1989:NLN

[TW89] G. Tunnicliffe-Wilson. Non-linear and non-stationary time series analysis. Journal of Time Series Analysis, 10(4):385–386, July 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Teles:2002:UAT

[TW02] Paulo Teles and William W. S. Wei. The use of aggregate time series in testing for Gaussianity. Journal of Time Series Analysis, 23(1):95–116, January 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tiku:2000:TSM

[TWVB00] M. L. Tiku, Wing-Keung Wong, David C. Vaughan, and Guorui Bian. Time series models in non-normal situations: symmetric innovations. Journal of Time Series Analysis, 21(5):571–596, September 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tschernig:2000:NLS

[TY00] Rolf Tschernig and Lijian Yang. Nonparametric lag selection for time series. Journal of Time Series Analysis, 21(4):457–487, July 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tang:2010:ATR

[TY10] Qi Tang and Danni Yan. Autoregressive trending risk function and exhaustion in random asset price movement. Journal of Time Series Analysis, 31(6):465–470, November 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tse:2002:VRT

[TZ02] Y. K. Tse and X. B. Zhang. The variance ratio test with stable Paretian errors. Journal of Time Series Analysis, 23(1):117–126, January 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Taqqu:2019:SNS

[TZ19] Murad S. Taqqu and Ting Zhang. A self-normalized semiparametric test to detect changes in the long memory parameter. Journal of Time Series Analysis, 40(4):411–424, July 2019. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Tan:2022:ALD

[TZ22] Songhua Tan and Qianqian Zhu. Asymmetric linear double autoregression. Journal of Time Series Analysis, 43(3):371–388, May 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ursu:2009:MDC

[UD09] Eugen Ursu and Pierre Duchesne. On modelling and diagnostic checking of vector periodic autoregressive time series models. Journal of Time Series Analysis, 30(1):70–96, January 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Usami:1995:ECT

[UH95] Yoshihiro Usami and Mituaki Huzii. Estimation of coefficients of time series regression with a nonstationary error process. Journal of Time Series Analysis, 16(1):105–118, January 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ula:1993:FMP

[Ula93] Taylan A. Ula. Forecasting of multivariate periodic autoregressive moving-average processes. Journal of Time Series Analysis, 14 (6):645–657, November 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Unnikrishnan:2004:BSM

[Unn04] N. K. Unnikrishnan. Bayesian subset model selection for time series. Journal of Time Series Analysis, 25(5):671–690, September 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ullah:1983:ELR

[USMS83] A. Ullah, V. K. Srivastava, L. Magee, and A. Srivastava. Estimation of linear regression model with autocorrelated disturbances. Journal of Time Series Analysis, 4(2):127–135, March 1983. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ursu:2012:PAM

[UT12] Eugen Ursu and Kamil Feridun Turkman. Periodic autoregressive model identification using genetic algorithms. Journal of Time Series Analysis, 33(3):398–405, May 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Vermaak:2004:RJM

[VADG04] J. Vermaak, C. Andrieu, A. Doucet, and S. J. Godsill. Reversible jump Markov chain Monte Carlo strategies for Bayesian model selection in autoregressive processes. Journal of Time Series Analysis, 25(6):785–809, November 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Viano:1995:LRD

[VDO95] M. C. Viano, Cl. Deniau, and G. Oppenheim. Long-range dependence and mixing for discrete time fractional processes. Journal of Time Series Analysis, 16(3):323–338, May 1995. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Velilla:1994:GFT

[Vel94] Santiago Velilla. A goodness-of-fit test for autoregressive movingaverage models based on the standardized sample spectral distribution of the residuals. Journal of Time Series Analysis, 15(6): 637–647, November 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Velasco:1999:GSE

[Vel99] Carlos Velasco. Gaussian semiparametric estimation of nonstationary time series. Journal of Time Series Analysis, 20(1):

87–127, January 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Velasco:2000:LCV

[Vel00] Carlos Velasco. Local cross-validation for spectrum bandwidth choice. Journal of Time Series Analysis, 21(3):329–361, May 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Velasco:2003:GSP

[Vel03] Carlos Velasco. Gaussian semi-parametric estimation of fractional cointegration. Journal of Time Series Analysis, 24(3):345–378, May 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Velasco:2007:PFP

[Vel07] Carlos Velasco. The periodogram of fractional processes. Journal of Time Series Analysis, 28(4):600–627, July 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Veres:1987:ADL

[Ver87] Sándor Veres. Asymptotic distributions of likelihood ratios for overparametrized ARMA processes. Journal of Time Series Analysis, 8(3):345–357, May 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

vanGarderen:1999:EGA

[vG99] Kees Jan van Garderen. Exact geometry of autoregressive models. Journal of Time Series Analysis, 20(1):1–21, January 1999. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Vidoni:2004:IPI

[Vid04] Paolo Vidoni. Improved prediction intervals for stochastic process models. Journal of Time Series Analysis, 25(1):137–154, January 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Vidoni:2009:SPC

[Vid09] Paolo Vidoni. A simple procedure for computing improved prediction intervals for autoregressive models. Journal of Time Series Analysis, 30(6):577–590, November 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Vijverberg:2006:TDC

[Vij06] Chu-Ping C. Vijverberg. Time deformation, continuous Euler processes and forecasting. Journal of Time Series Analysis, 27(6): 811–829, November 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Villani:2001:FBL

[Vil01] Mattias Villani. Fractional Bayesian lag length inference in multivariate autoregressive processes. Journal of Time Series Analysis, 22(1):67–86, January 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

VonSachs:2000:WBT

[VN00] Rainer Von Sachs and Michael H. Neumann. A wavelet-based test for stationarity. Journal of Time Series Analysis, 21(5):597–613, September 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Vogelsang:2017:EIL

[VN17] Timothy J. Vogelsang and Nasreen Nawaz. Estimation and inference of linear trend slope ratios with an application to global temperature data. Journal of Time Series Analysis, 38(5):640– 667, September 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Vogelsang:1999:TSP

[Vog99] Timothy J. Vogelsang. Two simple procedures for testing for a unit root when there are additive outliers. Journal of Time Series Analysis, 20(2):237–252, March 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Vollenbroker:2012:SSS

[Vol12] Bernd Vollenbröker. Strictly stationary solutions of ARMA equations with fractional noise. Journal of Time Series Analysis, 33(4): 570–582, July 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Valk:2012:TSC

[VP12] Marcio Valk and Aluísio Pinheiro. Time-series clustering via quasi U-statistics. Journal of Time Series Analysis, 33(4):608–619, July 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Vaillant:2011:STA

[VPWD11] Jean Vaillant, Gavino Puggioni, Lance A. Waller, and Jean Daugrois. A spatio-temporal analysis of the spread of sugarcane yellow leaf virus. Journal of Time Series Analysis, 32(4):396–406, July 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

vonSachs:1994:PIN

[vS94] Rainer von Sachs. Peak-insensitive non-parametric spectrum estimation. Journal of Time Series Analysis, 15(4):429–452, July 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

VanHecke:2018:FAS

[VVD18] Ria Van Hecke, Stanislav Volgushev, and Holger Dette. Fourier analysis of serial dependence measures. Journal of Time Series Analysis, 39(1):75–89, January 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Velasco:2015:JPT

[VW15] Carlos Velasco and Xuexin Wang. A joint portmanteau test for conditional mean and variance time-series models. Journal of Time Series Analysis, 36(1):39–60, January 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Velu:1987:NNS

[VWR87] Raja P. Velu, Dean W. Wichern, and Gregory C. Reinsel. A note on non-stationarity and canonical analysis of multiple time series models. Journal of Time Series Analysis, 8(4):479–487, July 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Vaninskii:1990:SPF

[VY90] K. L. Vaninskii and A. M. Yaglom. Stationary processes with a finite number of non-zero canonical correlations between future and past. Journal of Time Series Analysis, 11(4):361–375, July 1990. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Vogelsang:2016:ENU

[VY16] Timothy J. Vogelsang and Jingjing Yang. Exactly/ nearly unbiased estimation of autocovariances of a univariate time series with unknown mean. Journal of Time Series Analysis, 37(6):723–740, November 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wahlberg:1989:EAM

[Wah89] Bo Wahlberg. Estimation of autoregressive moving-average models via high-order autoregressive approximations. Journal of Time Series Analysis, 10(3):283–299, May 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wall:1987:ITV

[Wal87a] Kent D. Wall. Identification theory for varying coefficient regression models. Journal of Time Series Analysis, 8(3):359–371, May 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wallis:1987:TSA

[Wal87b] Kenneth F. Wallis. Time series analysis of bounded economic variables. Journal of Time Series Analysis, 8(1):115–123, January 1987. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Walker:1995:RPC

[Wal95] A. M. Walker. On results of Porat concerning asymptotic efficiency of sample covariances of Gaussian ARMA processes. Journal of Time Series Analysis, 16(2):237–248, March 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Walker:2000:SRC

[Wal00] A. M. Walker. Some results concerning the asymptotic distribution of sample Fourier transforms and periodograms for a discrete-time stationary process with a continuous spectrum. Journal of Time Series Analysis, 21(1):95–109, January 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Walker:2003:NEL

[Wal03] A. M. Walker. A note on estimation by least squares for harmonic component models. Journal of Time Series Analysis, 24(5):613– 629, September 2003. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wang:1993:ATE

[Wan93a] Xiaobao Wang. An AIC type estimator for the number of cosinusoids. Journal of Time Series Analysis, 14(4):433–440, July 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wang:1993:NSF

[Wan93b] Xiaobao Wang. Non-singularity of Fisher information for autoregressive moving-average processes. Journal of Time Series Analysis, 14(5):547–548, September 1993. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wang:2005:PES

[Wan05] Hai-Bin Wang. Parameter estimation and subset selection for separable lower triangular bilinear models. Journal of Time Series Analysis, 26(5):743–757, September 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wang:2008:NAM

[Wan08] Hai-Bin Wang. Nonlinear ARMA models with functional MA coefficients. Journal of Time Series Analysis, 29(6):1032–1056, November 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wang:2009:SIC

[Wan09] Xiao Wang. Semiparametric inference on a class of Wiener processes. Journal of Time Series Analysis, 30(2):179–207, March 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wang:2016:UMI

[Wan16] Fangfang Wang. An unbiased measure of integrated volatility in the frequency domain. Journal of Time Series Analysis, 37(2): 147–164, March 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wang:2021:BRS

[Wan21] Weining Wang. Book review: Statistical foundations of data science, by Jianqing Fan, Runze Li, Chun-Hui Zhang, Hui Zou. Published by Taylor and Francis Group. Total number of pages: 729. ISBN: 978-1-466-51084-5. Journal of Time Series Analysis, 42(3): 372–373, May 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wang:2024:NKM

[Wan24] Tao Wang. Nonlinear kernel mode-based regression for dependent data. Journal of Time Series Analysis, 45(2):189–213, March 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Watanabe:1985:NKF

[Wat85] N. Watanabe. Note on the Kalman filter with estimated parameters. Journal of Time Series Analysis, 6(4):269–278, July 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wang:2010:WCP

[WC10] Lihong Wang and Haiyan Cai. Wavelet change-point estimation for long memory non-parametric random design models. Journal of Time Series Analysis, 31(2):86–97, March 2010. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wu:2014:PDL

[WC14] Rongning Wu and Yunwei Cui. A parameter-driven logit regression model for binary time series. Journal of Time Series Analysis, 35 (5):462–477, September 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Woodward:1998:FGL

[WCG98] Wayne A. Woodward, Q. C. Cheng, and H. L. Gray. A k-factor GARMA long-memory model. Journal of Time Series Analysis, 19(4):485–504, July 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wei:2012:SBN

[WCK12] Lai Wei, Peter F. Craigmile, and Wayne M. King. Spectral-based non-central F mixed effect models, with application to otoacoustic emissions. Journal of Time Series Analysis, 33(5):850–862, September 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wu:2010:LAD

[WD10] Rongning Wu and Richard A. Davis. Least absolute deviation estimation for general autoregressive moving average time-series models. Journal of Time Series Analysis, 31(2):98–112, March 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Westerlund:2007:NIT

[WE07] Joakim Westerlund and David L. Edgerton. New improved tests for cointegration with structural breaks. Journal of Time Series Analysis, 28(2):188–224, March 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Weiss:1984:AMA

[Wei84] Andrew A. Weiss. Arma models with arch errors. Journal of Time Series Analysis, 5(2):129–143, March 1984. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Weiss:1985:SPC

[Wei85] Andrew A. Weiss. The stability of the AR(1) process with an AR(1) coefficient. Journal of Time Series Analysis, 6(3):181–186, May 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Weis:1986:SHP

[Wei86] Andrew A. Weis. On the stability of a heteroscedastic process. Journal of Time Series Analysis, 7(4):303–310, July 1986. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Westerlund:2013:CCU

[Wes13] Joakim Westerlund. A computationally convenient unit root test with covariates, conditional heteroskedasticity and efficient detrending. Journal of Time Series Analysis, 34(4):477–495, July 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Westerlund:2019:CBM

[Wes19a] Joakim Westerlund. Common breaks in means for cross-correlated fixed-T panel data. Journal of Time Series Analysis, 40(2):248– 255, March 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Westerlund:2019:EIH

[Wes19b] Joakim Westerlund. On estimation and inference in heterogeneous panel regressions with interactive effects. Journal of Time Series Analysis, 40(5):852–857, September 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wikle:2011:PNS

[WH11] Christopher K. Wikle and Scott H. Holan. Polynomial nonlinear spatio-temporal integro-difference equation models. Journal of Time Series Analysis, 32(4):339–350, July 2011. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wikle:2019:RAS

[WH19] Christopher K. Wikle and Scott H. Holan. Recent advances in spatio-temporal methodology. Journal of Time Series Analysis, 40(3):267–268, May 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Whittaker:1985:AEA

[Whi85] Joe Whittaker. Additive elements of ARMA models. Journal of Time Series Analysis, 6(2):135–140, March 1985. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Whittle:2002:ETF

[Whi02] P. Whittle. The estimation and tracking of frequency. Journal of Time Series Analysis, 23(5):627–628, September 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Whittle:2005:BRE

[Whi05] P. Whittle. Book review: The Estimation and Tracking of Frequency. Journal of Time Series Analysis, 26(4):625–626, July 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wu:2022:MEP

[WHY22] Yanfeng Wu, Jianqiang Hu, and Xiangyu Yang. Moment estimators for parameters of Lévy-driven Ornstein–Uhlenbeck processes. Journal of Time Series Analysis, 43(4):610–639, July 2022. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wied:2013:CTT

[Wie13] Dominik Wied. CUSUM-type testing for changing parameters in a spatial autoregressive model for stock returns. Journal of Time Series Analysis, 34(2):221–229, March 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wilson:2016:BRT

[Wil16] Granville Tunnicliffe Wilson. Book review: Time Series Analysis: Forecasting and Control, 5th Edition, by George E. P. Box, Gwilym M. Jenkins, Gregory C. Reinsel and Greta M. Ljung, 2015. Published by John Wiley and Sons Inc., Hoboken, New Jersey, pp. 712. ISBN: 978-1-118-67502-1. Journal of Time Series Analysis, 37(5): 709–711, September 2016. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wilson:2017:SEM

[Wil17] Granville Tunnicliffe Wilson. Spectral estimation of the multivariate impulse response. Journal of Time Series Analysis, 38(2): 381–391, March 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wilson:2018:TSR

[Wil18] Granville Tunnicliffe Wilson. Tata Subba Rao, 1942–2018. Journal of Time Series Analysis, 39(5):640, September 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wang:2011:LSD

[WJM11] Cheng Wang, Baisuo Jin, and Baiqi Miao. On limiting spectral distribution of large sample covariance matrices by VARMA(p, q). Journal of Time Series Analysis, 32(5):539–546, September 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wong:1998:NCA

[WL98] C. S. Wong and W. K. Li. A note on the corrected Akaike information criterion for threshold autoregressive models. Journal of Time Series Analysis, 19(1):113–124, January 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wong:2005:MPT

[WL05] Heung Wong and Shiqing Ling. Mixed portmanteau tests for timeseries models. Journal of Time Series Analysis, 26(4):569–579, July 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wang:2011:AFA

[WL11] Chao Wang and Wai Keung Li. On the autopersistence functions and the autopersistence graphs of binary autoregressive time series. Journal of Time Series Analysis, 32(6):639–646, November 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wang:2012:SSM

[WLC12] Jiabin Wang, Hua Liang, and Rong Chen. A state space model approach for HIV infection dynamics. Journal of Time Series Analysis, 33(5):841–849, September 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Westerlund:2022:FAA

[WNS22] Joakim Westerlund, Milda Norkute, and Ovidijus Stauskas. The factor analytical approach in trending near unit root panels. Journal of Time Series Analysis, 43(3):501–508, May 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wong:1997:FDT

[Won97] Woon Wong. Frequency domain tests of multivariate Gaussianity and linearity. Journal of Time Series Analysis, 18(2):181–194, March 1997. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Weiss:2014:BAP

[WP14] Christian H. Weiß and Philip K. Pollett. Binomial autoregressive processes with density-dependent thinning. Journal of Time Series Analysis, 35(2):115–132, March 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wang:2021:CAS

[WP21] Jiang Wang and Dimitris N. Politis. Consistent autoregressive spectral estimates: Nonlinear time series and large autocovariance matrices. Journal of Time Series Analysis, 42(5-6):580–596, September 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wilson:2008:SPC

[WR08] Granville Tunnicliffe Wilson and Marco Reale. The sampling properties of conditional independence graphs for I(1) structural VAR models. Journal of Time Series Analysis, 29(5):802–810, September 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wright:1995:SOM

[Wri95] J. H. Wright. Stochastic orders of magnitude associated with twostage estimators of fractional ARIMA systems. Journal of Time Series Analysis, 16(1):119–125, January 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wright:1998:TSB

[Wri98] Jonathan H. Wright. Testing for a structural break at unknown date with long-memory disturbances. Journal of Time Series Analysis, 19(3):369–376, May 1998. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wall:2002:SSA

[WS02] Kent D. Wall and David S. Stoffer. A state space approach to bootstrapping conditional forecasts in ARMA models. Journal of Time Series Analysis, 23(6):733–751, November 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wahlberg:2011:LSH

[WS11] Patrik Wahlberg and Peter J. Schreier. Locally stationary harmonizable complex improper stochastic processes. Journal of Time Series Analysis, 32(1):33–46, January 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wang:2020:ATU

[WS20] Xuexin Wang and Yixiao Sun. An asymptotic F test for uncorrelatedness in the presence of time series dependence. Journal of Time Series Analysis, 41(4):536–550, July 2020. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wang:2004:ETP

[WSS04] Dehui Wang, Lixin Song, and Ningzhong Shi. Estimation and testing for the parameters of ARCH(q) under ordered restriction. Journal of Time Series Analysis, 25(4):483–499, July 2004. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Willekens:1988:SSP

[WT88] E. Willekens and J. L. Teugels. Subordination of stationary processes. Journal of Time Series Analysis, 9(3):281–299, May 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wang:2019:CMT

[WT19] Yongning Wang and Ruey S. Tsay. Clustering multiple time series with structural breaks. Journal of Time Series Analysis, 40(2): 182–202, March 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wong:2017:NMN

[WTSL17] Shiu Fung Wong, Howell Tong, Tak Kuen Siu, and Zudi Lu. A new multivariate nonlinear time series model for portfolio risk measurement: The threshold copula-based TAR approach. Journal of Time Series Analysis, 38(2):243–265, March 2017. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wolf:2015:BJP

[WW15] Michael Wolf and Dan Wunderli. Bootstrap joint prediction regions. Journal of Time Series Analysis, 36(3):352–376, May 2015. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wagner:2017:CMC

[WW17] Martin Wagner and Dominik Wied. Consistent monitoring of cointegrating relationships: The US housing market and the subprime crisis. Journal of Time Series Analysis, 38(6):960–980, November 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wang:2009:AKF

[WWG09] Zhu Wang, Wayne A. Woodward, and Henry L. Gray. The application of the Kalman filter to nonstationary time series through time deformation. Journal of Time Series Analysis, 30(5):559– 574, September 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wen:2012:OGP

[WWW12] Qiuzi H. Wen, Augustine Wong, and Xiaolan L. Wang. Overlapped grouping periodogram test for detecting multiple hidden periodicities in mixed spectra. Journal of Time Series Analysis, 33(2): 255–268, March 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wu:2018:PTC

[WX18] Jilin Wu and Zhijie Xiao. A powerful test for changing trends in time series models. Journal of Time Series Analysis, 39(4): 488–501, July 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Wylomanska:2008:SMP

[Wyl08] Agnieszka Wylomańska. Spectral measures of PARMA sequences. Journal of Time Series Analysis, 29(1):1–13, January 2008. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Xia:1999:PPA

[XA99] Xingcun Xia and H. Z. An. Projection pursuit autoregression in time series. Journal of Time Series Analysis, 20(6):693–714, November 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Xia:2024:IHD

[XCG24] Jiaqi Xia, Yu Chen, and Xiao Guo. Inference for high-dimensional linear models with locally stationary error processes. Journal of Time Series Analysis, 45(1):78–102, January 2024. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Xia:2017:MAT

[XHN17] Qiang Xia, Kejun He, and Cuizhen Niu. A model-adaptive test for parametric single-index time series models. Journal of Time Series Analysis, 38(6):981–999, November 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Xiao:2001:TNH

[Xia01] Zhijie Xiao. Testing the null hypothesis of stationarity against an autoregressive unit root alternative. Journal of Time Series Analysis, 22(1):87–105, January 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Xiao:2002:NPC

[XL02] Zhijie Xiao and Oliver Linton. A nonparametric prewhitened covariance estimator. Journal of Time Series Analysis, 23(2):215– 250, March 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Xu:2023:HOA

[XLT23] Xiaofei Xu, Yan Liu, and Masanobu Taniguchi. Higher-order asymptotics of minimax estimators for time series. Journal of Time Series Analysis, 44(2):247–257, March 2023. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Xia:2010:BNT

[XPZL10] Qiang Xia, Jiazhu Pan, Zhiqiang Zhang, and Jinshan Liu. A Bayesian nonlinearity test for threshold moving average models. Journal of Time Series Analysis, 31(5):329–336, September 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Xie:2018:SRL

[XX18] Fang Xie and Zhijie Xiao. Square-root LASSO for high-dimensional sparse linear systems with weakly dependent errors. Journal of Time Series Analysis, 39(2):212–238, March 2018. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Xinjian:1989:SMN

[XY89] Gu Xinjian and Huang Yiyun. A simulation method for non-normal random processes. Journal of Time Series Analysis, 10(4):373–374, July 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Xu:2022:NGG

[XZ22] Yue Xu and Fukang Zhu. A new GJR–GARCH model for Z-valued time series. Journal of Time Series Analysis, 43(3):490–500, May 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Xia:2020:PTS

[XZL20] Qiang Xia, Zhiqiang Zhang, and Wai Keung Li. A portmanteau test for smooth transition autoregressive models. Journal of Time Series Analysis, 41(5):722–730, September 2020. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yabe:2012:LDS

[Yab12] Ryota Yabe. Limiting distribution of the score statistic under moderate deviation from a unit root in MA(1). Journal of Time Series Analysis, 33(4):533–541, July 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yajima:1985:APS

[Yaj85] Yoshihiro Yajima. Asymptotic properties of the sample autocorrelations and partial autocorrelations of a multiplicative ARIMA process. Journal of Time Series Analysis, 6(3):187–201, May 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yajima:1989:CLT

[Yaj89] Yoshihiro Yajima. A central limit theorem of Fourier transforms of strongly dependent stationary processes. Journal of Time Series Analysis, 10(4):375–383, July 1989. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yakowitz:1987:NNM

[Yak87] S. Yakowitz. Nearest-neighbour methods for time series analysis. Journal of Time Series Analysis, 8(2):235–247, March 1987. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yamaguchi:2011:ECP

[Yam11] Keiko Yamaguchi. Estimating a change point in the long memory parameter. Journal of Time Series Analysis, 32(3):304–314, May 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yao:2020:BRL

[Yao20] Jianfeng Yao. Book review: Large Covariance and Autocovariance Matrices, By Arup Bose and Monika Bhattacharjee. Published by Taylor and Francis Group, LLC, Boca Raton, London, New York, 2019. ISBN: 978-1-138-30386-7 (hardback). Journal of Time Series Analysis, 41(1):173–174, January 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yau:2012:ELL

[Yau12] Chun Yip Yau. Empirical likelihood in long-memory time series models. Journal of Time Series Analysis, 33(2):269–275, March 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yao:2006:GML

[YB06] Qiwei Yao and Peter J. Brockwell. Gaussian maximum likelihood estimation for ARMA models. I. Time series. Journal of Time Series Analysis, 27(6):857–875, November 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yau:2012:LID

[YD12] Chun Yip Yau and Richard A. Davis. Likelihood inference for discriminating between long-memory and change-point models. Journal of Time Series Analysis, 33(4):649–664, July 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yu:2014:NPE

[YFL^{+}14] Chao Yu, Yue Fang, Zeng Li, Bo Zhang, and Xujie Zhao. Nonparametric estimation of high-frequency spot volatility for Brownian semimartingale with jumps. Journal of Time Series Analysis, 35(6):572–591, November 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yang:1999:NAM

[YHN99] Lijian Yang, Wolfgang Härdle, and Jens Nielsen. Nonparametric autoregression with multiplicative volatility and additive mean. Journal of Time Series Analysis, 20(5):579–604, September 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yamamoto:2006:TLR

[YK06] Taku Yamamoto and Eiji Kurozumi. Tests for long-run Granger non-causality in cointegrated systems. Journal of Time Series Analysis, 27(5):703–723, September 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yu:1991:MSS

[YL91] Gwo-Hsing Yu and Yow-Chang Lin. A methodology for selecting subset autoregressive time series models. Journal of Time Series Analysis, 12(4):363–373, July 1991. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yang:2020:SWL

[YL20] Yaxing Yang and Dong Li. Self-weighted lad-based inference for heavy-tailed continuous threshold autoregressive models. Journal of Time Series Analysis, 41(1):163–172, January 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yang:2022:EER

[YL22] Xuanling Yang and Dong Li. Estimation of the empirical riskreturn relation: a generalized-risk-in-mean model. Journal of Time Series Analysis, 43(6):938–963, November 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yu:2012:NMS

[YLC12] Shu-Hui Yu, Chien-Chih Lin, and Hung-Wen Cheng. A note on mean squared prediction error under the unit root model with deterministic trend. Journal of Time Series Analysis, 33(2):276–286, March 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yang:2021:TMT

[YLC21] Lixiong Yang, Chingnun Lee, and I-Po Chen. Threshold model with a time-varying threshold based on Fourier approximation. Journal of Time Series Analysis, 42(4):406–430, July 2021. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Young:2006:EOB

[YP06] K. D. S. Young and L. I. Pettit. The effect of observations on Bayesian choice of an autoregressive model. Journal of Time Series Analysis, 27(1):41–50, January 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yuan:1992:CTU

[YR92] J. Yuan and T. Subba Rao. Classification of textures using secondorder spectra. Journal of Time Series Analysis, 13(6):547–562, November 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yap:1995:RET

[YR95] Sook Fwe Yap and Gregory C. Reinsel. Results on estimation and testing for a unit root in the nonstationary autoregressive movingaverage model. Journal of Time Series Analysis, 16(3):339–353, May 1995. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yuan:2022:NVM

[YSX^{+}22] Huiling Yuan, Yulei Sun, Lu Xu, Yong Zhou, and Xiangyu Cui. A new volatility model: GQARCH–Itô model. Journal of Time Series Analysis, 43(3):345–370, May 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yu:2007:HMP

[Yu07] Hao Yu. High moment partial sum processes of residuals in ARMA models and their applications. Journal of Time Series Analysis, 28(1):72–91, January 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yuan:2000:TGL

[Yua00a] J. Yuan. Testing Gaussianity and linearity for random fields in the frequency domain. Journal of Time Series Analysis, 21(6):723– 737, November 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Yuan:2000:TLS

[Yua00b] J. Yuan. Testing linearity for stationary time series using the sample interquartile range. Journal of Time Series Analysis, 21(6): 713–722, November 2000. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zaffaroni:2007:CAG

[Zaf07] Paolo Zaffaroni. Contemporaneous aggregation of GARCH processes. Journal of Time Series Analysis, 28(4):521–544, July 2007. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zaffaroni:2008:LSV

[Zaf08] Paolo Zaffaroni. Large-scale volatility models: theoretical properties of professionals’ practice. Journal of Time Series Analysis, 29(3):581–599, May 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zarepour:2002:NMA

[ZB02] M. Zarepour and D. Banjevic. A note on maximum autoregressive processes of order one. Journal of Time Series Analysis, 23(5):619– 626, September 2002. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhou:2005:MLE

[ZB05] J. Zhou and I. V. Basawa. Maximum likelihood estimation for a first-order bifurcating autoregressive process with exponential errors. Journal of Time Series Analysis, 26(6):825–842, November 2005. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zheng:2006:IPT

[ZBD06] Haitao Zheng, Ishwar V. Basawa, and Somnath Datta. Inference for p th-order random coefficient integer-valued autoregressive processes. Journal of Time Series Analysis, 27(3):411–440, May 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:2012:MLE

[ZC12] Rong-Mao Zhang and Ngai Hang Chan. Maximum likelihood estimation for nearly non-stationary stable autoregressive processes.

Journal of Time Series Analysis, 33(4):542–553, July 2012. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zadrozny:2019:WCF

[ZC19a] Peter A. Zadrozny and Baoline Chen. Weighted-covariance factor decomposition of Varma models applied to forecasting quarterly U.S. real GDP at monthly intervals. Journal of Time Series Analysis, 40(6):968–986, November 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:2019:ESC

[ZC19b] Bohai Zhang and Noel Cressie. Estimating spatial changes over time of Arctic sea ice using hidden 2×2 tables. Journal of Time Series Analysis, 40(3):288–311, May 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhao-Guo:1985:AEL

[ZG85] Chen Zhao-Guo. The asymptotic efficiency of a linear procedure of estimation for ARMA models. Journal of Time Series Analysis, 6(1):53–62, January 1985. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhao-Guo:1988:ACP

[ZG88] Chen Zhao-Guo. An alternative consistent procedure for detecting hidden frequencies. Journal of Time Series Analysis, 9(3):301–317, May 1988. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhao-Guo:1980:DPO

[ZGH80] Chen Zhao-Guo and E. J. Hannan. The distribution of periodogram ordinates. Journal of Time Series Analysis, 1(1):73–82, January 1980. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:1992:RAB

[Zha92] H.-C. Zhang. Reduction of the asymptotic bias of autoregressive and spectral estimators by tapering. Journal of Time Series Analysis, 13(5):451–469, September 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:2004:ALI

[Zha04] Xibin Zhang. Assessment of local influence in GARCH processes. Journal of Time Series Analysis, 25(2):301–313, March 2004. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:2013:BRS

[Zha13] Tusheng Zhang. Book review: Statistical Methods for Stochastic Differential Equations. Edited By, Mathieu Kessler, Alexander Lindner and Michael Sørensen. Publishers CRC Press, Taylor and Francis Group. London, ISBN 978-1-4398-4940-8. 483 Pages. Journal of Time Series Analysis, 34(3):422, May 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:2024:AAM

[Zha24] Hong-Fan Zhang. Additive autoregressive models for matrix valued time series. Journal of Time Series Analysis, 45(3):398–420, May 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zamani:2022:SFA

[ZHHH22] Atefeh Zamani, Hossein Haghbin, Maryam Hashemi, and Rob J. Hyndman. Seasonal functional autoregressive models. Journal of Time Series Analysis, 43(2):197–218, March 2022. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhou:1992:AEP

[Zho92] Guofu Zhou. Algorithms for estimation of possibly nonstationary vector time series. Journal of Time Series Analysis, 13(2):171–188, March 1992. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhou:2012:MND

[Zho12] Zhou Zhou. Measuring nonlinear dependence in time-series, a distance correlation approach. Journal of Time Series Analysis, 33(3): 438–457, May 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhou:2013:INS

[Zho13] Zhou Zhou. Inference for non-stationary time-series autoregression. Journal of Time Series Analysis, 34(4):508–516, July 2013. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhu:2011:NBI

[Zhu11] Fukang Zhu. A negative binomial integer-valued GARCH model. Journal of Time Series Analysis, 32(1):54–67, January 2011. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhu:2013:MPT

[Zhu13] Ke. Zhu. A mixed portmanteau test for ARMA-GARCH models by the quasi-maximum exponential likelihood estimation approach. Journal of Time Series Analysis, 34(2):230–237, March 2013. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zielinski:1999:MUE

[Zie99] Ryszard Zielinski. A median-unbiased estimator of the AR(1) coefficient. Journal of Time Series Analysis, 20(4):477–481, July 1999. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhu:2006:MCD

[ZJ06] Rong Zhu and Harry Joe. Modelling count data time series with Markov processes based on binomial thinning. Journal of Time Series Analysis, 27(5):725–738, September 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:2024:MCV

[ZJN24] Lin Zhang, Harry Joe, and Natalia Nolde. Margin-closed vector autoregressive time series models. Journal of Time Series Analysis, 45(2):269–297, March 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zambom:2022:VLM

[ZKG22] Adriano Zanin Zambom, Seonjin Kim, and Nancy Lopes Garcia. Variable length Markov chain with exogenous covariates. Journal of Time Series Analysis, 43(2):312–328, March 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:2022:WAM

[ZKP22] Chao Zhang, Piotr Kokoszka, and Alexander Petersen. Wasserstein autoregressive models for density time series. Journal of Time Series Analysis, 43(1):30–52, January 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:2012:LTG

[ZL12a] Rong-Mao Zhang and Zheng-Yan Lin. Limit theory for a general class of GARCH models with just barely infinite variance. Journal of Time Series Analysis, 33(1):161–174, January 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhu:2012:LRT

[ZL12b] Ke Zhu and Shiqing Ling. Likelihood ratio tests for the structural change of an AR(p) model to a threshold AR(p) model. Journal of Time Series Analysis, 33(2):223–232, March 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:2024:STM

[ZL24a] Xinyu Zhang and Dong Li. Smooth transition moving average models: Estimation, testing, and computation. Journal of Time Series Analysis, 45(3):463–478, May 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhou:2024:SJ

[ZL24b] Weilian Zhou and Soumendra Lahiri. Stationary jackknife. Journal of Time Series Analysis, 45(3):333–360, May 2024. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:2019:ABO

[ZLP19] Ting Zhang, Liliya Lavitas, and Qiao Pan. Asymptotic behavior of optimal weighting in generalized self-normalization for time series. Journal of Time Series Analysis, 40(5):831–851, September 2019. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:2020:TSE

[ZLSY20] Yuanyuan Zhang, Rong Liu, Qin Shao, and Lijian Yang. Two-step estimation for time varying arch models. Journal of Time Series Analysis, 41(4):551–570, July 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:2006:MGT

[ZLY06] Zhiqiang Zhang, Wai Keung Li, and Kam Chuen Yuen. On a Mixture GARCH Time-Series Model. Journal of Time Series Analysis, 27(4):577–597, July 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:2001:EHF

[ZM01] Hao Zhang and V. Mandrekar. Estimation of hidden frequencies for 2D stationary processes. Journal of Time Series Analysis, 22 (5):613–629, September 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:2006:CAD

[ZM06] Y. Zhang and A. I. McLeod. Computer algebra derivation of the bias of linear estimators of autoregressive models. Journal of Time Series Analysis, 27(2):157–165, March 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhao:2020:EBP

[ZP20a] Wenjie Zhao and Raquel Prado. Efficient Bayesian PARCOR approaches for dynamic modeling of multivariate time series. Journal of Time Series Analysis, 41(6):759–784, November 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhu:2020:HOA

[ZP20b] Tingyi Zhu and Dimitris N. Politis. Higher-order accurate spectral density estimation of functional time series. Journal of Time Series Analysis, 41(1):3–20, January 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhou:2021:ELT

[ZPZ21] Mo Zhou, Liang Peng, and Rongmao Zhang. Empirical likelihood test for the application of SWQMELE in fitting an ARMA– GARCH model. Journal of Time Series Analysis, 42(2):222–239, March 2021. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:2001:ASM

[ZS01] Jing Zhang and Robert A. Stine. Autocovariance structure of Markov regime switching models and model selection. Journal of Time Series Analysis, 22(1):107–124, January 2001. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhou:2017:LMT

[ZS17] Xingwu Zhou and Martin Solberger. A Lagrange multiplier-type test for idiosyncratic unit roots in the exact factor model. Journal of Time Series Analysis, 38(1):22–50, January 2017. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:2022:NPT

[ZSW22] Erhua Zhang, Xiaojun Song, and Jilin Wu. A non-parametric test for multi-variate trend functions. Journal of Time Series Analysis, 43(6):856–871, November 2022. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:1994:DAS

[ZT94] Guoqiang Zhang and Masanobu Taniguchi. Discriminant analysis for stationary vector time series. Journal of Time Series Analysis, 15(1):117–126, January 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:1997:PMN

[ZT97] Xichuan Zhang and R. Deane Terrell. Projection modulus: a new direction for selecting subset autoregressive models. Journal of Time Series Analysis, 18(2):195–212, March 1997. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhu:2006:ISR

[ZT06] Zhengyuan Zhu and Murad S. Taqqu. Impact of the sampling rate on the estimation of the parameters of fractional Brownian motion. Journal of Time Series Analysis, 27(3):367–380, May 2006. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:2018:TCT

[ZT18] Shibin Zhang and Xin M. Tu. Tests for comparing time-invariant and time-varying spectra based on the Pearson statistic. Journal of Time Series Analysis, 39(5):709–730, September 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:1994:NWE

[ZW94] Hu-Ming Zhang and Ping Wang. A new way to estimate orders in time series. Journal of Time Series Analysis, 15(5):545–559, September 1994. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhu:2008:EPM

[ZW08] Fukang Zhu and Dehui Wang. Estimation of parameters in the NLAR(p) model. Journal of Time Series Analysis, 29(4):619–628, July 2008. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zheng:2012:WSE

[ZW12] Lingyu Zheng and William W. S. Wei. Weighted scatter estimation method of the GO-GARCH models. Journal of Time Series Analysis, 33(1):81–95, January 2012. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhang:2011:ELI

[ZWZ11] Haixiang Zhang, Dehui Wang, and Fukang Zhu. Empirical likelihood inference for random coefficient INAR(p) process. Journal of Time Series Analysis, 32(3):195–203, May 2011. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zheng:2022:GAM

[ZXC22] Tingguo Zheng, Han Xiao, and Rong Chen. Generalized autoregressive moving average models with GARCH errors. Journal of Time Series Analysis, 43(1):125–146, January 2022. CODEN JT-SADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhao:2014:NPE

[ZZL14] Zhibiao Zhao, Yiyun Zhang, and Runze Li. Non-parametric estimation under strong dependence. Journal of Time Series Analysis, 35(1):4–15, January 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Zhu:2020:BIG

[ZZL20] Qianqian Zhu, Ruochen Zeng, and Guodong Li. Bootstrap inference for Garch models by the least absolute deviation estimation. Journal of Time Series Analysis, 41(1):21–40, January 2020. CO-DEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).